A Complete Bibliography of Statistica Neerlandica

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 October 2024 Version 1.00

Title word cross-reference

(s, Q) [1301]. (s, S) [785]. 2^{K} [2162]. 2n + 1 [743]. 2 × 2 [2301, 758, 1201, 1832, 1853, 1552, 2026, 1958, 733, 1749, 269, 393]. 2 × 2 × K [1844, 1552]. ^{1} [298, 1322]. _{1} [410]. a [1247, 1290, 1246, 804]. α [1394, 1902, 1555]. AR(p) [1572, 2061]. B [158]. β [800]. c = 0 [110, 110]. C[0, 1] [1156]. χ^{2} [1046, 1136, 142, 1073, 19, 375]. D [2062]. F [479, 158, 838]. GARCH(p, q) [2218]. h [2169]. ∞ [1583, 503]. J [1590, 696, 1927]. k [1887, 837, 2308, 2211, 801, 2202, 602, 1090, 575, 419]. L [1970, 2016, 1169]. L_{p} [2196]. M [1004, 1361, 1054, 2240, 1913, 1169, 1817, 1425]. N [1121, 992]. N(−1, l) [1024]. N(1, 1) [1024]. n = 12 [805]. n = 13(1)16 [858]. O [2085]. P [608, 1193, 1933, 2274, 1728, 1702, 1727, 2282, 1791]. P (X_{t} > Y_{t}) [1911]. p_{2} [2201, 1705]. ϕ [2210, 1516]. π [1066]. PINAR(p) [2214]. ∝ [1226]. R^{2} [1345]. R_{0} [1455]. R × C [1844, 1609, 787]. S [1499, 858]. S∗ [1502]. T [1137, 1424, 1503, 2081, 1735, 2266]. T_{1} [666]. U [1741]. \underline{S} [805]. X_{t} [1911]. Y_{t} [1911].

* [771, 802].

-class [801, 1137]. -content [800]. -correlations [1970]. -coverage [1791]. -dependent [676]. -distributions [2081]. -divergence-based [2210]. -divergences [1516]. -Estimates [1169, 1054]. -estimation [1004]. -estimators [1502, 1361, 1499, 1913, 1425]. -F-test [394]. -F-toets [394]. -fluctuation [1817]. -function [1927]. -likelihood [2169]. -methode [142]. -mixing [1066]. -model [593]. -moments [1970, 2016]. -monotone [1887]. -monotonicity [1290]. -Nearest [2202]. -optimal [2062]. -policy [992]. -proeven [628]. -ranking [1121]. -regularly [2085]. -sample [837, 1090]. -stable [1902]. -stage [696]. -staps [503]. -statistics [1741]. -step [2211, 503]. -tabel [269, 733]. -table [758, 393]. -technieken [925]. -test [1073, 19]. -tests [375]. -toets [479]. -unimodal [1394, 1555]. -Unimodality [1226, 1247, 1246]. -value [2282]. -values [1728, 1702, 1727, 1933]. -variate [2308].

1 [1585, 1013, 1219, 931, 1921, 1759, 1696, 992, 1270]. 10 [394]. 100-m [1918]. 1047 [230]. 13-de [332]. 13th [332]. 19 [2258, 2228, 2209]. 1946 [31, 33]. 1949 [105, 99]. 1951 [179]. 1952/1953 [204]. 1953/1954 [225]. 1957 [280]. 1959 [358].

2 [99, 1270]. 20 [31]. 2011 [1948]. 21 [699]. 26 [1998]. 26e [100]. 26th [100]. 2e [654]. 2nd [654].

3117 [278, 420].

5 [105]. 5-year [1888]. 50-year-old [1768]. 500-m [1910]. 53 [1567]. 56 [1662]. 57 [1689].

60 [522]. 61 [1818]. 62 [1850]. 63 [1884]. 67 [1998].

802.11s [1836].

A. [906]. Aalen [2146]. aan [540, 118, 217, 504, 131, 630, 291, 454]. aanbestedingen [447, 45]. Aanhangsel [5]. aanleiding [132]. aannemelijke [434]. aanpassen [217]. aanpassing [752, 131, 291]. aanpassingstoetsen [717]. aansluiting [500]. aantal [917, 107, 272, 265, 217, 246, 516]. aantallen [678]. Aantekeningen [436]. Aanvoer [448]. aanwijzen [377]. Aardal [1550]. aardolie [908]. aardolieprodukten [908]. Abnormal [1534]. Abraham [708]. Absolute [1121]. absorbing [2267]. Abstracts [463]. academia [1782]. Academic [80]. Academische [80]. accelerated [2173]. Accept [1724]. acceptance [1775, 841, 813]. accident [1786]. accidents [949, 270]. accommodation [137]. accomodatie [137]. according [327, 736, 302, 37, 793]. account [2258, 2002, 146]. accountantscontrole [353]. accountantscontroles [525, 706]. Accounting [2168, 2167, 2128]. accounts [2123, 1771, 2126, 2129, 148, 553, 2124, 177]. accrual [2007]. Accuracy [1600, 1601, 1030, 486, 42, 497]. Achtergronden [409]. across [706]. action [2181, 170]. activity [638, 2015]. actually [645]. actuarial [1645]. adaptations [1846]. adapting [1776, 131]. Adaptive [1127, 1809, 1659, 1178, 1420, 2013, 1045, 1655, 1874, 1898, 1656, 2187, 2300]. added [639]. Addendum [885]. addilog [1875]. Additional [745]. Additive [1622, 1788, 1657, 2298, 1378]. address [2293, 2243]. adequacy [1864]. adjusted [1898, 2129]. adjustment [1777, 2128, 717]. Adjustments [735, 2125, 762, 4]. administratieve [259]. administrative [1947, 1948, 1949, 1952, 259]. Admissibility [1019]. Advances [1685, 1558]. advantages [1626]. Advertentie [501]. Advertising [501, 281]. affecting [2028, 1038]. afgeleid [148]. Afleiding [710, 873, 29]. aflezen [109, 711]. afname [473]. Africa [1908, 113]. Afrika [113]. Afschatting [717]. afscheid [1001, 1029, 905, 749]. AFT [2169]. after [1731, 1458]. Afterthought [128]. afzetanalyse [604]. again [1467, 740]. Against [1183, 159, 246, 897, 516]. aged [1407]. ageing [1395]. aggregate [2024, 2116, 141]. aggregation [849]. agility [739]. Agreement [1173, 1857, 1063]. Agricultural [1148, 193, 1387, 598]. agriculture [1513]. aids [15]. airline [1660]. Albert [1276, 1357]. algebra [1376]. Algemeen [296, 308, 334]. algemene [288, 625, 425]. ALGOL [830, 522]. Algorithm [1165, 2048, 1379, 1378, 1448, 1790, 2295, 1704]. Algorithmic [1107]. algorithms [2172, 1995, 1793, 1317, 1035, 1843, 1034]. Aligned [1122]. aliter [1764]. alles [28]. allocation [1655, 1974, 696, 879, 901]. allocations [1222]. Almost [1636]. along [1986]. alternate [997]. alternatief [544]. Alternative [636, 470, 910, 1570, 801, 897]. alternatively [544]. Alternatives [1183, 1655, 1385, 1517, 1972, 1012]. ambiguity [2224]. America [178]. Amerika [178]. among [345, 1330, 2097, 929]. amount [355]. amounts [447, 2278, 45]. Amsterdam [244]. analgesics [457]. Analist [387, 433, 478, 517, 564, 609, 654, 647]. analyse [413, 426, 72, 500, 186, 224, 425, 629, 116, 197, 561, 215, 277, 2, 290, 889]. analyses [1045, 2208, 352]. Analysing [1682]. Analysis [1047, 700, 2279, 1610, 1089, 1135, 1174, 1611, 1807, 972, 1214, 1119, 1492, 1951, 2207, 1926, 1619, 1671, 1993, 1665, 1240, 1513, 1604, 1586, 1981, 2000, 2036, 427, 1841, 2225, 519, 1341, 413, 1060, 1324, 1771, 819, 1703, 2228, 2025, 827, 885, 250, 976, 1308, 1401, 1330, 1336, 1661, 1854, 426, 1716, 1535, 468, 72, 1030, 1495, 1830, 1387, 1390, 1095, 500, 470, 469, 908, 186, 916, 921, 1995, 864, 829, 224, 1909, 2081, 2162, 1699, 1709, 834, 1241, 1973, 1826, 61, 604, 2170, 2187, 1916, 425, 1328, 1642, 629, 116, 197, 157, 418, 1895, 1762, 1867, 561, 2292]. analysis [1784, 2055, 1243, 294, 295, 277, 2, 290, 889, 935, 280, 619, 1099]. Analyst [204, 225, 308, 609, 205, 204, 225, 267, 296, 308, 334, 387, 433, 478, 517, 564, 654, 647]. Analytic [2095, 846]. analytica [1764]. analytical [770, 1305]. analyze [2147]. Analyzing [1822]. anchored [2225]. Ancient [1218, 1924]. andere [475]. Anderson [2239]. AneuRisk65 [2111]. animal [586]. animals [44]. anisotropic [2025]. annealing [1265]. Announcement [613, 953, 1069, 1883, 868]. Announcements [89, 1453, 11, 20, 47, 55, 62, 82]. annual [2120, 710, 711, 553]. ANOVA [2153, 727]. Answers [1283]. anthropo [92]. anthropo-sociological [92]. anthropo-sociologische [92]. Anthropological [92]. Anthropologische [92]. anthropology [1744]. anti [1537]. anti-monotone [1537]. Antithetic [1112]. any [1986]. anything [28]. APARCH [2042]. Appendix [576, 5]. applicability [1798]. Application [605, 1084, 2103, 1121, 1604, 1005, 2267, 2302, 1391, 1747, 2225, 2223, 833, 1066, 1938, 1988, 902, 2066, 1977, 1446, 2228, 1685, 129, 1330, 1323, 1458, 709, 1790, 908, 788, 2046, 2071, 1868, 241, 61, 1746, 961, 1715, 1514, 2067, 2111, 2117, 2047, 573, 939, 642, 630, 813, 2055, 1445, 317, 746, 16, 162]. Applications [697, 1150, 988, 1169, 262, 1551, 2000, 2141, 390, 1710, 2129, 1379, 163, 1272, 926, 2295, 1337, 1234, 942, 1642, 853, 717, 136, 2014, 1581, 295, 1744, 511]. Applied [1168, 1030, 3, 1045, 1276, 1772, 1990, 1630, 601, 611, 769, 949, 595, 469, 13, 961, 2310, 823, 150, 189, 182, 417]. applying [71, 107, 77, 353]. Appointment [782]. Approach [1611, 1668, 2137, 554, 1657, 2083, 2152, 1989, 1486, 1401, 1781, 2097, 1792, 2169, 1810, 2195, 1372, 2290, 761, 2139, 1523, 975, 589]. Approaches [1603, 1784, 2206, 1901, 1642, 1821]. Appropriate [1188]. approval [31]. Approximate [1106, 1618, 1275, 2132, 1323, 1514, 1823]. approximated [1730]. Approximatie [376]. Approximating [1250, 745, 725]. Approximation [1160, 376, 1338, 2201, 1654, 687, 1650, 1739, 1370, 2017, 2029, 2051, 1843, 652, 838]. approximations [957, 744, 2311, 1917, 1987, 2056, 1328]. April [1281]. arbeidsprestaties [127]. arbeidsproduktiviteit [354, 350]. arbitrage [401]. arbitrary [1904, 995, 2065]. arch [1294]. arch-enemy [1294]. Archimedean [2292]. area [1632, 1070, 395]. areas [657, 2231, 2175]. argmax [1695]. argmax-functional [1695]. ARIMA [1042]. ARIMA-processen [1042]. Arising [1047, 1573, 916, 921]. arithmetic [704]. Arm [780]. ARMA [2159, 1285]. army [302]. arrangement [383]. array [1990]. array-CGH [1990]. arrester [707]. Arrival [1123, 496]. arrival-intervals [496]. article [41, 497, 132, 2273]. artifact [1924]. artificial [70]. artikel [41, 497, 132]. ascending [1396]. ascent [2048]. ascertainment [1993]. Aselect [441, 408]. aselecte [525]. aseptic [601]. aseptische [601]. aspect [1158]. aspecten [462, 175, 719, 247, 127, 294, 52]. Aspects [1157, 1675, 1782, 1973, 1375, 428, 462, 887, 1588, 175, 1077, 719, 1683, 23, 247, 592, 127, 294, 512, 52]. assay [933]. assays [1310]. assembly [1582]. assess [2195]. Assessing [2252, 1996, 2282, 1823, 1949]. assessment [1948, 2083, 1706, 1571, 2090, 2097, 1779]. assessments [42]. assets [2009]. Assign [1620]. Assignment [492, 1790, 1798, 814, 854, 866, 874, 881, 886, 890, 895, 898, 904, 918, 927, 932, 936, 940, 943, 946, 952, 958, 963, 967, 978, 984, 989, 994, 1000, 1007, 1028, 1037, 1044, 1050, 1056, 1062, 1068, 1021, 922, 1016]. assisted [1866]. associated [245, 1735, 562]. Association [31, 1610, 423, 1750, 1071, 2026, 2040, 782, 733, 422]. Assortimentsbepaling [438]. assumptions [2024]. asymmetric [1925, 2092, 2133]. Asymptotic [1424, 1048, 1714, 2311, 1941, 1115, 2305, 1195, 1435, 2265, 1040, 1499, 1516, 1337, 1286, 1121, 1074, 1070, 1405, 1012, 1362, 1848, 1102, 1897, 1728, 1058, 2096, 488, 1272, 1702, 842, 1890, 1476, 1036, 1096, 1295, 166]. Asymptotical [728]. Asymptotically [1454, 1498, 1512, 2214, 1404]. Asymptotics [2069, 2297]. Asymptotische [166, 488]. atpecten [23]. attachment [844]. Attempts [1107]. Attribute [337]. Attributenkeuring [337]. Attrition [1443, 246]. Attuning [843]. auction [762]. audit [1245, 1569]. Auditing [1282, 1283, 353]. audits [525, 706]. augmentation [1667]. Austrian [1948]. auteurs [345]. authentication [1879]. authors [345]. autocorrelated [864]. autocorrelations [2159, 796]. Automatic [617, 2253, 2295, 136, 16, 661]. Automatische [661, 136, 16]. autopark [703, 593]. autoregression [2072, 2055]. autoregressions [1923, 2074]. Autoregressive [1236, 2235, 1176, 2228, 1822, 1870, 1967, 2081, 2277, 1864, 1487, 2260, 2003]. Autoverkopen [703]. Auxiliary [329, 1847]. Average [2243, 2137, 651, 484, 265, 2293, 694, 2235, 110, 383, 1101, 704, 753, 2300]. averages [985, 116]. averaging [1963, 2012]. Award [1471, 1307]. axis [746].

B [2032]. B-splines [2032]. Back [691, 1837, 730, 730]. backfitting [1378]. Backgrounds [409]. Bacteriological [933]. Bad [1283]. Bahadur [941]. bake [1317]. Bakker [222]. balanced [1755, 1309, 2287, 2104, 838]. balancing [2125]. balls [1640]. bandwidth [2051, 1852]. bank [1712, 980]. banknotes [1078]. bare [97]. bars [1966]. Bartlett [2246, 2086, 2003]. base [755, 1235]. Based [1602, 1122, 1253, 2229, 2141, 2225, 1946, 1945, 1737, 2201, 2137, 1722, 1948, 605, 1863, 548, 2002, 1959, 987, 1716, 2210, 2291, 508, 1808, 770, 1667, 1778, 1796, 2107, 2031, 955, 2240, 1626, 1244, 2186, 231, 2111, 2154, 2121, 1982, 1260, 2052, 1496, 2215, 1422, 1799, 1983, 1787, 1950, 2194, 2239, 395, 1225, 1628]. baseline [2280]. basic [2248, 1388, 1505, 183]. Basil [1234]. basis [534, 239, 680]. basisperiode [755]. basketball [2133]. Batch [1584, 770, 1779]. Batch-service [1584]. Bayes [1385, 1019, 1703, 1633, 1652, 2237, 1564, 1036, 1024, 948, 1730, 897, 1085]. Bayesian [1491, 2032, 2270, 2085, 2236, 1656, 1301, 2216, 2083, 1706, 1844, 1729, 1473, 1930, 1318, 2113, 1726, 2153, 2237, 1922, 1810, 1627, 2133, 2174, 1668, 2245, 1609, 1129, 2188, 2067, 1936, 1563, 1647, 2260, 1933, 1962, 2012]. be [1985, 734, 1241, 473]. beads [848]. Bedaux [77]. Bedaux-systeem [77]. bediening [492, 555]. bedragen [45]. bedrijf [67, 241, 362, 455, 49, 658]. bedrijfsbeheer [57]. bedrijfseconometrie [461]. bedrijfseconomische [902]. bedrijfsleider [241]. bedrijfsleven [90]. Bedrijfsmodellen [662, 865]. bedrijfstakken [303]. Bedrijfsvergelijking [367]. beet [578]. begin [196]. beginning [196]. begrip [32]. begrippen [278]. behandeling [143]. Behavior [1115, 2147, 1812, 211, 2102, 2139]. Behavioral [1149]. behaviour [1023, 1476, 907, 1036, 1096]. behendigheid [739]. behoefte [540]. Behrens [892]. behulp [707, 616, 911, 370, 218, 619, 159, 298, 256, 593, 39]. beide [78]. beïnvloed [629]. bekeken [973]. belief [2148]. Belletjes [258]. belong [136]. Belsley [1020]. bemonsteringsschema’s [348]. benadering [687]. benaderingsformules [652]. benchmark [2126]. Benchmarking [2119, 2120, 2121]. benefit [1888]. Benelux [99, 99]. Benoeming [782]. beoor [42]. beoor-delingen [42]. bepaald [272, 265]. bepalen [316, 473, 138, 137]. Bepaling [303, 261, 680, 386, 42, 160, 444, 395]. berekenen [397]. berekening [272, 206, 290]. berekeningen [648]. berekeningsmethode [712]. Bergstrom [1823]. berichtenverkeer [340]. Bern [105]. Bernoulli [1655, 1167, 1426, 1231]. Bertus [1493]. beschouwingen [625, 740, 579, 237]. Beschrijvende [437]. Besliskunde [535, 607]. Beslissen [542, 534, 541, 256, 543]. beslissingsprobleem [371]. beslissingsproblemen [503]. beslist [645]. Best [1137, 800, 992, 1563, 418, 933]. bestand [591]. bestek [537]. bestelniveaus [473]. Bestimmung [88]. bestuur [679]. Beta [2213, 1135, 2057, 1853, 1934, 1535, 2029, 2234]. Beta-Binomial [1135, 1853]. betekenis [461, 90]. Betere [585]. betreffende [579, 648, 747]. Betrekkingen [151, 141]. betrekkingtussen [574]. betrouwbaarheid [7, 40, 382]. betrouwbaarheidsgebied [395]. betrouwbaarheidsgrenzen [333, 551]. Betrouwbaarheidsintervallen [265]. better [2293, 1871, 585, 585]. Between [1106, 1184, 1621, 1024, 354, 554, 688, 1298, 2231, 2126, 545, 574, 1397, 415, 880, 151, 44, 1403, 2151, 97, 951, 965, 466, 1823, 2040, 1579, 1857, 729, 96, 2045, 2309, 587, 141, 586]. bevolking [244]. bevoorradingsmagazijnen [498]. Beware [985]. bewegende [116]. bewerking [37, 78]. bewijs [312, 183]. bewijzen [28]. bezetting [651]. bezettingstheorie [162]. bezien [307]. bezoek [118]. Bias [2037, 643, 1941, 1803, 2071, 1310, 2253]. bias-corrected [2071]. bias-reducing [1310]. Bibliografie [424]. Bibliography [424]. Bibliotheek [422]. bieten. [578]. Bij [196, 1001, 1029, 905, 749, 31, 233, 345, 713, 912, 492, 555, 413, 84, 917, 145, 624, 57, 41, 14, 24, 71, 76, 107, 484, 70, 109, 394, 532, 525, 706, 151, 333, 498, 631, 44, 12, 56, 332, 317, 231, 110, 91, 742, 197, 768, 460, 925, 561, 753, 929, 630, 410, 542, 779, 77, 59, 259, 353, 436, 277, 454, 183, 290, 733, 889, 652, 27, 594, 112, 401, 95, 45]. bijdrage [741, 579, 679]. bijeenkomsten [179]. bijzonder [77]. bilateral [2269]. Binary [1670, 1182, 1809, 2262, 1446, 1807, 2181, 1790, 788, 2046, 2162, 2280, 2038, 1993, 1295]. binnenwerk [768]. binomal [951]. Binomial [1127, 1192, 1135, 1003, 2031, 872, 1832, 1853, 1987, 687, 850, 488, 745, 725, 2038, 998, 729, 2045, 2194]. binomiale [687, 488]. Bio [1106]. Bio-medical [1106]. bioequivalence [1570]. Biological [423, 1521, 1251, 499]. Biologische [499, 423]. Biometric [99, 106, 106]. biometrica [120]. biometrics [120]. Biometrische [99]. bioscopen [137]. Biostatistics [1151]. bird [1800, 1070, 1038]. Birnbaum [1916]. birth [1032, 1033]. birth-death [1032, 1033]. Bivariate [1166, 1113, 2157, 2145, 1711, 1573, 1891, 470, 1877, 2046, 1814, 857, 1497, 1477, 2194, 825, 869]. Blackjack [1505]. Blanco [51]. Blank [51]. blind [1881]. Block [838, 1755, 2050, 394, 2186, 1697]. block-diagonal [2050]. blocked [1780]. blocking [1232]. blockmodels [2187]. bloedproef [307]. Bloemena [404]. blokkenproef [394]. blood [307, 2310]. BLUF [873]. BLUF-schatter [873]. BLUS [1235]. Boekbeapreking [46]. Boekberpreking [121]. Boekbespreking [30, 53, 68, 73, 81, 93, 104, 133, 320, 705, 715, 723, 737, 748, 757, 789, 794, 799, 803, 815, 855, 867, 875, 891, 899, 903, 913, 923, 947, 10, 772]. Boekbesprekingen [111, 140, 153, 167, 185, 201, 209, 221, 226, 232, 248, 255, 263, 268, 274, 279, 285, 292, 299, 305, 311, 331, 341, 356, 373, 384, 392, 403, 416, 430, 440, 450, 464, 474, 482, 493, 505, 515, 523, 533, 547, 557, 570, 581, 599, 606, 633, 653, 663, 682, 689]. Boekbesprekingens [612]. Bogaard [490]. Bok [558]. Bolk [216]. Book [644, 674, 698, 731, 766, 1067, 1438, 809, 1020, 298, 30, 46, 53, 68, 73, 81, 93, 104, 111, 121, 133, 140, 144, 153, 167, 185, 201, 209, 221, 226, 232, 248, 255, 263, 268, 274, 279, 285, 292, 299, 305, 311, 331, 320, 341, 356, 373, 384, 392, 403, 416, 430, 440, 450, 464, 474, 482, 493, 505, 515, 523, 533, 547, 557, 570, 581, 599, 606, 612, 633, 653, 663, 682, 689, 705, 715, 723, 737, 748, 757, 789, 794, 799, 803, 815]. Book [855, 867, 875, 891, 899, 903, 913, 923, 947, 10, 772]. Bookbesprekingen [144]. Bootstrap [2172, 2218, 1548, 1728, 1208]. Bootstrapping [1169, 1575]. bosbouw [366, 350]. both [770, 2287, 78, 1911]. bound [879, 1476, 951]. boundary [1507, 1514]. bounded [1514]. Bounding [1957]. Bounds [998, 1887, 1801, 1796, 2059, 729]. bouwmarkt [480]. bovengrens [110]. Box [2061]. brain [2187]. branched [348]. branching [1640]. Brand [929]. Branley [962]. Brazil [2223]. bread [72, 95]. breakdown [1402]. Breast [1168]. breeding [231]. Bridging [1403]. brief [1777, 591]. brightness [676]. broken [467]. bron [413]. brood [72, 95]. Brownian [1013, 1632]. Brussels [421]. Bubbles [258]. budget [1288]. Building [1383, 969]. Bureau [630, 1030]. Buridanus [777]. burst [591]. Business [471, 2206, 554, 902, 461, 57, 470, 469, 402, 865, 90, 367, 662].

C [818, 1008, 1087, 1408, 648, 817]. cable [1756]. calculating [397]. calculation [712, 272, 147, 206, 2045, 290, 444]. calculations [734, 648, 289]. calculator [1094]. calculators [511, 614]. calculus [149]. Calcutta [380]. Calibration [1602, 2192]. Call [1239]. Calling [122]. Camp [108]. can [28]. Cancer [1168, 1605, 1095]. Candy [1674]. Canonical [1064, 1065]. capability [2092]. capacitated [1798]. capacities [1469]. car [703, 590, 593, 703]. car-following [590]. cards [228, 2177]. care [2166, 2097]. Carlo [2048, 1385, 2166, 372, 1965, 2057, 386, 1112, 621, 668]. carousel [1909]. carried [734]. carry [2279]. carry-over [2279]. carryover [2091]. cars [650, 649, 651]. Case [987, 1609, 1604, 575, 1443, 1022, 1341, 1786, 2083, 1712, 1486, 1832, 171, 2242, 1695, 2124, 323, 1435, 1667, 1095, 915, 458, 955, 950, 2009, 2290, 847, 1787, 456, 1038, 1383, 452, 1221, 322]. cases [1683, 1966, 455]. Cash [1688]. Casino’s [1505]. casual [2304]. Categorical [1207, 2109, 2312, 1487]. categories [1892]. Cauchy [701, 807]. causal [1989, 1707]. Causes [1177, 1367]. cell [1832, 1853, 1959, 2112]. Cells [1145]. Celsius [203]. Censored [1184, 2082, 2173, 1435, 1845, 1272, 2060, 2308, 2160, 1916, 2261, 1979, 1692, 1982, 1757, 643]. Censoring [1014, 1484, 1424, 1454, 1498, 2207, 2079, 2289, 1497]. censorship [1333]. Census [1948, 735]. censuses [836]. centered [2293, 2243]. Centraal [630]. Central [1494, 1361, 1030, 1156, 980, 1483, 2143, 2088, 124, 146, 2303, 672, 630]. centrale [124, 146]. Certain [1161, 1396, 795, 765, 476]. Certainty [690, 742]. CGH [1990]. chain [2257, 1965, 639, 2267]. Chains [1130, 560, 462, 902]. Chairman [234]. challenge [1920]. challenges [1946, 1945, 1774, 1952]. Chamber [975]. chance [738, 880, 257, 740, 739]. Chances [708]. Change [2228, 1182, 1873, 471, 628, 2024, 1336, 1908, 628]. Change-over [1182, 628, 628]. Change-point [2228]. changes [1356, 475]. changing [1841]. character [15, 1073]. Characteristic [1253, 1886]. characteristics [1904, 244, 538]. characterization [1297, 1364]. Characterizations [1314, 1887]. Characterizing [1170, 1480, 2132]. chart [1367, 860, 309]. charts [1776, 1856, 1778, 1779]. chauffeursploeg [364]. check [149]. checking [7, 1864, 2219]. checks [259]. chemical [2015]. Chen [2229, 2289]. Chi [1145, 683, 752, 1638, 672, 838]. chi-kwadraat-toets [752]. Chi-Square [1145, 752, 672, 838]. chi-squared [683, 1638]. Chief [1001, 1029, 749]. child [1407]. children [1459]. Chimeric [1517]. Choice [1145, 1446, 1965, 386, 258, 929, 1295, 276, 257]. cholera [1922]. Choosing [1235, 896, 542]. Christiaan [1408]. chromatograph [2266]. chromatography [1022]. chronic [1988]. cijfers [64, 72]. cinemas [137]. Circle [521]. circles [665]. circulation [2114, 626]. cirkulatie [626]. citation [1661]. cities [2231]. City [2124]. claim [2278]. claims [2116]. CLAR [2193]. Class [1863, 1200, 1156, 1169, 666, 1396, 852, 203, 801, 1137, 2182, 2156, 948, 1523, 962]. Class-specific [1863]. Classes [1046, 1658, 1506, 2309]. classical [1716, 1727, 562]. Classification [942, 988, 1024, 713, 2295, 2280, 2115, 664]. classified [1232]. cleaner [844]. cleanliness [555]. climate [1908]. climatological [79]. climatology [1264, 65]. clinic [489]. Clinical [1669, 1898, 1656, 2262, 2167, 325, 457]. clinische [325]. closed [1580, 2281]. closed-form [2281]. closing [2002, 1070]. clothes [1503]. clothing [347, 52]. cluster [2078, 1709, 2187]. Clustered [1619, 2298, 1807, 2241, 2078, 1741, 1826, 2047]. Clustering [987, 1359, 1969, 2295, 2074, 1936]. coarsening [1574]. coast [286]. Cochran [787]. Cochranen [320]. coderingsvraagstuk [162]. codes [591]. coding [1842, 162]. codispersion [2040]. coefficient [477, 957, 1953, 1002, 1923, 151, 2095, 1937, 160, 1899, 395]. coefficients [2028, 1703, 1957, 1926]. cognitive [2267]. Coherent [2061, 1940]. cohesion [2016]. coincidence [1924, 64, 783]. coincidental [109]. cointegrated [1717]. collapsibility [2312]. collection [901]. collision [1482].

Combination [1948]. combinations [2032, 1496]. Combinatorial [1205, 1451, 1551, 1523]. Combinatorics [1205]. combineren [747]. Combining [2066, 1931, 1691, 787, 991, 747]. Comment [1999, 1998, 912]. Commentaar [912, 740]. commentary [740]. Comments [1047, 1997, 1158, 600, 132, 272, 489, 25, 92, 77, 340, 917, 648]. Committee [278, 491]. committees [445]. common [2269, 182]. Communicatie [588, 587, 584, 585, 598]. Communication [586, 584, 585, 587, 598, 588, 587]. compactification [1358]. companion [452]. Company [12, 658, 461, 67, 241, 362, 147, 455, 49]. Comparatieve [460]. Comparative [2035, 460]. compared [1009, 98]. Comparing [957, 1333, 1072, 602, 2239, 207, 608]. Comparison [1515, 1193, 2059, 1172, 2047, 877, 1603, 1122, 2231, 1941, 837, 1728, 2305, 1716, 58, 367, 1328, 1982, 368, 621, 838, 1012, 679, 516]. Comparisons [1088, 1318, 2161, 2278, 1506, 1271, 962]. Competing [2298, 2207, 2097, 2217, 2261, 2289, 2292]. Competition [346]. compiled [1945]. Complementary [2049]. complete [786, 1220, 1697]. completely [1384]. complex [2172, 1784]. complexity [1794, 1039, 1962]. component [579, 469, 808]. Components [1005, 1308, 834, 1460, 1010]. Composed [1005]. composite [579, 1868, 2023, 2139]. compositions [2230]. Compound [1019, 1202]. comprehension [2115]. comprised [811]. Computation [1377]. computational [926, 1434, 512]. computations [1551]. Computer [1376, 511, 616, 619, 618, 630, 627, 760, 615]. computerized [1393]. computers [624]. Computing [1723, 1621, 1632, 830, 1375]. concave [1633]. Concentration [2102, 1842]. concept [1052, 1455, 2313, 32]. concepts [278]. concerning [870, 743, 671]. Concomitant [1653]. Concurrency [1583]. condities [316]. condition [824]. Conditional [1673, 2090, 393, 1711, 758, 2304, 1897, 2313, 2148, 2151, 2234, 1477, 686]. Conditionally [1880]. conditionals [1477]. Conditioning [1504, 997]. conditions [1769, 831, 316, 1843, 2286]. conduct [694, 2170]. Confidence [1106, 1160, 1801, 1184, 770, 245, 1882, 1114, 1424, 957, 1298, 2090, 2200, 333, 968, 2215, 2132, 1070, 551, 395, 265]. confirmatory [1330]. conflict [1330]. conflicts [912, 889]. confounding [2001]. congres [106, 105]. Congress [106, 105]. connecting [1986]. connection [971, 500, 386, 97, 556, 244]. Connections [2309]. Conover [2301]. consequences [2213, 1405, 1447]. consideration [1806]. Considerations [1019, 1622, 914, 625, 579, 961, 237]. Consistency [1848, 1887, 2127]. Consistent [1018, 1526]. Constance [321]. constant [2251, 2124, 131, 1413, 277, 290]. constante [277, 290]. constanten [182]. constants [182]. Constrained [1186, 1486]. constraint [1790]. constraints [2032, 1938, 2276, 1959, 1728, 2294, 880, 1889, 2056, 2128, 1304]. Constructie [36, 70]. Constructing [1806]. Construction [1219, 1136, 1602, 480, 2206, 70, 556, 1221, 36]. Constructive [1442]. consulting [1782]. Consument [284]. consumenten [327, 929, 90]. consumenten-onderzoek [90]. consumer [327, 594, 90, 284]. consumers [929]. consumption [502]. contact [1644]. contacts [788]. Contamination [1315]. contemporary [1794, 58]. Content [1624, 800, 72, 790]. Contents [1539, 1597]. contest [399]. contingency [1863, 1844, 1552, 787, 1318, 2026, 375, 1805]. continue [484, 752]. Continuity [930]. Continuous [1825, 931, 1671, 1820, 1656, 2262, 2276, 2251, 2020, 1822, 1828, 1695, 2183, 484, 2173, 752, 1912, 1808, 1824, 2062, 2186, 1821, 1085]. Continuous-Time [931, 1825, 1822, 1824, 2186]. contöle [132]. contrôle [72, 101, 259]. contrôle-analyse [72]. contractor [519]. contrast [1996]. contrasts [2287, 1972]. contribution [741, 579, 1876, 679]. contributions [1452]. Control [1583, 714, 1776, 1856, 1333, 1367, 363, 14, 24, 171, 1072, 34, 132, 72, 1778, 2282, 1779, 602, 992, 695, 847, 1412, 101, 1787, 6, 543, 1212]. controle [261, 6]. controle-interval [261]. contröle-materiaal [6]. controlegroep [602]. controleren [149]. controleurs [40]. controlled [1985, 472]. controls [1569]. Convenient [2009]. conventional [554]. Convergence [1228, 1843, 1061, 993, 2016, 1913, 1547, 1097]. conversion [1796, 203]. converting [78]. Convex [1676, 2213, 589, 1654, 1514]. convexity [692]. convolution [1853, 2314]. Convolutions [1555]. cookie [1500]. Cooperation [1330]. Coordination [21]. Coördinotie [21]. Copula [2048, 1815, 2073, 2292]. copula-driven [2073]. copulas [2000, 2147]. cords [617]. Cornelius [1281]. Corrected [1684, 2071]. Correctie [716, 713, 59]. Correcting [1935, 1546]. Correction [1247, 699, 970, 1491, 577, 714, 528, 1590, 2246, 2086, 2253, 59, 2003, 716, 713, 930]. correlated [2299, 2142, 2248, 1969, 1031, 1816, 1483, 2030, 2269, 1709, 2047, 1932]. correlatie [443]. correlatieberekingen [289]. correlatiecoëfficiënt [160]. correlatiediagram [36]. Correlatierekening [150, 149, 147, 146, 79]. Correlation [271, 477, 957, 149, 1002, 611, 2096, 36, 139, 2241, 200, 272, 443, 2095, 1871, 147, 289, 2154, 146, 79, 160, 2098, 397, 825, 150]. correlations [1970]. Correspondence [1099]. Corrigenda [1254, 1293, 1143, 1189]. Corrigendum [1804]. Corsten [905]. cost [414, 2011, 1579, 33, 449, 472, 680]. cost-controlled [472]. costs [1459]. Count [1174, 2229, 1711, 2064, 2279, 1994, 2061, 1979, 2235, 2132]. Counting [1178, 1800]. countries [540]. country [118]. counts [1832, 1853, 846, 1541]. coupled [2113]. courses [911]. Covariables [1177]. Covariance [1811, 1416, 688, 2050, 1499, 786, 1328, 1632]. covariate [1898, 1651, 2245, 1788, 2300]. covariate-adaptive [2300]. covariate-adjusted [1898]. Covariates [1684, 1985, 2046, 2015, 1705]. covariates-dependent [2015]. coverage [1791]. covering [1426]. covers [535]. COVID [2258, 2228, 2209]. COVID-19 [2258, 2228, 2209]. Cox [320, 1996, 2196, 1862, 1880, 2146]. Coxian [1270]. Coxian-2 [1270]. CPB [1766]. Cramér [1467, 2059]. Creating [538]. creation [123]. Credibility [2104]. Credit [1775]. crime [2231, 1681]. Criminal [1679]. criminology [1682]. criteria [1543, 1385, 917, 1942]. criterion [2221, 896]. criterium [896]. critical [1640]. Criticism [102]. Critiek [102]. critique [1503].

Crommelin [1837]. Cross [1615, 1602, 1608, 1748, 2279, 1627, 2052, 2041]. cross-over [2279]. cross-section [1608]. cross-sectional [1748]. cross-validation [2052, 2041]. crossing [507, 600, 1968, 572]. crossing-points [1968]. crossings [1637]. crossover [1809, 2091]. crossroad [1575]. crumble [1500]. cubic [1825]. cum [321]. cumulative [1338, 1805]. Cup [1919]. cure [2270, 2217]. currency [401]. Current [1148, 1622, 1777, 2078, 2160, 2155]. curse [2222]. curve [712]. Curved [1113]. curven [159, 152]. curves [159, 642, 798, 152]. cushion [2276]. customer [1580, 843]. customers [1488]. CUSUM [1967]. Cusums [2065]. Cutting [492, 555, 1881, 1810]. cytostatica [489]. cytotoxic [489].

D [1047, 374, 1020, 1837]. Dag [358, 280, 155, 773, 99]. dagen [89]. D’Agostino [877]. Daily [1686]. dairy [448]. Dalenius [1348]. dangerous [1786]. Dantzig [374, 1471, 1307, 781, 1017]. Dantzig-prijs [781, 1017]. Darling [2239]. Data [1183, 1670, 1608, 1680, 1184, 1200, 1437, 1120, 1174, 1182, 1207, 1622, 1619, 1259, 1240, 1049, 1434, 528, 1243, 1443, 2109, 1980, 1981, 2036, 2018, 2225, 912, 1533, 1898, 1737, 2255, 1948, 1711, 1755, 2206, 554, 2304, 2258, 1534, 1333, 1446, 2228, 2298, 2152, 2172, 1825, 2175, 2025, 2082, 2279, 2203, 1828, 1756, 1651, 2129, 1232, 1309, 1897, 1530, 1854, 2166, 1807, 2312, 1969, 1764, 1531, 1930, 2181, 1949, 1830, 2197, 1329, 1435, 2241, 2050, 1421, 1845, 1448, 1912, 1929, 2030, 1952, 1492, 2308, 1994, 470, 469, 2207, 1824, 2230, 1942, 2078, 916, 921, 1741, 2269, 1638]. data [2160, 1709, 834, 2155, 1241, 1522, 846, 243, 2127, 1826, 822, 541, 617, 1820, 2168, 1460, 1916, 2146, 2202, 2038, 1372, 1993, 2261, 1979, 1090, 2186, 2188, 1449, 2110, 2310, 2111, 1788, 1972, 1802, 1895, 2011, 1982, 1867, 2247, 2267, 2292, 1322, 2215, 2037, 1757, 476, 2132, 1983, 1787, 1950, 2055, 2266, 889, 1447]. Data-Analysis [1240]. data-analytic [846]. datasets [1718]. David [374]. Day [358, 280, 155, 773, 99]. Days [89]. DEA [1685]. Dead [537]. death [1032, 2155, 1033]. decade [1834, 1458]. December [31, 179, 33]. decentralized [734]. decides [645]. deciding [534, 541]. Decision [1019, 1153, 1385, 607, 1233, 371, 900, 1581, 325, 503, 364, 535]. decisions [955, 256, 542, 543, 449, 491, 637]. decks [2177]. decomposability [1468]. deconvoluting [2266]. deconvolution [1672, 1512, 2135, 1931]. Decreasing [1494]. deel} [654]. deficiency [914, 1091]. Defined [1183, 1090, 2051]. Defining [471]. definition [691, 1570, 730]. degree [203, 1571]. degrees [203]. dekt [535]. delay [507, 600, 1756, 572]. delayed [1991]. delicate [925]. delingen [42]. deliveries [604]. delivery [473, 243]. Delta [625]. Deltaplan [625]. Demand [1005, 1875, 847]. Deming [1087]. demografische [244]. demographic [244]. demonstratie [124]. demonstration [124]. densities [1554, 1388, 1032, 514, 2237, 857, 1427, 573]. Density [1494, 1676, 1673, 1672, 1170, 1887, 1560, 1592, 1561, 1334, 1978, 1750, 2079, 2295, 1852, 2041, 672]. Departure [1123]. dependence [1815, 2000, 1748, 2031, 1814, 1589, 2104, 825]. dependencies [1701]. dependency [2265]. Dependent [1195, 1177, 1396, 2254, 1645, 1880, 2168, 676, 2146, 496, 2015, 2023, 878]. deployment [604]. deployment-sales [604]. derde [84]. derivation [801, 873, 29, 442, 672, 710]. Derivative [1676]. derived [1693, 148]. Description [1542]. Descriptive [1372, 437]. Design [1174, 2011, 325, 1809, 1898, 1656, 2091, 1968, 1780, 1781, 1398, 489, 2005, 2026, 1697, 1310, 896, 1558, 1342, 491, 1749, 250, 325]. Designing [1159]. Designs [1201, 1618, 1612, 1391, 428, 315, 917, 1639, 251, 324, 312, 1806, 1862, 2162, 2153, 920, 2287, 2062, 838, 320]. desires [131]. destructive [2217]. detectie [323, 312]. Detecting [2039]. detection [1409, 1873, 323, 312, 1447]. determinant [2096]. Determination [870, 1677, 605, 386, 42, 160, 668, 88, 303, 261, 395, 680, 298]. determinations [532]. determine [147, 137, 326]. determined [265]. Determining [1161, 519, 316, 444, 473, 138, 438]. developing [540, 2168]. development [1771, 324, 253, 13, 736, 631, 1770, 439, 460, 202, 997, 254]. Developmental [1619]. Developments [1146, 188, 2123]. deviation [1465, 1011, 277, 290]. deviations [941]. Device [1161]. deze [218]. Diagnostic [1934, 1600, 2081]. Diagnostics [1916, 2020, 1020]. diagonal [2050, 1805]. diagonal-parameter [1805]. diagram [36]. diagrams [560, 260]. diameters [665]. DIC [1725]. Dice [750]. Dichotomous [955, 1324, 1694]. dichotomously [955]. dichotomy [1635]. did [1500]. dienst [2, 349]. Differ [1665]. Difference [1106, 1184, 1298, 1930, 1910, 2269, 1011, 2067, 466]. differences [2257, 1023, 2187, 2045, 545]. different [608, 2206, 1997, 577, 549, 1999, 1976, 1998]. Differentiable [1368, 1169, 673]. Differential [1170, 1988, 2174]. differentiation [853]. difficulties [71]. diffusion [1617, 1984, 1922, 593]. digitaal [504]. digitally [504]. dilution [1310, 1867, 642]. dimension [2108, 2242, 1868, 1035, 2303]. dimension-reduction [2242]. dimensional [1980, 2036, 1940, 1838, 1881, 1553, 2050, 508, 1740, 2047, 2253, 1425]. dimensionality [1706]. Dimensionering [610]. diptheria [115]. diptherie [115]. Direct [1561, 873, 1025]. directed [1730, 1705]. directeur [238]. Directions [1176, 78, 1412]. director [238]. directory [527]. Disaggregation [2119, 2123, 2122]. discharge [707]. discipline [1774]. disciplines [1835]. Disclosure [1353, 1349, 1352, 1553, 1544, 1350]. Disconnecting [1792]. discontinue [484]. discontinuity [1643]. discontinuous [223, 484]. Discounted [900]. discouraged [1241]. discreet [443]. Discrete [1226, 1687, 1677, 1247, 1394, 1290, 2018, 2259, 1648, 2183, 1990, 1814, 961, 893, 2007, 2062, 2186, 1246, 1802, 1823, 1586, 948, 641, 1555, 2132]. discrete-time [2007, 2062, 2186, 1586]. discretely [443]. Discretized [2266]. Discrimination [988, 1940, 942, 2111]. Discriminative [165, 479, 394, 143]. discriminatory [218]. Discussie [741]. Discussion [1949, 1952, 1628, 1240, 230, 1243, 1342, 741]. disease [1691, 115, 456]. diseases [1988, 456]. Disequilibrium [1767, 1929]. disintegration [1504].

Disorder [412, 405, 415, 411, 406]. Dispersion [1684, 1502, 670, 1593, 1934, 1023, 1421, 2057, 183]. Displaced [1084, 1168]. dispositions [2181]. Disruption [1795]. Distance [1610, 965, 1131, 1527, 651, 1397, 1716, 514, 1986, 951, 2154, 1641, 729]. distance-based [1716]. Distance-regular [1131]. distances [1518, 975]. distortion [2070, 2192]. distortions [2286]. distributed [1987, 333, 443, 473, 544]. distributes [152]. Distributing [706]. Distribution [1215, 1166, 1298, 1157, 1184, 1192, 1187, 1052, 1134, 1072, 1105, 2159, 1156, 306, 968, 1605, 1866, 669, 1211, 1227, 2302, 1253, 1316, 608, 390, 1396, 2259, 524, 744, 872, 1061, 1297, 1250, 1296, 1917, 687, 1091, 1891, 2254, 849, 1482, 1650, 1562, 2142, 1693, 1388, 2209, 2096, 1633, 72, 488, 203, 949, 508, 2281, 1808, 820, 272, 265, 1305, 2031, 1956, 1734, 665, 701, 2017, 1735, 1829, 2029, 1403, 2237, 1059, 2234, 858, 97, 385, 676, 807, 1338, 951, 1208, 2067, 995, 2289, 273, 466, 924, 1251, 1594, 29, 729, 326, 1477, 847, 1404, 1295]. distribution [2132, 743, 127, 805, 123, 395, 183, 732, 1270, 825, 804, 621, 637, 838, 442, 95, 1158, 43, 1371, 157, 551, 269]. Distribution-Free [1184, 306, 1052, 1072, 968, 2254, 1059, 621, 637, 838, 551, 157, 269]. Distributional [1215, 1236, 1428, 2248, 1564]. Distributions [1128, 1202, 931, 1003, 1429, 223, 1394, 2085, 2213, 1527, 1828, 677, 126, 1630, 828, 484, 870, 1873, 1970, 752, 508, 1808, 1314, 66, 806, 2081, 1814, 1516, 1480, 620, 1735, 2237, 1337, 1220, 966, 845, 2261, 496, 2156, 75, 808, 948, 717, 692, 641, 642, 1555, 1933, 125, 152, 393, 434, 444, 726, 27, 668]. disturbance [852]. disturbances [1816, 864, 786]. dit [122]. divergence [2210, 1097, 1563]. divergences [1516]. diverging [2101]. Diversification [1815, 2182]. Divisia [1731]. divisible [2145]. division [684]. DNA [1683]. Dnderzoek [45]. do [362, 1855]. dobbelspel [750]. Doctrine [708]. doeltreffendheid [545]. doen [362]. does [1888]. domain [1418, 2230]. dommekracht [780]. donkey [777]. Door [494, 495, 278, 350, 550, 504, 78, 629, 894, 585, 303, 238, 353]. doorgelaten [110]. Doornbos [1029]. dosage [44]. Dose [2046, 61, 1310]. doseeringsproblemen [25]. dosering [44]. doseringsletaliteitproef [61]. dosing [25]. double [1732, 2260]. Downdating [1621]. Dr [357, 374, 530, 775, 242, 558, 749, 782, 1408, 222, 550]. Dr. [1087, 475]. Draft [278, 230]. draw [753, 746, 733, 2280, 925]. drempelwaardemetingen [454]. Drie [50]. drift [2002, 1984]. driven [2137, 2073, 1718, 2116]. driver [949, 364]. driving [303, 364]. droge [72, 95]. droge-stofonderzoek [95]. Dropouts [1669, 1895]. drugs [489, 779]. dry [72, 95]. dual [1874]. Duals [2314]. Duel [646, 646]. dumbass [780]. dummy [1708]. dunne [448]. dunnen [578]. Dunnett [602]. Duplicability [198]. duplicate [532]. Dupliceerbaarheid [198]. duplobepalingen [532]. duration [2145, 1758]. Durations [1443, 638]. Durbin [2200, 838]. during [132, 72, 197, 101, 259]. Dutch [657, 707, 168, 608, 5, 30, 46, 53, 21, 11, 20, 47, 55, 1, 31, 68, 73, 81, 62, 82, 69, 93, 104, 111, 100, 121, 133, 140, 144, 153, 167, 185, 174, 201, 209, 204, 221, 226, 232, 225, 248, 255, 249, 233, 263, 268, 274, 267, 279, 285, 292, 299, 296, 297, 278, 305, 311, 331, 320, 308, 330, 341, 356, 373, 384, 334, 345, 346, 358, 344, 392, 403, 416, 390, 387, 388, 396, 389, 398, 424, 430, 440, 450, 464, 433, 445]. Dutch [446, 474, 482, 493, 505, 478, 515, 523, 533, 547, 517, 506, 510, 518, 526, 536, 557, 570, 581, 599, 564, 552, 606, 612, 633, 609, 653, 663, 654, 647, 682, 689, 705, 715, 723, 716, 737, 748, 757, 775, 773, 781, 789, 794, 799, 803, 815, 855, 814, 854, 867, 875, 866, 874, 881, 891, 899, 886, 890, 895, 898, 903, 913, 923, 904, 918, 927, 932, 936, 940, 947, 943, 946, 952, 958, 963, 967, 978, 984, 989, 994, 1000, 1007, 1028, 1017]. Dutch [1037, 1044, 1050, 1056, 1062, 1068, 713, 702, 912, 169, 284, 447, 492, 480, 537, 555, 560, 527, 354, 413, 467, 188, 646, 208, 84, 196, 135, 315, 408, 10, 43, 15, 18, 366, 350, 379, 462, 540, 1771, 917, 758, 60, 102, 902, 848, 184, 145, 494, 1685, 461, 741, 288, 545, 548, 700, 582, 624, 650, 649, 703, 57, 574, 380, 755, 632, 363, 377, 616, 710, 149, 722, 625, 911, 568, 7, 41, 14, 24, 71, 76, 107, 156, 171, 659, 774, 906, 712, 228, 129, 376, 687, 686, 479, 227, 677, 126, 365, 400]. Dutch [405, 777, 205, 651, 34, 497, 132, 426, 484, 534, 601, 369, 338, 591, 372, 74, 58, 64, 70, 105, 87, 118, 122, 179, 180, 235, 251, 324, 313, 337, 417, 740, 738, 661, 415, 211, 579, 623, 36, 72, 94, 142, 117, 139, 163, 175, 321, 323, 312, 441, 488, 739, 783, 1030, 752, 203, 349, 830, 563, 655, 772, 109, 115, 98, 22, 409, 170, 195, 210, 253, 193, 595, 301, 119, 597, 590, 272, 954, 394, 500, 532, 550, 370, 709, 218, 265, 280, 487, 391, 908, 260, 525, 607]. Dutch [706, 1797, 190, 176, 186, 217, 361, 521, 619, 499, 220, 327, 788, 316, 559, 65, 151, 333, 443, 13, 66, 67, 224, 420, 750, 504, 719, 736, 192, 458, 241, 778, 264, 498, 502, 678, 701, 99, 631, 535, 473, 620, 310, 243, 38, 120, 351, 343, 44, 61, 106, 159, 386, 12, 457, 604, 873, 367, 155, 177, 541, 97, 51, 42, 2280, 138, 143, 298, 56, 332, 520, 286, 658, 362, 214, 307, 300, 317, 425, 423, 489, 588, 602, 618, 431, 509, 571, 648, 1001, 1021, 1029, 19, 302, 439, 754, 411, 414, 131]. Dutch [50, 258, 231, 465, 565, 25, 28, 4, 75, 110, 91, 124, 147, 239, 256, 371, 347, 566, 556, 742, 309, 378, 40, 137, 289, 352, 381, 780, 213, 240, 199, 407, 383, 922, 37, 78, 629, 116, 197, 287, 148, 304, 768, 562, 108, 157, 283, 704, 711, 146, 257, 460, 656, 894, 718, 896, 717, 708, 29, 23, 79, 103, 160, 348, 585, 593, 303, 437, 553, 628, 925, 756, 561, 33, 660, 685, 136, 17, 35, 96, 181, 191, 583, 753, 9, 206, 326, 422, 459, 622, 884, 888, 335, 202, 929, 359]. Dutch [615, 905, 939, 236, 261, 455, 101, 567, 749, 782, 230, 630, 247, 578, 262, 319, 746, 410, 16, 238, 92, 150, 189, 456, 6, 49, 127, 125, 194, 234, 501, 542, 587, 551, 779, 325, 747, 973, 1016, 596, 77, 59, 294, 503, 645, 113, 215, 282, 128, 123, 162, 244, 259, 276, 291, 353, 435, 495, 178, 368, 412, 158, 237, 152, 402, 436, 793, 141, 166, 161, 187, 212, 252, 207, 246, 269, 395, 397, 295, 277, 865, 271, 454, 544, 32, 2, 48, 85, 172, 254, 342, 452, 776]. Dutch [627, 382, 355, 543, 183, 219, 290, 130, 448, 812, 475, 364, 434, 732, 733, 889, 652, 444, 626, 406, 318, 86, 27, 594, 675, 751, 134, 112, 511, 328, 401, 598, 679, 538, 52, 26, 114, 322, 314, 399, 584, 516, 182, 1456, 39, 614, 340, 662, 472, 165, 164, 481, 580, 90, 89, 80, 95, 198, 680, 438, 610, 45]. Dynamic [1614, 1144, 2097, 2098, 1359, 2123, 1925, 2183, 1830, 2113, 2071, 1344, 2188, 1101, 802, 939, 1983, 449]. dynamics [1879, 1870, 2184, 1405, 1447]. dynamische [939].

earnings [1453, 849, 1992]. echelon [1305]. Ecological [1609]. Econom [1681]. Econometric [1761, 1770, 1762, 519, 624, 1769, 769, 954, 1525, 1234, 137, 352]. Econometrics [1146, 1048, 697, 461, 57, 1321, 1767, 1762, 1763]. econometrie [57]. econometrisch [624, 137]. Econometrische [352]. Economic [1212, 720, 1766, 34, 1876, 190, 2122, 116, 283, 339, 257, 282, 865, 679, 369, 295]. economic-technical [865]. Economics [1261, 902]. economisch [34, 865, 679]. economisch-technische [865]. Economische [295, 190, 116, 283, 257, 282, 369]. economy [954]. edge [1590, 1840]. edge-reinforced [1840]. Edgeworth [1343]. editing [1434]. editor [1530, 506, 1450, 1374, 1001, 1029, 1240, 749]. Editor-in-Chief [1001, 1029, 749]. Editorial [1833, 1300, 1943, 1975, 2004, 1678, 2044, 2053, 2075, 2093, 2130, 2163, 1736, 1773, 1598, 1625, 1760, 1849, 1914, 2021, 2106, 1900, 1721, 1783, 1698, 38, 1819, 2204, 2226, 2249, 1663, 1789, 1568, 2264, 2165, 2283, 1577, 1634, 1883, 475, 445]. editors [345, 1945, 905, 1320, 1347]. education [540, 571, 302]. Edward [1087]. Eeden [321]. een- [117]. Eendracht [134]. Eenige [25]. eentoppige [484]. Eenvoudige [6, 29, 291]. eenzijdige [484]. effect [2258, 2091, 1896, 1908, 2200, 1707, 2169, 2001, 229, 1445, 1070, 1758]. Effective [1545, 2262, 2108, 1939]. effectiveness [545, 1700, 1697]. Effects [1670, 1049, 1041, 2109, 1639, 1997, 2276, 827, 885, 2279, 1701, 2166, 1935, 1816, 2030, 1999, 2224, 1976, 2240, 2187, 2005, 1998, 2267, 2300, 1983, 1730, 1705, 1405]. Efficacies [1113]. efficacité [821]. efficacy [489]. Efficiencies [1618, 1058]. Efficiency [914, 1622, 369, 1195, 370, 1709, 360, 306, 1114, 1121, 359, 1384, 1425, 1276, 1091, 941, 693, 1506, 821, 367, 813, 369, 370, 359]. Efficiency-onderzoek [370]. efficiencymeter [367]. efficient [235, 2281, 2134, 2295, 2214, 981, 2243, 363]. Efficiënte [363, 235]. egocentric [2181]. eigenlijk [645]. eigenschappen [707, 390, 58, 488, 562, 166]. eigenvalue [1458]. einfache [824]. Einflussgrösse [824]. Eisenhart [1335]. elasticities [148]. election [1935]. elections [9]. electricity [1685]. electronic [511, 614]. elektronische [511, 614]. Elementair [183]. elementaire [152]. elementary [1575, 937, 152]. elements [811]. elevator [780]. ellipses [36]. elliptic [1940, 1906]. Embedded [931, 1130]. embedding [1311]. Emeritus [1118]. emperor [1503]. Emphasis [1605]. Empirical [1019, 1688, 1866, 1188, 1024, 1703, 1466, 1453, 1278, 1296, 1661, 1842, 1652, 995, 1036, 948, 897, 1085]. employment [2001]. empty [1640]. encompassing [1726]. encrypto [215]. encrypto-analyse [215]. encryptoanalysis [215]. endogenous [1957, 1708]. endpoints [2007, 2062]. enemy [1294]. ener [309]. energieverbruik [502]. energy [502]. engineering [595, 4]. England [3]. Enige [911, 132, 272, 719, 631, 351, 431, 407, 127, 294, 865, 271, 114]. Enkale [23]. Enkele [390, 462, 625, 579, 623, 175, 66, 489, 562, 77, 141, 340, 16, 444]. enkelvoudig [76]. enkelvoudige [110]. Enough [1046]. enquêtes [327]. Enschede [249]. Entrepreneur [284]. entries [393]. entropies [1981]. Entropy [1842, 2121, 2294, 1104]. Entropy-based [2121]. environment [1585, 1513]. environmental [831]. Epidemic [2180, 1455, 1922]. epidemiology [2167]. equal [2059, 1971]. equality [758, 1031, 1390, 218, 966]. equally [218]. equals [378]. equating [2028, 2257]. Equation [1170, 1988, 1896, 1831, 1393, 786, 1327, 1101, 1708, 1865]. Equations [1345, 1005, 2172, 2071, 2174, 1546, 182]. equi [669]. equi-probable [669]. equipment [626]. Equivalence [931, 1912]. equivalent [1780, 1512, 1882]. equivalent-estimation [1780]. erelid [782]. Eretoespraak [775]. Erhaltung [826]. Erlang [1837]. ermee [562]. Errata [83, 546]. Erratum [54, 1745, 63]. Error [102, 720, 2077, 2176, 60, 1997, 2082, 2196, 1700, 591, 1999, 1740, 2010, 2159, 1976, 1393, 2282, 97, 1998, 1546, 2135, 1322, 1966, 652, 74]. error-correcting [1546]. Errors [1940, 1146, 1005, 1635, 1323, 109, 1329, 1344, 91, 1788, 2023, 1101, 1932, 2070, 2192, 1225]. errors-in-variables [1323, 1344]. Ervaringen [618, 630, 911, 631, 865, 85]. erythematosus [1988]. est [1764]. estimate [788, 1420]. estimated [266, 1096]. Estimates [1613, 1178, 1169, 1385, 1054, 1768, 1793, 1377, 761, 934]. Estimating [1747, 1676, 1590, 1947, 1157, 1964, 1959, 2181, 1371, 1421, 1929, 1824, 1707, 1615, 2277, 1679, 2052, 2267, 1708, 1865, 1717, 2172, 574, 1694, 514, 1715, 207, 1158, 540, 768]. Estimation [1127, 810, 1887, 1200, 1673, 1968, 1182, 1484, 1518, 2136, 2209, 1214, 1904, 1613, 665, 786, 1389, 1937, 1382, 1129, 1813, 1112, 1162, 1864, 1005, 1368, 1443, 1572, 2077, 1591, 480, 1991, 1436, 477, 1737, 1592, 2201, 1066, 688, 2125, 2251, 1724, 1963, 1903, 2082, 1004, 2103, 1996, 1486, 1844, 1643, 1561, 1651, 1334, 1978, 2002, 1984, 1323, 1990, 1562, 1454, 1498, 1803, 1780, 1930, 1672, 811, 139, 1750, 2197, 1896, 508, 1318, 1448, 840, 1571, 2265, 1524, 1335, 861, 1027, 2079, 1659, 1955, 2071, 1638, 2160, 2155, 2153, 1526, 2193, 2237, 2217, 1715, 1627, 807, 2038, 2009]. estimation [2214, 754, 1692, 1208, 2015, 982, 1427, 1511, 103, 1010, 1260, 1704, 1496, 2300, 1799, 1983, 1852, 2041, 643, 621, 467, 717, 753]. Estimator [1133, 1137, 1446, 2196, 1296, 1897, 767, 795, 2291, 1845, 2281, 2293, 2102, 873, 2237, 1244, 998, 924, 878, 418]. Estimators [1527, 1693, 1195, 1156, 1866, 666, 608, 1560, 2176, 1714, 1848, 1315, 1941, 1703, 1515, 852, 1091, 1552, 1388, 1502, 1361, 1512, 532, 693, 1525, 1506, 1499, 2143, 1459, 801, 1913, 1369, 981, 1971, 2051, 2288, 1497, 948, 1025, 2309, 2104, 2219, 432, 435, 1931, 1085, 1225, 1362, 1425]. ethical [426]. ethics [1851]. ethische [426]. Ethnocentrism [1823]. etr [1681]. Ettinger [1051]. Eucalyptus [1810]. EURANDOM [1834, 1540, 1839]. European [1834, 1453]. evaluatie [457]. evaluating [828]. Evaluation [1282, 1911, 1431, 1666, 790, 638, 1393, 457]. evaluations [499]. even [912, 218, 889]. event [2197, 1707, 2170, 2146, 1445]. evidence [1948, 1964, 312, 1933, 1985]. evolutie [98]. Evolution [1147, 1766, 98]. ewewydige [131]. Exact [1537, 1912, 838, 1810, 1802, 2045]. exam [267, 296, 334, 387, 433, 478, 517, 564, 654, 647, 205, 204, 225, 308, 609]. Examen [205, 267, 296, 334, 387, 433, 478, 517, 564, 654, 647, 204, 225, 308, 609]. examination [95]. examining [1768]. example [720, 832, 829, 1342]. examples [431]. Exceedance [18, 484, 117, 333, 386, 747, 59]. Excellency [775]. Excellentie [775]. Exceptional [1228]. exces [27]. excess [1979, 27]. Exchange [1617, 1453]. exchangeability [1289]. excursion [1013]. existence [1369, 869]. Existential [176]. Existentiële [176]. exogeneity [2030]. expansion [1499]. expansions [1343]. Expectation [986, 635, 272, 990, 678]. expectations [2311, 686, 327]. Expected [2173, 2294]. expeditiebedrijf [147]. expenditure [1875, 639]. experience [926]. experiences [735, 911, 631, 865, 85, 618, 630]. Experiment [282, 1997, 313, 1889, 1999, 1998, 181, 325, 318, 314, 282]. Experimental [668, 315, 917, 632, 324, 312, 1398, 896, 476, 1342, 586, 671, 251, 320]. Experimentally [316]. experimenteel [632, 316]. Experimenten [465, 632, 235]. experimenter [317]. Experiments [1159, 1198, 2279, 250, 632, 2305, 1780, 235, 1387, 1976, 465]. Expert [1872]. expertise [1872]. explained [2153]. Explanation [13]. explanatory [1018, 2138]. Explicit [1053]. exploratory [1372]. Exponential [1166, 1750, 1512, 1253, 1297, 2157, 1562, 611, 811, 993, 1808, 1652, 1955, 966, 1629, 1757, 1933, 1085, 726]. exponentialiteit [717]. exponentiality [717]. exponentially [2217]. exponentiated [2157, 2144, 2197]. exponentiated-exponential [2157]. expository [529]. Extended [1259, 2301, 1917]. extension [1525, 983]. Extensions [1805, 944, 1798, 2151, 962]. External [1165]. Extra [1120, 688]. Extra-Poisson [1120]. Extracting [1531]. Extrapolatie [622]. Extrapolation [622]. extreme [1941, 1918, 761, 96, 1404]. extreme-value [1918]. extremen [732]. extremes [1264, 1742, 2132, 937, 732]. eye [724, 454]. ezel [777].

F [132, 394]. fabricage [132, 101]. fabricagemethode [258]. face [742]. facets [623]. facetten [623]. facto [324]. Factor [700, 2024, 976, 1330, 1667, 824, 294, 295, 1888]. factoranalyse [294, 295, 700]. factoren [137]. factorial [1639, 1997, 1999, 2162, 1976, 1998, 1749]. Factors [2028, 1605, 519, 324, 2133, 137, 1855, 2098, 1730, 1749]. factory [239, 448]. facts [1388, 92]. Fahrenheit [203]. Failure [1177, 1926, 1707, 2097, 110, 1788, 1743]. False [1168]. Families [1113, 1808, 1652, 1220, 148, 692]. Family [1191, 1873, 2261, 573, 1085]. familywise [1740]. farewell [1001, 1029, 905, 749, 475]. Farkas [1391]. farmacologisch [317]. Fast [1485, 1379, 1990]. fatigue [840]. Fay [2230]. feasible [880]. features [1437]. februari [558]. February [483]. feiten [92]. Feller [1167]. Few [1157, 1158]. fiber [562]. Field [1677, 1940, 126, 1640, 231, 37, 78, 283]. fields [2074, 2310]. fifth [145]. Fifty [1761, 1713, 1456]. Fighting [1294]. figure [276]. figures [755, 710, 64, 72, 78]. files [1351]. filling [288]. filter [961, 1820, 2222]. Filtering [960]. Finals [1919]. finance [1530, 1645]. Financial [2000, 2252, 1588, 449]. finding [385]. Finite [1019, 2060, 1130, 1692, 1546, 696, 1323, 1694, 1564, 2267, 1653, 1447]. First [1236, 1157, 1195, 1661, 694, 1445, 1158]. First-Order [1236, 1195]. Fisher [2100, 892]. fishing [202]. Fit [1046, 1159, 909, 1136, 2229, 1517, 1410, 1478, 19, 640, 641, 1413, 1422, 1383, 726]. fitted [245, 796]. fitting [752, 217, 476, 291]. five [1772]. fixed [1639, 1968, 1986, 2026, 1035, 1983, 1034]. flag [535]. flame [707]. Flanders [1786, 1823]. flat [2156]. fleet [703, 593]. Flexible [1303]. flight [1795]. flow [560]. fluctuation [1817]. FMS [1303]. focus [1841]. Focused [1963]. follow [2245]. follow-up [2245]. following [590]. Football [788, 1071, 2067, 301]. force [1446]. Forecast [1030, 17, 35]. Forecasting [721, 2306, 2116, 2012, 1872, 623, 759, 2061, 2193, 283, 2042, 662, 938, 190, 593]. Forecasts [1030, 301, 486]. foregoing [971]. foreign [1956]. Forensic [1928]. forest [1810]. forestry [366, 350]. Foreword [816, 1, 233, 84, 145, 56, 112]. form [2281]. Formal [1241]. FORMASY [938]. formed [665]. Forms [1134, 986, 977, 990, 1064, 1065, 717]. formula [1840]. formulas [1416, 652]. formulations [2242]. FORTRAN [830]. Forty [1140]. Forward [1581]. forwarding [147]. Foundation [446, 238]. foundations [61]. Fourier [2137, 1379, 1641]. Fourier-based [2137]. fourth [112]. fout [652]. fouten [109]. Foutentheorie [74]. foutenuitbarsting [591]. foutenverdeling [97]. Foutenwet [102, 60]. fowl [326]. fractile [761]. fractional [1297, 1433, 1266]. fractions [1714]. fragmentation [1500]. frailty [2161, 2216, 1691, 1826]. framework [1532, 2125, 1869, 1514, 2154]. Franklin [2299]. Franses [1861, 1860]. Free [1184, 306, 1298, 1052, 2254, 1072, 163, 159, 968, 1059, 157, 1581, 1445, 269, 621, 637, 838, 551]. French [821]. frequencies [1959, 70, 1421, 1516]. Frequency [1610, 1747, 1532, 126, 1531, 2181, 203, 1904, 1790, 66, 806, 2306, 75, 466, 270, 123, 27, 125]. frequent [1907]. frequentie [27]. frequentie- [27]. frequenties [70]. frequentieverdeling [203, 123]. Frequentieverdelingen [125, 126, 66, 75]. Frequentist [1609, 2313, 793]. frequentistische [793]. Friedman [983, 1082, 1122]. Frisch [151]. fugit [2183]. functie [57, 228]. functie-ponskaarten [228]. functies [217]. Function [1161, 1156, 1866, 1590, 57, 1296, 228, 1845, 2308, 861, 2277, 385, 1328, 1497, 2104, 1927, 442, 672]. Functional [1466, 1027, 2302, 1897, 1695, 1969, 1571, 1275, 1954, 2202, 2111]. Functionals [1169, 1454, 1498, 1435, 2065]. Functions [1368, 1227, 2302, 1091, 1397, 1388, 1535, 1732, 1633, 217, 2148, 673, 2295, 1715, 1811, 857, 1642, 995, 2069, 668]. Fung [1491]. further [1777]. Future [451, 1765, 2240, 1375, 1770, 565, 1412].

G [320, 775, 132, 1013, 1219, 931, 1921, 1759, 1583, 992, 1270, 1129]. Galton [151]. Game [750, 1842]. games [911, 738, 928, 740]. Gamma [1157, 2198, 2144, 2297, 2281, 473, 1364, 1338, 2208, 2156, 1158]. gamma-distributed [473]. gamma-verdeeld [473]. gang [243]. gap [288, 1403]. GARCH [1925]. garen [562]. Gas [1743, 707]. gasontladingsbuis [707]. gate [1795]. gated [1696]. Gaussian [2048, 1658, 2147, 1996, 1828, 1409, 1824, 2134, 2100, 1880, 2310]. GCL [795]. GDP [1859, 2124, 1861, 1860]. ge [562]. ge-sponnen [562]. gebit [98]. gebroken [467]. Gebruik [129, 147, 240, 624, 149, 156, 117, 394, 116, 894, 652]. gecorreleerde [435]. gedachte [146]. gedachten [176]. gedeelte [308, 334, 296]. gedrag [211]. gegeven [203, 768]. gegevens [912, 243, 541, 889]. gegewens [131]. Gegist [537]. gehele [203]. geheugen [774]. gehouden [179]. gekregen [31]. gelaagde [353]. gelijken [378]. gelijkheid [218]. gemeen [182]. Gemeente [249]. gemeten [736]. gemiddeld [110]. Gemiddelde [973, 651, 484, 265, 383, 704, 183]. gemiddelden [116]. gemodificeerd [479]. Geneeskundige [349]. geneesmiddelen [499, 779]. General [1134, 1869, 1227, 1572, 1815, 288, 1553, 625, 852, 2210, 2107, 1890, 1709, 1814, 2182, 1064, 1514, 425, 1579, 296, 308, 334]. generalisations [1928]. generalised [666]. Generalization [575, 787, 861, 2031, 1082]. Generalizations [1719]. Generalized [1160, 1649, 1192, 1187, 1684, 1193, 1554, 2308, 1741, 1933, 1817, 2041, 1122, 2018, 1848, 2020, 1957, 1309, 1401, 1891, 1870, 1387, 1979, 2067, 2289, 1594, 1096, 1358]. Generating [1939, 1514]. generation [806]. generators [2231, 1363, 1094]. genetic [1790]. Geometric [1215, 2064, 2157, 1547]. geometry [642]. geostatistics [1513]. geprogrammeerde [632]. gepromoveerd [321]. gereduceerde [142]. gerekt [685]. German [826, 824, 88]. Gerrit [216]. Gertrude [320]. gesimuleerde [504]. Gesuperponeerde [152]. getal [458, 452]. getallen [455]. getijdebeweging [629]. geval [171, 323, 458, 452]. Gevallen [455]. gevoelig [707]. gewarde [394]. gewichten [276]. gewichtscontrole [363]. gewogen [276]. gezinnen [148]. Gezond [779]. Ghana [1922]. GI [1270]. GI/G/1 [1270]. Gibbs [2015]. Gibbsian [1839]. Gijs [1893]. Gill [1047, 1472]. giro [13]. giroverkeer [13]. Gissen [534, 541, 542, 543]. given [203, 272, 665, 768, 230, 869]. GJR [2014]. glasbuizen [258]. glass [258]. GLMM [2073]. Global [1172]. GM [2291]. go [1712]. goederen [197]. goedkeuring [31]. Goldbach [846]. Goldberger [1768]. Golden [1450]. Goodness [2229, 1046, 1517, 1136, 1413, 1410, 1478, 19, 640, 641, 1422, 1383, 726]. Goodness-of-fit [2229, 1410, 1422]. goods [197]. governance [679]. graad [203]. graden [203]. grafiek [309, 116]. grafieken [711]. Grafiekenpapier [227]. Grafische [160, 574, 291, 397]. Grandfather [1139]. Grandson [1139]. graph [116, 227]. graphical [2048, 574, 944, 2189, 2040, 291, 397, 160]. Graphs [1131, 1337, 1792, 711, 1730, 1705]. Gravenhage [33]. great [820]. Greenwood [1250]. Grenander [2196, 2143]. Grenander-type [2196, 2143]. groei [704]. groei-indicator [704]. groeipercentages [711]. groepen [545]. grond [548, 231, 395]. grondslag [534]. grondslagen [61]. grondwaterstanden [629]. groot [217, 459, 544]. groot-heden [544]. grootheden [333, 443]. grootte [484]. grote [70, 218, 678, 347]. groundwater [629]. Group [1613, 1468, 602, 2007, 1857, 1757]. group-censored [1757]. grouped [1737, 1334, 665, 637]. Grouping [822]. groups [2066, 545, 602, 2007]. growth [861, 1382, 704, 711]. guaranteed [840]. guessing [543, 534, 541]. Guest [1577, 1634, 1945, 1320, 1374, 1347]. Guidelines [206]. Gulf [2290].

H [775, 1057, 1408, 41, 242]. H. [818, 817]. haar [137]. Hachelijk [675]. half [1417, 701, 1244]. half-Cauchy [701]. Hall [1920]. halve [701]. Hamaker [818, 1008, 1087, 1408, 841, 817]. handle [1942]. Handling [1669]. handmade [1690]. Handsome [675, 1690]. handy [1690]. Hans [1861, 1860]. harmonic [1977, 1547]. Hastings [1724]. Hausdorff [1641]. Hausman [2090]. Hautreaktionen [88]. hazard [2082, 2308, 2079, 1955]. Hazards [1133, 1622, 2298, 2268, 1442, 1788]. Hazelwood [509]. health [2166]. Healthy [779]. Heavy [1589, 1585]. heavy-tailed [1585]. hebben [182]. heden [544]. hedendaagse [58]. hedonic [2009]. heeft [31, 203]. Heerde [525]. Heerden [525]. heid [497]. height [1396, 276]. held [179]. Hellinger [2143]. hem [26]. Hemelrijk [1118, 749, 782]. Hemelrijks [740]. Hendrik [216]. Henri [1616]. her [321]. herd [326]. Herinneringen [774]. herleiding [78, 79, 158]. Herriot [2230]. herstel [2]. herziening [436]. heterogeneity [2185]. Heterogeneous [1614, 2144, 1542, 2278]. heteroscedastic [1880]. heteroscedasticity [2149]. heteroskedasticity [1870, 1384]. Hetschatten [540]. Heuristic [1793, 1488, 1304]. Hierarchical [1969, 1922, 2190, 2156, 2011, 1704]. high [1980, 2036, 1532, 2231, 2002, 1531, 2050, 1740, 822, 2306, 2253]. high-dimensional [2036, 2050, 1740, 2253]. high-frequency [1531, 2306]. high-risk [2231]. higher [1981, 869]. higher-order [1981]. hij [26]. him [26]. histograms [1649]. Historical [1142, 188, 541, 291, 141]. historische [188, 541, 291, 141]. history [1691]. Hodges [1113]. Hoe [459, 26, 26]. hoenderstapel [326]. Hoesel [1550]. hoeveelheid [355]. Holland [1505]. Hollerith [184]. Hollerith-machines [184]. Hollerith-masjiene [184]. homogeneity [149, 1702]. homogeneous [2066]. homogeniseringsinstallaties [610]. homogeniteit [149]. homogenization [610]. honorary [782, 775]. Hoofdcommissie [278]. hoofdredacteur [1001, 1029, 749]. hoogte [276]. hora [2183]. Horvitz [666]. hospital [2167]. hospitals [411]. houdsdiensten [498]. hours [147]. household [148, 901]. housing [556, 48]. Hubertus [1281]. Hugo [1408, 1008, 1087]. huishoudrekeningen [148]. hulpmiddel [70, 460]. hulpmiddelen [14, 24].

Hulst [1142]. human [98, 2187, 724]. humbug [80]. hun [163]. hunger [539]. hurdle [1979]. hurdle-generalized [1979]. hybrid [2306, 2289, 1708]. Hyperbolic [60, 102]. Hyperbolische [60, 102]. hypercalcemia [1445]. hypergeometric [1917, 919, 683]. hypothesen [117]. hypotheses [117, 1727, 1064, 1065, 784]. hypothesis [2271, 897]. hypothetical [1516].

ics [1681]. idea [146]. ideal [562]. ideale [562]. ideas [883]. identical [1302]. identically [1987, 2017, 2029]. Identifiability [2010, 2286]. Identification [1393, 945, 1325, 1329, 1442, 2155, 1317, 1629, 1350, 1928]. Identificeren [548]. Identify [548]. Identifying [1717, 2231, 1786, 377, 1500]. identity [2229]. idleness [2088]. IEEE [1836]. If [1620, 1742, 378]. ignoring [532]. II [155, 1551, 43, 2262, 735, 24, 830, 2207, 921, 1065, 1581, 293, 979]. IIa [988]. iid [2271]. III [1100]. Ikeda [666, 1137, 728]. Ill [1335, 1561]. ill-posed [1561]. Illustration [2242]. image [1827, 1995]. images [1521]. imitation [372]. Immunität [88]. immuniteit [159]. immunity [159, 88]. impact [1925, 2090, 1979, 1855, 1888]. Imperfect [449]. implementation [625]. implications [449]. importance [1571, 709, 2134, 136]. improduktiviteit [346, 399]. improper [1511]. Improve [1601]. Improved [2162, 2300, 1652]. improvement [1779]. Improving [2113]. impulse [1535]. Imputation [1665, 1666, 1667, 1994, 1664, 2008, 1434]. imputed [1718]. Imputing [1828]. INAR [1236, 2064, 2194]. INARCH [2039]. incidence [1542, 1956, 539]. inclusion [1079]. Income [1200, 736, 780, 148]. incompatible [1664]. Incomplete [1671, 838, 1755, 1651, 1448, 1994, 2188, 1449]. inconsistencies [272]. inconsistenties [272]. increase [347, 704]. increasing [442, 350, 353]. increments [1466]. incubation [2209]. indelingsprobleem [713, 702]. independence [2304, 1297, 2246, 1709, 2151, 1234, 728, 2190, 1266, 2013, 183]. Independent [1615, 1987, 2144, 1930, 1433, 2017, 2029, 2280, 995, 863]. Indeterminacy [976]. Index [1567, 1662, 1689, 1720, 1753, 1850, 1915, 1944, 2221, 1941, 649, 1515, 1803, 2136, 2042, 1496, 2309, 2149, 276, 755]. indexcijfer [276]. Indexcijfers [755]. indexes [2009]. India [179, 439]. Indian [380, 380]. Indicator [1161, 1868, 704]. Indices [1602, 1731, 2092, 1729, 206, 436]. indien [378]. indirect [1875, 1638]. indirectly [1089]. individual [2181, 1570, 2187, 2139, 1405]. Individualism [1823]. induced [2216, 1746]. Industrial [1774, 1452, 1152, 250, 659, 821, 270, 254, 565, 567]. Industrie [238, 251]. industriële [659, 254, 567, 565]. industrielle [821]. industries [303]. Industry [238, 89, 114, 251]. ineffective [704]. inefficiency [2033]. Inequalities [1593, 1009, 1957, 1410, 1506, 1317]. Inequality [1126, 1938, 1959, 1728, 2056, 736, 108]. infallible [2215]. infants [410]. Inference [2268, 1286, 1609, 2261, 882, 2137, 2180, 2304, 2152, 2172, 1102, 1588, 1309, 1633, 2313, 876, 1628, 1428, 2107, 1275, 2087, 1626, 2114, 2174, 1564, 2186, 1647, 945, 1674, 1932, 2260, 1899, 2070, 293]. inferences [1989, 1718, 2241]. infinite [694, 1420, 785, 1425, 678]. infinitely [2145]. infinity [1040]. Inflated [2223, 1994, 1877, 2235]. inflation [1877, 2012]. influence [1315, 2020, 831, 1682, 2208, 383, 2055, 410, 276]. influenced [629]. influencing [824]. influential [2203]. informal [1241]. informatie [413, 597]. informatietheoretische [736]. informatietheorie [407]. Information [2034, 2043, 2054, 2058, 2063, 2068, 2076, 2080, 2084, 2089, 2094, 2099, 2105, 2118, 2131, 2140, 2150, 2158, 2164, 2171, 2178, 2191, 2199, 2205, 2212, 2220, 2227, 2232, 2238, 2244, 2250, 2256, 2263, 2272, 2284, 2296, 2307, 2225, 1832, 1114, 1162, 1015, 826, 2022, 2027, 413, 688, 1959, 2049, 1531, 2210, 597, 1942, 790, 2100, 736, 1847, 407, 826]. information-theoretic [736]. Informative [1670, 2302, 2166, 2078, 1101]. infrequently [2009]. ingeschreyen [45]. inhomogeneous [1591, 1984, 1854, 1973, 2015, 1927]. injuries [949]. inkomenselasticiteiten [148]. inkomenslift [780]. inkomensongelijkheid [736]. Inleiding [381, 732, 614, 425, 280]. inner [1910]. Input [1041, 763, 771, 636, 908, 2280]. input-independent [2280]. input-output [636, 908]. input-outputanalyse [908]. inschrijvingsbedragen [447]. insertion [1304]. inspection [337, 110, 1858, 197]. inspections [2173, 302]. inspector [831]. inspectors [40]. instability [1817]. installations [610]. instance [843]. instelling [616]. Institut [100]. Institute [1761, 1762, 281, 380, 105, 420]. Instituut [105, 420]. instructive [825]. Instrumental [1737, 1953, 2010]. instruments [1957]. insurance [1379, 86]. Integer [1494, 1236, 1144, 2228, 1297, 203, 2285, 1433, 2218, 1479, 1266]. Integer-Valued [1236, 2228, 2285, 2218]. Integrals [1014, 1902, 1632]. integrated [2215]. integrated-likelihood-ratio [2215]. integration [1385, 1938, 2127, 1950]. intelligence [302, 244]. intelligentie [302]. intensiteit [340]. Intensity [2015, 1518, 1880, 340]. intensity-dependent [1880]. intensive [2097]. interaction [1591, 1997, 1052, 1318, 1999, 1976, 1998, 1474]. Interarrival [1800]. intercept [1703, 1702, 1692, 2003]. intercept-only [1703]. Interdisciplinary [1621]. interest [519, 1520, 1746]. Interior [1458, 768]. interlokaal [651]. intermediate [1496]. intermodal [1939]. Internal [604]. Internationaal [568, 105]. International [100, 463, 179, 2066, 460, 568, 338, 105]. Internationale [338, 460, 179]. Interne [604]. interobserver [1063]. Interpolation [1269]. interpret [2109]. interpretation [976]. interpreteeren [6]. interpreting [1002, 1683, 6]. interrupties [648, 578]. interruptions [871, 648, 578]. intertemporal [849]. Interval [1106, 1184, 471, 1482, 1454, 1498, 1930, 2173, 1435, 2261, 728, 684, 261, 2215]. Intervals [1183, 1114, 1424, 957, 1298, 800, 2090, 2200, 770, 245, 265, 1882, 2240, 968, 496, 1090, 1070]. intrinsic [2310]. Introducing [2182, 1540, 1550]. Introduction [1678, 1607, 1265, 1833, 1300, 1736, 1773, 1530, 1760, 2106, 1900, 1721, 1783, 1698, 1819, 425, 1240, 1789, 1960, 937, 1888, 280, 381, 732, 614]. Invariance [1137, 1491, 1473, 1508, 754]. invariant [2239]. Invariantie [754]. Inventory [847, 1301, 1053, 785, 802, 1579, 627]. Inverse [2255, 569, 1655, 2198, 1561, 2035, 919, 1954, 1338, 998]. Inverse-probability-weighted [2255]. inversion [184, 1641]. Investeren [756]. Investigating [1979]. investigation [790]. Investing [756]. investment [531]. invloed [383, 410, 276]. involving [1645]. inzet [604]. inzet-afzetanalyse [604]. IPPS [1079]. Ir. [404, 1057]. irregular [1382, 1843]. irregularly [1825, 1903]. IRT [2028, 1729]. ISI [1977]. island [1928]. isolated [1857]. isometric [2230]. Isomoor [361]. isomor [361]. Isotonic [1388]. Issue [2022, 2027, 2034, 2043, 2054, 2058, 2063, 2068, 2076, 2080, 2084, 2089, 2094, 2099, 2105, 2118, 2131, 2140, 2150, 2158, 2164, 2171, 2178, 2191, 2199, 2205, 2212, 2220, 2227, 2232, 2238, 2244, 2250, 2256, 2263, 2272, 2284, 2296, 2307, 1599, 1441, 1945, 1530, 2179]. Issues [1148, 1794, 1699, 1784]. Itô [1902]. Italian [2123]. Italy [639]. items [955]. itemsets [1907]. iterated [1466, 1384]. iteratiemethode [574]. iteration [574, 900]. Iterative [1286, 1668, 1102, 1492, 1802]. IV [1536, 1983].

J [357, 1118, 906, 118, 550, 490, 749, 782]. J. [648]. j.l. [99]. Jaap [1894, 1578]. jaar [253, 252, 254]. jaargang [233, 84, 196, 145, 56, 332, 112]. jaarlijkse [711]. jaarrekeningen [553]. jaartotalen [710]. Jackknife [1552, 1803, 1208]. Jan [1876, 1051]. jaren [521]. Jasso [1829]. Jenkins [2061]. job [1758]. job-training [1758]. jobs [1302]. Johannes [1281]. Joint [1123, 2294, 949, 2097]. jointly [1714]. Jong [906]. journal [1977]. journals [1855, 1888]. Jubilee [1450]. judged [1063]. Judgment [675]. Juli [99]. July [374, 99]. Jump [1617, 1903, 1984, 1965]. Jump-diffusion [1617]. jumps [2288]. June [216]. Juran [118]. Just [91]. Justesen [1057].

K.B. [31]. Kalman [1448, 961]. kan [28]. Kans [739, 257]. kansellipsen [36]. kansen [43, 218, 487, 884, 888, 207, 246, 516]. kansspelen [738, 740]. kansverdeelingen [27]. kansverdelingen [444]. Kantorovich [1506]. Kaplan [1133]. karaktertype [15]. Karmarkar [1458]. Keen [1493]. keeping [1740]. Kendall [273, 397]. kenmerken [244, 538]. kernel [1091, 1750, 2051, 1647, 1799, 1852, 2041, 1931]. kernel-based [1799]. kernels [2151]. ketens [560, 902]. Keuls [479]. keuring [197]. keuringsschema [110]. Keus [257]. keuze [258, 542, 276]. keuzetest [386]. keys [657]. keystroke [1879]. kiezen [408, 896]. klassebreedte [203]. klassieke [562]. kleding [347, 52]. kleding-maatnormalisatie [52]. Klein [1768]. kleine [218]. kleinste [291, 219, 182]. Klimatologie [65]. klimatologische [79]. kliniek [489]. Klinische [457]. Kloek [154]. KMT [1410]. KMT-type [1410]. knoopsel [408]. knopen [408]. knot [408]. knots [408]. knowing [494, 500, 504]. knowledge [495]. Knut [37]. Kolmogorov [1409]. Koninklijke [31]. konsumenten [594]. konsumenten-ondenoek [594]. kort [591]. korte [283]. kosten [33, 746]. kostenbasis [680]. kostenbeheersende [472]. kostprijs [414]. Kotz [1829]. Koyck [530, 1713]. KPSS [1719]. KPSS-test [1719]. kraaltjes [848]. Kronecker [697]. Kuh [1020]. Kuiper [1478, 764, 797]. Kullback [1959, 1563, 2052].

Kumaraswamy [2223]. Kumasi [1922]. kunstmatige [70]. kursussen [911]. kurtosis [2016, 1208, 589]. kwadraat [752]. kwadraatsom [973]. kwadraten [291, 219, 182]. kwadratische [717]. kwaliteitsbaheersing [24]. Kwaliteitsbeheersing [171, 14]. Kwaliteitsbeleid [566, 567, 565]. kwaliteitscontrôle [34]. Kwaliteitsdienst [238]. Kwaliteitsnormen [236]. kwaliteitsprobleem [894]. Kwantielen [657]. Kwantificering [214, 213, 215, 580]. kwantitatieve [324]. kwantitatievestochastische [574]. kwartaalcijfers [710].

L [530, 1123]. L. [154]. label [2115]. labelling [1521]. labor [354, 350]. laboratorium [317]. laboratory [44, 317]. Labour [1005, 1446]. lack [1491, 1473]. ladder [1396]. lading [848, 535]. land [118]. landbouw [193]. landbouw-opbrengsten [193]. landbouwmethoden [598]. landen [540]. lane [1910]. lange [17, 35]. Language [2115, 215, 582]. Laplace [2201, 691, 730]. Large [1530, 1465, 1640, 1533, 1066, 1534, 941, 70, 217, 678, 2127, 1011, 1372, 347, 459]. large-data [2127]. largest [1507, 2144, 1426, 2278]. larynx [1095]. larynx-cancer [1095]. last [2290]. lateness [926]. Latent [1047, 1680, 949, 1321, 2185, 1822, 1626, 1326, 1325, 2098, 1288, 935]. later [1767]. Latijns [379]. Latin [379]. lattice [1426]. Lattices [1469, 1995]. laude [321]. Law [102, 2145, 1066, 60, 993, 590, 678, 740, 738]. ’lawfulnes [783]. laws [1648, 1466, 1465, 1364]. layer [1022]. layered [353]. layout [702]. layouts [1052]. laziness [1241]. League [788]. learning [1301, 2306, 2189, 780, 632]. Least [1572, 1186, 2221, 1066, 1848, 2242, 245, 693, 2059, 909, 2102, 1096, 2243, 2041, 291, 219, 182]. least-squares [2221, 291]. leemtevulling [288]. Leerprocessen [632]. left [1640]. legal [1683, 742]. legerkeuringen [302]. legit [1422]. Lehmann [1113]. Leibler [1959, 1563]. length [1424, 1417, 1078, 1986, 2167]. lengths [728]. leren [780]. letaliteit [44]. lethality [44, 61]. Lets [148]. Leve [1893]. level [758, 1990, 1806, 1637, 2133, 1847]. levels [1941, 1343, 192, 473, 629]. Levensbericht [1057]. Levensduur [650]. Levensduuronderzoek [659]. levensonderhoud [33, 436]. leveringen [604]. levertijd [473]. levertijden [243]. Lévy [2137, 1903, 1904]. Lévy-driven [2137]. lexical [1421]. lezers [122]. LIBOR [1901]. Library [422]. licht [707]. Liebler [2052]. life [650, 2173, 840, 820, 1078, 2211, 1578, 1757]. life-length [1078]. Lifespan [659]. lifetime [2216, 849]. lifetimes [1911]. light [707, 724]. light-sensitivity [724]. Lijnvereffening [39]. Likelihood [1113, 987, 1195, 1866, 1601, 1005, 2270, 2201, 2125, 1724, 1486, 1309, 2268, 1335, 2293, 2033, 2057, 2169, 2038, 1328, 2086, 948, 2215]. likelihood-centered [2293]. likelihoods [1723, 1647]. limietstelling [124]. Limit [1167, 1361, 2006, 1156, 724, 1636, 1013, 1483, 2143, 2088, 1364, 1905, 110, 124, 393]. limit-distributions [393]. limitation [498, 1544]. limitations [2242]. Limiting [1482, 931, 1310, 442]. limits [1543, 1838, 1367, 333, 2132, 551, 59]. Lin [1982]. line [527, 245, 1986, 1074, 1791, 39]. lineair [467]. Lineaire [889, 94, 908, 260, 151, 912]. Linear [1333, 1684, 1115, 1186, 1207, 1185, 1194, 1172, 1015, 2286, 666, 2018, 763, 771, 467, 1560, 1649, 1315, 2020, 1343, 1269, 1486, 1875, 852, 1308, 1309, 1401, 1323, 1728, 2312, 1870, 734, 94, 1435, 1387, 1912, 1039, 2056, 908, 260, 693, 1831, 151, 909, 1506, 2071, 2087, 2057, 920, 2240, 1389, 961, 2202, 2174, 1065, 1317, 857, 1344, 1813, 824, 1805, 1096, 1358, 1496, 573, 900, 1413, 1899, 2070, 2219, 2253, 291, 1384, 760, 912, 889]. lineare [824]. linearity [1362]. linearly [1840]. link [2139, 96]. Linkage [1951, 1929]. linked [2152, 2172]. Linking [2175]. links [2009]. Linnik [1648]. liquid [2266]. Lisman [1281, 475]. LISREL [1448, 1326]. List [1111, 1667]. literature [351]. literatuur [351]. litigation [1681]. Littlewood [1436]. Liu [2291]. Liu-type [2291]. living [33]. LM [1546]. load [848, 535]. loading [840]. loans [1712]. Local [1592, 1732, 1272, 2055, 1231, 1548, 1636, 2202, 2208, 340]. Locally [2214, 1369]. locating [1397]. Location [1502, 1991, 1023, 1058, 1499, 1377, 2261, 934]. location-scale [1058]. locations [1786]. Log [1207, 2233, 2145, 1996, 1844, 1650, 1728, 2312, 2087, 1880, 1916, 2261, 642]. log-Birnbaum [1916]. log-Gaussian [1996]. Log-Linear [1207, 1728, 2312]. log-location-scale [2261]. log-normal [642]. log-returns [2145]. Log-symmetric [2233, 2087]. logarithm [1466]. Logarithmic [2192, 1314, 711]. logaritmisch [711]. logical [1430]. Logistic [2275, 1604, 1257, 1277, 2176, 1651, 1258, 712, 861, 2133, 1787, 1383, 1012, 1360]. Logistics [1300]. logistische [712]. Logit [1261, 1260]. loglikelihood [1524]. loglog [1465]. Lognormal [1605, 390]. lognormale [390]. logrank [2255]. logratio [2230]. lokaal [340]. loketten [481]. Lomax [1316]. Long [2290, 2216, 2091, 1635, 651, 845, 1589, 17, 35]. long-distance [651]. long-memory [1635]. long-range [1589]. long-tailed [845]. long-term [2216, 2091, 17, 35]. Longitudinal [1670, 2206, 1669, 1448, 1671, 1441, 1981, 1718, 2166, 2097, 1612, 2245, 2007, 2062, 1895]. look [2002, 2148, 846]. Loor [483]. loose [774]. loss [1832, 2143, 2069, 2104]. losse [774]. Loten [925]. lotingsas [746]. lots [925]. lottery [1495]. low [1532, 2002, 822, 2306, 1855]. low- [2306]. lower [1887, 1796, 2059, 879, 869]. LP [1796]. LP-based [1796]. LR [2083]. Lupus [1988].

M [320, 357, 530, 1123, 1585, 1997, 118, 550, 1696, 490, 1918, 1910, 1123, 1585, 1013, 1219, 931, 1837, 1921, 1759, 1696, 1583, 992]. M/D/s [1837]. M/G/ [1583]. M/G/1 [1013, 1219, 931, 1921, 1759, 992]. M/M/1 [1585, 1696]. M/M/L [1123]. maakt [134]. maatnormalisatie [52, 347]. maatschappelijke [253]. maatstaven [736]. machinaal [578]. machine [492, 555, 616, 926, 2306, 239, 355]. machinebouw [4]. machinefabriek [239]. machinegereedschap [355]. machinery [513]. machines [184, 1302]. Macht [210, 134]. Macro [1938, 2127]. Macro-integration [1938, 2127]. macromoleculen [94]. macromolecules [94]. made [42, 89].

Madedelingen [89]. magazine [122]. magic [1840]. magnitude [2145]. magnitudes [1251]. Main [278]. maintenance [498, 513]. major [820]. make [256, 542]. maken [538]. making [2152, 1989, 535, 542, 543]. malaria [1908]. male [1095]. Management [1763, 57, 1795, 911, 1992, 119, 1660, 627, 449]. manager [241]. managers [256]. MAnpower [938]. Mantel [2301]. manufacture [132, 101]. Manufacturing [1303, 258]. many [839, 1271]. map [2040]. mapping [1695]. March [154]. margarine [363]. Marginal [2312, 1723, 2298, 1724, 828, 508, 768, 1416]. marginally [1807]. marginals [857]. markers [2047]. Market [568, 661, 480, 284, 280, 2290]. marketing [1533, 1710, 1530, 1830]. markets [2252]. marking [1880]. Markov [560, 1680, 902, 1573, 1965, 2072, 2074, 2151, 1130, 639, 2186, 2310, 2267, 900, 1581, 928, 1153]. Markov-ketens [560, 902]. Markovian [1053]. Markowitz [1252]. Marktanalyse [280]. Marktonderzoek [568, 284]. Marktregulering [661]. martingalen [686]. martingales [686, 1416]. masjiene [184]. mass [258, 2266]. mass-produced [258]. massaal [91, 90]. massaproduct [258]. massive [91, 2247, 90]. match [2083, 2174]. matchen [199]. matches [1814]. Matching [1668, 1985, 220, 2008, 199]. materiaal [6]. material [2295, 1966, 6]. materials [1630]. materieel [626]. math [6]. Mathematical [959, 1974, 1920, 74, 175, 391, 1545, 2114, 143, 681, 592, 77, 141, 252]. mathematical-statistical [77]. mathematics [1379, 407, 1294, 607]. mathematisch [391, 77]. mathematisch-statistische [77]. mathematische [175, 141]. mathode [6]. Matrices [1130, 673]. Matriks [184]. Matriks-omkering [184]. Matrix [697, 1974, 2096, 2224, 786, 673, 1937, 768, 853, 184, 569]. matter [72, 95]. Maxima [484, 1061, 2254, 484]. maximal [1593, 1882]. maximin [862]. Maximum [2201, 2125, 987, 1195, 1104, 2038, 1005, 1446, 1486, 2294, 993, 926, 1335, 2293, 2033, 2057, 863]. may [1712]. MCAR [2086]. MCMC [1723, 1626, 2114, 1664, 1628]. ‘MCMC-based [1628]. Mean [1536, 2176, 1100, 1003, 996, 964, 979, 1980, 1801, 1812, 1977, 2251, 1507, 1917, 2142, 2248, 1397, 2294, 1631, 529, 2010, 764, 1847, 797, 924, 878, 2008, 183, 973]. mean-reverting [2010]. Means [1162, 2311, 1822, 1518, 870, 1054, 862, 1547, 59, 1383]. measurable [538]. measure [1475, 1563, 1474]. Measured [1126, 2082, 1089, 736, 1986]. Measurement [1198, 714, 1026, 1322, 169, 1329, 170, 2010, 1393, 351, 343, 1810, 1382, 1788, 695, 2070, 2192, 495, 1456, 159]. measurements [2255, 1390, 42, 2245, 454]. measures [1981, 1593, 2049, 2016, 915, 1480]. Measuring [2024, 1600, 1459, 1063, 494, 1351, 597, 504]. mechanical [4, 578]. mechanism [1798]. medal [1008, 1087]. Mededeelingen [11, 20]. Mededeling [69, 506, 135]. Mededelingen [47, 55, 62, 82]. Media [281, 501]. Median [1100, 1003, 1536, 744, 996, 1059, 964, 743, 979]. medians [1011]. mediation [2236, 2274, 2170]. Medical [349, 423, 1599, 1106, 426, 214, 588, 411, 587, 300, 618]. medical-ethical [426]. Medicine [1151, 427]. medicines [499]. medicus [587]. medisch [426, 214, 423]. medisch-ethische [426]. medische [588, 411, 587, 300, 618]. meer [492, 555, 182, 481]. meer-loketten [481]. meer-machine-bediening [492, 555]. meerjarige [755]. meervoudige [382]. Meest [434]. meetbare [538]. Meetings [179]. Meets [421, 1467]. mega [1531]. mega-panels [1531]. Meidell [108]. Meier [1133]. melk [448]. melkproduktenfabriek [448]. member [782]. Memoriam [1894, 154, 357, 374, 404, 1754, 1893, 1281, 216, 483, 530, 1408, 1493, 222, 242, 1357, 1051, 490, 1181]. Memories [774]. Memory [2290, 1635, 774, 906]. men [28, 26]. Mendes [1997, 1998]. mengproblemen [125]. mens [98]. Merkkeuze [929]. messaging [340]. meten [494, 597, 495, 504]. meter [367]. metgezel [452]. Method [735, 1361, 1831, 511, 288, 2103, 574, 836, 1245, 712, 132, 142, 1896, 514, 1252, 1524, 1638, 2277, 258, 2051, 37, 2117, 1096, 791, 1025, 101, 1799, 6, 291, 269, 397, 219, 444, 760, 88, 1866, 182]. Method-of-moment [1831]. Methode [182, 288, 132, 142, 37, 101, 291, 269, 397, 219]. Methoden [207, 377, 601, 163, 370, 525, 520, 157, 303, 77]. methodiek [436]. Methodological [1946, 1157, 1685, 1675, 1158, 1945]. methodology [1803, 436]. Methods [1048, 1669, 987, 828, 959, 1601, 1107, 1043, 1666, 1963, 944, 1766, 377, 601, 163, 1458, 370, 525, 916, 921, 620, 1559, 2170, 2306, 2168, 1544, 2114, 2189, 520, 982, 1310, 2121, 157, 1427, 303, 2300, 77, 594, 598, 207]. Methorst [242]. Meting [159, 343]. metingen [42]. metoden [594]. metric [485]. Metropolis [1724]. Mexico [2124]. mg [410]. microaggregation [1813]. microarray [1841, 1784]. microdata [1353, 1351, 1349, 1352, 1350]. middel [303]. middelen [457]. middle [2251, 1407]. middle-aged [1407]. Midzuno [666, 1137]. Militair [349]. Militaire [262]. Military [349, 262]. milk [448]. millennium [1568]. Minimal [1161, 975]. Minimax [605, 1214, 1967, 1073, 1887, 1592, 1362]. minimization [720, 449]. Minimizing [926]. minimum [1996, 1027, 1638, 2154]. Mining [1907]. Minkowski [1391]. Misclassification [1127, 1940, 1324]. misclassified [2038]. Mises [1361, 793]. missing [2255, 1832, 1853, 1942, 1522, 2202, 1802]. Missingness [1670]. Misspecification [1285]. mistakes [542]. mittels [88]. Mixed [1207, 2166, 2262, 2276, 1309, 1401, 1387, 1667, 1492, 1442, 1524, 1370, 2057, 2187, 1064, 1065, 2219]. Mixed-effects [2166, 2276]. mixed-outcome [1667]. Mixing [1025, 1066, 125]. Mixture [1648, 2270, 1710, 920, 1522, 2074, 1653, 573, 2260, 2266]. mixtures [2060, 1364]. ML [1651]. ML- [1651]. Mode [1100, 1003, 1536, 1083, 1897, 964, 979]. Model [1045, 1670, 1261, 1617, 1174, 1622, 1965, 1778, 1185, 1194, 1878, 1605, 1866, 1583, 1129, 1178, 1679, 1005, 2219, 1730, 2299, 2077, 2270, 2236, 912, 2221, 2161, 1436, 1722, 1711, 1503, 1710, 1786, 2231, 2276, 2258, 2216, 1703, 1446, 2298, 1825, 1963, 1822, 852, 2196, 1853, 1309, 2157, 980, 1892, 1323, 1990, 935, 1691, 1672, 919, 811, 2197, 1387, 1448, 2030, 1667, 1702, 1492, 954, 1726, 1967, 1824, 1442, 391, 2230, 1524, 2246, 1335, 1525, 1877, 443, 2138, 2081, 1961, 2071, 2061, 2160, 2088, 1768, 2057, 502, 1742, 2168, 2217, 1715, 2114, 2133, 2146, 1064, 1979, 439]. model [1692, 2156, 2177, 2289, 878, 2135, 2019, 785, 802, 1823, 1864, 1096, 1674, 593, 2139, 1260, 2014, 1413, 1422, 1960, 2260, 578, 2266, 1288, 962, 1705, 2003, 889, 1384, 1962, 2012, 481, 1420]. Model-assisted [1866]. Model-based [1778]. Modeling [1533, 2185, 1612, 1885, 2042, 2109, 1532, 2145, 2064, 1988, 1588, 1785, 1896, 1398, 2207, 2097, 1922, 1820, 1901, 1821, 2101, 1342]. Modellen [559]. Modelling [1680, 1700, 1614, 1908, 1686, 1071, 1956, 1814, 2110, 1049, 2085, 1825, 1920, 2112, 2046, 1541, 2067]. Models [1684, 1200, 1610, 1756, 1675, 1766, 1120, 1207, 1159, 1259, 1285, 1665, 1326, 1449, 1015, 1966, 1242, 1441, 2048, 1572, 2018, 1431, 2236, 2201, 1411, 1848, 1315, 2198, 2180, 1657, 1953, 2251, 2020, 883, 2185, 2103, 1486, 1651, 1588, 1701, 1308, 1401, 2242, 1729, 2183, 1473, 1728, 2166, 1807, 2136, 1378, 969, 1816, 1329, 949, 1272, 2060, 2056, 2134, 693, 1942, 1882, 2107, 2224, 1890, 1506, 1707, 2159, 2100, 2087, 920, 2240, 1393, 1690, 1522, 1327, 1389, 2074, 961, 1626, 2169, 1460, 1234, 796, 2277, 1916, 2174, 2182, 1065, 2190, 2214, 2235, 2233, 1344, 2188, 2042, 1895]. models [1805, 1647, 1864, 1579, 1358, 727, 1704, 2292, 2167, 1653, 1708, 1295, 1960, 1932, 2309, 1899, 1983, 1787, 2070, 2149, 2192, 2219, 2286, 2253, 1730, 1445, 2116, 1383, 865, 1865, 901, 1405, 1447, 662, 559, 1491]. moderate [877]. modes [1978]. modification [1729]. Modified [2134, 479]. modulo [1404, 804]. moeilijkheden [71]. Mogelijkheden [71, 136]. mogelijkheid [77]. mogen [394]. Moivre [708]. Moment [1990, 1913, 957, 1957, 1002, 1739, 1831, 13, 694]. momentopnamen [240]. Moments [1494, 1157, 1593, 767, 795, 1970, 2016, 977, 1547, 1225, 869, 1158]. Monitoring [1631, 1847, 1779, 261]. Monographs [336]. Monotone [1522, 1085, 2270, 1887, 2221, 2145, 1388, 2136, 1537, 2079, 2086, 948]. Monotonicity [1213, 1290, 2268, 1033]. monotonized [1036]. Monotonizing [948]. monsternemen [49]. monsterneming [136]. Monte [2048, 1385, 2166, 372, 1965, 2057, 386, 1112, 621, 668]. Monte-Carlo [386]. Monte-Carlo-bepaling [386]. monthly [1449]. Monty [1920]. Morgen [566, 567]. mortality [1395, 2097]. mortgages [1797]. Mortibus [1251]. mosaics [790]. most [806, 434]. motivated [1364]. motley [933]. movement [2112, 1485, 629]. movements [735]. mover [847]. moving [2137, 2235, 116, 1101]. moving-average [2137]. Mozart [750]. Mr [41]. MRAS [1420]. MSE [1276]. Multi [1584, 1889, 1951, 2187, 1385, 1886, 1582, 755, 1330, 1023, 1305, 2115, 240, 481]. multi-criteria [1385]. multi-echelon [1305]. multi-label [2115]. Multi-Linkage [1951]. multi-momentopnamen [240]. Multi-sample [1889, 1330, 1023]. Multi-server [1584]. multi-snapshots [240]. multi-stage [1582]. multi-station [481]. Multi-subject [2187]. multi-type [1886]. multi-year [755]. multicollinearity [2293, 2243]. multidimensional [2185]. multifold [2052]. multilateral [2285]. multilevel [1835, 2236, 1703, 2274, 1701, 1896, 1699]. Multilinear [1739]. Multimodality [573, 1978]. Multinomial [1630, 1368, 1693, 2305, 375, 1422]. Multiple [1088, 1177, 1994, 1665, 1310, 1982, 555, 477, 1948, 1666, 1508, 468, 1667, 1937, 1664, 1101, 1708, 1558, 382]. Multiple-dose [1310]. multiple-objective [1558]. Multiplicative [827, 2159, 878, 2247, 2192, 2219, 2286, 885]. Multiplicity [1794, 1157, 1784, 1158]. multiply [1718]. multisample [1801]. multistate [2078]. multistation [481]. Multivariate [1906, 1502, 1619, 1665, 1412, 1980, 2161, 1592, 1812, 819, 1503, 1710, 1925, 2092, 1974, 1738, 2305, 1716, 1870, 1542, 2281, 2230, 2246, 1275, 1742, 1735, 1937, 1377, 1642, 924, 1427, 1864, 2098, 1221, 1744, 1065]. Municipality [249]. Murphy [1185]. music [66, 224, 247]. Musical [750]. Must [1620, 1742]. mutual [2210, 368]. muziek [66, 224, 247]. Muzikale [750]. mvu [1369].

’n [131]. naam [379]. Naar [601, 235, 479, 132, 489, 410, 45]. Nabeschouwing [128]. nabootsing [372]. nagedachtenis [906]. name [379]. Naschrift [41, 38]. Nationaal [501]. National [2124, 177, 2123, 1771, 2129, 501]. Nationale [177]. Nationalism [1823]. nations [1330]. natural [494]. Nature [161]. natuur [161]. natuurkunde [50]. Natuurwetenschap [494]. nauwkeurigheid [42]. nearer [1655]. Nearest [2202, 1863, 1518]. Nearly [2045]. Nebenbedingung [824]. necessary [107, 59]. Nederland [278, 650, 703, 908, 137]. Nederlands [420]. need [540, 2222]. Neerlandica [1997, 1450, 1998, 699, 1567, 1662, 1689, 1818, 1850, 1884, 1001, 1029, 2179, 1407, 905, 2273]. negatieve [443]. Negative [1192, 2305, 443, 375, 729, 798, 2194]. Negatives [1168]. negende [233]. Neglected [1447, 1546, 1405]. neighbor [1863, 1518]. neighboring [2258]. neighboring-provinces [2258]. neighbors [2202]. neighbourhood [1682]. nemen [525]. NEN [420]. neoclassical [2024]. nequal [1389]. Nested [1664, 1875, 1707, 2184]. nests [1070, 1038]. Net [91, 98, 218, 602, 452]. Netherlands [278, 999, 1006, 1075, 1080, 1109, 1116, 1124, 1154, 1196, 1203, 1216, 1223, 1237, 1255, 1262, 1267, 1273, 1279, 1291, 1312, 1331, 1339, 1354, 1365, 1414, 1439, 1489, 1509, 1528, 1556, 1595, 1606, 636, 650, 703, 1163, 1452, 1380, 1140, 1209, 1229, 1461, 908, 420, 2127, 1179, 137, 1092, 1055, 1444, 1248]. nets [1586]. Network [638, 2184, 2179, 1587, 2201, 2183, 2181, 1580, 2187, 2186, 2188, 1589]. networks [2180, 2185, 1756, 1336, 2113, 1796, 1337, 2190, 1733]. Neumann [389, 561]. neural [1587]. news [1925]. next [1445]. Neyman [1734, 879, 1136, 784]. Neymanian [2162]. Niet [435, 722, 677, 538]. niet-normale [677]. Nieuwe [388, 563, 649, 309, 779, 598]. ninth [233]. NMES [1664]. No [31, 471]. No-Change [471]. node [2185, 2184]. Nogmaals [740]. noise [1658, 2258, 1409, 87]. noisy [1903]. nomografisch [138]. Nomogram [479, 19, 333, 711, 479, 19]. nomogrammen [344, 388, 563, 394]. nomograms [344, 388, 563, 394]. nomographically [138]. Non [1591, 1903, 1186, 1133, 2078, 1668, 857, 1213, 1368, 1980, 1658, 2147, 1875, 1828, 1363, 1987, 1897, 1695, 677, 1970, 1824, 2134, 693, 909, 2071, 2017, 2029, 1802, 2019, 1895, 1839, 2045, 1581, 1932, 1383, 733, 1865, 538, 760, 672]. Non- [1591]. non-association [733]. non-central [672]. Non-Differentiable [1368]. non-Gaussian [2147, 1828, 1824, 2134]. non-Gibbsian [1839]. non-identically [1987, 2017, 2029]. Non-Iterative [1668, 1802]. Non-Linear [1186, 693, 909, 2071, 760]. non-measurable [538]. non-nested [1875]. Non-normal [857, 677, 1970, 1865]. non-normality [1980]. Non-parametric [1903, 2078, 1213, 1897, 1932]. Non-proportional [1133]. non-random [1895]. non-randomized [2019, 2045]. non-standard [1581]. non-uniform [1363]. non-unique [1695]. noncentral [2234]. none [950]. none-simple [950]. nonhomogeneous [2113]. nonignorable [2303]. nonindependent [2159]. Nonlinear [1048, 1821, 2149, 1043, 1066, 2231, 1275, 1704, 1523]. Nonlinearity [915, 1546]. nonnegative [2277, 1097]. Nonparametric [1183, 1643, 1675, 1334, 1484, 1633, 2160, 2155, 1559, 670, 1208, 1497, 1856, 1801, 1635, 837, 1023, 1807, 2210, 2265, 2107, 1826, 2195, 1090, 1858, 2135, 1704, 1799, 1474]. nonredundant [1317]. nonresponse [2303]. nonstationary [1811]. noodzakelijke [107, 59]. Norm [420]. normaal [333, 473]. normaal-verdeeld [473]. Normal [1856, 1113, 1134, 986, 1170, 1227, 608, 2225, 744, 1507, 2073, 266, 1738, 677, 2142, 2248, 2305, 870, 1031, 1970, 683, 265, 977, 990, 1327, 620, 857, 924, 1563, 573, 642, 1852, 743, 152, 395, 183, 1865, 668]. normale [608, 677, 265, 620, 152, 395, 183]. normaliratie [120]. Normalisatie [278, 371, 420]. Normalisatie-instituut [420]. normaliteit [718, 717]. Normality [1195, 1980, 1848, 1435, 1272, 764, 1074, 797, 792, 877, 718, 717, 671, 825]. normalization [52, 371, 347]. normally [333, 473]. Normontwerp [278]. North [286]. Note [477, 1614, 1861, 759, 806, 1798, 1286, 1175, 1246, 1860, 1368, 1142, 1443, 1980, 506, 688, 507, 600, 1289, 1508, 2142, 572, 767, 795, 880, 996, 217, 1370, 2088, 2151, 1986, 1220, 951, 983, 1420, 476, 672]. Notes [1252, 980, 436, 277]. Notice [69, 135]. NPMLE [1435]. nu [776]. null [1727, 858, 805]. number [917, 1978, 107, 1892, 272, 265, 217, 458, 2007, 1094, 2101, 456, 1070, 1038, 863, 246, 452, 399, 516]. number-of-modes [1978]. numbers [1066, 156, 678, 455]. numerical [980, 623, 893, 496, 1919]. numerieke [623]. Numismatics [1218]. nummertje [346, 399]. nurturing [303]. nut [130]. nutritional [126].

O [654, 222]. O.R. [654]. O.R.-Analist [654]. obesity [2001]. Obituary [1616]. object [22]. Objective [2189, 1558]. observable [97]. observation [1465, 282]. observational [217, 2168, 91]. Observations [1195, 1177, 1991, 1940, 208, 2066, 548, 149, 1334, 107, 2240, 198]. observed [1904, 665, 1516, 822, 2001, 2289, 2111]. observer [916, 921]. observers [1063]. obtain [800, 734]. Obtained [1195, 321]. occasions [810, 432]. occupancy [651, 162]. occur [378]. occurrence [841, 286]. occurring [724, 454]. odds [1844, 1650]. ODE [2103]. oetan [98]. off [230, 229]. official [1353]. oftewel [780]. Old [1901, 1952, 996, 1768, 1342]. OLS [1768]. Olympic [1910]. om [149]. omkering [184]. OMNITAB [830]. omtrent [626]. Omvang [364, 76, 70]. Omzetting [203]. omzetvergelijking [368]. onafhankelijkheid [183]. onbekende [207]. Once [740]. ondenoek [594]. onder [327, 498, 780]. onder-houdsdiensten [498]. Onderlinge [368]. Ondernemer [284]. onderneming [461]. onderscheidingsver [394]. onderscheidingsver-mogen [394]. onderscheidingsvermogen [479, 218, 143, 165]. Onderwijs [571, 540]. onderzoek [21, 624, 400, 72, 22, 370, 214, 91, 137, 23, 630, 410, 501, 779, 90]. onderzoekingen [489]. ondoelmatige [704]. one [1060, 2103, 1409, 484, 972, 117, 203, 926, 13, 2088, 1742, 2038, 1271, 1933, 1085, 1749, 26, 182]. one- [117]. one-parameter [1085]. one-sample [2038]. one-sided [1409, 484, 2088, 1271, 1933]. one-step [2103]. one-top [484]. one-way [1060]. Oneindige [678]. ongeassocieerdheid [733]. ongelijke [753, 884, 888]. ongelijkheid [108]. only [1703]. onmacht [210]. ons [118]. ontmengingsmethoden [620]. ontstaan [123]. ontwikkeling [324, 253, 736, 631, 202, 254]. ontwikkelingen [188]. ontwikkelings [540]. ontwikkelings-landen [540]. ontwikkelingsmodel [439]. ontwikkelingsplanning [460]. onze [345]. onzekerheid [646, 742, 756]. oog [454]. oogenblik [13]. ook [912, 889]. Oordeel [675, 241]. Oorlog [778]. oorsprong [379]. op [912, 447, 480, 208, 102, 548, 616, 126, 534, 740, 109, 500, 13, 99, 61, 231, 347, 383, 283, 711, 717, 17, 35, 16, 162, 276, 395, 733, 182, 680]. opbouw [793]. opbrengstcijfers [78]. opbrengsten [193]. Open [287, 287]. openbare [447, 45]. Opening [1348, 234]. Openingswoord [234]. operation [555]. Operational [256, 256, 615]. Operationele [615]. Operations [463, 1300, 1140, 1751, 980, 1907, 1772, 262]. Opgavenrubriek [814, 854, 866, 874, 881, 886, 890, 895, 898, 904, 918, 927, 932, 936, 940, 943, 946, 952, 958, 963, 967, 978, 974, 984, 989, 994, 1000, 1007, 1028, 1037, 1044, 1050, 1056, 1062, 1068, 1021, 922, 1016]. opinie [21, 22, 23]. opinie-onderzoek [21, 22, 23]. opinieonderzoek [9]. opinion [21, 1872, 22, 241, 23, 9]. Opleiding [302]. Opleidingen [446]. oplossing [260, 220, 894]. Oplossingen [974]. opmerking [217]. Opmerkingen [917, 648, 277, 132, 272, 489, 25, 92, 77, 340]. opnieuw [973]. opponent [161]. opportunities [1774, 1777, 487, 246, 401, 516]. Oproep [122]. opslag [448]. optical [1796]. Optimal [696, 1780, 1580, 1618, 1797, 1862, 920, 1954, 2007, 1487, 2247, 859, 527, 1941, 1269, 1454, 1498, 1806, 316, 1810, 2280, 2062, 1505, 261, 491, 492, 555, 616, 908, 382, 355, 322]. Optimale [492, 555, 616, 908, 382, 355, 322, 527, 316, 261]. optimaliseren [348]. optimaliteitscriteria [917]. optimality [1874, 917, 1967, 1040, 992, 785]. Optimistic [1583]. Optimization [1205, 1733, 1451, 1551, 1989, 2117, 2011, 1523]. optimizing [673, 348]. Optimum [1974, 415, 734]. optimumprobleem [415]. Options [1453, 1534]. optreden [378]. opzet [489]. orang [98]. orangutan [98]. Orde [405, 411, 406, 415, 412]. Order [1236, 1128, 1105, 1195, 1253, 1981, 882, 2213, 872, 1886, 2025, 2144, 1854, 1554, 415, 2285, 864, 1973, 473, 851, 1696, 1496, 1404, 2003, 1225, 869, 405, 411, 412, 406]. Ordered [223, 1183, 1892, 2287, 1972, 2042]. Ordering [2144, 1202, 2254]. ordinal [2109, 2203, 1492, 2110]. ordinary [2174, 2243]. organ [2097]. Organisatie [567, 584, 380]. Organization [567, 584, 380]. orientations [1986, 1026]. oriented [2183]. origin [379, 1426]. ornithology [1747]. Ornstein [2006, 2088]. ORTEC [1772]. Orthogonal [2287, 1780, 287, 1972, 39]. orthogonale [287, 39]. Other [1131, 1593, 780, 475]. our [345, 118]. outcome [1667]. Outcomes [1671, 217, 2046, 2162, 2170]. outer [1910]. outlier [1491, 1473, 532]. Outlier-ignoring [532]. Outliers [1185, 1194, 377, 856, 972, 1629]. Output [1041, 636, 1723, 908]. outputanalyse [908]. Outstanding [487]. overdispersion [2064, 1541, 2039]. Overflow [956]. overlapping [2231]. overschrijding [484]. overschrijdingsgrenzen [59]. Overschrijdingskans [18, 484]. overschrijdingskansen [117, 333, 386, 747]. overtime [481]. Overview [1660, 1717, 1838, 1963, 591, 925, 550, 620, 202]. overwerk [481]. Overzicht [550, 620, 202, 591, 925].

p [699]. P. [490]. Paap [1628]. packaging [25]. packets [363]. packing [616, 859]. pair [860, 2139]. paired [1559, 962]. pairs [669]. Pairwise [1088, 1722, 1063, 1982]. pakjes [363]. pakmachine [616]. Palm [1403]. Panel [1748, 2005, 1049, 1443, 1446, 1825, 1816, 1329, 1448, 2030, 1824, 1820, 1449, 1821, 1322, 1983, 1447]. Panels [1614, 1785, 1531]. Paper [1047, 971, 208, 1997, 129, 227, 1398, 1550, 711, 130]. Papers [1239]. papier [208, 711]. parallel [1302, 1488, 1304, 2019, 382]. parallelle [382]. Parameter [1984, 2018, 1991, 2103, 1996, 1323, 163, 1638, 1937, 676, 1805, 1260, 2069, 1817, 1085]. parameter-free [163]. parameters [1436, 2312, 2113, 138, 807, 2174, 761, 1864, 2101]. parametervrije [163]. parametric [1591, 2276, 1903, 1897, 2078, 1926, 1459, 2214, 1213, 1704, 1932, 1383, 163]. Parametrische [163]. parent [1407, 668]. Pareto [2132]. part [308, 334, 1454, 1498, 942, 654, 916, 921, 1064, 1065, 988, 1581]. Partial [1214, 2274, 2242, 1790, 1311, 2070]. Partially [1953, 2134, 1664, 2289, 2111, 1899]. participation [1446]. particles [665]. particular [77]. particulate [1630, 1693]. partitioned [728, 853]. partitions [1516, 1936]. parts [1297, 1433, 1266]. Parus [820]. passage [1032]. passed [110]. passenger [650, 649, 651, 1751]. passingen [4]. passingsprobleem [220]. Passive [1484]. Past [721, 451, 1765, 1375, 1770]. Path [1219, 1722, 1986, 1793]. path-based [1722]. Patient [365]. patienten [365]. patients [1095, 2097]. pattern [1522]. pattern-mixture [1522]. Patterns [1613, 1924, 1591, 1341, 1333, 1527, 790, 1922, 1474]. Payments [1688]. Pays [1139]. PCA [2111]. PCA-based [2111]. peace [778]. Pearson [2095, 784]. Peck [509]. penalization [1963]. Penalized [2169, 1601, 2048, 2036, 2018, 2230, 2293, 2102, 2243, 2069]. penalties [1659].

Penalty [2077]. penetrations [1449]. percentiel [608]. percentile [608]. Percentiles [744]. Perception [211]. percolation [1838, 1995, 1419, 1526]. Performance [1703, 1185, 1194, 2170, 1779, 1024, 909, 1909, 2306, 2195, 127]. period [1809, 2091, 755, 785, 1445]. periodic [1923]. periodically [1864]. periods [2145]. permutatietoets [544]. permutation [2013, 544, 1221]. personeelssterkte [680]. personeelsterkte [309]. personeelsterkte-grafiek [309]. personenauto [650, 649, 651]. perspective [1907]. PERT [638]. Petri [1586]. petroleum [908]. Ph.D [1489]. Ph.D. [999, 1006, 1075, 1080, 1109, 1116, 1124, 1154, 1196, 1203, 1216, 1223, 1237, 1255, 1262, 1267, 1273, 1279, 1291, 1312, 1331, 1339, 1354, 1365, 1414, 1439, 1509, 1528, 1556, 1565, 1595, 1606, 1163, 1380, 1209, 1229, 1461, 1179, 1092, 1055, 1248]. pharmacological [317]. pharmacologist [586]. phase [2262, 2251, 1778]. PhD [321]. phenomenological [2258]. phenomenon [87, 13]. Philip [1861, 1860]. Philosophers [1836]. phylogenetic [2182]. physical [1744]. physicians [587]. physics [50]. physiology [603]. piece [7]. piecewise [1560, 1649]. pijnstillende [457]. PIT [12]. Pitman [813]. Pitman-efficiency [813]. Pivoting [1165]. plaats [67]. plaatsen [287]. places [287]. Plan [625]. plane [665]. Planning [1030, 1757, 513, 980, 409, 638, 2282, 950, 520, 239, 460, 660, 472]. plans [850, 1575]. plant [231]. plantenveredeling [231]. plants [2276]. plausible [434]. players [264]. plot [1174, 1780]. plots [1372]. Plotting [208]. PLS [1326]. Plug [2051, 1702]. Plug-in [2051, 1702]. pocket [1094]. Point [1173, 1591, 1341, 2228, 1886, 2025, 1854, 1518, 1873, 1458, 1973, 1403, 2193, 2015, 1035, 1434, 1474, 1927, 1034]. pointer [109]. points [1968, 1986, 1514, 407]. pointwise [1592]. poison [745, 725]. Poisson [2229, 1160, 1711, 2064, 1710, 1084, 1168, 1187, 1120, 841, 687, 1891, 1518, 1930, 850, 1542, 2060, 2246, 1877, 2293, 1638, 1370, 745, 725, 2217, 676, 951, 648, 1979, 2156, 2067, 466, 1594, 1679, 29, 685, 729, 1477, 578, 2055]. Poisson-gamma [2156]. Poisson-proces [648, 578]. Poisson-process [676]. Poissonproces [685]. Police [1679]. policies [785]. Policy [566, 567, 1301, 1771, 1766, 1876, 1919, 992, 257, 900, 565]. polinome [131]. poliomyelitis [318]. politiek [257]. poll [9]. Pollaczek [1837]. polls [1935]. Pólya [524]. Polyhedra [1131]. Polyhedral [1451, 1551, 1205, 1792]. polymers [1457]. polynomia [131, 39]. polynomial [2251]. polynomials [1575]. Polytomous [1258]. ponskaarten [228]. Pooling [1830]. Poor [780]. popular [1495]. populatiegemiddelde [753]. Population [1162, 1679, 2221, 1507, 696, 1694, 1956, 1564, 1563, 753, 244]. populations [1955, 1277, 1360]. Portfolio [2154, 1535, 1252, 2117]. portfolios [2278]. portmanteau [2271]. posed [577, 549, 1561]. position [109, 67]. positive [1652, 824]. positiver [824]. possibilities [136, 71]. possibility [1993, 77]. possible [637]. Posterior [1722, 1727, 2274, 1650, 1728, 1702, 2237]. posteriors [2095]. Postscript [41, 38]. Power [1902, 165, 1997, 2311, 479, 1732, 394, 218, 1999, 1976, 143, 1328, 1998, 877, 134, 668, 1012, 210]. powerful [2045]. powerlessness [210]. Powers [1122, 1991, 788, 798, 621]. pp [1997]. Practical [1282, 892, 1427, 432, 322, 290]. practice [1245, 171, 337, 990]. Practische [290]. Practitioners [281]. praktijk [171, 337]. praktijkgeval [322]. pre [1935]. pre-election [1935]. precision [2198, 2224, 879, 1222]. preclassified [2066]. Predicting [1691, 188, 193, 192, 243, 191, 194]. Prediction [1600, 2240, 1599, 1445, 1269, 1738, 2230, 2097, 639, 1350, 1885, 2135, 2014, 679, 189, 187]. predictions [636, 449]. Predictive [1601, 2008, 545, 1728, 1702, 1727, 531, 2237]. predictor [2062]. Preface [1320, 1450, 1347, 1374, 332]. prehistory [1099]. preliminary [1572]. premature [410]. prepayment [1797]. preponderance [2203]. presence [1004, 2279, 1023, 2268, 2207, 2155, 2001]. Present [451, 1765, 1375, 1770, 273]. preservation [826]. preserving [692]. presmoothed [1845]. pressure [2310]. pretest [2090, 2200]. prevalence [1988, 1908]. preventable [115]. Price [605, 1041, 480, 1731, 762, 649, 1939, 2009, 414, 436, 327]. price-effective [1939]. prices [2002, 2124]. pricing [480]. priesthood [275, 293]. prijs [781, 1017]. prijsindex [649]. prijsindexcijfers [436]. prijsraming [480]. Prijsverwachtingen [327]. prijsvorming [480]. Prijsvraag [346, 399]. Principal [469, 834]. Principal-component [469]. Principles [664, 681, 1465]. Prior [2262, 1015, 2302, 2156]. prioriteitsregeling [383]. priority [529, 871, 383]. priors [1722, 1726, 2310, 1936]. privacy [1806]. Prize [781, 971, 1017]. Probabilistic [1951, 2172, 2242, 1495]. probabilistically [2152]. Probabilité [821]. Probabilities [1127, 1620, 1615, 1112, 1368, 720, 1722, 1711, 43, 758, 941, 2294, 117, 218, 876, 1921, 333, 1520, 386, 1759, 1627, 385, 1805, 753, 888, 747, 207]. Probability [18, 1200, 1150, 808, 1655, 2255, 208, 1411, 1977, 1079, 129, 1920, 484, 835, 36, 1376, 2148, 821, 2295, 571, 692, 884, 573, 798, 578, 16, 793, 130, 444, 27]. probable [806, 669]. probiotica [1851]. probleem [34, 65, 264, 131, 91, 49, 481]. Problem [1106, 971, 1076, 1086, 1093, 1098, 1103, 1110, 1117, 1125, 1132, 1138, 1155, 1164, 1171, 1180, 1190, 1204, 1210, 1217, 1224, 1249, 1256, 1263, 1268, 1274, 1280, 1287, 1292, 1299, 1313, 1319, 1366, 1373, 1386, 1392, 1400, 1406, 1415, 1423, 1432, 1440, 1462, 1470, 1481, 1490, 1501, 1510, 1519, 1529, 1538, 1549, 1557, 1566, 1576, 1596, 1623, 492, 1167, 1381, 1081, 1303, 1306, 1332, 1340, 1346, 1355, 1197, 1230, 1238, 914, 713, 702, 554, 507, 600, 577, 549, 1881, 1836, 1561, 849, 34, 1920, 1058, 572, 1739, 415, 734, 830, 1371]. problem [1458, 842, 1790, 391, 569, 996, 220, 65, 264, 1954, 893, 1233, 131, 2115, 91, 371, 1928, 992, 892, 894, 1511, 1791, 49, 162, 975, 603, 454, 1931, 512, 481]. problemen [562, 114, 779]. Problems [281, 1019, 988, 1665, 1213, 1218, 763, 771, 1276, 1801, 1409, 976, 1645, 2294, 1912, 1349, 260, 909, 1798, 2127, 942, 25, 562, 1802, 1523, 1933, 592, 16, 125, 77, 503, 114, 779]. Procedure [1122, 666, 1301, 910, 787, 991, 788, 1868, 1885, 1802]. Procedures [1286, 1137, 1024, 1874, 1052, 1102, 1318, 529, 673, 1500, 1982, 813]. proces [578, 648]. Procesnauwkeurig [497]. Process [1236, 1160, 1123, 1084, 1168, 931, 1188, 1176, 1800, 2064, 2092, 2228, 1886, 2025, 1881, 1777, 1278, 841, 1573, 1644, 1984, 1854, 2166, 1518, 1904, 2285, 1631, 1778, 1637, 2088, 1403, 676, 1779, 1847, 2015, 995, 1311, 562, 1412, 578, 2194, 1212, 497, 648, 685]. processen [504, 431, 1042]. Processes [1746, 1178, 1153, 2137, 1906, 1886, 1903, 1902, 1996, 1466, 611, 1542, 2006, 1032, 504, 1973, 1395, 1880, 2218, 1465, 431, 724, 2186, 1578, 2039, 900, 1581, 1033, 1927, 632]. processing [617, 37, 78]. processor [1835]. processor-sharing [1835]. produced [258]. Product [697, 957, 1002, 2142, 2248, 981, 258, 982, 853, 1025]. product-moment [957]. productie [206, 194]. productie-indices [206]. productie-resultaten [194]. productieplanning [472]. Production [1356, 354, 1794, 1881, 228, 1305, 1810, 950, 206, 194, 472, 543]. production/distribution [1305]. Productiviteit [168, 172]. Productiviteitsmeting [169, 170]. productivity [2024, 354, 350, 351, 343, 352, 353, 172, 168, 345, 169, 170, 347, 342]. Products [986, 659, 1483, 908, 977, 990, 448, 438]. produkten [659]. produktie [354]. Produktiebesturing [543]. Produktiviteit [345, 342, 343]. produktiviteitsanalyses [352]. produktiviteitsmeting [351]. Produktiviteitsverhoging [347, 353]. proefdieren [44]. proefgroepen [602]. proefopzet [896]. Proefopzetten [251, 315, 917, 324, 312]. proefschema’s [287, 894]. proefveld [231, 78]. proefveld-resultaten [231]. Proeve [309]. proeven [51, 628]. Prof [749, 222]. Prof. [357, 374, 530, 1408, 118, 242, 558, 749, 782, 483, 550]. professional [823]. Professor [1118]. profetie [187]. profiling [1683]. profit [401]. profitability [662, 401]. Prognose [593]. Prognostic [1605, 1602, 2168]. program [632, 631, 839, 1763]. programering [939]. programma [632, 630]. programme [380]. programmed [632]. Programmering [556, 908]. programmeringsproblemen [260]. programmes [1758]. Programming [1144, 959, 667, 1043, 763, 771, 1974, 734, 1039, 908, 260, 950, 1479, 761, 681, 900, 939, 556]. programs [909, 630]. progress [243]. progression [516]. Progressive [528, 2207, 2261, 2289]. progressively [2173, 1757]. Project [520, 243, 520]. Project-planning [520]. Projection [1121]. Proof [931, 1311, 992, 183]. propensity [1985]. Properties [1215, 1236, 1059, 1286, 2032, 707, 1253, 390, 1714, 1417, 58, 488, 770, 2265, 1692, 1546, 562, 1674, 166, 825]. property [924]. prophecy [187]. proportion [1694, 2038, 2215]. proportional [1079, 1133, 2308, 2268, 1442, 879, 1520, 975]. proportions [2276, 1387, 2269, 2045]. proposal [120]. protection [1806, 1349, 2059, 1971]. proteomics [2047]. prove [28]. provinces [2258]. Pseudo [1614, 1866]. pseudopopulations [2141]. psychologie [199, 150, 189]. psychology [199, 150, 189]. psychometrics [1765]. public [447, 519, 1351, 45, 48]. Publications [818, 550, 1762]. publikaties [550]. published [278]. pull [1582]. pulmonary [2114]. punch [228]. punten [407]. purchase [1542, 473]. pure [1903]. Push [1582]. puzzle [1920].

Quadratic [1134, 986, 1227, 1654, 977, 990, 418, 717, 449]. quadrature [2100]. Qualitative [1502, 88]. qualitativen [88]. Quality [565, 238, 1948, 2206, 605, 1977, 1452, 14, 24, 34, 1949, 617, 2005, 556, 894, 171, 566, 236, 567]. Quantal [1259]. quantielen [333]. Quantification [214, 213, 215, 580]. Quantifying [2276]. Quantile [1397, 1642, 2036, 2276, 1845, 2265, 2072, 1480, 2233, 2023, 442]. Quantile-locating [1397]. Quantiles [1106, 1184, 1298, 1397, 333, 2117, 657]. quantitatieve [44]. Quantitative [1049, 574, 324, 44]. quantities [333, 443, 1526, 544]. quantity [1731]. Quantum [2014]. quarterly [2123, 710, 2129]. Quasi [1309, 1612, 1805, 1647]. Quasi-likelihood [1309]. Quasi-longitudinal [1612]. quasi-symmetry [1805]. question [722, 1961, 2177]. questionnaires [1430]. Questions [1283, 925]. Queue [1123, 1219, 931, 1585, 1013, 1837, 1921, 1759, 1696, 1270]. Queueing [1147, 1583, 1402, 887, 871, 893, 992, 481]. Queues [1846, 1573, 765]. queuing [1580, 481]. quotënt [561]. quotient [389, 743, 389, 561].

R [404, 654, 1020, 1628]. R. [1047, 793]. radio [87]. radiobuizen [87]. railway [1751]. Rain [939]. Raised [758, 758]. Random [1494, 1670, 1128, 1923, 1105, 1677, 1433, 684, 1094, 1457, 1131, 1253, 1396, 1940, 1424, 408, 688, 1714, 1585, 2258, 1703, 1289, 1363, 1974, 1701, 1987, 1232, 156, 2144, 2254, 2142, 2248, 1630, 910, 2096, 1468, 634, 1816, 840, 2030, 1702, 525, 485, 2224, 2079, 2240, 1840, 2017, 2029, 2074, 1574, 1337, 2169, 2202, 851, 1696, 995, 1097, 2310, 1895, 1497, 181, 2267, 798, 1730, 1445, 1705, 804, 70, 441, 259]. Random-coefficient [1923]. random-effect [2169]. random-effects [1701, 2224]. randomization [1398, 2300, 1342]. Randomized [1206, 1995, 2299, 2035, 1806, 394, 1697, 1971, 2019, 925, 2045, 1749]. randomly [1272, 728]. randomness [907]. randtotalen [768]. Rangcorrelatie [139]. rangcorrelatieschema [272]. Rangcorrelatieschema’s [271]. rangcorrelatietoet [397]. Range [1194, 744, 1589, 743, 438]. Rank [1201, 1115, 1232, 1193, 2241, 1213, 1122, 271, 1639, 1333, 1343, 1706, 1052, 1058, 1272, 1912, 200, 272, 991, 1040, 1271, 1082, 397, 668, 139]. rank-interaction [1052]. ranked [1755, 1972]. Ranking [1977, 1786, 1121]. ranks [1977]. rantsoenering [908]. Rao [2059]. rare [1988, 2083]. Rasch [1260]. rat [2147]. rate [2082, 1061, 841, 993, 1740, 2217, 1954, 1311, 110, 704, 1743]. rater [1857]. raters [1857]. Rates [1617, 1930, 1750, 1955, 1746, 2005, 711]. rather [830]. Ratio [1113, 1650, 1716, 835, 2031, 981, 1328, 2086, 982, 948, 1025, 2215]. rational [211]. rationalisatieonderroek [350]. Rationalisatieonderzoek [366]. rationalization [350, 366]. rationeel [211]. rationing [908]. Ratios [1162, 1844, 1958]. rats [2139]. Ray [1157, 1158]. rayon [617]. Re [1768, 1492]. Re-examining [1768]. re-weighted [1492]. Reaction [1861, 1860]. reactions [88]. readers [122]. Readership [281]. reading [109, 711]. ready [1571]. real [97, 1791]. realisatie [679]. realistic [601]. realistische [601]. realization [679]. received [31]. Recente [283, 324]. Receptuur [297, 330, 396, 389, 467, 310, 509, 378, 304, 335, 551, 395, 419]. rechtlijnige [291]. rechtstreekse [873]. Rechtszekerheid [742]. Recidivism [1680, 1824]. Recipe [297, 330, 396, 389, 467, 310, 509, 378, 304, 335, 551, 395]. recipes [892]. reciprocal [1822]. reckoning [537]. Reconciliation [2119, 2129, 2126]. Record [1215, 1316]. Records [1679, 1918]. recourse [950]. recovery [2303, 2]. recreation [1038]. Recruitment [938]. rectangles [1739]. Recurrent [1146, 2146]. recurrent-event [2146]. recursion [1581]. Recursive [1199, 960, 2051]. redacteuren [905]. Redactie [38, 345, 506, 475]. Redactiecommissies [445]. Reduced [1803, 1941, 1706, 142, 804]. Reduced-bias [1803, 1941]. reducing [1310]. Reduction [1188, 1043, 2242, 529, 1868, 79, 2037, 2303, 158]. reeks [149, 131, 291]. reeksen [79]. referees [1111]. Reference [1178, 1420, 1045, 2225, 1978, 1852]. reference-based [2225]. refined [1516]. reflected [2006, 2088]. reflections [740]. refuse [901]. Reg [1160]. regarding [625, 132, 579, 648, 747]. regelbreedte [527]. Regen [939]. regime [2251]. region [2231, 2092]. regional [2016, 553]. regionale [553]. regions [880, 1514]. register [1946, 1945, 1948, 1950]. register-based [1946, 1945, 1948, 1950]. Registration [1484]. regressie [912, 413, 561, 277, 290, 889]. regressie-analyse [413, 561, 277, 290, 889]. regressieanalyse [61, 157]. regressiecoëfficiënten [151]. Regression [1236, 1324, 1083, 1675, 1214, 1159, 1951, 1282, 1345, 824, 1604, 1679, 1867, 1005, 1283, 1257, 1225, 1045, 2032, 2302, 2036, 912, 413, 2201, 2176, 1315, 2198, 1503, 2276, 2251, 2298, 2082, 1706, 1635, 1643, 1258, 2157, 1968, 1934, 1892, 468, 1378, 2197, 2200, 2060, 2268, 1942, 2078, 1525, 864, 151, 1926, 2293, 1862, 1699, 2087, 61, 1937, 961, 1916, 2202, 857, 1979, 2233, 1692, 1813, 2051, 1788, 1546, 2023, 157, 418, 2135, 1895, 1096, 561, 2243, 2247, 2167, 1932, 2069, 1787, 2070, 2149, 2192, 2219, 2286, 2101, 1752, 277, 290, 889, 1020]. regressions [2223, 2123, 2208]. Regressor [1701, 1957]. regressors [1324, 1708]. Regular [1131]. regularly [2085]. Regulation [661]. reinforced [1840]. reject [1724]. Rejoinder [1284]. rekenautomaten [614]. rekening [793]. Rekeningen [177]. rekenkundig [704]. rekenmachines [511]. rekenwijze [444]. Related [1199, 2077, 1739, 1469, 385, 642, 798, 237]. Relating [2139, 1593, 2024]. Relations [1041, 1875, 470, 141, 151]. relationship [354, 467, 554, 574, 880, 1027, 44, 2151, 1579]. relationships [1822, 2314, 1275, 339, 1823]. Relative [1091, 1574, 1114, 2135, 1552, 1691, 1058, 788, 879, 1222, 1986]. relevant [2124, 1740]. Reliability [811, 1436, 883, 1879, 7, 1882, 916, 921, 40, 592, 382]. remark [745]. remarks [835, 1329, 1398, 743, 234, 402]. Remembering [1463]. REML [1492]. removable [992]. removing [1792]. ren [324]. renewal [842, 956]. rentabiliteitsprognoses [662]. Reparameterized [2198]. Repeated [1981, 1610, 1611, 856, 1613, 1569, 1107, 1390, 1869, 1627]. Replacement [1175, 965, 733]. replicability [2282]. Reply [1399]. report [1243]. reporting [1930]. Representation [727, 1845, 975]. representations [1648]. reproduceerbaarheid [198]. reproducibility [198]. resampling [2141, 1575]. Research [281, 336, 463, 1300, 1148, 1140, 1751, 1621, 1464, 21, 1946, 1945, 366, 350, 624, 659, 126, 1700, 769, 72, 370, 1907, 214, 1821, 23, 1412, 630, 1784, 779, 86, 594, 90, 284, 1772, 568, 256, 615, 262, 45]. reservation [391]. reserveringsprobleem [391]. resident [1956]. Residual [2208, 1983, 2159, 418, 1911, 863]. Residual-based [1983]. Residuals [829, 1199, 796, 1295, 1422, 413]. Residuen [413]. resistant [591]. resolution [1482]. Resolving [2224]. resp [473]. respect [1902]. respectively [473]. Respects [1139]. respondents [2059, 1971]. Response [1874, 1206, 1120, 1259, 2299, 1700, 1535, 1806, 1781, 2046, 1813, 2005, 2110, 1971, 2019, 925, 1749]. Responses [1135, 1809, 1656, 2262, 2020]. responsibilities [212]. Restricted [1335, 1717, 2033, 2070, 429]. restriction [2042]. restrictions [882, 1389]. result [1011, 399]. resultaten [231, 78, 283, 194]. Results [1166, 1133, 1835, 1290, 1402, 1703, 976, 1897, 2144, 1323, 788, 602, 496, 231, 78, 283, 802, 1412, 194, 671, 1012]. retrial [1402, 1573, 1759]. Retropolating [2124]. retrospective [1777]. return [1532]. returns [2145, 1534]. revenue [1660]. Revenues [1686]. reversed [2161]. reversible [1965]. reverting [2010]. Review [320, 1166, 1206, 1088, 1119, 30, 46, 53, 68, 73, 81, 93, 104, 121, 133, 705, 715, 723, 737, 748, 757, 789, 794, 799, 803, 815, 855, 867, 875, 891, 899, 903, 913, 923, 947, 1067, 10, 1418, 944, 1964, 772, 916, 921, 1525, 2168, 809, 1257]. Reviews [674, 698, 731, 766, 1438, 111, 140, 144, 153, 167, 185, 201, 209, 221, 226, 232, 248, 255, 263, 268, 274, 279, 285, 292, 299, 305, 311, 331, 341, 356, 373, 384, 392, 403, 416, 430, 440, 450, 464, 474, 482, 493, 505, 515, 523, 533, 547, 557, 570, 581, 599, 606, 612, 633, 653, 663, 644, 682, 689]. revised [1859, 1861, 1860, 791]. revisietijd [322]. revision [436, 322]. revisions [2126]. Revisited [1104, 1602, 2301, 1734, 1752, 973]. Reweighted [1524]. rich [780]. Richard [1472]. richtingen [78]. Richtlijnen [206]. Ridge [1129, 1752, 2176, 2243]. Ridit [1119]. right [2160, 1429, 1979, 1982, 1581]. right-censored [2160]. rijk [780]. rijtijdenbesluit [364]. Risk [1024, 1353, 2231, 1534, 1552, 1351, 1535, 2207, 2269, 2097, 1954, 2052, 1958]. Risks [1202, 2298, 1645, 1352, 2217, 2261, 2289, 2292, 2104]. road [836, 949, 596]. Robust [1359, 1048, 2020, 1562, 2291, 2230, 934, 1852, 1142, 2276, 1524, 1377, 1244, 2288]. Robustness [1363, 1502, 1699, 1327, 2085, 878, 1867]. ROC [2047]. role [2141, 119]. roller [1881]. rolling [107]. Ronald [293]. rondom [318]. root [1748]. roots [1732, 1837]. Rotatable [315]. Rotation [1613, 1869]. Roteerbare [315]. Rotterdam [1762, 1763]. Rough [1571]. Rough-and-ready [1571]. Rounded [1618, 2037]. rounding [2141, 230, 229]. rounding-off [230]. routes [1939]. routing [1580, 1488, 1304]. row [995, 933]. royal [31]. ruïering [264]. ruis [87]. rule [1978, 1488, 897]. Rules [1600, 969, 1544, 1036, 230]. run [762]. RV [1167].

S [1997, 1057, 558, 1160, 33, 1837, 802]. ‘s-Gravenhage [33]. S. [728]. Saddlepoint [1917, 1987, 2017, 2029]. safeguard [1856]. Sales [528, 703, 604, 368]. same [811, 1063]. samengestelde [579]. samenhangende [562]. Samenvatting [497, 195]. Sample [1128, 1105, 180, 1228, 937, 1443, 1253, 2262, 744, 872, 2175, 837, 76, 1330, 2254, 1323, 1023, 1133, 1058, 870, 2210, 993, 910, 1896, 2241, 1912, 1889, 265, 1741, 1073, 2087, 2282, 694, 1059, 851, 2038, 1011, 728, 1692, 1090, 761, 2086, 273, 1546, 1101, 326, 459, 2045, 329, 743, 59, 395, 419, 1221, 621, 637, 668, 1012, 442]. sampled [1903]. Samples [1168, 1618, 575, 744, 1486, 70, 218, 525, 706, 1559, 51, 1629, 808, 877, 229, 259, 269, 643, 1447, 701]. Sampling [1282, 1520, 1735, 1137, 1175, 1858, 1283, 453, 666, 1391, 635, 810, 1775, 1655, 1724, 1079, 1974, 1245, 696, 76, 107, 2035, 1630, 1693, 850, 1537, 58, 770, 709, 2134, 1222, 1646, 1869, 2237, 138, 143, 1514, 1971, 982, 965, 843, 1802, 348, 136, 884, 813, 59, 432, 353, 183, 888, 49]. sanitation [2223]. Sanov [1467]. SAS [1666]. satisfaction [1790, 1868]. Saunders [1916]. SCAD [2036]. SCAD-penalized [2036]. scalar [673]. Scale [1173, 1991, 1004, 1562, 1058, 1873, 1244, 2261, 347, 538]. Scaling [1838, 2310, 594, 1456, 594]. scan [1485]. scatter [1372]. sceptical [2282]. Schaafsma [1001]. schaal [347, 538]. schatten [574, 768, 207]. schatter [873]. schatters [608, 532, 435]. Schatting [753, 467]. schattingsproef [139, 103]. schattingstheorie [754]. Scheduling [1302, 1303, 1835, 1794, 1795]. Scheefheid [27]. Scheefheidstoetsen [328]. Scheffer [905]. schema [143, 260]. schema’s [58]. scheme [1079, 272, 138, 143]. Schemes [271, 58, 110, 287, 894, 348]. scholars [1769]. school [1700, 109, 1444]. Schweitzer [577, 549]. science [494, 595, 607, 256, 1966, 2179, 1763]. Sciences [1149, 213, 495, 86, 1456]. Scientific [293]. Score [1684, 1446, 1985, 2097, 1787, 2116]. score-driven [2116]. scored [955]. scores [1071, 2067]. Scoring [1712, 791]. scrambled [2299]. scrapping [650]. Screening [1168, 839]. Search [1740]. Seasonal [2128, 735, 2159, 1864, 438]. seasonally [2129]. Seating [137]. Second [2025, 1854, 106, 864, 1973, 56, 2003, 72, 975]. Second-order [2025, 1854, 864, 1973, 2003]. secondary [824]. Sectie [423]. Sectiegedeelte [510, 518, 526, 536, 552]. Secties [174]. Section [1103, 1110, 1117, 1125, 1132, 1138, 1155, 1164, 1171, 1180, 1190, 1204, 1210, 1217, 1224, 1249, 1256, 1263, 1268, 1274, 1280, 1287, 1292, 1299, 1313, 1319, 1366, 1373, 1386, 1392, 1400, 1406, 1415, 1423, 1432, 1440, 1462, 1470, 1481, 1490, 1501, 1510, 1519, 1529, 1538, 1549, 1557, 1566, 1576, 1596, 1623, 1381, 423, 1306, 1332, 1340, 1346, 1355, 1197, 1230, 1238, 518, 526, 536, 552, 814, 854, 866, 874, 881, 886, 890, 895, 898, 904, 918, 927, 932, 936, 940, 943, 946, 952, 958, 963, 967, 978]. section [984, 989, 994, 1000, 1007, 1028, 1037, 1044, 1050, 1056, 1062, 1068, 1076, 1086, 1093, 1098, 1608, 1081, 665, 1021, 922, 1016, 296, 518, 526, 536, 552]. sectional [1748]. Sections [510, 1615, 1627, 174]. sector [1685, 502, 660]. sectormodel [502]. Securities [2290]. security [1511]. seeding [1919].

Seemingly [1345, 1005, 1932, 1899]. sees [26]. segmentation [1990]. segregation [1863]. seizoenprodukten [438]. Selectie [231]. Selection [2203, 1603, 1277, 2036, 2018, 2236, 1018, 408, 1507, 1726, 1252, 1878, 2169, 2245, 2154, 2023, 1563, 2069, 1730, 1445, 1360, 1725, 398, 379, 298, 231, 319, 291, 397, 434]. Selective [453]. selector [2051]. Selectuur [398, 379, 298, 319, 291, 397, 434]. Self [1776, 1468, 1589]. Self-adapting [1776]. self-decomposability [1468]. self-similar [1589]. SEM [1612, 1820]. Semaphone [1211]. semi [1591, 1807, 1468, 1926, 711, 2292, 900]. semi-competing [2292]. semi-group [1468]. semi-logarithmic [711]. semi-logaritmisch [711]. semi-Markov [900]. semi-nonparametric [1807]. semi-parametric [1591, 1926]. semicontinuous [2008]. semimartingales [1905]. Semiparametric [1431, 1525, 1826, 1895, 2101, 2303, 1651, 1459, 1692]. semivariance [2154]. Sen [1137]. Senegal [2184]. sensitive [707]. Sensitivity [1671, 2225, 1694, 724, 1895]. separate [688]. Sept [105]. September [374, 105, 1754]. sequence [71, 426, 425, 197]. sequences [669, 181]. Sequente [197, 426, 425]. Sequential [427, 1991, 635, 428, 1507, 2120, 835, 1646, 2097, 1233, 429, 637]. sequentially [2113]. sequentietest [71]. serial [2013, 1867]. serially [1932]. serie [563]. Series [2119, 1611, 1686, 1615, 2229, 1608, 1812, 2064, 1657, 2304, 2126, 2120, 1897, 1785, 2124, 563, 1314, 1890, 765, 2061, 2074, 2306, 796, 2122, 2235, 131, 1647, 1864, 1487, 79, 553, 622, 798, 2101, 2116, 291]. Server [1107, 1584, 1573, 992]. servers [1402, 1488]. Service [238, 1584, 1846, 871, 1696, 1579, 2, 1270, 650, 349]. services [498]. session [100]. set [266, 149, 508, 2184, 2111, 182]. sets [1533, 1714, 1534, 1530, 1740, 1372, 1344]. setting [1543, 616, 1738, 1858]. Setup [1303, 1740]. seven [2184]. seventh [196]. several [1072, 880, 909, 1506, 1955, 1222, 1328]. sexual [2181]. Shake [1317]. Shake-and-bake [1317]. shape [2032, 1377]. Shapiro [877, 718]. shares [639]. sharing [1835]. sharpening [1827]. Shewhart [1008]. shift [1012]. shifting [1990]. Shoe [1167]. Short [762, 1614, 2290, 1424, 1905, 283]. Short-run [762]. Short-Term [2290, 283]. shortest [1417, 1793, 1244]. should [1985, 1241]. Showing [1120]. shrink [2222]. Shrinkage [2197, 1604, 1602, 1603, 1601]. sib [1929]. sick [779]. side [426, 1304, 219]. sided [1409, 484, 117, 2088, 1271, 1933]. sieve [1245]. sightings [2147]. sign [991, 429]. Signaalstroomdiagrammen [560]. signal [1409, 560]. Signed [1494, 1193, 991, 1040]. Signed-Rank [1193]. Significance [1343, 266, 758, 461, 2222, 90]. similar [1589]. similarity [828]. simpele [574]. Simple [2109, 1108, 1115, 931, 1136, 574, 1401, 1889, 864, 1544, 950, 1233, 1311, 824, 1546, 1475, 992, 29, 291, 897, 442, 513, 6]. Simplex [1165]. Simplicial [1034]. Simplified [807]. Simulated [1265, 504, 2117]. Simulatie [632, 616, 719]. simulaties [746]. Simulating [1970]. Simulation [1718, 1122, 554, 1985, 616, 2166, 1077, 719, 1971, 1012, 632]. simulations [1966, 746]. simultane [182]. Simultaneous [720, 1088, 1955, 954, 1831, 1393, 1395, 339, 182]. since [1713]. singen [136]. Single [1107, 2221, 1723, 76, 1573, 2136, 110, 2309, 2149]. single-index [2136, 2309, 2149]. single-server [1573]. singular [2224]. Sir [293]. site [1995]. situation [190, 13]. Size [1128, 1105, 931, 1679, 1253, 2262, 744, 872, 1079, 76, 484, 870, 1896, 2241, 2078, 1709, 2282, 851, 1101, 2045, 743, 52, 442, 347, 364]. sized [218]. sizes [70, 2007]. Sizing [610]. skating [1910]. Skew [2073, 798, 2266]. skew- [2266]. Skew-normal [2073]. Skewness [1191, 2252, 2049, 2016, 1208, 1475, 589, 27, 328]. skin [88]. skip [1581]. skip-free-to-the-right [1581]. Slaapmiddelen [15]. slanted [1429]. Sleeping [15]. slippage [837]. sloop [650]. Slovakia [1758]. slow [847]. slow-mover [847]. smaak [400]. smaak-onderzoek [400]. Small [1418, 1023, 1133, 1618, 2087, 2175, 1466, 2241, 218, 877]. Small-sample [1133, 2087]. smallest [266, 2144, 1458]. Smirnov [1409]. Smooth [2079, 1136, 1091, 1454, 1498, 2174]. smooth-and-match [2174]. smoothed [1446]. smoother [1448]. Smoothing [528, 1635, 611, 1571, 2069, 39]. smoothness [2032]. snapshots [240]. snijsnekheid [555]. snijsnelheid [492]. soccer [1814, 1919]. Sociability [2184]. Social [1149, 2180, 253, 1337, 213, 1511, 660, 495, 1456]. sociale [213, 660, 495]. Society [106]. sociological [630, 92, 244]. sociologie [580]. sociologisch [630]. sociologische [92, 244]. sociology [580]. SOFA [2097]. soft [1659]. software [1436, 1770]. soil [808]. sojourn [1835]. Solution [1170, 1344, 2270, 577, 549, 220, 893, 1053, 894]. solutions [1349]. solve [892]. Solving [2127, 760, 830, 569, 260, 2071]. Some [1047, 1290, 1166, 1157, 1703, 1417, 600, 887, 883, 831, 1675, 944, 1588, 1379, 1897, 1133, 1645, 835, 1329, 697, 306, 670, 1234, 1271, 496, 802, 1603, 592, 743, 402, 671, 726, 1744, 901, 512, 1491, 1658, 2213, 2092, 2228, 1777, 1828, 1363, 2242, 1023, 1473, 1058, 2124, 1398, 529, 926, 1544, 845, 1213, 562, 16, 444, 390, 462, 625, 911, 132, 579, 623, 175, 272, 66, 719, 631, 351, 489, 431, 25, 407, 562, 23, 127, 77, 294, 402, 141, 865, 271, 114, 340, 1158]. Somermeijer [41]. Something [148]. song [346]. soorten [50]. sorting [1813]. source [413]. sources [1948]. South [113]. Space [1612, 2185, 1611, 2134, 485, 1922, 1715, 1811, 1513, 2303]. space-time [1922]. spaced [1825]. spaces [727]. Spacings [1687]. Spain [2231, 2133]. sparse [2117]. Spatial [2074, 2077, 2147, 1341, 1863, 1527, 1996, 2265, 2285, 2057, 2040, 1474]. spatially [2231, 1969, 1816]. spatio [1854, 1973]. spatio-temporal [1854, 1973]. spatiotemporal [2025, 1996, 1956]. Spearman [858, 805]. Special [1599, 1441, 1945, 1367, 1530, 839, 2179]. species [2182]. specific [1863]. specificaties [585]. specification [2092, 1477]. specifications [470, 585]. specified [1807, 1970]. spectra [2239]. spectrometry [2266]. Spectrum [1677]. Speech [1348, 775, 238]. Speed [492, 1846, 1910]. spelers [264]. spellende [545]. spherical [665]. spike [2046]. spinning [562]. spinproces [562]. Spline [2082, 2107, 1825, 2069]. Spline-based [2107]. splines [2032]. Split [1174, 1780]. Split-plot [1174, 1780]. splitsing [152]. splitting [152]. sponnen [562]. spot [2288]. spread [447, 265]. spreiding [447, 272, 265, 127, 183, 95]. spun [562]. Square [1145, 379, 752, 1426, 1995, 1805, 838, 672]. squared [477, 2176, 683, 1638, 1295, 1422]. Squares [1186, 1572, 2221, 1066, 1848, 2242, 245, 693, 909, 2102, 1096, 2243, 2041, 973, 291, 219, 182]. squares-centered [2243]. Staatsbedrijf [12]. Stability [1468]. stable [1902, 1364, 2003]. Staff [38, 309, 680]. stage [1898, 1060, 1657, 1582, 696, 2211, 1059, 1479]. stand [109]. standaard [631]. standaard-toetsingsprogramma [631]. standaardafwijking [277, 290]. standard [2270, 744, 631, 1581, 230, 277, 290, 278, 420]. Standardization [278, 420, 120]. standards [736, 236]. standarized [1278]. staps [503]. STAR [1848]. stars [676]. STAT25 [773]. State [1019, 1611, 12, 2134, 2107, 1612, 1715, 1053]. state-switching [2107]. states [1839, 2267]. static [1580]. station [481]. stationarity [1719, 1746]. stationary [1714, 611, 1631, 1637, 1403, 1811]. statisisch [105]. Statistic [1194, 575, 1250, 2297, 1478, 983, 1208, 2239, 1082, 397, 838, 172]. Statistica [43, 1764, 120, 43, 1997, 1998, 699, 1567, 1662, 1689, 1818, 1850, 1884, 8, 18, 1450, 1001, 1029, 2179, 1407, 905, 2273, 32]. Statistical [204, 225, 267, 296, 308, 334, 358, 336, 387, 433, 478, 517, 564, 609, 647, 882, 1341, 2137, 1148, 1777, 1553, 380, 1336, 2112, 105, 179, 1600, 1095, 1077, 1282, 876, 521, 486, 155, 1544, 1668, 2186, 1683, 1218, 1932, 1169, 1899, 2070, 1283, 275, 293, 720, 388, 773, 354, 833, 2126, 1588, 534, 58, 87, 323, 969, 1495, 830, 563, 508, 1876, 1952, 1782, 370, 1778, 916, 921, 1995, 65, 790, 719, 1375, 961, 1779, 784, 25, 239, 664, 945, 23, 303, 1412, 1950, 6, 77, 294, 114, 88, 707, 278, 344, 377]. Statistical [14, 24, 205, 34, 72, 312, 280, 220, 224, 2179, 247, 237, 2, 85, 86, 52, 89]. statistically [307]. statistician [67, 823, 212, 26]. Statistick [172, 168]. Statistics [1943, 1975, 2004, 1264, 2053, 2075, 1128, 1115, 1140, 1644, 2021, 1851, 338, 1827, 1105, 451, 1161, 1621, 1149, 2226, 2249, 658, 1213, 565, 1240, 1152, 1568, 1599, 437, 412, 1142, 1253, 2301, 1353, 554, 1938, 1639, 872, 2093, 2130, 2163, 1977, 1418, 582, 1774, 1452, 944, 2144, 1849, 1883, 1914, 1554, 601, 74, 417, 175, 349, 655, 210, 253, 119, 1376, 1637, 1741, 1545, 241, 1241, 1327, 360, 120, 851, 1372, 2204, 362, 317, 588, 411, 414, 50, 28, 273, 339, 1855, 1496, 1404, 1950, 1243, 587, 282, 141, 252, 48, 254, 342, 130, 1888]. statistics [3, 1744, 2283, 31, 446, 369, 313, 94, 280, 176, 2127, 12, 300, 423, 618, 656, 660, 583, 422, 359, 615, 630, 113, 178, 776, 318, 314]. statisticus [67, 212, 26]. Statistiek [313, 94, 280, 12, 565, 437, 660, 583, 615, 113, 178, 776, 318, 314, 601, 74, 417, 175, 349, 655, 210, 253, 241, 362, 317, 588, 411, 414, 50, 28, 587, 282, 141, 252, 48, 254, 342, 31, 446, 369, 176, 658, 300, 423, 618, 1141, 656, 422, 359, 630, 412]. Statistieken [338]. Statistique [100]. Statistisch [204, 225, 308, 609, 72, 205, 87, 65, 307, 6, 267, 296, 334, 387, 433, 478, 517, 564, 647]. Statistische [707, 278, 344, 377, 14, 24, 312, 521, 220, 224, 247, 237, 85, 86, 114, 773, 354, 34, 534, 58, 323, 563, 370, 719, 25, 239, 23, 303, 77, 294, 358, 388, 179, 280, 155, 89]. statistisches [88]. Statistischo [52]. statistitche [2]. Status [1622, 1777, 2078, 2160, 2155]. stay [2167]. steady [1053]. steekproef [76, 58, 143, 326, 459, 395]. steekproef-schema’s [58]. steekproefbreedtes [265]. steekproefgemiddelden [59]. Steekproefonderzoek [180]. steekproefschema [138]. steekproefsysteem [76]. Steekproeftrekking [884, 888]. Steekproeven [701, 259, 107, 70, 218, 525, 706, 353, 269, 183]. Steel [602]. Steer [655, 658, 656]. Steerneman [1754]. Stein [1572, 2229, 924]. Stein-type [1572]. stel [182]. stemming [1353]. Stemmingen [751]. step [2103, 2173, 1889, 2207, 1485, 2211, 1646, 2086, 1757, 503]. step-and-scan [1485]. step-stress [2173, 1889, 1757]. stepper [1485]. Stepwise [1564]. Steriel [722]. sterile [722, 722]. Stichting [446, 238]. stijgingspercentage [704]. Stochastic [1521, 2161, 1988, 1014, 2278, 1144, 950, 667, 1176, 1033, 574, 1974, 2254, 611, 2138, 638, 2187, 431, 724, 1479, 2051, 562, 1586, 1578, 2098, 1843, 376, 503]. stochastics [1834]. Stochastische [376, 503, 431, 562]. Stock [1443, 1881, 2042, 2055]. stofgehalte [72]. stofonderzoek [95]. stopping [1507]. storage [448]. storms [286]. straight [245, 1986]. strata [508]. Strategies [1351, 2077, 1582, 1485, 1937, 1505, 1558]. stratification [696]. Stratified [1168, 1958, 2255, 1974, 2269, 1222, 59]. stream [507, 600, 572]. strength [309, 680]. stress [2173, 1889, 2207, 1757]. stretched [685]. strict [2030]. strijd [912, 415, 889]. Stringer [1476]. Strong [1845]. Structural [1535, 1896, 1707, 786, 1327, 2189, 1958, 1865]. Structure [1047, 935, 949, 2050, 834, 1814, 2187, 1328, 1919, 2186, 1430, 2104, 793]. structured [1973]. structures [2190]. struggle [415]. student [1868, 2081]. Studentization [1738, 934]. Studentized [1194]. studie [626]. Studiekring [521]. Studies [1619, 916, 921, 1878, 2245, 489, 724, 460, 1867, 1958]. Study [987, 1688, 1604, 1022, 1341, 1786, 1712, 1985, 1453, 400, 2166, 1667, 1095, 915, 955, 2217, 1382, 1919, 2290, 1971, 137, 410, 77, 1038, 1383, 626, 521]. stukscontrôle [7]. Sturen [655, 658, 656]. stuwende [303]. styles [2110]. subaggregate [2024]. subannual [2120]. subcohort [2245]. Subcommissie [180]. Subcommittee [180]. subdistribution [2298]. subintervals [728]. Subject [1443, 1333, 1930, 2128, 2187]. subjective [876, 821, 821]. subjects [1063]. submission [1661]. subpopulaties [706]. subpopulations [706]. subsampling [2247]. subset [1360]. subsistence [436]. Substitute [289]. Substituut [289]. Substituut-variabelen [289]. success [1655, 2133]. successes [1426]. Successive [1198, 810, 843, 432]. sufficiency [2313]. Sufficient [1161]. Suggestie [747]. Suggestion [747]. Suid [113]. Suid-Afrika [113]. sum [1486, 1987, 1027, 2017, 2029, 2026, 1311, 1295, 973]. sum-constrained [1486]. Summary [173, 497, 195]. sums [1645, 1483, 1097]. sunspots [96]. Superimposed [152]. superiority [1276]. Superpositie [75]. Superposition [75]. Supplementary [1162]. supplier [843]. supplier-customer [843]. supply [2223, 462, 498, 448]. sure [1636]. surface [1781, 2295]. surfaces [1925, 2047]. surgical [1885]. surrounding [318]. Surveillance [1812]. Survey [281, 1147, 1282, 1685, 2183, 1058, 1935, 1694, 22, 1458, 1077, 529, 1909, 1869, 1905, 1479, 91, 501, 180]. Surveys [1611, 1613, 471, 327, 531, 402]. Survival [1605, 2225, 1898, 2255, 2216, 2298, 1691, 1095, 2308, 1709, 2097, 1826, 1460, 2007, 2062, 1982, 1497]. survivors [2216]. Swedish [2129]. switching [2107, 2072]. symbol [361]. symbolen [120]. symbols [120]. symbool [361]. symmetric [2087, 2233, 1575, 812]. symmetrical [845]. symmetrietoets [297, 378]. symmetrische [812]. symmetry [914, 297, 991, 378, 1805]. systeem [228, 77, 472]. System [1303, 1178, 1420, 1005, 1211, 763, 771, 1875, 76, 228, 811, 830, 871, 786, 1053, 1743, 77, 472, 2124]. systematic [58, 1824, 109]. Systematische [109, 58]. systemen [382]. systemic [1988]. Systems [1584, 2119, 1107, 1045, 1402, 1717, 1582, 1537, 1830, 1305, 1831, 1909, 1480, 843, 339, 1101, 1579, 1911, 382, 859, 760, 938].

T [558, 1129]. taak [300]. taal [215]. Taalstatistiek [582]. Tabel [711, 269, 733]. tabelkunde [381]. Tabellen [394, 509, 344, 388, 563]. table [758, 2125, 269, 393, 733, 711]. table-balancing [2125]. Tables [1610, 200, 509, 2301, 344, 388, 1863, 1553, 1844, 1832, 1853, 1552, 563, 787, 1318, 529, 385, 983, 2026, 375, 1805, 1958, 329, 394]. Tabulation [381]. Tail [1496, 1515, 1296, 1803, 2297]. tailed [1585, 845]. tails [745, 725, 1589]. taking [2258, 836, 2002, 910, 525]. Target [1617]. task [300]. taste [400, 309]. tau [1882]. tau-equivalent [1882]. taught [1782]. Tax [1686]. te [380, 149, 105, 33, 244]. team [2133, 364]. technical [865]. techniek [614, 521]. technieken [925]. technique [1694, 569]. Techniques [1206, 1604, 1451, 1551, 2127, 1820, 1660, 1208, 925, 1242, 668]. technische [865]. technology [833, 614, 521]. teeth [98]. tegen [591, 159, 246, 516]. tegenspeler [161]. telecommunication [1733]. telefonie [16]. telefoonboek [298]. telefoongids [527]. telephone [527, 298]. telephony [16]. Temporal [2123, 2119, 2122, 1854, 1973]. Tempus [2183]. Ten [253, 252, 254]. Tendency [471, 402]. tender [447]. tendered [45]. tenders [447, 519, 45]. tentative [790]. Ter [906, 260, 326, 444]. Term [2290, 2216, 2091, 283, 17, 35, 1413, 277, 290]. termen [278]. termijn [283, 17, 35]. terminal [2146]. termination [2173]. terms [278, 2159]. terrein [126, 283]. Terry [962, 621, 668]. teruglegging [733]. Test [1183, 1655, 1046, 1108, 1193, 1185, 1114, 1172, 1122, 165, 720, 1572, 1543, 297, 1060, 2255, 554, 2271, 1997, 2274, 1052, 1409, 71, 479, 2210, 835, 139, 752, 1719, 394, 218, 470, 469, 1999, 991, 2246, 862, 1652, 764, 1073, 1976, 61, 386, 1059, 966, 307, 602, 19, 1328, 1090, 1998, 2086, 231, 273, 378, 907, 797, 792, 78, 1546, 877, 1982, 894, 640, 641, 103, 1422, 1757, 329, 1787, 542, 1383, 166, 246, 397, 544, 637, 838, 1012, 164, 218, 718, 733, 394]. Testing [2257, 2304, 1875, 1992, 1728, 1870, 1031, 1816, 2050, 1390, 1702, 991, 2269, 1213, 2288, 1430, 1980, 467, 758, 1717, 1997, 1968, 856, 117, 1929, 1999, 2211, 2087, 1976, 631, 486, 1234, 784, 1998, 1972, 1933, 2253, 207, 897]. Tests [1755, 1113, 1201, 1684, 1877, 306, 1145, 670, 1136, 1121, 2239, 2229, 1748, 1997, 1863, 1333, 1343, 1367, 837, 1232, 841, 1072, 1023, 1508, 1058, 800, 870, 1732, 2173, 972, 1272, 1912, 2030, 1517, 1999, 1410, 1040, 1976, 1064, 1065, 1271, 602, 1998, 2026, 375, 37, 78, 2013, 717, 2045, 1817, 269, 429, 1221, 671, 726, 328, 516, 621, 668]. textiel [237]. textielindustrie [114, 89]. Textile [89, 114]. textiles [237]. Thank [1472]. thans [31]. Theil [577, 549, 1616, 1126]. their [2312, 163, 529, 1986, 796, 449]. Themes [1146]. Theorem [1167, 1361, 1391, 1276, 1695, 1483, 2143, 2088, 1234, 124]. theorems [1636, 1013, 2006, 1364, 1905, 724]. theoretic [736]. Theoretical [428, 219]. theoretische [219]. theorie [902, 337, 257]. Theorms [1156]. Theory [1451, 1048, 1147, 893, 1150, 1411, 833, 902, 887, 883, 831, 1918, 1573, 2248, 611, 74, 337, 1455, 842, 1305, 1890, 2046, 1419, 1327, 571, 754, 784, 407, 257, 79, 956, 202, 1950, 16, 150, 162, 793]. these [621]. Theses [999, 1006, 1075, 1080, 1109, 1116, 1124, 1154, 1196, 1203, 1216, 1223, 1237, 1255, 1262, 1267, 1273, 1279, 1291, 1312, 1331, 1339, 1354, 1365, 1414, 1439, 1489, 1509, 1528, 1556, 1565, 1595, 1606, 1163, 1380, 1229, 1461, 1179, 1092, 1055, 1248, 1209]. Thin [1022, 448]. thinned [2314, 685]. thinning [578, 2194]. thinning-based [2194]. third [84, 1886, 1225]. third-order [1886]. Thompson [666]. those [602]. thoughts [176]. Three [2091, 1684, 2231, 2251, 1553, 972, 1971, 2047, 637, 50]. three-dimensional [1553, 2047]. Three-period [2091]. three-phase [2251]. Threshold [2251, 1926, 2138, 454]. thresholding [1658, 1659]. Thresholds [1925]. tickets [1495]. tidal [629]. tie [2183]. tie-oriented [2183]. Tien [253, 252, 254]. ties [1755, 273]. Tight [1796]. tijdens [132, 72, 101]. Tijdreeksen [553, 622]. tijdschrift [122]. tijdstudieproblemen [77]. Time [1608, 2119, 1611, 1177, 1219, 931, 1686, 1615, 1605, 1211, 1176, 1835, 2229, 1812, 2064, 1657, 2304, 2126, 1825, 1822, 1897, 1984, 1785, 2183, 2209, 2197, 1889, 1032, 1824, 1921, 1926, 1890, 1707, 2061, 473, 2074, 1759, 2170, 1922, 1820, 2306, 2168, 796, 1811, 676, 2122, 496, 2007, 2062, 2235, 2186, 1821, 1513, 383, 1788, 1647, 1864, 1487, 1586, 622, 798, 2104, 2239, 2101, 77, 1288, 2116, 603, 1270, 322, 2012, 553]. time-dependent [2168]. time-invariant [2239]. Time-series [1608, 1897, 1890]. time-to-event [2197, 1707, 2170]. time-varying [1707, 2239, 2012]. Times [1231, 1800, 365, 2173, 529, 243, 1810, 1382, 1885, 863, 364, 421]. Tinbergen [1876]. tire [617]. tits [820]. Tobit [2030]. today [776]. toegepast [13, 182]. Toegepaste [417, 601, 595, 150, 189]. toekomst [565]. toepas [136]. toepas-singen [136]. toepassen [71]. Toepassing [317, 746, 16, 162, 902, 107, 129, 709, 908, 241, 61, 939, 630, 77, 59, 353]. toepassingen [390, 163, 717, 262, 295]. toepassingsmogelijkheden [511]. Toespraak [238]. toestand [190, 13]. Toets [218, 718, 733, 479, 752, 542, 166, 246, 165, 164, 394]. toetsen [657, 117, 602, 207, 269, 516]. toetsing [467]. toetsingsgrootheid [397]. toetsingsprogramma [631]. Toeval [783]. toevallige [109]. Toevals [64, 70]. Toevals-cijfers [64]. Toevals-frequenties [70]. toevalscijfers [156]. toevalsreeksen [181]. tolerance [800, 1781, 2132, 657]. tolerances [497, 4]. Tolerantiegebieden [657]. toleranties [497, 4]. Tomorrow [566, 567]. Ton [1754]. tool [1771, 70, 119, 460, 2040]. toolkit [1666]. tools [14, 24, 1934, 355]. top [484, 2133]. top-level [2133]. Topics [1950]. topologies [1469]. Tore [1348]. tot [494, 288, 122, 579, 280, 458, 120, 425, 381, 782, 456, 495, 158, 452, 679]. total [2024, 1452, 951, 2026, 729]. totals [2120, 710, 768]. tournament [1919]. Toxiciteitsbepaling [298]. Toxicity [1619, 1878, 298]. Toxoplasmose [159]. Toxoplasmosis [159]. tracer [642]. tracer-dilution [642]. tracking [2065]. trade [1938, 2066]. traded [2009]. traffic [507, 600, 836, 651, 572, 13, 1589, 583, 956, 16, 596, 595]. Training [446, 1758]. trajectory [1822]. transaction [2055]. transform [2213, 1379]. Transformatie [677, 151, 812]. transformation [864, 151, 812, 677]. transformations [2230, 1487, 1865, 589]. Transforms [385]. Transition [1670, 1610, 1615, 1627, 1444]. transitions [1581]. transmission [1756]. transport [1938]. transportation [2294, 1751, 569]. Treatment [1758, 1809, 2091, 1707, 2168, 143, 2300]. Treatments [1193, 1333, 1072, 839, 2287]. trees [2182]. trekking [733]. trekkingskansen [753]. trenches [2122]. Trend [1613, 1172, 1972, 291, 2003]. trends [1824]. trendtest [419]. trial [287]. Trials [1182, 1669, 1898, 1656, 2262, 1426, 2007, 2062, 628]. triangulations [1035]. Tribute [817]. trimloss [512]. Trimmed [1054]. Trivariate [2147]. truck [901]. True [1126, 1971]. Truncated [1679, 2225, 2259, 811, 1525]. truth [2108]. tube [707]. tubes [87, 258]. Tuning [2018]. tussen [354, 545, 415, 151, 44, 97, 96, 587, 141]. Twee [126, 608, 218, 620, 75, 157, 269]. tweede [72, 106, 56]. tweetal [652, 516]. tweevoudige [160]. tweezijdige [117]. Twenty [1772, 521]. Twenty-five [1772]. twin [286]. Twintig [521]. Two [1046, 2206, 1173, 2030, 1185, 1194, 1646, 1479, 1005, 1176, 429, 608, 810, 1809, 1898, 1060, 688, 758, 1657, 1838, 2091, 2126, 827, 885, 1881, 1052, 837, 987, 1058, 1930, 2210, 972, 117, 508, 1318, 1912, 1808, 1512, 218, 1741, 1073, 620, 1986, 1059, 966, 2177, 2086, 75, 273, 808, 843, 466, 157, 2040, 1632, 1966, 573, 2309, 329, 1911, 269, 652, 1221, 516, 621, 637, 668, 1012, 126]. two-component [808]. two-dimensional [1838, 1881, 508]. two-period [1809]. Two-Point [1173]. two-sample [1058, 2210, 1912, 1741, 1073, 1059, 1221, 621, 637, 668, 1012]. two-sided [117]. Two-stage [1479, 1898, 1060, 1657, 1059]. Two-step [1646, 2086]. two-treatment [1809, 2091]. two-way [827, 885, 1052, 1318]. twofold [160]. Type [1113, 1572, 15, 2083, 1886, 2196, 841, 1091, 2291, 1410, 2143, 765, 1734, 1757, 643, 2173, 2207, 1999, 498, 1976, 2282, 2261, 1997, 1998]. Type-I [2173, 2282, 2261]. Type-II [2207]. Typebeperking [498]. types [880, 50, 375, 2273, 476]. typically [1859, 1861, 1860]. Typologie [26]. Typology [26].

Ueber [824]. Uhlenbeck [2006, 2088]. uit [710, 625, 171, 774, 265, 13, 701, 243, 407, 148, 562, 183]. uit-voering [625]. Uitbijternegerende [532]. uitbijters [377]. Uitgedund [685]. uitgegeven [278]. uitkomsten [602]. Uitschietende [487]. Uitslag [399]. uitvalpercentage [110]. uitzetten [208]. Ultimate [1918]. ultrasonic [2139]. ultraviolet [707]. unbalanced [1308, 1524, 2287]. unbiased [477]. uncertain [2230]. Uncertainty [695, 646, 2193, 2110, 742, 756, 1960, 690, 714]. Unconditional [2026, 2151]. uncorrelated [2159, 435]. under-reporting [1930]. undergraduate [2195]. Underlining [634]. underlying [2133]. underreported [2215]. Undersize [1371]. undersized [1486]. Understanding [1905, 845, 1871]. unemployment [1758]. unequal [1333, 1004, 753, 888, 884]. uni [2269]. uni- [2269]. Unification [2092, 1587]. unified [1532]. uniform [1548, 1066, 1363, 1302, 1672, 2017, 1404, 1931]. Uniformity [1108]. unimodal [1394, 1555]. Unimodality [1226, 1192, 1187, 1247, 1032, 1246]. unique [1695, 2009]. Uniqueness [1419]. unit [1748, 1732, 2208]. units [1869]. Unity [134]. univariate [819, 1064]. univariate- [819]. Universe [32]. university [1868]. Universum [32]. unknow [847]. unknown [764, 797, 207, 1384]. unmixing [620]. unobserved [2001]. Unordered [1137, 2287]. unproductiveness [346, 399]. Unrelated [1345, 2177, 1005, 1932, 1899]. unseen [1747]. updated [1077]. upon [1070]. Upper [729, 951, 110]. upwards [1859, 1861, 1860]. use [624, 149, 156, 1351, 800, 117, 200, 1952, 394, 2168, 116, 894, 1511, 1749, 129, 147]. used [909]. Useful [1161, 919]. usefulness [130]. user [1879]. Using [1613, 1866, 1972, 1162, 1679, 707, 1981, 2085, 1655, 1711, 688, 1385, 2172, 1957, 1825, 1343, 616, 911, 1691, 1930, 2210, 1535, 2181, 1965, 830, 1448, 1667, 1702, 370, 218, 1726, 1824, 2230, 1942, 619, 1926, 1707, 2061, 1814, 2057, 2127, 2097, 159, 298, 2038, 1847, 256, 1563, 1823, 1427, 1647, 593, 2292, 1730, 59, 2116, 1865, 652, 39, 2012, 88, 240]. Usual [2254]. Uven [357, 550].

V [230]. vacancies [1074]. vaccin [318]. vaccine [318]. vacuum [844]. vacuum-cleaner [844]. Valencian [2231]. Validate [1159]. Validation [1675, 1602, 1867, 2052, 2041]. validity [1947, 1548]. value [1941, 2274, 1918, 2294, 2282, 531, 639, 454]. Valued [1236, 2228, 1468, 2285, 2218]. Values [1215, 1316, 1728, 1702, 1727, 822, 385, 1933]. valuta [401]. valuta-arbitrage [401]. Varangot [139]. variabele. [579]. variabelen [545, 289, 157]. variabelen. [574]. variability [2028, 1513, 1251]. Variable [1035, 2023, 1603, 2036, 1737, 579, 2010, 2057, 1520, 1626, 2169, 1813, 878, 1725]. Variables [1494, 1146, 1134, 1367, 986, 850, 223, 2299, 1947, 1018, 688, 1953, 545, 2203, 1289, 574, 1987, 1089, 2144, 2254, 1323, 2142, 2248, 1468, 1031, 634, 1321, 485, 1740, 2138, 977, 990, 1433, 2017, 2029, 1344, 1326, 995, 1097, 289, 1858, 157, 2040, 2008, 1653, 1325, 1225, 1865, 804]. Variance [1135, 1228, 1460, 1060, 827, 885, 1917, 1089, 1308, 2002, 1031, 770, 500, 1631, 2056, 619, 764, 834, 2153, 998, 1208, 797, 2154, 1420, 418, 1010, 186]. variances [266, 1716, 870, 1390, 2059]. variant [981]. Variantie [186, 500]. Variantie-analyse [186, 500]. Variantieanalyse [619]. variate [1363, 2308, 385]. Variates [1112, 466]. variatiecoefficient [395]. variaties [45]. variatiestatistiek [130]. Variation [1120, 1902, 951, 729, 395, 130]. variations [45]. variety [1543]. variograms [1799]. various [2069]. varying [2085, 2198, 1953, 1934, 1926, 1707, 1864, 1899, 2239, 2012]. vaststellen [147]. vaststelling [326]. VCG [1798]. vector [1572, 1235, 1870, 2081, 727]. Vectors [1227, 727]. vehicle [1304]. veldproeven [37, 78]. ver [542]. verantwoordelijkheden [212]. verband [354, 467, 625, 44, 386, 556, 96, 244, 237]. Verbeek [1357]. verbond [97]. verdeeld [473]. verdeelde [333, 443, 544]. verdeeling [29]. verdelen [706]. verdeling [608, 390, 687, 72, 488, 265, 701, 97, 326, 395, 183, 732, 43]. verdelingen [677, 484, 752, 620, 717, 434]. Verdelingsvrije [157, 269, 551]. Vereeniging [31, 1141]. Vereniging [423, 782, 422]. Verfahren [88]. vergeleken [98]. vergelijken [608, 602, 207]. Vergelijking [679, 516, 58]. vergelijkingen [182]. Verhoging [350]. Veringa [775]. verkeer [651, 583]. Verkeerskunde [595]. verkeersproblemen [16]. verkiezingen [9]. Verklaring [13]. verloop [13, 246, 516]. vermijdbare [115]. verpakkingsen [25]. verricht [42]. verschijnsel [87, 13]. Verschillen [545]. verschillende [608]. Versus [1126, 1172, 1875, 1872, 1780, 1820, 1326, 1240, 1589]. vertakte [348]. vertices [1792]. vervaardiging [228]. Vervaat [1463, 1464]. verwachting [272, 678]. verwachtingen [686]. verzamelende [141]. verzekeringswetenschappen [86]. verzorgende [303]. vezels [562]. via [1722, 2298, 2074, 1664, 2023, 1416, 975]. vierde [112]. vierkant [379]. view [1831, 2122, 1434, 1962]. viewed [307]. viewpoint [956]. vijfde [145]. Vik [37]. villages [2184]. vine [2000]. Virtual [1219]. virus [1867]. Visions [1765]. visit [118]. visiting [2166]. visserijtheorie [202]. visualize [2040]. vitality [2276]. vitamin [410]. vitamine [410]. vlag [535]. vlambeveiliging [707]. vocalizations [2139]. voedingsonderzoek [126]. voering [625]. Voetbalprognoses [301]. volat [2183]. volatility [1532, 2138, 1746, 2288, 2042, 2098]. volatility-induced [1746]. volgens [327, 736, 302, 37, 291, 793]. Volkshuisvesting [48]. Volume [1539, 1597, 1624, 233, 1915, 1944, 84, 1714, 145, 56, 112]. Voor [1141, 657, 707, 608, 31, 278, 917, 461, 545, 574, 377, 712, 687, 479, 484, 180, 752, 394, 218, 265, 391, 525, 361, 521, 333, 443, 620, 423, 602, 439, 147, 116, 562, 711, 718, 896, 717, 206, 422, 630, 578, 238, 6, 59, 291, 207, 246, 269, 395, 397, 544, 130, 538, 516, 662, 90, 89]. voor-spellende [545]. vooraf [332]. voorbeelden [431]. voorkomende [454]. Voorraadbeheer [627]. voorraadketens [462]. Voorspellen [188, 193, 190, 192, 191, 189, 194, 243]. Voorspelling [187, 679]. voorspellingen [283]. voorstel [120]. voort [243]. voort-gang [243]. voortschrijdende [107]. Voorwaardelijke [686]. Voorwoord [1, 233, 84, 145, 56, 112]. Voorzitter [234]. vormen [717]. Vos [1398]. voting [1411, 751]. vraag [722]. vragen [925]. vrede [778]. vrije [159]. vroeggeborenen [410].

W [1020, 1087, 41, 242]. W. [1491]. waardebepalingen [499]. waarmede [42]. waarneem [97]. waarneem-bare [97]. Waarneming [211, 282]. waarnemingen [208, 548, 149, 107, 198]. waarnemingsfouten [91]. waarnemingsuitkomsten [217]. waarschijnlijkheids [208, 578, 793]. waarschijnlijkheids-model [578]. waarschijnlijkheids-papier [208]. waarschijnlijkheids-rekening [793]. waarschijnlijkheidspapier [129]. waarschijnlijkheidsrekening [571, 16]. waarshijnlijkheidspapier [130]. waaryoor [45]. Wachten [596]. wachtprobleem [454]. wachttijd [383]. Wachttijden [365]. Waerden [621, 668]. wafer [1485]. Waiting [1219, 1921, 1759, 1211, 365, 529, 496, 383, 603, 863, 454, 1270, 596]. Waiting-time [1759, 496]. Wald [767]. Walk [1231]. Walks [1131, 1396, 1840, 798]. Walsh [812]. Wang [1982]. wanneer [473]. Wanorde [412, 405, 415, 411, 406]. War [72, 778]. warehouse [1909]. warehouses [498]. Wat [362]. watching [1800]. water [2223, 192]. waterstanden [192]. Watson [2200]. wavelength [1796]. Wavelet [1658, 1427]. way [1060, 827, 885, 1052, 734, 1318, 2243]. ways [2109]. weak [1066]. weather [623, 17, 35, 191]. weer [191]. Weersverwachting [17, 35]. weersverwachtingen [623]. wegverkeer [596]. Weibull [1191, 2197]. weight [363, 1371]. Weighted [1639, 1202, 1188, 2255, 1492, 1478, 2217, 995, 1295, 1422, 276]. weighting [2152]. weights [2141, 276]. wel [31]. welke [707, 203]. Welsch [1020]. wereld [776]. wereldoorlog [72]. werk [180]. werkelijke [97]. werking [170]. werkprogramma [380]. werktijden [147]. Werkwijze [511]. werkzaamheid [489]. Wessels [1894, 1578]. West [286]. western [1908]. wet [678, 740, 738]. Weten [500, 504, 494, 495]. wetenschap [595, 256]. wetenschappen [213, 495]. wetmatigheid [783]. where [1749]. Which [1985, 203, 42, 1961]. white [1409]. Who [645]. Wicksell [1371]. width [527, 203]. widths [265]. Wie [645]. Wiener [1640]. wij [362]. wijzer [109]. wijzigingen [475]. Wilcoxon [297, 273, 378, 575, 329, 166, 621, 668, 1012, 165, 164]. Wilk [877, 718]. Willemze [132]. William [320]. Wim [1463, 1464]. winnaar [751]. winner [751, 2222]. winning [1505]. Winstgevendheid [401]. winstkansen [401]. winters [96]. Wiskunde [607, 407]. wiskundige [74, 143, 252]. within [1881]. Without [1175, 1590, 2024, 1318, 2046, 965, 1966, 1413, 542, 543, 733, 751]. WLS [1384]. Wolff [525]. woningbouw [556]. woningkwaliteit [556]. Woord [332]. word [1421]. worden [42, 629]. words [780]. wordt [45]. work [1777, 380, 180, 728, 1026, 127, 1444]. working [1769, 1709, 147]. works [1837]. World [72, 1919, 1918, 776]. writing [1545]. Wrong [1283, 1712, 1961, 1993, 1960].

yarn [562]. yarns [617]. Yates [2301]. year [196, 755, 1768, 332, 1888].

Years [1140, 1731, 1772, 1766, 1761, 1713, 1765, 253, 521, 1767, 2290, 252, 254, 1456]. yield [78]. yields [193]. Yigit [1997, 1998].

Zaat [648]. zero [2142, 2248, 1994, 1877, 2046, 2235, 1958]. zero-inflated [1994, 1877, 2235]. zero-mean [2142]. zeros [2203, 1979]. zevende [196]. ziek [779]. ziekenhuizen [411]. ziekte [115]. ziektegeval [456]. ziektengetal [456]. ziet [26]. zijde [426, 219]. zijn [591, 473, 708]. Zijne [775]. Zipf [2259]. Zitplaats [137]. Zitplaats-accomodatie [137]. zitting [100]. zonder [542, 543, 733, 751]. Zone [1617]. zonnevlekken [96]. zur [88].

References

Anonymous:1946:VDF

[1] Anonymous. Voorwoord. (Dutch) [Foreword]. Statistica (Vereniging voor Statistiek), 1(1):1–2, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1946.tb00039.x.

vanEttinger:1946:SAD

[2] J. van Ettinger. De statistitche analyse in dienst van het herstel. (Dutch) [Statistical analysis in the service of recovery]. Statistica (Vereniging voor Statistiek), 1(1):3–5, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1946.tb00040.x.

vanRest:1946:ASE

[3] E. D. van Rest. Applied statistics in England. Statistica (Vereniging voor Statistiek), 1(1):5–11, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1946.tb00041.x.

Sittig:1946:TPM

[4] J. Sittig. Over toleranties on passingen in den machinebouw. (Dutch) [About tolerances and adjustments in mechanical engineering]. Statistica (Vereniging voor Statistiek), 1(1):11–24, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00042.x.

Anonymous:1946:ADA

[5] Anonymous. Aanhangsel. (Dutch) [Appendix]. Statistica (Vereniging voor Statistiek), 1(1):24–27, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1946.tb00043.x.

deJong:1946:EMV

[6] A. J. de Jong. Eenvoudige mathode voor het interpreteeren van statistisch contröle-materiaal. (Dutch) [Simple math method for interpreting statistical control material]. Statistica (Vereniging voor Statistiek), 1(1): 28–39, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00044.x.

Enters:1946:BVS

[7] J. H. Enters. De betrouwbaarheid van stukscontrôle. (Dutch) [The reliability of piece checking]. Statistica (Vereniging voor Statistiek), 1(1): 40–44, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00045.x.

Bok:1946:S

[8] S. T. Bok. Statistica. Statistica (Vereniging voor Statistiek), 1(1):45– 46, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00046.x.

Vorangot:1946:OVD

[9] W. Vorangot. Het opinieonderzoek en de verkiezingen. (Dutch) [The opinion poll and the elections]. Statistica (Vereniging voor Statistiek), 1(1):47–50, February 1946. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1946.tb00047.x.

Bok:1946:BDB

[10] S. T. Bok. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 1(1):50–52, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1946.tb00048.x.

Anonymous:1946:MDAa

[11] Anonymous. Mededeelingen. (Dutch) [Announcements]. Statistica (Vereniging voor Statistiek), 1(1):52, February 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00049.x.

Neher:1946:SBS

[12] L. Neher. Statistiek bij het Staatsbedrijf der PIT. (Dutch) [Statistics at the state company of the PIT]. Statistica (Vereniging voor Statistiek), 1 (2):53–55, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00050.x.

Lisman:1946:VVV

[13] J. H. C. Lisman. Verklaring van het verloop van een verschijnsel uit den toestand op één oogenblik, toegepast op het giroverkeer. (Dutch) [Explanation of the development of a phenomenon from the situation at one moment, applied to giro traffic]. Statistica (Vereniging voor Statistiek), 1 (2):55–63, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00051.x.

Enters:1946:SHBa

[14] J. H. Enters. Statistische hulpmiddelen bij de kwaliteitsbeheersing I. (Dutch) [Statistical tools for quality control I]. Statistica (Vereniging voor Statistiek), 1(2):63–78, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1946.tb00052.x.

Bok:1946:SKD

[15] S. T. Bok. Slaapmiddelen en karaktertype. (Dutch) [Sleeping aids and character type]. Statistica (Vereniging voor Statistiek), 1(2):79– 90, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00053.x.

deBruyn:1946:TWO

[16] M. van Dobben de Bruyn. Toepassing der waarschijnlijkheidsrekening op enkele verkeersproblemen in de automatische telefonie. (Dutch) [Application of probability theory to some traffic problems in automatic telephony]. Statistica (Vereniging voor Statistiek), 1(2):91– 97, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00054.x.

Visser:1946:WOLa

[17] S. W. Visser. Weersverwachting op lange termijn. (Dutch) [Longterm weather forecast]. Statistica (Vereniging voor Statistiek), 1(2):

98–101, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00055.x.

Bok:1946:SOD

[18] S. T. Bok. Statistica: De overschrijdingskans. (Dutch) [Statistica: The exceedance probability]. Statistica (Vereniging voor Statistiek), 1(2):102– 107, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00056.x.

S:1946:NVT

[19] J. S. Nomogram van de χ^{2}-test of goodness of fit. (Dutch) [Nomogram of the χ^{2} test of goodness of fit]. Statistica (Vereniging voor Statistiek), 1(2): 107–111, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00057.x.

Anonymous:1946:MDAb

[20] Anonymous. Mededeelingen. (Dutch) [Announcements]. Statistica (Vereniging voor Statistiek), 1(2):112, April 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00058.x.

Anonymous:1946:COO

[21] Anonymous. Coördinotie von het opinie-onderzoek. (Dutch) [Coordination of the opinion research]. Statistica (Vereniging voor Statistiek), 1(3): 113–114, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00059.x.

Hugos:1946:OVO

[22] A. Hugos. Het object van het opinie-onderzoek. (Dutch) [The object of the opinion survey]. Statistica (Vereniging voor Statistiek), 1(3):115– 118, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00060.x.

Varangot:1946:ESA

[23] V. Varangot. Enkale statistische atpecten van het opinie-onderzoek. (Dutch) [Some statistical aspects of opinion research]. Statistica (Vereniging voor Statistiek), 1(3):119–123, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00061.x.

Enters:1946:SHBb

[24] J. H. Enters. Statistische hulpmiddelen bij de kwaliteitsbaheersing II. (Dutch) [Statistical tools at the quality control II]. Statistica (Vereniging voor Statistiek), 1(3):124–130, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1946.tb00062.x.

Sittig:1946:ESO

[25] J. Sittig. Eenige statistische opmerkingen over verpakkingsen doseeringsproblemen. (Dutch) [Some statistical comments about packaging and dosing problems]. Statistica (Vereniging voor Statistiek), 1(3):131– 135, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00063.x.

vanderBurg:1946:TVS

[26] A. R. van der Burg. Typologie van den statisticus: Hoe men hem ziet en hoe hij. (Dutch) [Typology of the statistician: How one sees him and how he]. Statistica (Vereniging voor Statistiek), 1(3):136– 144, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00064.x.

vanUven:1946:SEB

[27] M. J. van Uven. Scheefheid en exces bij frequentie- en kansverdeelingen. (Dutch) [Skewness and excess in frequency and probability distributions]. Statistica (Vereniging voor Statistiek), 1(3):145–155, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00065.x.

Sittig:1946:MSK

[28] J. Sittig. Met statistiek kan men alles bewijzen. (Dutch) [You can prove anything with statistics]. Statistica (Vereniging voor Statistiek), 1(3): 156–160, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00066.x.

Varangot:1946:EAV

[29] V. Varangot. Een eenvoudige afleiding van de verdeeling van Poisson. (Dutch) [A simple derivation of the Poisson distribution]. Statistica (Vereniging voor Statistiek), 1(3):161, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00067.x.

Anonymous:1946:BDBa

[30] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 1(3):162–163, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00068.x.

Anonymous:1946:VVS

[31] Anonymous. De Vereeniging voor Statistiek heeft thans de Koninklijke goedkeuring gekregen, en wel bij K.B. van 20 December 1946, no 72. (Dutch) [The Association for Statistics has now received royal approval, by K.B. of 20 December 1946, no. 72]. Statistica (Vereniging voor Statistiek), 1(3):164–170, June 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1946.tb00069.x.

vanEttinger:1946:BUS

[32] J. van Ettinger. Het begrip Universum in de Statistica. (Dutch) [The concept of Universe in Statistica]. Statistica (Vereniging voor Statistiek), 1(4–5):171–175, August 1946. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1946.tb00596.x.

Vermeulen:1946:KVL

[33] A. J. M. Vermeulen. De kosten van levensonderhoud te ‘s-Gravenhage in December 1946. (Dutch) [The cost of living in ‘s-Gravenhage in December 1946]. Statistica (Vereniging voor Statistiek), 1(4–5):176– 191, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00597.x.

Geiss:1946:SKE

[34] W. Geiss. De statistische kwaliteitscontrôle als economisch probleem (I). (Dutch) [Statistical quality control as an economic problem (I)]. Statistica (Vereniging voor Statistiek), 1(4–5):192–196, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00598.x.

Visser:1946:WOLb

[35] S. W. Visser. Weersverwachting op lange termijn. (Dutch) [Long-term weather forecast]. Statistica (Vereniging voor Statistiek), 1(4–5):197– 202, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00599.x.

Hekker:1946:CVK

[36] Th. Hekker. Constructie van kansellipsen in een correlatiediagram. (Dutch) [Construction of probability ellipses in a correlation diagram]. Statistica (Vereniging voor Statistiek), 1(4–5):203–208, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00600.x.

Stoffels:1946:BVV

[37] A. Stoffels. De bewerking van veldproeven volgens de methode van Knut Vik. (Dutch) [The processing of field tests according to the method of Knut Vik]. Statistica (Vereniging voor Statistiek), 1(4–5):209–216, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00601.x.

N:1946:NVR

[38] C. A. G. N. Naschrift van de redactie. (Dutch) [Postscript from the editorial staff]. Statistica (Vereniging voor Statistiek), 1(4–5):216–218, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00602.x.

vanderLoan:1946:LMB

[39] E. van der Loan and Ir J. G. W. Ignatius. Lijnvereffening met behulp van orthogonale polynomia. (Dutch) [Line smoothing using orthogonal polynomia]. Statistica (Vereniging voor Statistiek), 1(4–5):219– 230, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00603.x.

Somermeijer:1946:BVC

[40] W. H. Somermeijer. De betrouwbaarheid van controleurs. (Dutch) [The reliability of inspectors]. Statistica (Vereniging voor Statistiek), 1(4–5): 231–235, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00604.x.

Enters:1946:NBA

[41] J. H. Enters. Naschrift bij het artikel van Mr W. H. Somermeijer. (Dutch) [Postscript to the article by Mr W. H. Somermeijer]. Statistica (Vereniging voor Statistiek), 1(4–5):236–239, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00605.x.

Posthumus:1946:BVN

[42] G. Posthumus. De bepaling van de nauwkeurigheid, waarmede metingen of beoor-delingen worden verricht. (Dutch) [The determination of the accuracy with which measurements or assessments are made]. Statistica (Vereniging voor Statistiek), 1(4–5):240–248, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00606.x.

Bok:1946:SIV

[43] S. T. Bok. Statistica II. Verdeling van kansen. (Dutch) [Statistica II. Distribution of probabilities]. Statistica (Vereniging voor Statistiek), 1 (4–5):249–256, August 1946. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1946.tb00607.x.

Nass:1946:QVT

[44] C. A. G. Nass. Het quantitatieve verband tussen dosering en letaliteit bij proefdieren. (Dutch) [The quantitative relationship between dosage and lethality in laboratory animals]. Statistica (Vereniging voor Statistiek), 1(4–5):257–266, August 1946. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1946.tb00608.x.

vonderWal:1946:DNV

[45] J. J. von der Wal and J. Sittig. Dnderzoek naar de variaties in de bedragen waaryoor bij openbare aanbestedingen wordt ingeschreyen.

(Dutch) [Research into the variations in the amounts tendered for in public tenders]. Statistica (Vereniging voor Statistiek), 1(4–5):267– 272, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00609.x.

Anonymous:1946:BDBb

[46] Anonymous. Boekbeapreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 1(4–5):273–278, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00610.x.

Anonymous:1946:MDAc

[47] Anonymous. Mededelingen. (Dutch) [Announcements]. Statistica (Vereniging voor Statistiek), 1(4–5):278, August 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00611.x.

vanEttinger:1946:VSD

[48] J. van Ettinger. Volkshuisvesting en statistiek. (Dutch) [Public housing and statistics]. Statistica (Vereniging voor Statistiek), 1(6):279–292, December 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00343.x.

deJong:1946:MPB

[49] A. J. de Jong. Het monsternemen als probleem in het bedrijf. (Dutch) [Sampling as a problem in the company]. Statistica (Vereniging voor Statistiek), 1(6):293–303, December 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1946.tb00344.x.

Schwarz:1946:DSS

[50] N. Schwarz, Jr. Drie soorten statistiek in de natuurkunde?. (Dutch) [Three types of statistics in physics?]. Statistica (Vereniging voor Statistiek), 1(6):304–316, December 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1946.tb00345.x.

Post:1946:BPD

[51] J. J. Post, Jr. Blanco proeven. (Dutch) [Blank samples]. Statistica (Vereniging voor Statistiek), 1(6):317–321, December 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00346.x.

vanderBurg:1946:SAV

[52] A. R. van der Burg. Statistischo aspecten van kleding-maatnormalisatie. (Dutch) [Statistical aspects of clothing size normalization]. Statistica (Vereniging voor Statistiek), 1(6):322–326, December 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00347.x.

Anonymous:1946:BDBc

[53] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 1(6):327–331, December 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00348.x.

Anonymous:1946:E

[54] Anonymous. Erratum. Statistica (Vereniging voor Statistiek), 1(6):331– 332, December 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1946.tb00349.x.

Anonymous:1946:MDAd

[55] Anonymous. Mededelingen. (Dutch) [Announcements]. Statistica (Vereniging voor Statistiek), 1(6):332–333, December 1946. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1946.
tb00350.x.

Redactie:1948:VBT

[56] De Redactie. Voorwoord bij de tweede jaargang. (Dutch) [Foreword to the second volume]. Statistica (Vereniging voor Statistiek), 2(1–2): 1, February 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00351.x.

Dalmulder:1948:FVE

[57] J. J. J. Dalmulder. De functie van de econometrie bij het bedrijfsbeheer. (Dutch) [The function of econometrics in business management]. Statistica (Vereniging voor Statistiek), 2(1–2):2–18, February 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00352.x.

Hamaker:1948:SVV

[58] H. C. Hamaker. Een systematische vergelijking van de statistische eigenschappen van hedendaagse steekproef-schema’s. (Dutch) [A systematic comparison of the statistical properties of contemporary sampling schemes]. Statistica (Vereniging voor Statistiek), 2(1–2):19–39, February 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00353.x.

deLeeuw:1948:NCV

[59] Th. A. de Leeuw. Een noodzakelijke correctie van de overschrijdingsgrenzen voor steekproefgemiddelden bij toepassing van stratified sampling. (Dutch) [A necessary correction of the exceedance limits for sample means when using stratified sampling]. Statistica (Vereniging voor Statistiek), 2(1–2):40–54, February 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1948.tb00354.x.

Brethouwer:1948:HFD

[60] D. H. G. Brethouwer. Hyperbolische foutenwet. (Dutch) [Hyperbolic error law]. Statistica (Vereniging voor Statistiek), 2(1–2):55–73, February 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00355.x.

Nass:1948:GRM

[61] C. A. G. Nass. De grondslagen der regressieanalyse, met een toepassing op een doseringsletaliteitproef. (Dutch) [The foundations of regression analysis, with an application to a dose lethality test]. Statistica (Vereniging voor Statistiek), 2(1–2):74–94, February 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1948.
tb00356.x.

Anonymous:1948:MDAa

[62] Anonymous. Mededelingen. (Dutch) [Announcements]. Statistica (Vereniging voor Statistiek), 2(1–2):95, February 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1948.
tb00357.x.

deJong:1948:E

[63] A. J. de Jong. Erratum. Statistica (Vereniging voor Statistiek), 2(1–2): 95, February 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00358.x.

Hamaker:1948:TCD

[64] H. C. Hamaker. Toevals-cijfers. (Dutch) [Coincidence figures]. Statistica (Vereniging voor Statistiek), 2(3):97–106, June 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1948.
tb00359.x.

Levert:1948:SPK

[65] C. Levert. Een statistisch probleem in de Klimatologie. (Dutch) [A statistical problem in Climatology]. Statistica (Vereniging voor Statistiek), 2 (3):107–119, June 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00360.x.

Lisman:1948:EFM

[66] J. H. C. Lisman. Enkele frequentieverdelingen in de muziek. (Dutch) [Some frequency distributions in music]. Statistica (Vereniging voor Statistiek), 2(3):120–123, June 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1948.tb00361.x.

Lisman:1948:PVS

[67] J. H. C. Lisman. De plaats van de statisticus in het bedrijf. (Dutch) [The position of the statistician in the company]. Statistica (Vereniging voor Statistiek), 2(3):124–126, June 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1948.tb00362.x.

Anonymous:1948:BDBa

[68] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 2(3):127–128, June 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1948.
tb00363.x.

Anonymous:1948:MDN

[69] Anonymous. Mededeling. (Dutch) [Notice]. Statistica (Vereniging voor Statistiek), 2(3):128, June 1948. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1948.tb00364.x.

Hamaker:1948:TFH

[70] H. C. Hamaker. Toevals-frequenties, een hulpmiddel bij de constructie van kunstmatige steekproeven van grote omvang. (Dutch) [Random frequencies, a tool for the construction of artificial samples of large sizes]. Statistica (Vereniging voor Statistiek), 2(4):129–137, August 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00365.x.

Enters:1948:MMB

[71] J. H. Enters. Mogelijkheden en moeilijkheden bij het toepassen van de sequentietest. (Dutch) [Possibilities and difficulties in applying the sequence test]. Statistica (Vereniging voor Statistiek), 2(4):138–154, August 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00366.x.

Hekker:1948:SOV

[72] Th. Hekker and Ir J. Straub. Statistisch onderzoek van contrôle-analyse cijfers: (De verdeling van het droge stofgehalte van het brood tijdens de tweede wereldoorlog). (Dutch) [Statistical research of control analysis figures: (The distribution of the dry matter content of bread during the Second World War)]. Statistica (Vereniging voor Statistiek), 2(4):155– 171, August 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00367.x.

Anonymous:1948:BDBb

[73] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 2(4):172–173, August 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1948.
tb00368.x.

Hamaker:1948:FWS

[74] H. C. Hamaker. Foutentheorie en wiskundige statistiek. (Dutch) [Error theory and mathematical statistics]. Statistica (Vereniging voor Statistiek), 2(5–6):177–205, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1948.tb00369.x.

Sittig:1948:SVT

[75] J. Sittig. Superpositie van twee frequentieverdelingen. (Dutch) [Superposition of two frequency distributions]. Statistica (Vereniging voor Statistiek), 2(5–6):206–227, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1948.tb00370.x.

Enters:1948:OVS

[76] J. H. Enters. De omvang van de steekproef bij een enkelvoudig steekproefsysteem. (Dutch) [The size of the sample in a single sampling system]. Statistica (Vereniging voor Statistiek), 2(5–6):228–234, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00371.x.

deLeeuw:1948:EOM

[77] Th. A. de Leeuw. Enkele opmerkingen over de mogelijkheid van toepassing van mathematisch-statistische methoden bij tijdstudieproblemen en in het bijzonder bij het Bedaux-systeem. (Dutch) [Some comments on the possibility of applying mathematical-statistical methods to time study problems and in particular to the Bedaux system]. Statistica (Vereniging voor Statistiek), 2(5–6):235–241, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00372.x.

Stoffels:1948:BVR

[78] A. Stoffels. De bewerking van de resultaten van veldproeven door herleiding van de opbrengstcijfers in beide richtingen van het proefveld. (Dutch) [The processing of the results of field tests by converting the yield figures in both directions of the test field]. Statistica (Vereniging voor Statistiek), 2(5–6):242–249, October 1948. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1948.tb00373.x.

Varangot:1948:CHV

[79] V. Varangot. Over correlatierekening en de herleiding van klimatologische reeksen. (Dutch) [On correlation theory and the reduction of climatological series]. Statistica (Vereniging voor Statistiek), 2(5–6):250– 252, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00374.x.

vdBurg:1948:AHD

[80] A. R. v. d. Burg. Academische humbug. (Dutch) [Academic humbug]. Statistica (Vereniging voor Statistiek), 2(5–6):253–255, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00375.x.

Anonymous:1948:BDBc

[81] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 2(5–6):256–260, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1948.
tb00376.x.

Anonymous:1948:MDAb

[82] Anonymous. Mededelingen. (Dutch) [Announcements]. Statistica (Vereniging voor Statistiek), 2(5–6):260, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1948.
tb00377.x.

Anonymous:1948:E

[83] Anonymous. Errata. Statistica (Vereniging voor Statistiek), 2(5–6): 260, October 1948. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1948.tb00378.x.

Bestuur:1949:VBD

[84] Het Bestuur. Voorwoord bij de derde jaargang. (Dutch) [Foreword to the third volume]. Statistica (Vereniging voor Statistiek), 3(1):1– 2, February 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00379.x.

vanEttinger:1949:SED

[85] J. van Ettinger. Statistische ervaringen in 1948. (Dutch) [Statistical experiences in 1948]. Statistica (Vereniging voor Statistiek), 3(1):3– 7, February 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00380.x.

vanTulder:1949:SRV

[86] J. J. M. van Tulder. Statistische research in de verzekeringswetenschappen. (Dutch) [Statistical research in insurance sciences]. Statistica (Vereniging voor Statistiek), 3(1):8–18, February 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00381.x.

Hamaker:1949:RRS

[87] H. C. Hamaker. De ruis in radiobuizen als statistisch verschijnsel. (Dutch) [Noise in radio tubes as a statistical phenomenon]. Statistica (Vereniging voor Statistiek), 3(1):19–30, February 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00382.x.

vonNoss:1949:SVQ

[88] C. A. G. von Noss. Ein statistisches Verfahren zur qualitativen Bestimmung der Immunität mittels Hautreaktionen. (German) [A statistical method for the qualitative determination of immunity using skin reactions]. Statistica (Vereniging voor Statistiek), 3(1):31–41, February 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00383.x.

vdB:1949:MSD

[89] A. R. v. d. B. Madedelingen: Statistische dagen voor de textielindustrie. (Dutch) [Made announcements: Statistical days for the textile industry]. Statistica (Vereniging voor Statistiek), 3(2):42–44, April 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00384.x.

vanderWaf:1949:BVM

[90] G. van der Waf. De betekenis van massaal consumenten-onderzoek voor het bedrijfsleven. (Dutch) [The significance of massive consumer research for business]. Statistica (Vereniging voor Statistiek), 3(2):45– 48, April 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00385.x.

Sittig:1949:NPW

[91] L. Sittig. Net probleem der waarnemingsfouten bij een massaal onderzoek. (Dutch) [Just the problem of observational errors in a massive survey]. Statistica (Vereniging voor Statistiek), 3(2):49–68, April 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00386.x.

deFroe:1949:AOA

[92] A. de Froe. Anthropologische opmerkingen over anthropo-sociologische feiten. (Dutch) [Anthropological comments on anthropo-sociological facts]. Statistica (Vereniging voor Statistiek), 3(2):69–74, April 1949. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00387.x.

Anonymous:1949:BDBa

[93] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 3(2):75–76, April 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00388.x.

Hekker:1949:SVL

[94] Th. Hekker. Statistiek van lineaire macromoleculen. (Dutch) [Statistics of linear macromolecules]. Statistica (Vereniging voor Statistiek), 3(2): 77–83, April 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00389.x.

vdBurg:1949:SBD

[95] L. v. d. Burg. Over de spreiding bij het droge-stofonderzoek van brood. (Dutch) [About the distribution in the dry matter examination of bread]. Statistica (Vereniging voor Statistiek), 3(3):85– 91, June 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00390.x.

Visser:1949:VTZ

[96] S. W. Visser. Een verband tussen zonnevlekken en extreme winters. (Dutch) [A link between sunspots and extreme winters]. Statistica (Vereniging voor Statistiek), 3(3):92–99, June 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00391.x.

Picard:1949:VTW

[97] H. C. Picard. Het verbond tussen werkelijke verdeling, foutenverdeling en waarneem-bare verdeling. (Dutch) [The connection between real distribution, error distribution and observable distribution]. Statistica (Vereniging voor Statistiek), 3(3):101–107, June 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1949.tb00392.x.

Hooijer:1949:EVN

[98] D. A. Hooijer. De evolutie van net gebit van de orang oetan vergeleken met de evolutie van het gebit van de mens. (Dutch) [The evolution of orangutan teeth compared to the evolution of human teeth]. Statistica (Vereniging voor Statistiek), 3(3):108–112, June 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00393.x.

Mijsberg:1949:BBD

[99] W. A. Mijsberg. De Benelux Biometrische Dag op 2 Juli j.l.. (Dutch) [The Benelux Biometric Day on 2 July 1949]. Statistica (Vereniging voor Statistiek), 3(3):113–116, June 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1949.tb00394.x.

Anonymous:1949:ZVI

[100] Anonymous. De 26e zitting van het Institut International de Statistique. (Dutch) [The 26th session of the Institut International de Statistique]. Statistica (Vereniging voor Statistiek), 3(3):116–117, June 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00395.x.

Willemze:1949:MVC

[101] F. G. Willemze. Een methode van contrôle tijdens de fabricage. (Dutch) [A method of control during manufacture]. Statistica (Vereniging voor Statistiek), 3(4):119–132, August 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1949.tb00396.x.

Brethouwer:1949:COH

[102] D. H. G. Brethouwer and H. C. Hamaker. Critiek op de “Hyperbolische Foutenwet”. (Dutch) [Criticism of the “Hyperbolic Error Law”]. Statistica (Vereniging voor Statistiek), 3(4):133–143, August 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00397.x.

Varangot:1949:SDE

[103] V. Varangot. Een schattingsproef. (Dutch) [An estimation test]. Statistica (Vereniging voor Statistiek), 3(4):144–153, August 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00398.x.

Anonymous:1949:BDBb

[104] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 3(4):154–158, August 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00399.x.

Hamaker:1949:ISI

[105] H. C. Hamaker. Het “Internationaal statisisch Instituut” en het Congres te Bern van 5–10 Sept. 1949. (Dutch) [The “International Statistical Institute” and the Congress in Bern of 5–10 September. 1949]. Statistica (Vereniging voor Statistiek), 3(5–6):159–172, October 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00400.x.

Nass:1949:TCV

[106] C. A. G. Nass. Het tweede congres van de Biometric Society. (Dutch) [The Second Congress of the Biometric Society]. Statistica (Vereniging voor Statistiek), 3(5–6):173–178, October 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00401.x.

Enters:1949:ANW

[107] J. H. Enters. Het aantal noodzakelijke waarnemingen bij toepassing van voortschrijdende steekproeven. (Dutch) [The number of observations necessary when applying rolling sampling]. Statistica (Vereniging voor Statistiek), 3(5–6):179–200, October 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1949.tb00402.x.

Theil:1949:OVC

[108] H. Theil. Over de ongelijkheid van Camp en Meidell. (Dutch) [On the inequality of Camp and Meidell]. Statistica (Vereniging voor Statistiek), 3(5–6):201–208, October 1949. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1949.tb00403.x.

Homaker:1949:STF

[109] H. C. Homaker. Systematische en toevallige fouten bij het aflezen van de stand van een wijzer op een school. (Dutch) [Systematic and coincidental errors in reading the position of a pointer in a school]. Statistica (Vereniging voor Statistiek), 3(5–6):209–223, October 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00404.x.

Sittig:1949:BVG

[110] J. Sittig. De bovengrens van het gemiddeld doorgelaten uitvalpercentage bij enkelvoudige keuringsschema’s met c = 0. (Dutch) [The upper limit of the average passed failure rate for single inspection schemes with c = 0]. Statistica (Vereniging voor Statistiek), 3(5–6):224–226, October 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1949.tb00405.x.

Anonymous:1949:BDBc

[111] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 3(5–6):227–232, October 1949. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1949.
tb00406.x.

vandeBestuur:1950:VBV

[112] Het van de Bestuur. Voorwoord bij de vierde jaargang. (Dutch) [Foreword to the fourth volume]. Statistica (Vereniging voor Statistiek), 4(1–2):1– 2, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00407.x.

deLoor:1950:SSA

[113] B. de Loor. Statistiek in Suid-Afrika. (Dutch) [Statistics in South Africa]. Statistica (Vereniging voor Statistiek), 4(1–2):3–6, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00408.x.

vanderBurg:1950:ESP

[114] A. R. van der Burg. Enige statistische problemen in de textielindustrie. (Dutch) [Some statistical problems in the textile industry]. Statistica (Vereniging voor Statistiek), 4(1–2):7–19, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00409.x.

Hoogendoorn:1950:DVZ

[115] D. Hoogendoorn. Is diptherie een vermijdbare ziekte?. (Dutch) [Is diptheria a preventable disease?]. Statistica (Vereniging voor Statistiek), 4(1–2): 20–36, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00410.x.

Storch:1950:GVB

[116] J. M. Storch. Het gebruik van bewegende gemiddelden voor de analyse van een economische grafiek. (Dutch) [The use of moving averages for the analysis of an economic graph]. Statistica (Vereniging voor Statistiek), 4(1–2):37–53, February 1950. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1950.tb00411.x.

Hemelrijk:1950:GVT

[117] J. Hemelrijk and H. R. van der Voor. Het gebruik van een- en tweezijdige overschrijdingskansen voor het toetsen van hypothesen. (Dutch) [The use of one- and two-sided exceedance probabilities for testing hypotheses]. Statistica (Vereniging voor Statistiek), 4(1–2):54–66, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00412.x.

Hamaker:1950:BVP

[118] H. C. Hamaker. Het bezoek van Prof. J. M. Juran aan ons land. (Dutch) [The visit of Prof. J. M. Juran to our country]. Statistica (Vereniging voor Statistiek), 4(1–2):67–68, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1950.tb00413.x.

Juran:1950:RST

[119] J. M. Juran. The role of statistics as a tool of management. (Dutch) [The role of statistics as a tool of management]. Statistica (Vereniging voor Statistiek), 4(1–2):69–79, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1950.tb00414.x.

NC7:1950:VTN

[120] Normalisatie Commissie 73. Een voorstel tot normaliratie van de symbolen in de “statistica” en de “biometrica”. (Dutch) [A proposal for standardization of symbols in “statistics” and “biometrics”]. Statistica (Vereniging voor Statistiek), 4(1–2):80–85, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00415.x.

Anonymous:1950:BDBa

[121] Anonymous. Boekberpreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 4(1–2):86–88, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1950.
tb00416.x.

Hamaker:1950:OTL

[122] H. C. Hamaker. Oproep tot de lezers van dit tijdschrift. (Dutch) [Calling the readers of this magazine]. Statistica (Vereniging voor Statistiek), 4(1– 2):88, February 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00417.x.

deWolff:1950:OVF

[123] P. de Wolff. Het ontstaan van een frequentieverdeling. (Dutch) [The creation of a frequency distribution]. Statistica (Vereniging voor Statistiek), 4(3–4):89–92, June 1950. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1950.tb00418.x.

Sittig:1950:DVC

[124] J. Sittig. Een demonstratie van de centrale limietstelling. (Dutch) [A demonstration of the central limit theorem]. Statistica (Vereniging voor Statistiek), 4(3–4):93–100, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1950.tb00419.x.

deJong:1950:FMD

[125] A. J. de Jong. Frequentieverdelingen en mengproblemen. (Dutch) [Frequency distributions and mixing problems]. Statistica (Vereniging voor Statistiek), 4(3–4):101–105, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1950.tb00420.x.

Fortuin:1950:TFO

[126] G. J. Fortuin. Twee frequentieverdelingen op het terrein van het voedingsonderzoek. (Dutch) [Two frequency distributions in the field of nutritional research]. Statistica (Vereniging voor Statistiek), 4(3–4):106– 109, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00421.x.

deJong:1950:EAV

[127] J. R. de Jong. Enige aspecten van de spreiding van arbeidsprestaties. (Dutch) [Some aspects of the distribution of work performance]. Statistica (Vereniging voor Statistiek), 4(3–4):110–118, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1950.
tb00422.x.

deWolff:1950:NDA

[128] P. de Wolff. Nabeschouwing. (Dutch) [Afterthought]. Statistica (Vereniging voor Statistiek), 4(3–4):119–120, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1950.
tb00423.x.

Exalto:1950:GTV

[129] L. J. H. Exalto. Gebruik en toepassing van waarschijnlijkheidspapier. (Dutch) [Use and application of probability paper]. Statistica (Vereniging voor Statistiek), 4(3–4):121–128, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1950.tb00424.x.

vanIterson:1950:NVW

[130] G. van Iterson, Jr. Het nut van waarshijnlijkheidspapier voor de variatiestatistiek. (Dutch) [The usefulness of probability paper for variation statistics]. Statistica (Vereniging voor Statistiek), 4(3–4):129– 157, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00425.x.

Schumann:1950:PVA

[131] T. E. W. Schumann. Die probleem van die aanpassing van polinome aan ’n reeks ewewydige gegewens. (Dutch) [The problem of adapting polynomia to a series of constant desires]. Statistica (Vereniging voor Statistiek), 4(3–4):158–167, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1950.tb00426.x.

Goosens:1950:EON

[132] L. Goosens. Enige opmerkingen naar aanleiding van het artikel, “Een methode van contöle tijdens de fabricage” van F. G. Willemze. (Dutch) [Some comments regarding the article, “A method of control during manufacture” by F. G. Willemze]. Statistica (Vereniging voor Statistiek), 4(3– 4):168–176, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00427.x.

Anonymous:1950:BDBb

[133] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica (Vereniging voor Statistiek), 4(3–4):177–180, June 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1950.
tb00428.x.

vandeBestuur:1950:EMM

[134] Het van de Bestuur. Eendracht maakt macht!. (Dutch) [Unity is power!]. Statistica (Vereniging voor Statistiek), 4(5):181–182, October 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00429.x.

Bestuw:1950:MDN

[135] Het Bestuw. Mededeling. (Dutch) [Notice]. Statistica (Vereniging voor Statistiek), 4(5):183, October 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1950.tb00430.x.

Visman:1950:BVA

[136] J. Visman. Het belong van automatische monsterneming; mogelijkheden en toepas-singen. (Dutch) [The importance of automatic sampling; possibilities and applications]. Statistica (Vereniging voor Statistiek), 4(5): 184–194, October 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00431.x.

Somermeijer:1950:ZAV

[137] W. H. Somermeijer and H. C. Hamaker. Zitplaats-accomodatie van bioscopen in Nederland, en de factoren die haar bepalen; een econometrisch onderzoek. (Dutch) [Seating accommodation of cinemas in The Netherlands, and the factors that determine it; an econometric study]. Statistica (Vereniging voor Statistiek), 4(5):195–203, October 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb00432.x.

Quintelier:1950:NBV

[138] G. Quintelier and H. C. Hama. Het nomografisch bepalen van de parameters van een steekproefschema. (Dutch) [Determining the parameters of a sampling scheme nomographically]. Statistica (Vereniging voor Statistiek), 4(5):204–215, October 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1950.tb00433.x.

Hemelrijk:1950:RSV

[139] J. Hemelrijk. Rangcorrelatie en de schattingsproef van Varangot. (Dutch) [Rank correlation and the Varangot estimation test]. Statistica (Vereniging voor Statistiek), 4(5):216–225, October 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1950.
tb00434.x.

Anonymous:1950:BDBc

[140] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 4(5):226–232, October 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1950.
tb00435.x.

vanDantzig:1950:EHB

[141] D. van Dantzig. Enkele historische betrekkingen tussen mathematische en verzamelende statistiek. (Dutch) [Some historical relations between mathematical and aggregate statistics]. Statistica (Vereniging voor Statistiek), 4(6):233–247, December 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1950.tb01050.x.

Hemelrijk:1950:GMD

[142] J. Hemelrijk. De gereduceerde χ^{2}-methode. (Dutch) [The reduced χ^{2} method]. Statistica (Vereniging voor Statistiek), 4(6):249–258, December 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1950.tb01051.x.

Quintelier:1950:WBV

[143] G. Quintelier. De wiskundige behandeling van het onderscheidingsvermogen van een steekproef schema. (Dutch) [The mathematical treatment of the discriminative power of a sampling scheme]. Statistica (Vereniging voor Statistiek), 4(6):259–271, December 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1950.tb01052.x.

Anonymous:1950:BDBd

[144] Anonymous. Bookbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 4(6):273–282, December 1950. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1950.
tb01053.x.

BvVvS:1951:VBV

[145] Het Bestuur van de Vereniging voor Statistiek. Voorwoord bij de vijfde jaargang. (Dutch) [Foreword to the fifth volume]. Statistica (Vereniging voor Statistiek), 5(1–2):1–2, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1951.tb00567.x.

Tinbergen:1951:CGV

[146] J. Tinbergen. De centrale gedachte van de correlatierekening. (Dutch) [The central idea of the correlation account]. Statistica (Vereniging voor Statistiek), 5(1–2):3–6, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1951.tb00568.x.

Sittig:1951:GVC

[147] J. Sittig. Gebruik van correlatierekening voor het vaststellen van werktijden in het expeditiebedrijf. (Dutch) [Use of correlation calculation to determine working hours in the forwarding company]. Statistica (Vereniging voor Statistiek), 5(1–2):7–16, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1951.tb00569.x.

Stuvel:1951:LIA

[148] G. Stuvel. Lets over inkomenselasticiteiten, afgeleid uit huishoudrekeningen van gezinnen. (Dutch) [Something about income elasticities, derived from household accounts of families]. Statistica (Vereniging voor Statistiek), 5(1–2):17–22, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1951.tb00570.x.

Drion:1951:GVC

[149] E. F. Drion. Het gebruik van correlatierekening, om de homogeniteit van een reeks waarnemingen te controleren. (Dutch) [The use of correlation calculus, to check the homogeneity of a set of observations]. Statistica (Vereniging voor Statistiek), 5(1–2):23–25, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00571.x.

deGroot:1951:CTP

[150] A. D. de Groot. Correlatierekening in de toegepaste psychologie. (Dutch) [Correlation theory in applied psychology]. Statistica (Vereniging voor Statistiek), 5(1–2):26–32, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1951.tb00572.x.

Levert:1951:BTR

[151] C. Levert. Betrekkingen tussen de regressiecoëfficiënten van Galton en Frisch bij lineaire transformatie. (Dutch) [Relations between the regression coefficient of Galton and Frisch in a linear transformation]. Statistica (Vereniging voor Statistiek), 5(1–2):33–46, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00573.x.

vanAlphen:1951:GND

[152] G. W. H. M. van Alphen. Gesuperponeerde normale distributes en splitsing in elementaire curven. (Dutch) [Superimposed normal distributions and splitting into elementary curves]. Statistica (Vereniging voor Statistiek), 5(1–2):47–53, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1951.tb00574.x.

Anonymous:1951:BDBa

[153] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 5(1–2):54–74, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1951.
tb00575.x.

Anonymous:1951:MKM

[154] Anonymous. In memoriam: L. Kloek (1895–3 march 1951). Statistica (Vereniging voor Statistiek), 5(1–2):74, February 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1951.
tb00576.x.

Oomens:1951:SDI

[155] C. A. Oomens. De Statistische Dag 1951 (II). (Dutch) [The Statistical Day 1951 (II)]. Statistica (Vereniging voor Statistiek), 5(3–4):75– 80, June 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00577.x.

Enters:1951:GVT

[156] J. H. Enters. Het gebruik van toevalscijfers. (Dutch) [The use of random numbers]. Statistica (Vereniging voor Statistiek), 5(3–4):81– 96, June 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00578.x.

Theil:1951:VMR

[157] H. Theil. Verdelingsvrije methoden in de regressieanalyse van twee variabelen. (Dutch) [Distribution-free methods in the regression analysis of two variables]. Statistica (Vereniging voor Statistiek), 5(3–4): 97–118, June 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00579.x.

vIJzeren:1951:HVT

[158] J. v. IJzeren. Over de herleiding van B tot F. (Dutch) [About the reduction of B to F]. Statistica (Vereniging voor Statistiek), 5(3–4): 119–122, June 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00580.x.

Nass:1951:MVI

[159] C. A. G. Nass. Meting van immuniteit tegen Toxoplasmose met behulp van vrije curven. (Dutch) [Measurement of immunity against Toxoplasmosis using free curves]. Statistica (Vereniging voor Statistiek), 5(3–4): 123–144, June 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00581.x.

Varangot:1951:GBV

[160] V. Varangot. Grafische bepaling van de tweevoudige correlatiecoëfficiënt. (Dutch) [Graphical determination of the twofold correlation coefficient]. Statistica (Vereniging voor Statistiek), 5(3–4):145–147, June 1951. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1951.
tb00582.x.

vanDantzig:1951:NTD

[161] D. van Dantzig. De natuur als tegenspeler. (Dutch) [Nature as an opponent]. Statistica (Vereniging voor Statistiek), 5(5–6):149–159, October 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00583.x.

deWolff:1951:TVB

[162] P. de Wolff. Toepassing van de bezettingstheorie op een coderingsvraagstuk. (Dutch) [Application of occupancy theory to a coding problem]. Statistica (Vereniging voor Statistiek), 5(5–6):161–170, October 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00584.x.

Hemelrijk:1951:PPM

[163] J. Hemelrijk. Parametrische en parametervrije methoden en hun toepassingen. (Dutch) [Parametric and parameter-free methods and their applications]. Statistica (Vereniging voor Statistiek), 5(5–6):171–184, October 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00585.x.

vanderVaart:1951:TVW

[164] H. R. van der Vaart and D. van Dantzig. De toets van Wilcoxon. (Dutch) [Wilcoxon’s test]. Statistica (Vereniging voor Statistiek), 5(5– 6):185, October 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00586.x.

vanderVaart:1951:OVT

[165] H. R. van der Vaart. Over het onderscheidingsvermogen van de toets van Wilcoxon. (Dutch) [On the discriminative power of Wilcoxon’s test]. Statistica (Vereniging voor Statistiek), 5(5–6):186–190, October 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00587.x.

vanDantzig:1951:AEV

[166] D. van Dantzig and J. Hemelrijk. Asymptotische eigenschappen van Wilcoxon’s toets. (Dutch) [Asymptotic properties of Wilcoxon’s test]. Statistica (Vereniging voor Statistiek), 5(5–6):191–194, October 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1951.tb00588.x.

Anonymous:1951:BDBb

[167] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 5(5–6):195–215, October 1951. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1951.
tb00589.x.

Albregts:1952:PSD

[168] A. H. M. Albregts. Productiviteit en Statistick. (Dutch) [Productivity and Statistick]. Statistica (Vereniging voor Statistiek), 6(1):1– 4, March 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00971.x.

Bakker:1952:PDP

[169] O. Bakker. Productiviteitsmeting. (Dutch) [Productivity measurement]. Statistica (Vereniging voor Statistiek), 6(1):5–8, March 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1952.
tb00972.x.

Idenburg:1952:PWD

[170] Ph. J. Idenburg. De productiviteitsmeting in werking. (Dutch) [Productivity measurement in action]. Statistica (Vereniging voor Statistiek), 6 (1):9–12, March 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00973.x.

Enters:1952:KGU

[171] J. H. Enters. Kwaliteitsbeheersing: een geval uit de praktijk. (Dutch) [Quality control: a case from practice]. Statistica (Vereniging voor Statistiek), 6(1):13–18, March 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1952.tb00974.x.

vanEttinger:1952:SPD

[172] ir J. van Ettinger. Statistick en productiviteit. (Dutch) [Statistic and productivity]. Statistica (Vereniging voor Statistiek), 6(1):19–28, March 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00975.x.

Anonymous:1952:S

[173] Anonymous. Summary. Statistica (Vereniging voor Statistiek), 6(1): 29–30, March 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00976.x.

Anonymous:1952:SDS

[174] Anonymous. Secties. (Dutch) [Sections]. Statistica (Vereniging voor Statistiek), 6(1):31, March 1952. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1952.tb00977.x.

Hemelrijk:1952:EAM

[175] J. Hemelrijk. Enkele aspecten der mathematische statistiek. (Dutch) [Some aspects of mathematical statistics]. Statistica (Vereniging voor Statistiek), 6(2):31–47, June 1952. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1952.tb00978.x.

Kuiler:1952:EGS

[176] H. C. Kuiler. Existentiële gedachten over Statistiek. (Dutch) [Existential thoughts on Statistics]. Statistica (Vereniging voor Statistiek), 6(2): 49–58, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00979.x.

Oomens:1952:NRD

[177] C. A. Oomens. De Nationale Rekeningen. (Dutch) [The National Accounts]. Statistica (Vereniging voor Statistiek), 6(2):59–75, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (elec-

tronic).  URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00980.x.

tenHave:1952:SAD

[178] O. ten Have. Statistiek in Amerika. (Dutch) [Statistics in America]. Statistica (Vereniging voor Statistiek), 6(2):77–87, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1952.
tb00981.x.

Hamaker:1952:ISB

[179] H. C. Hamaker. De Internationale Statistische bijeenkomsten in India gehouden in December 1951. (Dutch) [The international statistical meetings in India held in December 1951]. Statistica (Vereniging voor Statistiek), 6(2):89–96, June 1952. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1952.tb00982.x.

Hamaker:1952:WVS

[180] H. C. Hamaker. Het werk van de UNO Subcommissie voor Steekproefonderzoek. (Dutch) [The work of the UNO Subcommittee on Sample Survey]. Statistica (Vereniging voor Statistiek), 6(2):97–100, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00983.x.

Visser:1952:EMT

[181] S. W. Visser and W. van der Bijl. Een experiment met toevalsreeksen. (Dutch) [An experiment with random sequences]. Statistica (Vereniging voor Statistiek), 6(2):101–105, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1952.tb00984.x.

vanderHeiden:1952:MVK

[182] J. A. van der Heiden. Methode van de kleinste kwadraten, toegepast op een stel simultane vergelijkingen die één of meer constanten gemeen hebben. (Dutch) [Method of least squares, applied to a set of simultaneous equations that have one or more constants in common]. Statistica (Vereniging voor Statistiek), 6(2):107–112, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1952.
tb00985.x.

vanIJzeren:1952:EBV

[183] J. van IJzeren. Elementair bewijs van de onafhankelijkheid van gemiddelde en spreiding bij steekproeven uit een normale verdeling. (Dutch) [Basic proof of the independence of mean and dispersion when sampling from a normal distribution]. Statistica (Vereniging voor Statistiek), 6(2): 113–119, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00986.x.

Burger:1952:MOM

[184] A. P. Burger. Matriks-omkering met Hollerith-masjiene. (Dutch) [Matrix inversion with Hollerith-machines]. Statistica (Vereniging voor Statistiek), 6(2):121–133, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1952.tb00987.x.

Anonymous:1952:BDB

[185] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 6(2):135–148, June 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1952.
tb00988.x.

Kuiper:1952:VAD

[186] N. H. Kuiper. Variantie-analyse. (Dutch) [Variance analysis]. Statistica (Vereniging voor Statistiek), 6(3):149–194, September 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1952.
tb00989.x.

vanDantzig:1952:VPD

[187] D. van Dantzig. Voorspelling en profetie. (Dutch) [Prediction and prophecy]. Statistica (Vereniging voor Statistiek), 6(4):195–203, December 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00990.x.

Baudet:1952:VVH

[188] H. Baudet. Voorspellen van historische ontwikkelingen. (Dutch) [Predicting historical developments]. Statistica (Vereniging voor Statistiek), 6(4):205–216, December 1952. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1952.tb00991.x.

deGroot:1952:VTP

[189] A. D. de Groot. Voorspellen in de toegepaste psychologie. (Dutch) [Prediction in applied psychology]. Statistica (Vereniging voor Statistiek), 6(4):217–226, December 1952. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1952.tb00992.x.

Kuiken:1952:VVE

[190] L. C. Kuiken. Voorspellen van de economische toestand. (Dutch) [Forecasting the economic situation]. Statistica (Vereniging voor Statistiek), 6(4):227–234, December 1952. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1952.tb00993.x.

Visser:1952:VVW

[191] S. W. Visser. Voorspellen van het weer. (Dutch) [Predicting the weather]. Statistica (Vereniging voor Statistiek), 6(4):235–244, December 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00994.x.

Mazure:1952:VVW

[192] ir J. P. Mazure. Voorspellen van waterstanden. (Dutch) [Predicting water levels]. Statistica (Vereniging voor Statistiek), 6(4):245–257, December 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00995.x.

Ignatius:1952:VVL

[193] ir J. G. W. Ignatius. Voorspellen van landbouw-opbrengsten. (Dutch) [Predicting agricultural yields]. Statistica (Vereniging voor Statistiek), 6(4):259–264, December 1952. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1952.tb00996.x.

deJong:1952:VVP

[194] A. J. de Jong. Voorspellen van productie-resultaten. (Dutch) [Predicting production results]. Statistica (Vereniging voor Statistiek), 6(4):265– 272, December 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1952.tb00997.x.

Idenburg:1952:SDS

[195] Ph. J. Idenburg. Samenvatting. (Dutch) [Summary]. Statistica (Vereniging voor Statistiek), 6(4):273–279, December 1952. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1952.
tb00998.x.

Bestuur:1953:BBV

[196] Het Bestuur. Bij het begin van de zevende jaargang. (Dutch) [At the beginning of the seventh year]. Statistica (Vereniging voor Statistiek), 7 (1):1–2, March 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1953.tb01054.x.

Storch:1953:SAB

[197] J. M. Storch. Sequente analyse bij keuring van goederen. (Dutch) [Sequence analysis during inspection of goods]. Statistica (Vereniging voor Statistiek), 7(1):3–14, March 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1953.tb01055.x.

vdHeiden:1953:DRV

[198] J. A. v. d. Heiden. Dupliceerbaarheid en reproduceerbaarheid van waarnemingen. (Dutch) [Duplicability and reproducibility of observations]. Statistica (Vereniging voor Statistiek), 7(1):15–22, March 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (elec-

tronic).  URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1953.tb01056.x.

Spitz:1953:MPD

[199] J. C. Spitz. Het matchen in de psychologie. (Dutch) [Matching in psychology]. Statistica (Vereniging voor Statistiek), 7(1):23–40, March 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1953.tb01057.x.

Kaarsemaker:1953:TUR

[200] L. Kaarsemaker and A. van Wijngaarden. Tables for use in rank correlation. Statistica (Vereniging voor Statistiek), 7(1):41–54, March 1953. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1953.
tb01058.x.

Anonymous:1953:BDBa

[201] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 7(1):55–69, March 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1953.
tb01059.x.

Westenberg:1953:OVO

[202] J. Westenberg. Overzicht van de ontwikkeling der visserijtheorie. (Dutch) [Overview of the development of fishing theory]. Statistica (Vereniging voor Statistiek), 7(2):71–95, June 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1953.tb00815.x.

Heyna:1953:OVG

[203] B. Heyna. Omzetting van een in gehele graden Fahrenheit gegeven frequentieverdeling in een, welke een klassebreedte van een graad Celsius heeft. (Dutch) [Conversion of a frequency distribution given in integer degrees Fahrenheit into one which has a class width of one degree Celsius]. Statistica (Vereniging voor Statistiek), 7(2):97–104, June 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1953.tb00816.x.

Anonymous:1953:ESA

[204] Anonymous. Het Examen Statistisch Analyst 1952/1953. (Dutch) [The Statistical Analyst Exam 1952/1953]. Statistica (Vereniging voor Statistiek), 7(2):105–106, June 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1953.tb00817.x.

Gedeelte:1953:ESA

[205] Algemeen Gedeelte. Examen statistisch analyst. (Dutch) [Statistical analyst exam]. Statistica (Vereniging voor Statistiek), 7(2):107– 125, June 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1953.tb00818.x.

Vos:1953:RVB

[206] J. W. E. Vos. Richtlijnen voor de berekening van productie-indices. (Dutch) [Guidelines for the calculation of production indices]. Statistica (Vereniging voor Statistiek), 7(3):127–140, September 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1953.
tb00819.x.

vanEeden:1953:MVV

[207] Constance van Eeden. Methoden voor het vergelijken, toetsen en schatten van onbekende kansen. (Dutch) [Methods for comparing, testing and estimating unknown probabilities]. Statistica (Vereniging voor Statistiek), 7(3):141–162, September 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1953.tb00820.x.

Benard:1953:UVW

[208] A. Benard and E. C. Bos-Levenbach. Het uitzetten van waarnemingen op waarschijnlijkheids-papier. (Dutch) [Plotting observations on probability paper]. Statistica (Vereniging voor Statistiek), 7(3):163–173, September 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1953.tb00821.x.

Anonymous:1953:BDBb

[209] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 7(3):175–181, September 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1953.
tb00822.x.

Idenburg:1953:MOV

[210] Ph. J. Idenburg. Macht en onmacht van de statistiek. (Dutch) [Power and powerlessness of statistics]. Statistica (Vereniging voor Statistiek), 7(4):183–191, December 1953. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1953.tb00999.x.

Hartog:1953:WRG

[211] J. A. Hartog. Waarneming en rationeel gedrag. (Dutch) [Perception and rational behavior]. Statistica (Vereniging voor Statistiek), 7(4):193– 198, December 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1953.tb01000.x.

vanDantzig:1953:VVS

[212] D. van Dantzig. De verantwoordelijkheden van de statisticus. (Dutch) [The responsibilities of the statistician]. Statistica (Vereniging voor Statistiek), 7(4):199–208, December 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1953.tb01001.x.

Somermeyer:1953:KSW

[213] W. H. Somermeyer. Kwantificering in de sociale wetenschappen. (Dutch) [Quantification in the social sciences]. Statistica (Vereniging voor Statistiek), 7(4):209–222, December 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1953.tb01002.x.

Rumke:1953:KMO

[214] Chr. L. Rümke. Kwantificering in medisch onderzoek. (Dutch) [Quantification in medical research]. Statistica (Vereniging voor Statistiek), 7(4):223–232, December 1953. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1953.tb01003.x.

deVries:1953:KTE

[215] M. de Vries. Kwantificering in taal encrypto-analyse. (Dutch) [Quantification in language encryptoanalysis]. Statistica (Vereniging voor Statistiek), 7(4):233–256, December 1953. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1953.tb01004.x.

Anonymous:1953:MHG

[216] Anonymous. In memoriam: Hendrik Gerrit Bolk (1896–5 June 1954). Statistica (Vereniging voor Statistiek), 7(4):257, December 1953. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1953.
tb01005.x.

Kuiper:1954:OAV

[217] N. H. Kuiper. Een opmerking over het aanpassen van functies aan een groot aantal waarnemingsuitkomsten. (Dutch) [A note on fitting functions to a large number of observational outcomes]. Statistica (Vereniging voor Statistiek), 8(1):1–6, March 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1954.tb00436.x.

Klerk-Grobben:1954:TVG

[218] Gerda Klerk-Grobben and H. J. Prins. Toets voor de gelijkheid van twee kleine kansen met behulp van even grote steekproeven en net onderscheidingsvermogen van deze toets. (Dutch) [Test for the equality of two small probabilities using equally sized samples and the net discriminatory power of this test]. Statistica (Vereniging voor Statistiek), 8(1): 7–20, March 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1954.tb00437.x.

vanIJzeren:1954:TZV

[219] J. van IJzeren. De theoretische zijde van de methode der kleinste kwadraten. (Dutch) [The theoretical side of the method of least squares]. Statistica (Vereniging voor Statistiek), 8(1):21–45, March 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1954.tb00438.x.

Landman:1954:SOV

[220] H. J. Landman. Statistische oplossing van een passingsprobleem. (Dutch) [Statistical solution of a matching problem]. Statistica (Vereniging voor Statistiek), 8(1):47–52, March 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1954.tb00439.x.

Anonymous:1954:BDBa

[221] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 8(1):53–57, March 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1954.
tb00440.x.

Idenburg:1954:MPD

[222] Ph. J. Idenburg. In memoriam prof, dr O. Bakker. Statistica (Vereniging voor Statistiek), 8(2):59–60, June 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1954.tb00441.x.

Abdel-Aty:1954:OVD

[223] S. H. Abdel-Aty. Ordered variables in discontinuous distributions. Statistica (Vereniging voor Statistiek), 8(2):61–82, June 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1954.
tb00442.x.

Lisman:1954:SAV

[224] J. H. C. Lisman. Statistische analyse van muziek. (Dutch) [Statistical analysis of music]. Statistica (Vereniging voor Statistiek), 8(2):83– 91, June 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1954.tb00443.x.

Anonymous:1954:ESA

[225] Anonymous. Het Examen Statistisch Analyst 1953/1954. (Dutch) [The Statistical Analyst Exam 1953/1954]. Statistica (Vereniging voor Statistiek), 8(2):93–112, June 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1954.tb00444.x.

Anonymous:1954:BDBb

[226] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 8(2):113–121, June 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1954.
tb00445.x.

Ferro:1954:GDG

[227] H. Ferro. Grafiekenpapier. (Dutch) [Graph paper]. Statistica (Vereniging voor Statistiek), 8(3–4):123–154, September 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1954.
tb00446.x.

Erlee:1954:VVS

[228] Th. J. D. Erlee and A. A. Koene. De vervaardiging van een systeem functie-ponskaarten. (Dutch) [The production of a system of function punch cards]. Statistica (Vereniging voor Statistiek), 8(3–4):155–167, September 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1954.tb00447.x.

Wise:1954:ERS

[229] M. E. Wise. The effect of rounding off in samples. Statistica (Vereniging voor Statistiek), 8(3–4):169–173, September 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1954.
tb00448.x.

Wise:1954:DRR

[230] M. E. Wise and H. C. Hamaker. A discussion of the rules for rounding-off as given in the Dutch draft standard V 1047. Statistica (Vereniging voor Statistiek), 8(3–4):175–178, September 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1954.tb00449.x.

Sieben:1954:SBP

[231] J. W. Sieben. Selectie bij de plantenveredeling op grond van proefveldresultaten. (Dutch) [Selection in plant breeding based on test field results]. Statistica (Vereniging voor Statistiek), 8(3–4):179–189, September 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1954.tb00450.x.

Anonymous:1954:BDBc

[232] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica (Vereniging voor Statistiek), 8(3–4):191–196, September 1954. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1954.
tb00451.x.

Anonymous:1955:VBN

[233] Anonymous. Voorwoord bij de negende jaargang. (Dutch) [Foreword to the ninth volume]. Statistica Neerlandica, 9(1–2):1–2, March 1955. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1955.
tb00735.x.

deJong:1955:OVV

[234] A. J. de Jong. Openingswoord van de Voorzitter. (Dutch) [Opening remarks by the Chairman]. Statistica Neerlandica, 9(1–2):3–5, March 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00736.x.

Hamaker:1955:NEE

[235] H. C. Hamaker. Naar efficiënte experimenten. (Dutch) [Towards efficient experiments]. Statistica Neerlandica, 9(1–2):7–25, March 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1955.
tb00737.x.

Wijtenburg:1955:KDQ

[236] J. M. Wijtenburg. Kwaliteitsnormen. (Dutch) [Quality standards]. Statistica Neerlandica, 9(1–2):27–35, March 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1955.tb00738.x.

vanAf:1955:SBV

[237] J. W. van Af. Statistische beschouwingen in verband met textiel. (Dutch) [Statistical considerations related to textiles]. Statistica Neerlandica, 9(1– 2):37–42, March 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00739.x.

deFremery:1955:TDD

[238] J. D. N. de Fremery. Toespraak door de directeur van de Stichting “Kwaliteitsdienst voor de Industrie”. (Dutch) [Speech by the director of the Foundation “Quality Service for the Industry”]. Statistica Neerlandica, 9(1–2):43–46, March 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1955.tb00740.x.

Sittig:1955:SBV

[239] J. Sittig. De statistische basis van de planning in een machinefabriek. (Dutch) [The statistical basis of planning in a machine factory]. Statistica Neerlandica, 9(1–2):47–69, March 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1955.tb00741.x.

Sparnaay:1955:GVM

[240] K. F. Sparnaay. Gebruik van multi-momentopnamen. (Dutch) [Using multi-snapshots]. Statistica Neerlandica, 9(1–2):71–77, March 1955. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1955.
tb00742.x.

Meijs:1955:OVB

[241] L. Th. J. Meijs. Het oordeel van een bedrijfsleider over de toepassing van de statistiek in het bedrijf. (Dutch) [The opinion of a manager on the application of statistics in the company]. Statistica Neerlandica, 9(1–2): 79–83, March 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00743.x.

Idenburg:1955:MPD

[242] Ph. J. Idenburg. In memoriam Prof. Dr H. W. Methorst (1868– 1955). Statistica Neerlandica, 9(3):85–88, September 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1955.
tb00287.x.

Muller:1955:VVL

[243] H. L. Muller. Het voorspellen van levertijden uit gegevens over de voort-gang van een project. (Dutch) [Predicting delivery times from data about the progress of a project]. Statistica Neerlandica, 9(3):89– 99, September 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00288.x.

deWolff:1955:ITA

[244] P. de Wolff and J. Meerdink. De intelligence te Amsterdam in verband met demografische en sociologische kenmerken van de bevolking. (Dutch) [The intelligence in Amsterdam in connection with demographic and sociological characteristics of the population]. Statistica Neerlandica, 9 (3):101–123, September 1955. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1955.tb00289.x.

Kerrich:1955:CIA

[245] J. E. Kerrich. Confidence intervals associated with a straight line fitted by least squares. Statistica Neerlandica, 9(3):125–129, September 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00290.x.

vanEeden:1955:TTV

[246] Constance van Eeden. Een toets tegen verloop voor een aantal kansen. (Dutch) [A test against attrition for a number of opportunities]. Statistica Neerlandica, 9(3):131–159, September 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1955.tb00291.x.

Yntema:1955:SAM

[247] L. Yntema. Statistische aspecten der muziek. (Dutch) [Statistical aspects of music]. Statistica Neerlandica, 9(3):161–171, September 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00292.x.

Anonymous:1955:BDBa

[248] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 9(3):173–187, September 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1955.tb00293.x.

Anonymous:1955:GED

[249] Anonymous. Gemeente Enschede. (Dutch) [Municipality of Enschede]. Statistica Neerlandica, 9(3):188, September 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1955.
tb00294.x.

Davies:1955:DAI

[250] O. L. Davies, F. J. van Dun En, and H. C. Hamaker. “Design and analysis of industrial experiments”. Statistica Neerlandica, 9(4):189– 207, December 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00295.x.

Hamaker:1955:PID

[251] H. C. Hamaker. Proefopzetten in de Industrie. (Dutch) [Experimental designs in the Industry]. Statistica Neerlandica, 9(4):209–232, December 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00296.x.

vanDantzig:1955:TJW

[252] D. van Dantzig. Tien jaar wiskundige statistiek. (Dutch) [Ten years of mathematical statistics]. Statistica Neerlandica, 9(4):233–242, December 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1955.tb00297.x.

Idenburg:1955:TJO

[253] Ph. J. Idenburg. Tien jaar ontwikkeling van de maatschappelijke statistiek. (Dutch) [Ten years of development of social statistics]. Statistica Neerlandica, 9(4):243–259, December 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1955.
tb00298.x.

vanEttinger:1955:TJO

[254] J. van Ettinger. Tien jaar ontwikkeling van de industriële statistiek. (Dutch) [Ten years of development of industrial statistics]. Statistica Neerlandica, 9(4):261–269, December 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1955.tb00299.x.

Anonymous:1955:BDBb

[255] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 9(4):271–280, December 1955. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1955.tb00300.x.

Sittig:1956:ORM

[256] J. Sittig. Operational Research — managers beslissen met behulp van de wetenschap. (Dutch) [Operational Research — managers make decisions using science]. Statistica Neerlandica, 10(1):1– 17, March 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1956.tb00001.x.

Tinbergen:1956:KKT

[257] J. Tinbergen. Keus en kans in de theorie der economische politiek. (Dutch) [Choice and chance in the theory of economic policy]. Statistica Neerlandica, 10(1):19–25, March 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1956.tb00002.x.

Seligmann:1956:BGK

[258] P. S. Seligmann. Belletjes in glasbuizen — de keuze van de fabricagemethode van een massaproduct. (Dutch) [Bubbles in glass tubes — the choice of manufacturing method of a mass-produced product]. Statistica Neerlandica, 10(1):27–33, March 1956. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1956.tb00003.x.

deWolff:1956:SBA

[259] P. de Wolff. Steekproeven bij administratieve contrôle. (Dutch) [Random samples during administrative checks]. Statistica Neerlandica, 10 (1):35–44, March 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1956.tb00004.x.

Kriens:1956:STO

[260] J. Kriens. Schema’s ter oplossing van lineaire programmeringsproblemen. (Dutch) [Diagrams for solving linear programming problems]. Statistica Neerlandica, 10(1):45–53, March 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1956.tb00005.x.

Wijvekate:1956:BVO

[261] M. L. Wijvekate. Bepaling van het optimale controle-interval. (Dutch) [Determination of the optimal monitoring interval]. Statistica Neerlandica, 10(1):55–63, March 1956. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1956.tb00006.x.

Zoutendijk:1956:MTV

[262] G. Zoutendijk. Militaire toepassingen van Operations Research. (Dutch) [Military applications of Operations Research]. Statistica Neerlandica, 10 (1):65–73, March 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1956.tb00007.x.

Anonymous:1956:BDBa

[263] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 10(1):75–86, March 1956. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1956.tb00008.x.

Meinesz:1956:PVR

[264] M. Meinesz. Het probleem van de ruïering der spelers. (Dutch) [The problem of the ruïering of the players]. Statistica Neerlandica, 10(2): 87–97, June 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1956.tb00009.x.

Klerk-Grobben:1956:BVS

[265] Gerda Klerk-Grobben and H. D. Sandberg. Betrouwbaarheidsintervallen voor de spreiding van een normale verdeling bepaald uit het gemiddelde van een aantal steekproefbreedtes. (Dutch) [Confidence intervals for the spread of a normal distribution determined from the average of a number of sample widths]. Statistica Neerlandica, 10(2): 99–115, June 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1956.tb00010.x.

Doornbos:1956:SSS

[266] R. Doornbos. Significance of the smallest of a set of estimated normal variances. Statistica Neerlandica, 10(2):117–126, June 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1956.
tb00011.x.

Anonymous:1956:ESA

[267] Anonymous. Het examen Statistisch Analyst 1955. (Dutch) [The Statistical Analyst exam 1955]. Statistica Neerlandica, 10(2):127–139, June 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1956.tb00012.x.

Anonymous:1956:BDBb

[268] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 10(2):141–155, June 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1956.tb00013.x.

vanEeden:1956:VTV

[269] Constance van Eeden. Verdelingsvrije toetsen voor twee steekproeven en de methode der 2 × 2-tabel. (Dutch) [Distribution-free tests for two samples and the method of the 2 × 2 table]. Statistica Neerlandica, 10 (3–4):157–162, September 1956. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1956.tb00590.x.

Thomissen:1956:FIA

[270] F. X. Thomissen. The frequency of industrial accidents. Statistica Neerlandica, 10(3–4):163–176, September 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1956.tb00591.x.

vanElteren:1956:ERD

[271] Ph. van Elteren and W. F. van Peype. Enige rangcorrelatieschema’s. (Dutch) [Some rank correlation schemes]. Statistica Neerlandica, 10(3– 4):177–195, September 1956. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1956.tb00592.x.

Kesten:1956:EOB

[272] H. Kesten and J. Th. Runnenburg. Enige opmerkingen over de berekening van de verwachting en de spreiding van het aantal inconsistenties in een bepaald rangcorrelatieschema. (Dutch) [Some comments on the calculation of the expectation and the distribution of the number of inconsistencies in a given rank correlation scheme]. Statistica Neerlandica, 10 (3–4):197–204, September 1956. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1956.tb00593.x.

Smid:1956:DTS

[273] L J. Smid. On the distribution of the test statistics of Kendall and Wilcoxon’s two sample test when ties are present. Statistica Neerlandica, 10(3–4):205–214, September 1956. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1956.tb00594.x.

Anonymous:1956:BDBc

[274] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 10(3–4):215–228, September 1956. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1956.tb00595.x.

vanDantzig:1957:SP

[275] D. van Dantzig. Statistical priesthood. Statistica Neerlandica, 11(1): 1–16, March 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00014.x.

deWolff:1957:IVK

[276] P. de Wolff. De invloed van de keuze der gewichten op de hoogte van een gewogen indexcijfer. (Dutch) [The influence of the choice of weights on the height of a weighted index figure]. Statistica Neerlandica, 11(1): 17–22, March 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00015.x.

vanEijk:1957:OSV

[277] C. J. van Eijk and H. Moerman. Opmerkingen over de standaardafwijking van de constante term bij regressie-analyse. (Dutch) [Notes on the standard deviation of the constant term in regression analysis]. Statistica Neerlandica, 11(1):23–30, March 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1957.tb00016.x.

Anonymous:1957:STB

[278] Anonymous. Statistische termen en begrippen: (Normontwerp 3117, uitgegeven door de Hoofdcommissie voor de Normalisatie in Nederland). (Dutch) [Statistical terms and concepts: (Standard Draft 3117, published by the Main Committee for Standardization in the Netherlands)]. Statistica Neerlandica, 11(1):31–39, March 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1957.tb00017.x.

Anonymous:1957:BDBa

[279] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 11(1):41–53, March 1957. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1957.tb00018.x.

Kohnstamm:1957:SMI

[280] G. A. Kohnstamm. Statistiek en Marktanalyse: Inleiding tot de Statistische Dag 1957. (Dutch) [Statistics and Market Analysis: Introduction to Statistical Day 1957]. Statistica Neerlandica, 11(2):55– 62, June 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00019.x.

Abrams:1957:PMR

[281] M. A. Abrams. Problems of media research: The readership survey of the Institute of Practitioners in Advertising. Statistica Neerlandica, 11 (2):63–76, June 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00020.x.

deVries:1957:EWE

[282] M. de Vries. Experiment en waarneming in de economische statistiek. (Dutch) [Experiment and observation in economic statistics]. Statistica Neerlandica, 11(2):77–87, June 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1957.tb00021.x.

Theil:1957:RRO

[283] H. Theil. Recente resultaten op het terrein van economische voorspellingen op korte termijn. (Dutch) [Recent results in the field of short-term economic forecasting]. Statistica Neerlandica, 11(2):89– 96, June 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00022.x.

Bakker:1957:OMC

[284] A. Bakker. Ondernemer — Marktonderzoek — Consument. (Dutch) [Entrepreneur — Market Research — Consumer]. Statistica Neerlandica, 11 (2):97–108, June 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00023.x.

Anonymous:1957:BDBb

[285] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 11(2):109–120, June 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1957.tb00024.x.

Rijkoort:1957:OTS

[286] P. J. Rijkoort and J. Hemelrijk. The occurrence of “twin” storms from the North West on the Dutch coast. Statistica Neerlandica, 11(3):121– 130, September 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00025.x.

Storch:1957:OPO

[287] J. M. Storch. Open plaatsen in orthogonale proefschema’s. (Dutch) [Open places in orthogonal trial schemes]. Statistica Neerlandica, 11 (3):131–139, September 1957. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1957.tb00026.x.

Corsten:1957:AMT

[288] L. C. A. Corsten. Een algemene methode tot leemtevulling. (Dutch) [A general method for gap filling]. Statistica Neerlandica, 11(3):141– 151, September 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00027.x.

Somermeijer:1957:SVC

[289] W. H. Somermeijer. Substituut-variabelen in correlatieberekingen. (Dutch) [Substitute variables in correlation calculations]. Statistica Neerlandica, 11(3):153–160, September 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1957.tb00028.x.

vanIJzeren:1957:PBV

[290] J. van IJzeren. Practische berekening van de standaardafwijking van de constante term bij regressie-analyse. (Dutch) [Practical calculation of the standard deviation of the constant term in regression analysis]. Statistica Neerlandica, 11(3):161–167, September 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1957.tb00029.x.

deWolff:1957:SEG

[291] P. de Wolff. Selectuur: Een eenvoudige grafische methode voor de aanpassing van een rechtlijnige trend volgens de methode van de kleinste kwadraten aan een historische reeks. (Dutch) [Selection: a simple graphical method for fitting a linear least-squares trend to a historical series]. Statistica Neerlandica, 11(3):169–170, September 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00030.x.

Anonymous:1957:BDBc

[292] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 11(3):171–183, September 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1957.tb00031.x.

vanDantzig:1957:SPI

[293] D. van Dantzig. Statistical Priesthood II: Sir Ronald on Scientific Inference. Statistica Neerlandica, 11(4):185–200, December 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1957.
tb00032.x.

deLeve:1957:ESA

[294] G. de Leve. Enige statistische aspecten van de factoranalyse. (Dutch) [Some statistical aspects of the factor analysis]. Statistica Neerlandica, 11(4):201–209, December 1957. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1957.tb00033.x.

vanEijk:1957:ETV

[295] C. J. van Eijk. Economische toepassingen van factoranalyse. (Dutch) [Economic applications of factor analysis]. Statistica Neerlandica, 11 (4):211–213, December 1957. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1957.tb00034.x.

Anonymous:1957:ESA

[296] Anonymous. Het examen Statistisch Analyst (Algemeen Gedeelte) 1956. (Dutch) [The Statistical Analyst exam (General Section) 1956]. Statistica Neerlandica, 11(4):215–230, December 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1957.tb00035.x.

Anonymous:1957:RSV

[297] Anonymous. Receptuur: De symmetrietoets van Wilcoxon. (Dutch) [Recipe: The symmetry test of Wilcoxon]. Statistica Neerlandica, 11 (4):231–235, December 1957. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1957.tb00036.x.

R:1957:STM

[298] Chr. L. R. Selectuur: Toxiciteitsbepaling met behulp van een “telefoonboek”^{1}. (Dutch) [Selection: Toxicity determination using a “telephone book”^{1}]. Statistica Neerlandica, 11(4):237, December 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1957.tb00037.x.

Anonymous:1957:BDBd

[299] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 11(4):239–247, December 1957. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1957.tb00038.x.

Rumke:1958:TVM

[300] Chr. L. Rümke. De taak van de Medische Statistiek. (Dutch) [The task of Medical Statistics]. Statistica Neerlandica, 12(1):1–16, March 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00823.x.

Jochems:1958:VDF

[301] D. B. Jochems. Voetbalprognoses. (Dutch) [Football forecasts]. Statistica Neerlandica, 12(1):17–31, March 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1958.tb00824.x.

Sandee:1958:OIV

[302] J. Sandee and R. Ruiter. Opleiding en intelligentie volgens de legerkeuringen. (Dutch) [Education and intelligence according to army inspections]. Statistica Neerlandica, 12(1):33–43, March 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1958.
tb00825.x.

Venekamp:1958:BVS

[303] P. E. Venekamp. Bepaling van stuwende en verzorgende bedrijfstakken door middel van statistische methoden. (Dutch) [Determination of driving and nurturing industries by statistical methods]. Statistica Neerlandica, 12(1):45–53, March 1958. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1958.tb00826.x.

Tas:1958:RDR

[304] L. Tas. Receptuur. (Dutch) [Recipe]. Statistica Neerlandica, 12(1): 55, March 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00827.x.

Anonymous:1958:BDBa

[305] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 12(1):57–62, March 1958. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1958.tb00828.x.

Noether:1958:ESD

[306] Gottfried E. Noether. The efficiency of some distribution-free tests. Statistica Neerlandica, 12(2):63–73, June 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1958.tb00829.x.

Rumke:1958:BSB

[307] Chr. L. Rumke and A. R. Bloemena. De bloedproef statistisch bezien. (Dutch) [The blood test is viewed statistically]. Statistica Neerlandica, 12 (2):75–82, June 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00830.x.

Anonymous:1958:ESA

[308] Anonymous. Het Examen Statistisch Analyst 1957 (Algemeen gedeelte). (Dutch) [The Statistical Analyst Exam 1957 (General part)]. Statistica Neerlandica, 12(2):83–92, June 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1958.tb00831.x.

Sliggers:1958:PEN

[309] B. C. Sliggers. Proeve ener nieuwe personeelsterkte-grafiek. (Dutch) [Taste a new staff strength chart]. Statistica Neerlandica, 12(2):93– 95, June 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00832.x.

Muilwijk:1958:RDR

[310] J. Muilwijk. Receptuur. (Dutch) [Recipe]. Statistica Neerlandica, 12(2): 97–98, June 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00833.x.

Anonymous:1958:BDBb

[311] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 12(2):99–109, June 1958. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1958.tb00834.x.

Hemelrijk:1958:SPB

[312] J. Hemelrijk. Statistische proefopzetten: bewijs en detectie. (Dutch) [Statistical experimental designs: evidence and detection]. Statistica Neerlandica, 12(3):111–118, September 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1958.tb00835.x.

Hamaker:1958:SED

[313] H. C. Hamaker. Statistiek en experiment. (Dutch) [Statistics and experiment]. Statistica Neerlandica, 12(3):119–130, September 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1958.
tb00836.x.

vanderBurg:1958:SED

[314] A. R. van der Burg. Statistiek of experiment. (Dutch) [Statistics or experiment]. Statistica Neerlandica, 12(3):131–134, September 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00837.x.

Bloemena:1958:RPD

[315] A. R. Bloemena. Roteerbare proefopzetten. (Dutch) [Rotatable experimental designs]. Statistica Neerlandica, 12(3):135–142, September 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00838.x.

Leppink:1958:EBV

[316] G. J. Leppink. Het experimenteel bepalen van optimale condities. (Dutch) [Experimentally determining optimal conditions]. Statistica Neerlandica, 12(3):143–148, September 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1958.tb00839.x.

Rumke:1958:TVS

[317] Chr. L. Rümke. Toepassing van statistiek bij experimenter! in een farmacologisch laboratorium. (Dutch) [Application of statistics to experimenter! in a pharmacological laboratory]. Statistica Neerlandica, 12 (3):149–158, September 1958. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1958.tb00840.x.

vanStrik:1958:SER

[318] R. van Strik. Statistiek en experiment rondom het poliomyelitis vaccin. (Dutch) [Statistics and experiment surrounding the poliomyelitis vaccine]. Statistica Neerlandica, 12(3):159–169, September 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00841.x.

and:1958:SDS

[319] and. Selectuur. (Dutch) [Selection]. Statistica Neerlandica, 12(3):170– 172, September 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00842.x.

Anonymous:1958:BED

[320] Anonymous. Boekbespreking: Experimental Designs, William G. Cochranen Gertrude M. Cox. (Dutch) [Book review]. Statistica Neerlandica, 12(3):173, September 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1958.tb00843.x.

Hemelrijk:1958:CVE

[321] J. Hemelrijk. Constance van Eeden cum laude gepromoveerd. (Dutch) [Constance van Eeden obtained her PhD cum laude]. Statistica Neerlandica, 12(4):175–176, December 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1958.tb00844.x.

vanderBurg:1958:ORP

[322] A. R. van der Burg. Optimale revisietijd: Een praktijkgeval. (Dutch) [Optimal revision time: a practical case]. Statistica Neerlandica, 12 (4):177–190, December 1958. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1958.tb00845.x.

Hemelrijk:1958:GVS

[323] J. Hemelrijk and C. G. G. van Herk. Een geval van statistische detectie. (Dutch) [A case of statistical detection]. Statistica Neerlandica, 12(4):191–200, December 1958. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1958.tb00846.x.

Hamaker:1958:ROP

[324] H. C. Hamaker. De recente ontwikkeling in proefopzetten met kwantitatieve facto ren. (Dutch) [The recent development in experimental designs with quantitative factors]. Statistica Neerlandica, 12(4):201–212, December 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00847.x.

deJongh:1958:DDC

[325] D. K. de Jongh. Design for decision in het clinische experiment. (Dutch) [Design for decision in the clinical experiment]. Statistica Neerlandica, 12(4):213–229, December 1958. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1958.tb00848.x.

Vos:1958:STV

[326] J. W. E. Vos. Een steekproef ter vaststelling van de verdeling van de hoenderstapel. (Dutch) [A sample to determine the distribution of the fowl herd]. Statistica Neerlandica, 12(4):231–242, December 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1958.tb00849.x.

Lange:1958:PVE

[327] H. Lange and E. Kaptein. Prijsverwachtingen volgens enquêtes onder consumenten. (Dutch) [Price expectations according to consumer surveys]. Statistica Neerlandica, 12(4):243–253, December 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1958.
tb00850.x.

vandePanne:1958:SDS

[328] C. van de Panne and H. Lange. Scheefheidstoetsen. (Dutch) [Skewness tests]. Statistica Neerlandica, 12(4):255–264, December 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1958.
tb00851.x.

Zaalberg:1958:ATW

[329] J. Zaalberg. Auxiliary tables for Wilcoxon’s two sample test. Statistica Neerlandica, 12(4):265–273, December 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1958.tb00852.x.

Anonymous:1958:RDR

[330] Anonymous. Receptuur. (Dutch) [Recipe]. Statistica Neerlandica, 12 (4):275–280, December 1958. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1958.tb00853.x.

Anonymous:1958:BDBc

[331] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 12(4):281–296, December 1958. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1958.tb00854.x.

Redactie:1959:WVB

[332] De Redactie. Woord vooraf bij de 13-de jaargang. (Dutch) [Preface to the 13th year]. Statistica Neerlandica, 13(1):1–2, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb01006.x.

Levert:1959:NVB

[333] C. Levert. Een nomogram voor betrouwbaarheidsgrenzen van quantielen en overschrijdingskansen bij normaal verdeelde grootheden. (Dutch) [A nomogram for confidence limits of quantiles and exceedance probabilities for normally distributed quantities]. Statistica Neerlandica, 13(1): 3–14, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01007.x.

Anonymous:1959:ESA

[334] Anonymous. Het examen Statistisch Analyst 1958 (Algemeen gedeelte). (Dutch) [The Statistical Analyst exam 1958 (General part)]. Statistica Neerlandica, 13(1):15–30, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1959.tb01008.x.

Wallin:1959:RDR

[335] J. F. E. Wallin. Receptuur. (Dutch) [Recipe]. Statistica Neerlandica, 13 (1):31–33, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01009.x.

Anonymous:1959:SRM

[336] Anonymous. Statistical research monographs. Statistica Neerlandica, 13 (1):34–35, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01010.x.

Hamaker:1959:ATP

[337] H. C. Hamaker. Attributenkeuring in theorie en praktijk. (Dutch) [Attribute inspection in theory and practice]. Statistica Neerlandica, 13(1): 37–58, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01011.x.

Goudswaard:1959:ISD

[338] G. Goudswaard. Internationale statistieken. (Dutch) [International statistics]. Statistica Neerlandica, 13(1):59–64, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb01012.x.

Thell:1959:SSS

[339] H. Thell and T. Kloek. The statistics of systems of simultaneous economic relationships. Statistica Neerlandica, 13(1):65–89, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01013.x.

vanderPutten:1959:EOI

[340] A. van der Putten and H. C. Swaagman. Enkele opmerkingen over de intensiteit van lokaal berichtenverkeer. (Dutch) [Some comments on the intensity of local messaging]. Statistica Neerlandica, 13(1):91– 100, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01014.x.

Anonymous:1959:BDBa

[341] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 13(1):101–110, March 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1959.tb01015.x.

vanEttinger:1959:PSD

[342] J. van Ettinger. Produktiviteit en statistiek. (Dutch) [Productivity and statistics]. Statistica Neerlandica, 13(2):111–114, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb00855.x.

Nanninga:1959:MVP

[343] H. G. C. Nanninga. De meting van de produktiviteit. (Dutch) [The measurement of productivity]. Statistica Neerlandica, 13(2):115–141, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00856.x.

Anonymous:1959:STN

[344] Anonymous. Statistische tabellen en nomogrammen. (Dutch) [Statistical tables and nomograms]. Statistica Neerlandica, 13(2):143– 146, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00857.x.

Anonymous:1959:PBO

[345] Anonymous. Produktiviteit bij onze auteurs en redactie. (Dutch) [Productivity among our authors and editors]. Statistica Neerlandica, 13(2): 147–149, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00858.x.

Anonymous:1959:PNI

[346] Anonymous. Prijsvraag: Een nummertje improduktiviteit!. (Dutch) [Competition: a song about unproductiveness!]. Statistica Neerlandica, 13(2):151–154, June 1959. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1959.tb00859.x.

Sittig:1959:POG

[347] J. Sittig. Produktiviteitsverhoging op grote schaal: Maatnormalisatie van kleding. (Dutch) [Productivity increase on a large scale: Size normalization of clothing]. Statistica Neerlandica, 13(2):155– 169, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00860.x.

Veen:1959:OVV

[348] B. Veen. Het optimaliseren van vertakte bemonsteringsschema’s. (Dutch) [Optimizing branched sampling schemes]. Statistica Neerlandica, 13(2): 171–186, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00861.x.

Hoeflake:1959:SVM

[349] G. J. Hoeflake. De statistiek van de Militair Geneeskundige Dienst. (Dutch) [The statistics of the Military Medical Service]. Statistica Neerlandica, 13(2):187–196, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1959.tb00862.x.

Bol:1959:VVA

[350] M. Bol, A. G. Gerritsen, and J. van Soest. Verhoging van de arbeidsproduktiviteit in de bosbouw door rationalisatieonderroek. (Dutch) [Increasing labor productivity in forestry through rationalization research]. Statistica Neerlandica, 13(2):197–202, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1959.tb00863.x.

Nanninga:1959:ELP

[351] H. G. C. Nanninga. Enige literatuur over produktiviteitsmeting. (Dutch) [Some literature on productivity measurement]. Statistica Neerlandica, 13(2):203–206, June 1959. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1959.tb00864.x.

Somermeijer:1959:EPD

[352] W. H. Somermeijer. Econometrische produktiviteitsanalyses. (Dutch) [Econometric productivity analyses]. Statistica Neerlandica, 13(2):207– 213, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00865.x.

deWolff:1959:PBA

[353] P. de Wolff. Produktiviteitsverhoging bij accountantscontrole door toepassing van gelaagde steekproeven. (Dutch) [Increasing productivity in auditing by applying layered sampling]. Statistica Neerlandica, 13(2): 215–232, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00866.x.

Barten:1959:SVT

[354] A. P. Barten. Het statistische verband tussen arbeidsproduktiviteit en produktie. (Dutch) [The statistical relationship between labor productivity and production]. Statistica Neerlandica, 13(2):233–242, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00867.x.

vanHees:1959:OHM

[355] R. N. van Hees and H. W. van den Meerendonk. Optimale hoeveelheid machinegereedschap. (Dutch) [Optimal amount of machine tools]. Statistica Neerlandica, 13(2):243–248, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1959.tb00868.x.

Anonymous:1959:BDBb

[356] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 13(2):249–253, June 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1959.tb00869.x.

Anonymous:1959:MPDa

[357] Anonymous. In memoriam Prof. Dr. M. J. van Uven 1878–1959. Statistica Neerlandica, 13(3):255–258, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1959.tb00870.x.

Anonymous:1959:SDD

[358] Anonymous. De Statistische Dag 1959. (Dutch) [The Statistical Day 1959]. Statistica Neerlandica, 13(3):259, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb00871.x.

Wiggers:1959:ESD

[359] B. G. Wiggers. Efficiency in de Statistiek. (Dutch) [Efficiency in Statistics]. Statistica Neerlandica, 13(3):261–280, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb00872.x.

Moroney:1959:ES

[360] M. J. Moroney. Efficiency by statistics. Statistica Neerlandica, 13(3):281– 294, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00873.x.

Kuiper:1959:SVI

[361] N. H. Kuiper. Een symbool voor Isomoor. (Dutch) [A symbol for isomor]. Statistica Neerlandica, 13(3):295–317, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb00874.x.

Root:1959:WDW

[362] W. J. Root. Wat doen wij in een bedrijf met statistiek?. (Dutch) [What do we do with statistics in a company?]. Statistica Neerlandica, 13 (3):321–322, September 1959. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1959.tb00875.x.

Doornbos:1959:EGV

[363] R. Doornbos. Efficiënte gewichtscontrole van pakjes margarine. (Dutch) [Efficient weight control of packets of margarine]. Statistica Neerlandica, 13(3):323–328, September 1959. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1959.tb00876.x.

vanMaare:1959:OCR

[364] J. van Maare. Omvang chauffeursploeg en rijtijdenbesluit. (Dutch) [Size of driver team and driving times decision]. Statistica Neerlandica, 13 (3):329–338, September 1959. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1959.tb00877.x.

Fortuin:1959:WVP

[365] G. J. Fortuin. Wachttijden van patienten. (Dutch) [Patient waiting times]. Statistica Neerlandica, 13(3):339–346, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb00878.x.

Bol:1959:RBD

[366] M. Bol and A. G. Gerritsen. Rationalisatieonderzoek in de bosbouw. (Dutch) [Rationalization research in forestry]. Statistica Neerlandica, 13(3):347–351, September 1959. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1959.tb00879.x.

Nonninga:1959:BED

[367] H. G. C. Nonninga. Bedrijfsvergelijking als efficiencymeter. (Dutch) [Business comparison as an efficiency meter]. Statistica Neerlandica, 13 (3):353–354, September 1959. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1959.tb00880.x.

tenHave:1959:OOD

[368] O. ten Have. Onderlinge omzetvergelijking. (Dutch) [Mutual sales comparison]. Statistica Neerlandica, 13(3):355–362, September 1959. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb00881.x.

Goudswaard:1959:EES

[369] G. Goudswaard. Efficiency en Economische Statistiek. (Dutch) [Efficiency and Economic Statistics]. Statistica Neerlandica, 13(3):363–379, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00882.x.

Klaassen:1959:EOM

[370] L H. Klaassen. Efficiency-onderzoek met behulp van statistische methoden. (Dutch) [Efficiency research using statistical methods]. Statistica Neerlandica, 13(3):381–388, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1959.tb00883.x.

Sittig:1959:NBD

[371] J. Sittig. Normalisatie als beslissingsprobleem. (Dutch) [Normalization as a decision problem]. Statistica Neerlandica, 13(3):389–394, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb00884.x.

Haitsma:1959:MCN

[372] A. Haitsma. Monte Carlo en nabootsing. (Dutch) [Monte Carlo and imitation]. Statistica Neerlandica, 13(3):395–406, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb00885.x.

Anonymous:1959:BDBc

[373] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 13(3):407–414, September 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1959.tb00886.x.

Anonymous:1959:MPDb

[374] Anonymous. In memoriam Prof. Dr. D. [David] van Dantzig (23 September 1900–22 July 1959). Statistica Neerlandica, 13(4):416–430, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01016.x.

Steyn:1959:TCT

[375] H. S. Steyn. On χ^{2}-tests for contingency tables of negative multinomial types. Statistica Neerlandica, 13(4):433–444, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb01017.x.

Fabius:1959:SAD

[376] J. Fabius. Stochastische approximatie. (Dutch) [Stochastic approximation]. Statistica Neerlandica, 13(4):445–452, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1959.
tb01018.x.

Doornbos:1959:SMV

[377] R. Doornbos. Statistische methoden voor het aanwijzen van uitbijters. (Dutch) [Statistical methods for identifying outliers]. Statistica Neerlandica, 13(4):453–462, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1959.tb01019.x.

Smid:1959:RSV

[378] L. J. Smid. Receptuur: De symmetrietoets van Wilcoxon, indien “gelijken” optreden. (Dutch) [Recipe: The symmetry test of Wilcoxon, if “equals” occur]. Statistica Neerlandica, 13(4):463–464, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01020.x.

Boogaerdt:1959:SOV

[379] J Boogaerdt. Selectuur: De oorsprong van de naam “Latijns vierkant”. (Dutch) [Selection: The origin of the name “Latin square”]. Statistica Neerlandica, 13(4):465–466, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1959.tb01021.x.

Derksen:1959:ISI

[380] J. B. D. Derksen. Het Indian Statistical Institute te Calcutta Organisatie en werkprogramma. (Dutch) [The Indian Statistical Institute at Calcutta Organization and work programme]. Statistica Neerlandica, 13(4):469– 475, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01022.x.

Somermeijer:1959:ITT

[381] W. H. Somermeijer. Inleiding tot de tabelkunde. (Dutch) [Introduction to tabulation]. Statistica Neerlandica, 13(4):477–485, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01023.x.

vanHees:1959:OBV

[382] R. N. van Hees and H. W. v. d. Meerendonk. Optimale betrouwbaarheid van parallelle meervoudige systemen. (Dutch) [Optimal reliability of parallel multiple systems]. Statistica Neerlandica, 13(4):487– 501, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01024.x.

Steutel:1959:IVP

[383] F. W. Steutel. De invloed van een prioriteitsregeling op de gemiddelde wachttijd. (Dutch) [The influence of a priority arrangement on the average waiting time]. Statistica Neerlandica, 13(4):503–513, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1959.tb01025.x.

Anonymous:1959:BDBd

[384] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 13(4):515–526, December 1959. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1959.tb01026.x.

Prins:1960:TFP

[385] H. J. Prins. Transforms for finding probabilities and variate values of a distribution function in tables of a related distribution function. Statistica Neerlandica, 14(1):1–17, March 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00887.x.

Nass:1960:MCB

[386] C. A. G. Nass. Een Monte-Carlo-bepaling van overschrijdingskansen, in verband met een keuzetest. (Dutch) [A Monte-Carlo determination of exceedance probabilities, in connection with a choice test]. Statistica Neerlandica, 14(1):19–31, March 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00888.x.

Anonymous:1960:ESA

[387] Anonymous. Het examen Statistisch Analist 1959. (Dutch) [The Statistical Analyst exam 1959]. Statistica Neerlandica, 14(1):33–46, March 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00889.x.

Anonymous:1960:NST

[388] Anonymous. Nieuwe Statistische tabellen en nomogrammen. (Dutch) [New statistical tables and nomograms]. Statistica Neerlandica, 14(1): 47–49, March 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00890.x.

Anonymous:1960:RQV

[389] Anonymous. Receptuur: Het quotiënt van von Neumann. (Dutch) [Recipe: The quotient of von Neumann]. Statistica Neerlandica, 14(1): 51–56, March 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00891.x.

Anonymous:1960:EET

[390] Anonymous. Enkele eigenschappen en toepassingen van de lognormale verdeling. (Dutch) [Some properties and applications of the lognormal distribution]. Statistica Neerlandica, 14(1):59–83, March 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb00892.x.

Kosten:1960:MMV

[391] L Kosten. Een mathematisch model voor een reserveringsprobleem. (Dutch) [A mathematical model for a reservation problem]. Statistica Neerlandica, 14(1):85–94, March 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00893.x.

Anonymous:1960:BDBa

[392] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 14(1):95–109, March 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00894.x.

vanEeden:1960:CLD

[393] Constance van Eeden and J. Th. Runnenburg. Conditional limitdistributions for the entries in a 2×2-table. Statistica Neerlandica, 14(2): 111–126, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00895.x.

Keuls:1960:TNV

[394] M. Keuls. Tabellen en nomogrammen voor het onderscheidingsver-mogen van de 5 en 10%-F-toets voor het gebruik bij de gewarde blokkenproef. (Dutch) [Tables and nomograms for the discriminative power of the 5% and 10%-F-test for use in the randomized block test]. Statistica Neerlandica, 14(2):127–150, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00896.x.

vanEedens:1960:RBV

[395] Constance van Eedens. Receptuur: Bepaling van een betrouwbaarheidsgebied voor de variatiecoefficient van een normale verdeling op grond van een steekproef. (Dutch) [Recipe: Determination of a confidence area for the coefficient of variation of a normal distribution based on a sample]. Statistica Neerlandica, 14(2):151–154, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00897.x.

Anonymous:1960:RDR

[396] Anonymous. Receptuur. (Dutch) [Recipe]. Statistica Neerlandica, 14 (2):155, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00898.x.

vanEedens:1960:SGM

[397] Constance van Eedens. Selectuur: Een grafische methode voor het berekenen van de toetsingsgrootheid van Kendall’s rangcorrelatietoet. (Dutch) [Selection: a graphical method for calculating the test statistic of Kendall’s rank correlation test]. Statistica Neerlandica, 14(2): 157–160, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00899.x.

Anonymous:1960:SDS

[398] Anonymous. Selectuur. (Dutch) [Selection]. Statistica Neerlandica, 14(2): 161–164, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00900.x.

vanderBurg:1960:UPN

[399] A. R. van der Burg. Uitslag prijsvraag “Een nummertje improduktiviteit”. (Dutch) [Result of the contest “A number of unproductiveness”]. Statistica Neerlandica, 14(2):165–172, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb00901.x.

Fortuin:1960:SOD

[400] G. J. Fortuin and A. van Beek. Een smaak-onderzoek. (Dutch) [A taste study]. Statistica Neerlandica, 14(2):175–185, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb00902.x.

vandePanne:1960:WWB

[401] C. van de Panne and A. Stranders. Winstgevendheid en winstkansen bij valuta-arbitrage. (Dutch) [Profitability and profit opportunities in currency arbitrage]. Statistica Neerlandica, 14(2):187–204, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00903.x.

vanBeeck:1960:SRB

[402] J. G. van Beeck, G. E. Engberts, and J. H. C. Lisman. Some remarks on business tendency surveys. (Dutch) [Some remarks on business tendency surveys]. Statistica Neerlandica, 14(2):205–222, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00904.x.

Anonymous:1960:BDBb

[403] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 14(2):223–229, June 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00905.x.

Anonymous:1960:MIR

[404] Anonymous. In memoriam Ir. A. R. Bloemena (1928–1960). Statistica Neerlandica, 14(3–4):232–234, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb01027.x.

Freudenthal:1960:OWD

[405] H. Freudenthal. Orde en wanorde. (Dutch) [Order and disorder]. Statistica Neerlandica, 14(3–4):235–247, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb01028.x.

vanSoest:1960:OWD

[406] J. L. van Soest. Orde en wanorde. (Dutch) [Order and disorder]. Statistica Neerlandica, 14(3–4):249–258, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb01029.x.

Stam:1960:EPU

[407] A. J. Stam. Enige punten uit de wiskunde der informatietheorie. (Dutch) [Some points from the mathematics of information theory]. Statistica Neerlandica, 14(3–4):259–265, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb01030.x.

Bloemena:1960:AKV

[408] A. R. Bloemena. Het aselect kiezen van knopen in een knoopsel. (Dutch) [The random selection of knots in a knot]. Statistica Neerlandica, 14 (3–4):267–274, September 1960. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1960.tb01031.x.

Hulshof:1960:AVP

[409] A. H. Hulshof and J. J. Wouterse. Achtergronden van planning. (Dutch) [Backgrounds to planning]. Statistica Neerlandica, 14(3–4):277–285, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb01032.x.

deBruijne:1960:ONI

[410] J. I. de Bruijne and En A. R. Bloemena. Een onderzoek naar de invloed van 2 mg vitamine K_{1} bij vroeggeborenen. (Dutch) [A study into the influence of 2 mg vitamin K_{1} in premature infants]. Statistica Neerlandica, 14(3–4):287–294, September 1960. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1960.tb01033.x.

Scheijde:1960:OWM

[411] E. H. Scheijde. Orde en wanorde in de medische statistiek in ziekenhuizen. (Dutch) [Order and disorder in medical statistics in hospitals]. Statistica Neerlandica, 14(3–4):295–307, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb01034.x.

tenHave:1960:OWS

[412] O. ten Have. Orde — Wanorde — Statistiek. (Dutch) [Order — disorder — statistics]. Statistica Neerlandica, 14(3–4):309–317, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb01035.x.

Barten:1960:RBV

[413] A. P. Barten. Residuen als bron van informatie bij regressie-analyse. (Dutch) [Residuals as a source of information in regression analysis]. Statistica Neerlandica, 14(3–4):319–341, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb01036.x.

Schmidt:1960:KSD

[414] P. Schmidt. De kostprijs en statistiek. (Dutch) [The cost price and statistics]. Statistica Neerlandica, 14(3–4):343–351, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb01037.x.

Harteveld:1960:STO

[415] R. B. Harteveld. De strijd tussen orde en wanorde, een optimumprobleem. (Dutch) [The struggle between order and disorder, an optimum problem]. Statistica Neerlandica, 14(3–4):353–371, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb01038.x.

Anonymous:1960:BDBc

[416] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 14(3–4):373–378, September 1960. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1960.tb01039.x.

Hamaker:1961:TSD

[417] H. C. Hamaker. Toegepaste statistiek. (Dutch) [Applied statistics]. Statistica Neerlandica, 15(1):1–18, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1961.tb00906.x.

Theil:1961:BQE

[418] H. Theil and A. Schweitzer. The best quadratic estimator of the residual variance in regression analysis. Statistica Neerlandica, 15(1):19– 23, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00907.x. See different solution [549, 577].

vanEedens:1961:RST

[419] Constance van Eedens and Chr. L. Rümke. Receptuur: a k sample trendtest. Statistica Neerlandica, 15(1):25–29, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00908.x.

Lisman:1961:NNV

[420] J. H. C. Lisman. Norm NEN 3117 van het Nederlands Normalisatieinstituut. (Dutch) [Standard NEN 3117 of the Netherlands Standardization Institute]. Statistica Neerlandica, 15(1):31–32, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00909.x.

Anonymous:1961:TMB

[421] Anonymous. Times meets in Brussels in 1961. Statistica Neerlandica, 15 (1):34, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00910.x.

Vos:1961:BVV

[422] J. W. E. Vos. De Bibliotheek van de Vereniging voor Statistiek. (Dutch) [The Library of the Association for Statistics]. Statistica Neerlandica, 15 (1):35–36, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00911.x.

Rumke:1961:MBS

[423] Chr. L. Rümke. De Medisch–Biologische Sectie van de Vereniging voor Statistiek. (Dutch) [The Medical–Biological Section of the Association for Statistics]. Statistica Neerlandica, 15(1):39, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00912.x.

Anonymous:1961:BDBa

[424] Anonymous. Bibliografie (1946–1960). (Dutch) [Bibliography (1946– 1960)]. Statistica Neerlandica, 15(1):41–46, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00913.x.

Rumke:1961:AIT

[425] Chr. L. Rümke. Een algemene inleiding tot de sequente analyse. (Dutch) [A general introduction to sequence analysis]. Statistica Neerlandica, 15 (1):47–54, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00914.x.

Goslings:1961:MEZ

[426] J. Goslings. De medisch-ethische zijde van de sequente analyse. (Dutch) [The medical-ethical side of sequence analysis]. Statistica Neerlandica, 15 (1):55–72, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00915.x.

Armitage:1961:SAM

[427] P. Armitage. Sequential analysis in medicine. Statistica Neerlandica, 15 (1):73–82, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00916.x.

Armitage:1961:TAS

[428] P. Armitage. Theoretical aspects of sequential designs. Statistica Neerlandica, 15(1):83–89, March 1961. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1961.tb00917.x.

vanElteren:1961:TRS

[429] Ph. van Elteren. Two restricted sequential sign tests. Statistica Neerlandica, 15(1):91–96, March 1961. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1961.tb00918.x.

Anonymous:1961:BDBb

[430] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 15(1):97–105, March 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00919.x.

Runnenburg:1961:EVV

[431] J. Th. Runnenburg. Enige voorbeelden van stochastische processen. (Dutch) [Some examples of stochastic processes]. Statistica Neerlandica, 15(2):107–122, June 1961. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1961.tb00920.x.

deVries:1961:PES

[432] M. de Vries. Practical estimators for sampling on successive occasions. Statistica Neerlandica, 15(2):123–133, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00921.x.

Anonymous:1961:ESA

[433] Anonymous. Het examen Statistisch Analist 1960. (Dutch) [The Statistical Analyst exam 1960]. Statistica Neerlandica, 15(2):135–145, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00922.x.

vanMeurs:1961:SMA

[434] J. van Meurs and J. H. C. Lisman. Selectuur: Meest aannemelijke verdelingen. (Dutch) [Selection: Most plausible distributions]. Statistica Neerlandica, 15(2):147–151, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00923.x.

deWolff:1961:NGS

[435] P. de Wolff. Niet gecorreleerde schatters. (Dutch) [Uncorrelated estimators]. Statistica Neerlandica, 15(2):153–155, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00924.x.

vanBraam:1961:ABH

[436] A. van Braam. Aantekeningen bij de herziening van de methodiek der prijsindexcijfers van het levensonderhoud. (Dutch) [Notes on the revision of the methodology of the price indices of subsistence]. Statistica Neerlandica, 15(2):159–170, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00925.x.

Venekamp:1961:BSD

[437] P. E. Venekamp. Beschrijvende statistiek. (Dutch) [Descriptive statistics]. Statistica Neerlandica, 15(2):171–176, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00926.x.

vdMeerendonk:1961:AVS

[438] H. W. v. d. Meerendonk and G. G. j. Bos. Assortimentsbepaling van seizoenprodukten. (Dutch) [Determining the range of seasonal products]. Statistica Neerlandica, 15(2):177–187, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00927.x.

Sandee:1961:OVI

[439] J. Sandee. Een ontwikkelingsmodel voor India. (Dutch) [A development model for India]. Statistica Neerlandica, 15(2):189–210, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00928.x.

Anonymous:1961:BDBc

[440] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 15(2):211–223, June 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00929.x.

Hemelrijk:1961:ADR

[441] J. Hemelrijk. Aselect. (Dutch) [Random]. Statistica Neerlandica, 15 (3):225–237, September 1961. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1961.tb00452.x.

vanderVaart:1961:SDL

[442] H. R. van der Vaart. A simple derivation of the limiting distribution function of a sample quantile with increasing sample size. Statistica Neerlandica, 15(3):239–242, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00453.x.

Levert:1961:MVD

[443] C. Levert. Een model voor discreet verdeelde grootheden met negatieve correlatie. (Dutch) [A model for discretely distributed quantities with negative correlation]. Statistica Neerlandica, 15(3):243–246, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00454.x.

vanOostrum:1961:RTB

[444] H. R. van Oostrum. Een rekenwijze ter bepaling van enkele kansverdelingen. (Dutch) [A calculation method for determining some probability distributions]. Statistica Neerlandica, 15(3):247–250, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00455.x.

Anonymous:1961:RDE

[445] Anonymous. Redactiecommissies:. (Dutch) [Editorial committees:]. Statistica Neerlandica, 15(3):251, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00456.x.

Anonymous:1961:SOS

[446] Anonymous. Stichting Opleidingen Statistiek. (Dutch) [Foundation Training Statistics]. Statistica Neerlandica, 15(3):252, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (elec-

tronic).  URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00457.x.

Bakker:1961:SVI

[447] E. Bakker. De spreiding van inschrijvingsbedragen op openbare aanbestedingen. (Dutch) [The spread of tender amounts for public tenders]. Statistica Neerlandica, 15(3):253–265, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00458.x.

vanKreveld:1961:AOV

[448] A. van Kreveld. Aanvoer en opslag van dunne melk in een melkproduktenfabriek. (Dutch) [Supply and storage of thin milk in a dairy products factory]. Statistica Neerlandica, 15(3):267–291, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00459.x.

vandePanne:1961:IMD

[449] C. van de Panne and G. J. Aeyclts Averink. Imperfect management decisions and predictions and their financial implications in dynamic quadratic cost minimization. Statistica Neerlandica, 15(3):293–318, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00460.x.

Anonymous:1961:BDBd

[450] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 15(3):319–331, September 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00461.x.

Kendallo:1961:SPP

[451] M. G. Kendallo. Statistics, past, present and future. Statistica Neerlandica, 15(4):333–349, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00930.x.

vanEttinger:1961:VGT

[452] J. van Ettinger. Van geval tot geval met als metgezel net getal. (Dutch) [From case to case with net number as companion]. Statistica Neerlandica, 15(4):351–361, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00931.x.

deVries:1961:SS

[453] M. de Vries. Selective sampling. Statistica Neerlandica, 15(4):363–383, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00932.x.

vanElteren:1961:WVB

[454] Ph. van Elteren and H. J. M. Gerrits. Een wachtprobleem voorkomende bij drempelwaardemetingen aan het oog. (Dutch) [A waiting problem occurring in threshold value measurements on the eye]. Statistica Neerlandica, 15(4):385–401, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00933.x.

Wijvekate:1961:GVG

[455] M. L. Wijvekate. Gevallen van getallen in een bedrijf. (Dutch) [Cases of numbers in a company]. Statistica Neerlandica, 15(4):405–417, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00934.x.

deGroot:1961:VZT

[456] Meindert J. W. de Groot. Van ziektegeval tot ziektengetal. (Dutch) [From case of disease to number of diseases]. Statistica Neerlandica, 15(4):419– 429, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00935.x.

Nelemans:1961:KEV

[457] F. A. Nelemans, W. G. Zelvelder, and G. J. Leppink. Klinische evaluatie van pijnstillende middelen. (Dutch) [Clinical evaluation of analgesics]. Statistica Neerlandica, 15(4):431–438, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00936.x.

Meerdink:1961:VGT

[458] J. Meerdink. Van geval tot getal. (Dutch) [From case to number]. Statistica Neerlandica, 15(4):439–449, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00937.x.

Vos:1961:HGS

[459] J. W. E. Vos. Hoe groot is de steekproef?. (Dutch) [How large is the sample?]. Statistica Neerlandica, 15(4):451–460, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00938.x.

Tinbergen:1961:CIS

[460] J. Tinbergen. Comparatieve internationale studies als hulpmiddel bij de ontwikkelingsplanning. (Dutch) [Comparative international studies as a tool for development planning]. Statistica Neerlandica, 15(4):461– 465, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00939.x.

Claus:1961:BVB

[461] H. Claus. De betekenis van de bedrijfseconometrie voor de onderneming. (Dutch) [The significance of business econometrics for the company]. Statistica Neerlandica, 15(4):467–487, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1961.
tb00940.x.

Bos:1961:EAV

[462] G. G. J. Bos. Enkele aspecten van voorraadketens. (Dutch) [Some aspects of supply chains]. Statistica Neerlandica, 15(4):489–500, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1961.tb00941.x.

Anonymous:1961:IAO

[463] Anonymous. International abstracts in operations research. Statistica Neerlandica, 15(4):501–502, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1961.tb00942.x.

Anonymous:1961:BDBe

[464] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 15(4):503–509, December 1961. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1961.tb00943.x.

Sieben:1962:EDE

[465] J. W. Sieben. Experimenten. (Dutch) [Experiments]. Statistica Neerlandica, 16(1):1–16, March 1962. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1962.tb01181.x.

Strackee:1962:FDD

[466] J. Strackee and J. J. Denier van der Gon. The frequency distribution of the difference between two Poisson variates. Statistica Neerlandica, 16 (1):17–23, March 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01182.x.

Barten:1962:RST

[467] A. P. Barten. Receptuur: Schatting en toetsing van een gebroken lineair verband. (Dutch) [Recipe: Estimation and testing of a broken linear relationship]. Statistica Neerlandica, 16(1):25–28, March 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01183.x.

Hamaker:1962:MRA

[468] H. C. Hamaker. On multiple regression analysis. Statistica Neerlandica, 16(1):31–56, March 1962. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1962.tb01184.x.

Kloek:1962:PCA

[469] T. Kloek and R. Bannink. Principal-component analysis applied to business test data. Statistica Neerlandica, 16(1):57–69, March 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1962.
tb01185.x.

Kloek:1962:ASB

[470] T. Kloek and D. S. Jochems. Alternative specifications of bivariate relations in the analysis of business test data. Statistica Neerlandica, 16(1): 71–87, March 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01186.x.

Sandee:1962:DNC

[471] J. Sandee and J. H. C. Lisman. Defining the no-change interval in business tendency surveys. Statistica Neerlandica, 16(1):89–97, March 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01187.x.

vanderSloot:1962:SVK

[472] J. de Boer en J. van der Sloot. Een systeem van kostenbeheersende productieplanning. (Dutch) [A system of cost-controlled production planning]. Statistica Neerlandica, 16(1):99–112, March 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1962.
tb01188.x.

Moonen:1962:BVB

[473] H. J. M. Moonen. Het bepalen van bestelniveaus wanneer afname en levertijd gamma-verdeeld resp. normaal-verdeeld zijn. (Dutch) [Determining order levels when purchase and delivery time are gamma-distributed respectively. be normally distributed]. Statistica Neerlandica, 16(1):113– 120, March 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01189.x.

Anonymous:1962:BDBa

[474] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 16(1):121–138, March 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1962.tb01190.x.

vanLisman:1962:AVD

[475] Afscheid van Lisman. Afscheid van Dr. Lisman: en andere wijzigingen in de redactie. (Dutch) [Farewell of Dr. Lisman: and other editorial changes]. Statistica Neerlandica, 16(2):141–142, June 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1962.
tb01060.x.

Wynn:1962:NFC

[476] P. Wynn. A note on the fitting of certain types of experimental data. Statistica Neerlandica, 16(2):143–150, June 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1962.tb01061.x.

Barten:1962:NUE

[477] A. P. Barten. Note on unbiased estimation of the squared multiple correlation coefficient. Statistica Neerlandica, 16(2):151–164, June 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01062.x.

Anonymous:1962:ESA

[478] Anonymous. Het examen Statistisch Analist 1961. (Dutch) [The Statistical Analyst exam 1961]. Statistica Neerlandica, 16(2):165–176, June 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01063.x.

Ferguson:1962:NVO

[479] J. H. A. Ferguson. Nomogram voor het onderscheidingsvermogen van de F -toets gemodificeerd naar Keuls. (Dutch) [Nomogram for the discriminative power of the F test modified to Keuls]. Statistica Neerlandica, 16 (2):177–181, June 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01064.x.

Bakker:1962:PPO

[480] E. Bakker. Over prijsraming en prijsvorming op de bouwmarkt. (Dutch) [About price estimation and pricing on the construction market]. Statistica Neerlandica, 16(2):183–194, June 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1962.tb01065.x.

vanderVeen:1962:MLP

[481] B. van der Veen. Een meer-loketten probleem met overwerk: a multistation queueing model with overtime. (Dutch) [A multi-station problem with overtime: a multistation queuing model with overtime]. Statistica Neerlandica, 16(2):195–204, June 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1962.tb01066.x.

Anonymous:1962:BDBb

[482] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 16(2):205–212, June 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1962.tb01067.x.

Anonymous:1962:MPL

[483] Anonymous. In memoriam prof. de Loor (1901–9 February 1962). Statistica Neerlandica, 16(3):213, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1962.tb01068.x.

Goudriaan:1962:MVE

[484] J. Goudriaan. Maxima voor de eenzijdige overschrijdingskans en voor de grootte van de gemiddelde overschrijding bij discontinue en continue, eentoppige verdelingen. (Dutch) [Maxima for the one-sided exceedance probability and for the size of the average exceedance in discontinuous and continuous, one-top distributions]. Statistica Neerlandica, 16(3):215– 230, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01069.x.

Kuiper:1962:MSR

[485] Nicolaas H. Kuiper. On a metric in the space of random variables. Statistica Neerlandica, 16(3):231–235, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1962.
tb01070.x.

Nagar:1962:STA

[486] A. L. Nagar. Statistical testing of the accuracy of forecasts. Statistica Neerlandica, 16(3):237–247, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1962.tb01071.x.

Korswagen:1962:UKD

[487] Catharina Korswagen. Uitschietende kansen. (Dutch) [Outstanding opportunities]. Statistica Neerlandica, 16(3):249–253, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01072.x.

Hemelrijk:1962:AEV

[488] J. Hemelrijk. De asymptotische eigenschappen van de binomiale verdeling. (Dutch) [The asymptotic properties of the binomial distribution]. Statistica Neerlandica, 16(3):255–259, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1962.
tb01073.x.

Rumke:1962:EOO

[489] Chr. L. Rümke. Enkele opmerkingen over de opzet van onderzoekingen naar de werkzaamheid van cytostatica in de kliniek. (Dutch) [Some comments on the design of studies into the efficacy of cytotoxic drugs in the clinic]. Statistica Neerlandica, 16(3):261–268, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01074.x.

Theil:1962:MPJ

[490] H. Theil. In memoriam P. J. M. van den Bogaard. Statistica Neerlandica, 16(3):269, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1962.tb01075.x.

vandenBogaard:1962:DOC

[491] P. J. M. van den Bogaard and J. Versluis. The design of optimal committee decisions. Statistica Neerlandica, 16(3):271–289, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01076.x.

Bakker:1962:OSB

[492] W. Bakker. Optimale snijsnelheid bij meer-machine-bediening. (Dutch) [Optimal cutting speed and the machine assignment problem]. Statistica Neerlandica, 16(3):291–302, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1962.tb01077.x.

Anonymous:1962:BDBc

[493] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 16(3):303–316, September 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1962.tb01078.x.

Casimir:1962:DMT

[494] H. B. G. Casimir. Door meten tot weten in de Natuurwetenschap. (Dutch) [By measuring to knowing in natural science]. Statistica Neerlandica, 16(4):317–320, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1962.tb01079.x.

deWolff:1962:DMT

[495] P. de Wolff. Door meten tot weten in de sociale wetenschappen. (Dutch) [Through measurement to knowledge in the social sciences]. Statistica Neerlandica, 16(4):321–336, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1962.tb01080.x.

Runnenburg:1962:SNR

[496] J. Th. Runnenburg. Some numerical results on waiting-time distributions for dependent arrival-intervals. Statistica Neerlandica, 16(4):337– 347, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01081.x.

Gelling:1962:SVA

[497] H. Gelling and A. H. Schaafsma. Samenvatting van het artikel “Procesnauwkeurig heid en toleranties”. (Dutch) [Summary of the article “Process accuracy and tolerances”]. Statistica Neerlandica, 16(4):349– 351, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01082.x.

Melessen:1962:TBB

[498] H. K. J. Melessen. Typebeperking bij bevoorradingsmagazijnen van onder-houdsdiensten. (Dutch) [Type limitation for supply warehouses of maintenance services]. Statistica Neerlandica, 16(4):353–360, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01083.x.

Lammers:1962:BWV

[499] W. Lammers. Biologische waardebepalingen van geneesmiddelen. (Dutch) [Biological evaluations of medicines]. Statistica Neerlandica, 16 (4):361–371, December 1962. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1962.tb01084.x.

Keuls:1962:WAO

[500] M. Keuls. Weten in aansluiting op een variantie-analyse. (Dutch) [Knowing in connection with a variance analysis]. Statistica Neerlandica, 16 (4):373–387, December 1962. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1962.tb01085.x.

deJonge:1962:NAM

[501] W. J. de Jonge. Nationaal Advertentie Media onderzoek. (Dutch) [National Advertising Media survey]. Statistica Neerlandica, 16(4):389–402, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01086.x.

Middelhoek:1962:SVE

[502] A. J. Middelhoek. Een sectormodel van het energieverbruik. (Dutch) [A sector model of energy consumption]. Statistica Neerlandica, 16(4):403– 431, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01087.x.

deLeve:1962:SSB

[503] G. de Leve. Stochastische ∞-staps beslissingsproblemen. (Dutch) [Stochastic ∞-step decision problems]. Statistica Neerlandica, 16(4):433– 448, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1962.tb01088.x.

Lombaers:1962:WDM

[504] H. J. M. Lombaers. Weten door Meten aan digitaal gesimuleerde processen. (Dutch) [Knowing by Measuring digitally simulated processes]. Statistica Neerlandica, 16(4):449–458, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1962.
tb01089.x.

Anonymous:1962:BDBd

[505] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 16(4):459–468, December 1962. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1962.tb01090.x.

Anonymous:1963:MVR

[506] Anonymous. Mededeling van de redactie. (Dutch) [Editor’s note]. Statistica Neerlandica, 17(1):1, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb00301.x.

Cohen:1963:NDP

[507] J. W. Cohen. A note on the delay problem in crossing a traffic stream. Statistica Neerlandica, 17(1):3–11, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1963.
tb00302.x.

Ivanovic:1963:ETD

[508] B. Ivanovic. The estimation of the two-dimensional distribution of a statistical set, based on the marginal distributions of its strata. Statistica Neerlandica, 17(1):13–24, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb00303.x.

Runnenburg:1963:RTV

[509] J. Th. Runnenburg and F. W. Steutel. Receptuur: De tabellen van Peck en Hazelwood. (Dutch) [Recipe: The tables of Peck and Hazelwood]. Statistica Neerlandica, 17(1):25–26, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1963.tb00304.x.

Anonymous:1963:S

[510] Anonymous. Sectiegedeelte. (Dutch) [sections]. Statistica Neerlandica, 17 (1):27–28, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00305.x.

vandeBosch:1963:CWT

[511] J. P. van de Bosch. De computer: Werkwijze en toepassingsmogelijkheden van elektronische rekenmachines. (Dutch) [The computer: Method and applications of electronic calculators]. Statistica Neerlandica, 17(1): 29–36, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00306.x.

vandenMeerendonk:1963:SCA

[512] H. W. van den Meerendonk, J. A. G. M. Kerbosch, P. Medema, and J. H. Schouten. Some computational aspects of a trimloss problem. Statistica Neerlandica, 17(1):37–47, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb00307.x.

vdHeiden:1963:PMS

[513] J. A. v. d. Heiden. Planning the maintenance of simple machinery. Statistica Neerlandica, 17(1):49–69, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1963.tb00308.x.

Keuls:1963:DME

[514] M. Keuls, H. J. Over, and C. T. de Wit. The distance method for estimating densities. Statistica Neerlandica, 17(1):71–91, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1963.
tb00309.x.

Anonymous:1963:B

[515] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 17(1):93–103, March 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb00310.x.

vanderGon:1963:VVT

[516] J. J. Denier van der Gon. Vergelijking van een tweetal toetsen tegen verloop voor een aantal kansen. (Dutch) [Comparison of two tests against progression for a number of opportunities]. Statistica Neerlandica, 17(2): 105–112, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00311.x.

Anonymous:1963:ESA

[517] Anonymous. Het examen Statistisch Analist 1962. (Dutch) [The Statistical Analyst exam 1962]. Statistica Neerlandica, 17(2):113–121, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00312.x.

Anonymous:1963:SDSb

[518] Anonymous. Sectiegedeelte. (Dutch) [Section section]. Statistica Neerlandica, 17(2):123, June 1963. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1963.tb00313.x.

Bakker:1963:EAF

[519] E. Bakker. An econometric analysis of factors determining contractors’ interest in public tenders. Statistica Neerlandica, 17(2):125– 136, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00314.x.

Reith:1963:PPM

[520] P. F. Reith. Project-planning methoden. (Dutch) [Project planning methods]. Statistica Neerlandica, 17(2):137–173, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1963.
tb00315.x.

Kuiper:1963:TJS

[521] N. H. Kuiper. Twintig jaren Studiekring voor Statistische Techniek. (Dutch) [Twenty years of Study Circle for Statistical Technology]. Statistica Neerlandica, 17(2):175–185, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1963.tb00316.x.

Dijkstra:1963:A

[522] E. W. Dijkstra. ,,Algol 60“. Statistica Neerlandica, 17(2):187–188, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00317.x.

Anonymous:1963:BDBb

[523] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 17(2):189–197, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb00318.x.

Bosch:1963:PD

[524] A. J. Bosch. The Pólya distribution. Statistica Neerlandica, 17(3):201– 213, September 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb01040.x.

Kriens:1963:MVW

[525] J. Kriens. De methoden van De Wolff en Van Heerden voor het nemen van aselecte steekproeven bij accountantscontroles. (Dutch) [The methods of De Wolff and Van Heerde for taking random samples in audits]. Statistica Neerlandica, 17(3):215–231, September 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1963.tb01041.x.

Anonymous:1963:SDSc

[526] Anonymous. Sectiegedeelte. (Dutch) [Section section]. Statistica Neerlandica, 17(3):233, September 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb01042.x.

Balk:1963:ORT

[527] W. Balk and C. Kool. De optimale regelbreedte in de telefoongids. (Dutch) [The optimal line width in the telephone directory]. Statistica Neerlandica, 17(3):235–242, September 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1963.tb01043.x.

deBruyn:1963:SSD

[528] C. S. van Dobben de Bruyn. Smoothing sales data by progressive correction. Statistica Neerlandica, 17(3):243–265, September 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1963.
tb01044.x.

Koerts:1963:MWT

[529] J. Koerts. On mean waiting times and their reduction by priority procedures: an expository survey and some tables. Statistica Neerlandica, 17(3):267–283, September 1963. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1963.tb01045.x.

Anonymous:1963:MPD

[530] Anonymous. In memoriam Prof. Dr. L. M. Koyck. Statistica Neerlandica, 17(3):285–286, September 1963. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1963.tb01046.x.

Mouchart:1963:PVI

[531] M. Mouchart, H. Theil, and J. I. Vorst. On the predictive value of investment surveys. Statistica Neerlandica, 17(3):287–297, September 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb01047.x.

Keuls:1963:USB

[532] M. Keuls. Uitbijternegerende schatters bij duplobepalingen. (Dutch) [Outlier-ignoring estimators for duplicate determinations]. Statistica Neerlandica, 17(3):299–317, September 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1963.tb01048.x.

Anonymous:1963:BDBc

[533] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 17(3):319–327, September 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb01049.x.

Goudriaan:1963:GBO

[534] J. Goudriaan. Gissen en beslissen op statistische grondslag. (Dutch) [Guessing and deciding on a statistical basis]. Statistica Neerlandica, 17(4):331–362, December 1963. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1963.tb00319.x.

Monhemius:1963:BVL

[535] W. Monhemius. Besliskunde, een vlag die de lading dekt?. (Dutch) [Decision making, a flag that covers the load?]. Statistica Neerlandica, 17 (4):363–382, December 1963. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1963.tb00320.x.

Anonymous:1963:SDSd

[536] Anonymous. Sectiegedeelte. (Dutch) [Section section]. Statistica Neerlandica, 17(4):383–384, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb00321.x.

Bakker:1963:GBD

[537] A. Bakker. Gegist bestek. (Dutch) [Dead reckoning]. Statistica Neerlandica, 17(4):385–394, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb00322.x.

vandenToorn:1963:MVS

[538] J. W. van den Toorn. Het maken van een schaal voor niet meetbare kenmerken. (Dutch) [Creating a scale for non-measurable characteristics]. Statistica Neerlandica, 17(4):395–400, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1963.
tb00323.x.

Sukhatme:1963:IH

[539] P. V. Sukhatme. The incidence of hunger. Statistica Neerlandica, 17 (4):401–424, December 1963. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1963.tb00324.x.

Bos:1963:HVB

[540] H. C. Bos. Hetschatten van de behoefte aan onderwijs in ontwikkelingslanden. (Dutch) [Estimating the need for education in developing countries]. Statistica Neerlandica, 17(4):425–429, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1963.
tb00325.x.

Oppers:1963:GBM

[541] V. M. Oppers. Gissen en beslissen met historische gegevens. (Dutch) [Guessing and deciding with historical data]. Statistica Neerlandica, 17 (4):431–448, December 1963. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1963.tb00326.x.

deJonge:1963:BZV

[542] H. de Jonge. Beslissen zonder (ver)gissen bij de keuze van een toets. (Dutch) [Make decisions without making mistakes when choosing a test]. Statistica Neerlandica, 17(4):449–477, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1963.
tb00327.x.

vanHees:1963:PBZ

[543] R. N. van Hees. Produktiebesturing: beslissen zonder gissen?. (Dutch) [Production control: making decisions without guessing?]. Statistica Neerlandica, 17(4):479–486, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1963.tb00328.x.

vanElteren:1963:PVA

[544] Ph. van Elteren. Een permutatietoets voor alternatief verdeelde grootheden. (Dutch) [A permutation test for alternatively distributed quantities]. Statistica Neerlandica, 17(4):487–505, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1963.
tb00329.x.

Corsten:1963:VDT

[545] L C. A. Corsten. Verschillen in doeltreffendheid tussen groepen van voor-spellende variabelen. (Dutch) [Differences in effectiveness between groups of predictive variables]. Statistica Neerlandica, 17(4):507–512, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00330.x.

Anonymous:1963:E

[546] Anonymous. Errata. Statistica Neerlandica, 17(4):513, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1963.tb00331.x.

Anonymous:1963:BDBd

[547] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 17(4):515–520, December 1963. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1963.tb00332.x.

Corsten:1964:IOG

[548] L. C. A. Corsten. Identificeren op grond van waarnemingen. (Dutch) [Identify based on observations]. Statistica Neerlandica, 18(1):1–14, March 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00333.x.

Corsten:1964:DSP

[549] L C. A. Corsten. A different solution of a problem posed by Theil and Schweitzer. Statistica Neerlandica, 18(1):15–18, March 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.

tb00334.x. See [418] and correction [577].

Keuls:1964:OVP

[550] M. Keuls and L. R. Verdooren. Overzicht van de publikaties door prof. dr. M. J. van Uven. (Dutch) [Overview of publications by Prof. Dr. M. J. van Uven]. Statistica Neerlandica, 18(1):19–25, March 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00335.x.

deJonge:1964:RVB

[551] H. de Jonge. Receptuur: Verdelingsvrije betrouwbaarheidsgrenzen. (Dutch) [Recipe: Distribution-free confidence limits]. Statistica Neerlandica, 18(1):27–33, March 1964. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1964.tb00336.x.

Anonymous:1964:SDS

[552] Anonymous. Sectiegedeelte. (Dutch) [Section section]. Statistica Neerlandica, 18(1):35, March 1964. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1964.tb00337.x.

Venekamp:1964:TRJ

[553] P. E. Venekamp. Tijdreeksen en regionale jaarrekeningen. (Dutch) [Time series and regional annual accounts]. Statistica Neerlandica, 18(1):37– 44, March 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00338.x.

Boas:1964:SAP

[554] J. Boas and J. Koerts. A simulation approach to the problem of the relationship between business test data and conventional statistics. Statistica Neerlandica, 18(1):45–59, March 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1964.tb00339.x.

Bakker:1964:OSB

[555] W. Bakker. Optimale snijsnekheid bij meer-machine-bediening. (Dutch) [Optimal cutting cleanliness with multiple machine operation]. Statistica Neerlandica, 18(1):61–64, March 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1964.tb00340.x.

Sittig:1964:PVW

[556] J. Sittig. Programmering van de woningbouw in verband met de woningkwaliteit. (Dutch) [Programming of housing construction in connection with housing quality]. Statistica Neerlandica, 18(1):65–73, March 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00341.x.

Anonymous:1964:BDBa

[557] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 18(1):75–94, March 1964. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1964.tb00342.x.

Nass:1964:PDB

[558] C. A. G. Nass. Prof. Dr. S. T. Bok — 4 februari 1964. Statistica Neerlandica, 18(2):97–98, June 1964. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1964.tb00499.x.

Leppink:1964:MDM

[559] G. J. Leppink. Modellen. (Dutch) [Models]. Statistica Neerlandica, 18 (2):99–109, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00500.x.

Bakker:1964:SMK

[560] W. Bakker. Signaalstroomdiagrammen en Markov-ketens. (Dutch) [Signal flow diagrams and Markov chains]. Statistica Neerlandica, 18(2): 111–128, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00501.x.

Vermetten:1964:QVN

[561] J. B. Vermetten and J. H. C. Lisman. Het quotiënt van von Neumann bij regressie-analyse. (Dutch) [The quotënt of von Neumann in regression analysis]. Statistica Neerlandica, 18(2):129–138, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00502.x.

Terpstra:1964:ESE

[562] T. J. Terpstra. Enkele stochastische eigenschappen van een uit vezels ge-sponnen garen voor het ideale klassieke spinproces en enkele ermee samenhangende problemen. (Dutch) [Some stochastic properties of a fiber spun yarn for the ideal classical spinning process and some associated problems]. Statistica Neerlandica, 18(2):139– 149, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00503.x.

Hogendijk:1964:NSS

[563] M. J. Hogendijk. Nieuwe serie statistische tabellen en nomogrammen. (Dutch) [New series of statistical tables and nomograms]. Statistica Neerlandica, 18(2):151–152, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1964.tb00504.x.

Anonymous:1964:ESA

[564] Anonymous. Het examen Statistisch Analist 1963. (Dutch) [The Statistical Analyst exam 1963]. Statistica Neerlandica, 18(2):153–161, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00505.x.

Sieben:1964:ISK

[565] J. W. Sieben. Industriële Statistiek en Kwaliteitsbeleid in de toekomst. (Dutch) [Industrial Statistics and Quality Policy in the future]. Statistica Neerlandica, 18(2):165–173, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1964.tb00506.x.

Sittig:1964:KMD

[566] J. Sittig. Kwaliteitsbeleid morgen. (Dutch) [Quality policy tomorrow]. Statistica Neerlandica, 18(2):175–187, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00507.x.

Willemze:1964:KIO

[567] F. G. Willemze. Kwaliteitsbeleid en Industriele Organisatie “Morgen”. (Dutch) [Quality policy and Industrial Organization “Tomorrow”]. Statistica Neerlandica, 18(2):189–200, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1964.tb00508.x.

Emanuel:1964:IMD

[568] H. Emanuel. Internationaal marktonderzoek. (Dutch) [International Market Research]. Statistica Neerlandica, 18(2):201–211, June 1964. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00509.x.

Kosten:1964:IMT

[569] L. Kosten. Inverse Matrix technique for solving the transportation problem. Statistica Neerlandica, 18(2):213–224, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00510.x.

Anonymous:1964:BDBb

[570] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 18(2):225–238, June 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1964.tb00511.x.

Runnenburg:1964:OWD

[571] J. Th. Runnenburg. Onderwijs en waarschijnlijkheidsrekening. (Dutch) [Education and probability theory]. Statistica Neerlandica, 18(3):241– 257, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00512.x.

Goudriaan:1964:AND

[572] J. Goudriaan. Another note on the delay problem in crossing a traffic stream. Statistica Neerlandica, 18(3):259–265, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00513.x.

Wessels:1964:MFP

[573] J. Wessels. Multimodality in a family of probability densities, with application to a linear mixture of two normal densities. Statistica Neerlandica, 18(3):267–282, September 1964. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1964.tb00514.x.

Delver:1964:SGI

[574] A. Delver. Een simpele grafische iteratiemethode voor het schatten van een betrekkingtussen kwantitatievestochastische variabelen.. (Dutch) [A simple graphical iteration method for estimating a relationship between quantitative stochastic variables]. Statistica Neerlandica, 18(3):283–292, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00515.x.

Yen:1964:GWS

[575] Elizabeth H. Yen. Generalization of Wilcoxon statistic for the case of k samples. Statistica Neerlandica, 18(3):293–302, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00516.x.

vanderVaart:1964:A

[576] H. R. van der Vaart. Appendix. Statistica Neerlandica, 18(3):303–305, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00517.x.

Corsten:1964:CDS

[577] L C. A. Corsten. Correction to “A different solution of a problem posed by Theil and Schweitzer”. Statistica Neerlandica, 18(3):307, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00518.x. See [549].

Zaat:1964:PPM

[578] J. C. A. Zaat. Poisson-proces met interrupties, een waarschijnlijkheidsmodel voor het machinaal dunnen van bieten.. (Dutch) [Poisson process with interruptions, a probability model for mechanical beet thinning]. Statistica Neerlandica, 18(3):311–323, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.

tb00519.x. See remark [648].

Hazewinkel:1964:EBB

[579] A. Hazewinkel. Enkele beschouwingen betreffende de bijdrage van een component tot een samengestelde variabele.. (Dutch) [Some considerations regarding the contribution of a component to a composite variable]. Statistica Neerlandica, 18(3):325–339, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00520.x.

vanderVoort:1964:KSD

[580] E. M. Blok van der Voort. Kwantificering in de sociologie. (Dutch) [Quantification in sociology]. Statistica Neerlandica, 18(3):341–346, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00521.x.

Anonymous:1964:BDBc

[581] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 18(3):347–352, September 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1964.tb00522.x.

Corstius:1964:TDL

[582] H. Brandt Corstius. Taalstatistiek. (Dutch) [Language statistics]. Statistica Neerlandica, 18(4):353–367, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00523.x.

Volmuller:1964:SVD

[583] J. Volmuller. Statistiek en verkeer. (Dutch) [Statistics and traffic]. Statistica Neerlandica, 18(4):369–379, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00524.x.

vanderBurg:1964:COD

[584] A. R. van der Burg. Communicatie en organisatie. (Dutch) [communication and organization]. Statistica Neerlandica, 18(4):383–390, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00525.x.

Veen:1964:BSD

[585] B. Veen. Betere specificaties door betere communicatie. (Dutch) [Better specifications through better communication]. Statistica Neerlandica, 18 (4):391–402, December 1964. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1964.tb00526.x.

vanNoordwijk:1964:CBE

[586] J. van Noordwijk. Communication between the experimental animal and the pharmacologist. Statistica Neerlandica, 18(4):403–416, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00527.x.

deJonge:1964:CVM

[587] H. de Jonge. Communicatie van medische statistiek: de communicatie tussen medicus en statistiek. (Dutch) [Communication of medical statistics: communication between physicians and statistics]. Statistica Neerlandica, 18(4):417–424, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1964.tb00528.x.

Rumke:1964:CMS

[588] Chr. L. Rümke. Communicatie en medische statistiek. (Dutch) [Communication and medical statistics]. Statistica Neerlandica, 18(4):425– 431, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00529.x.

vanZwet:1964:CTN

[589] W. R. van Zwet. Convex transformations: a new approach to skewness and kurtosis. Statistica Neerlandica, 18(4):433–441, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00530.x.

Kamphorst:1964:CFL

[590] H. Kamphorst. De “car-following law”. (Dutch) [The “car-following law”]. Statistica Neerlandica, 18(4):443–445, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00531.x.

Gross:1964:KOV

[591] Alan J. Gross. Een kort overzicht van codes die tegen foutenuitbarsting bestand zijn. (Dutch) [A brief overview of error burst resistant codes]. Statistica Neerlandica, 18(4):447–451, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00532.x.

deBoer:1964:SMA

[592] J. de Boer. Some mathematical aspects of reliability problems. Statistica Neerlandica, 18(4):453–461, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1964.tb00533.x.

Vemtetten:1964:PVA

[593] J. B. Vemtetten. Prognose van het autopark met behulp van een “diffusion”-model. (Dutch) [Forecasting the car fleet using a “diffusion” model]. Statistica Neerlandica, 18(4):463–471, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00534.x.

vanWestendorp:1964:SMB

[594] P. H. van Westendorp. Scaling metoden bij konsumenten-ondenoek. (Dutch) [Scaling methods in consumer research]. Statistica Neerlandica, 18(4):473–477, December 1964. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1964.tb00535.x.

Jacobsen:1964:VTW

[595] L H. Jacobsen. Verkeerskunde als toegepaste wetenschap. (Dutch) [Traffic engineering as an applied science]. Statistica Neerlandica, 18(4):479– 484, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1964.tb00536.x.

deLange:1964:WWD

[596] S. J. de Lange. Wachten in het wegverkeer. (Dutch) [Waiting in road traffic]. Statistica Neerlandica, 18(4):485–493, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00537.x.

Justesen:1964:MVI

[597] S. H. Justesen. Het meten van informatie. (Dutch) [Measuring information]. Statistica Neerlandica, 18(4):495–496, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1964.
tb00538.x.

vandenBan:1964:CVN

[598] A. W. van den Ban. De communicatie van nieuwe landbouwmethoden. (Dutch) [The communication of new agricultural methods]. Statistica Neerlandica, 18(4):497–508, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1964.tb00539.x.

Anonymous:1964:BDBd

[599] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 18(4):509–512, December 1964. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1964.tb00540.x.

Cohen:1965:SCA

[600] J. W. Cohen. Some comments on: “Another note on the delay problem in crossing a traffic stream”. Statistica Neerlandica, 19(1):1– 2, March 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00462.x.

Goudriaan:1965:NAR

[601] J. Goudriaan. Naar aseptische en realistische methoden in de toegepaste statistiek. (Dutch) [Towards aseptic and realistic methods in applied statistics]. Statistica Neerlandica, 19(1):3–13, March 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1965.
tb00463.x.

Rumke:1965:TVD

[602] Chr. L. Rümke and H. de Jonge. De toetsen van Dunnett en van Steel voor net vergelijken van de uitkomsten van k proefgroepen met die van een controlegroep. (Dutch) [The Dunnett and Steel tests for comparing the results of k test groups with those of a control group]. Statistica Neerlandica, 19(1):15–24, March 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1965.tb00464.x.

tenHoopen:1965:WTP

[603] M. ten Hoopen and H. A. Reuver. On a waiting time problem in physiology. Statistica Neerlandica, 19(1):27–34, March 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1965.
tb00465.x.

Neudecker:1965:ILI

[604] H. Neudecker. Interne leveringen in de inzet-afzetanalyse. (Dutch) [Internal deliveries in the deployment-sales analysis]. Statistica Neerlandica, 19 (1):35–38, March 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00466.x.

Boot:1965:PDB

[605] J. C. G. Boot. Price determination based on quality: An application of minimax. Statistica Neerlandica, 19(1):41–53, March 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1965.
tb00467.x.

Anonymous:1965:BDBa

[606] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 19(1):55–61, March 1965. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1965.tb00468.x.

Kriens:1965:WBD

[607] J. Kriens. Wiskunde en besliskunde. (Dutch) [Mathematics and decision science]. Statistica Neerlandica, 19(2–3):65–80, June 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1965.
tb00944.x.

Alma:1965:VVT

[608] D. Alma. Het vergelijken van twee verschillende schatters voor een percentiel P van een normale verdeling. (Dutch) [Comparing two different estimators for a percentile P of a normal distribution]. Statistica Neerlandica, 19(2–3):81–91, June 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1965.tb00945.x.

Anonymous:1965:ESA

[609] Anonymous. Het Examen Statistisch Analist 1964. (Dutch) [The Statistical Analyst Exam 1964]. Statistica Neerlandica, 19(2–3):93– 103, June 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00946.x.

vonderMooren:1965:DVH

[610] A. L. von der Mooren. Dimensionering van homogeniseringsinstallaties. (Dutch) [Sizing of homogenization installations]. Statistica Neerlandica, 19(2–3):107–127, June 1965. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1965.tb00947.x.

Grunwald:1965:CTS

[611] H. Grunwald. The correlation theory for stationary stochastic processes applied to exponential smoothing. Statistica Neerlandica, 19(2–3):129– 138, June 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00948.x.

Anonymous:1965:BDBb

[612] Anonymous. Boekbesprekingens. (Dutch) [Book reviews]. Statistica Neerlandica, 19(2–3):139–162, June 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1965.tb00949.x.

Anonymous:1965:A

[613] Anonymous. Announcement. Statistica Neerlandica, 19(2–3):165, June 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00950.x.

vanderPoel:1965:ITV

[614] W. L. van der Poel. Inleiding in de techniek van elektronische rekenautomaten. (Dutch) [Introduction to the technology of electronic calculators]. Statistica Neerlandica, 19(4):167–172, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00951.x.

Wiggers:1965:SOR

[615] B. G. Wiggers. Statistiek, Operationele Research en Computer. (Dutch) [Statistics, Operational Research and Computer]. Statistica Neerlandica, 19(4):173–199, December 1965. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1965.tb00952.x.

Doornbos:1965:OIV

[616] R. Doornbos. Optimale instelling van een pakmachine met behulp van simulatie op een computer. (Dutch) [Optimal setting of a packing machine using simulation on a computer]. Statistica Neerlandica, 19(4):201– 212, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00953.x.

Ottens:1965:APQ

[617] F. Ottens. Automatic processing of quality data of rayon tire yarns and cords. Statistica Neerlandica, 19(4):213–225, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1965.
tb00954.x.

Runke:1965:EMC

[618] Chr. L. Rünke. Ervaringen met een computer in de Medische Statistiek. (Dutch) [Experiences with a computer in Medical Statistics]. Statistica Neerlandica, 19(4):227–240, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1965.tb00955.x.

Kuipers:1965:VMB

[619] H. A. Kuipers. Variantieanalyse met behulp van een computer. (Dutch) [Analysis of variance using a computer]. Statistica Neerlandica, 19(4): 241–247, December 1965. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1965.tb00956.x.

Motenaar:1965:OVO

[620] W. Motenaar. Overzicht van ontmengingsmethoden voor twee normale verdelingen. (Dutch) [Overview of unmixing methods for two normal distributions]. Statistica Neerlandica, 19(4):249–263, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00957.x.

vanderLaan:1965:MCE

[621] P. van der Laan and J. Oosterhoff. Monte Carlo estimation of the powers of the distribution-free two-sample tests of Wilcoxon, van der Waerden and Terry and comparison of these powers. Statistica Neerlandica, 19 (4):265–275, December 1965. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1965.tb00958.x.

Vos:1965:EVT

[622] J. W. E. Vos. Extrapolatie van tijdreeksen. (Dutch) [Extrapolation of time series]. Statistica Neerlandica, 19(4):277–288, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00959.x.

Heijna:1965:EFV

[623] B. Heijna. Enkele facetten van numerieke weersverwachtingen. (Dutch) [Some facets of numerical weather forecasting]. Statistica Neerlandica, 19(4):289–292, December 1965. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1965.tb00960.x.

Cramer:1965:GVC

[624] J. S. Cramer. Het gebruik van computers bij econometrisch onderzoek. (Dutch) [The use of computers in econometric research]. Statistica Neerlandica, 19(4):293–296, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1965.tb00961.x.

Dronkers:1965:EAB

[625] J. J. Dronkers. Enkele algemene beschouwingen in verband met de uitvoering van het Deltaplan. (Dutch) [Some general considerations regarding the implementation of the Delta Plan]. Statistica Neerlandica, 19 (4):297–304, December 1965. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1965.tb00962.x.

vanRees:1965:SOC

[626] J. van Rees. Een studie omtrent de cirkulatie van materieel. (Dutch) [A study on the circulation of equipment]. Statistica Neerlandica, 19(4):305– 312, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00963.x.

vanHam:1965:VCD

[627] Ir. C. J. G. J. van Ham. Voorraadbeheer en computer. (Dutch) [Inventory management and computer]. Statistica Neerlandica, 19(4):313– 322, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00964.x.

Verdooren:1965:CPD

[628] L R. Verdooren. Change-over“-proeven. (Dutch) [Change-over“ trials]. Statistica Neerlandica, 19(4):323–333, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1965.
tb00965.x.

Stol:1965:AVG

[629] Ir. Ph. Th. Stol. Een analyse van grondwaterstanden die beïnvloed worden door de getijdebeweging. (Dutch) [An analysis of groundwater levels influenced by tidal movement]. Statistica Neerlandica, 19(4):335– 343, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00966.x.

Wit:1965:EMT

[630] J. W. W. A. Wit. Ervaringen met toepassing van computer programma’s bij het sociologisch onderzoek aan het Centraal Bureau voor de Statistiek. (Dutch) [Experiences with the application of computer programs in sociological research at the Central Bureau of Statistics]. Statistica Neerlandica, 19(4):345–349, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1965.tb00967.x.

Mokken:1965:EEB

[631] R. J. Mokken. Enige ervaringen bij de ontwikkeling van een standaardtoetsingsprogramma. (Dutch) [Some experiences in the development of a standard testing program]. Statistica Neerlandica, 19(4):351–354, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1965.tb00968.x.

Dirkzwoger:1965:SVL

[632] A. Dirkzwoger. Simulatie van Leerprocessen: geprogrammeerde experimenten; een experimenteel programma. (Dutch) [Simulation of Learning Processes: programmed experiments; an experimental program]. Statistica Neerlandica, 19(4):355–374, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1965.
tb00969.x.

Anonymous:1965:BDBc

[633] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 19(4):375–395, December 1965. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1965.tb00970.x.

Hemelrijk:1966:URV

[634] J. Hemelrijk. Underlining random variables. Statistica Neerlandica, 20 (1):1–7, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00488.x.

Andrews:1966:ESS

[635] F. C. Andrews and J. R. Blum. On expectation in sequential sampling. Statistica Neerlandica, 20(1):9–18, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1966.
tb00489.x.

Beerens:1966:AIO

[636] G. A. C. Beerens and C. B. Tilanus. Alternative input-output predictions for The Netherlands, 1948–1958. Statistica Neerlandica, 20(1): 19–29, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00490.x.

vanderLaan:1966:SDF

[637] P. van der Laan. A sequential distribution-free two-sample grouped test with three possible decisions. Statistica Neerlandica, 20(1):31– 41, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00491.x.

Lootsma:1966:NPS

[638] F. A. Lootsma. Network planning with stochastic activity durations, an evaluation of PERT. Statistica Neerlandica, 20(1):43–69, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00492.x.

Rey:1966:MCP

[639] Guido Rey. A Markov chain prediction of value added and expenditure shares, Italy 1861–1956. Statistica Neerlandica, 20(1):71–86, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00493.x.

Vandewiele:1966:TGFa

[640] Georges Vandewiele and Paul de Witte. A test of goodness of fit. Statistica Neerlandica, 20(1):87–105, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1966.tb00494.x.

Vandewiele:1966:TGFb

[641] Georges Vandewiele. A test of goodness of fit for discrete distributions. Statistica Neerlandica, 20(1):107–118, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1966.
tb00495.x.

Wise:1966:GLN

[642] M. E. Wise. The geometry of log-normal and related distributions and an application to tracer-dilution curves. Statistica Neerlandica, 20(1): 119–142, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00496.x.

vanZwet:1966:BET

[643] W. R. van Zwet. Bias in estimation from type I censored samples. Statistica Neerlandica, 20(1):143–148, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1966.tb00497.x.

Anonymous:1966:BR

[644] Anonymous. Book reviews. Statistica Neerlandica, 20(1):149–154, March 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00498.x.

deLeve:1966:WBE

[645] G. de Leve. Wie beslist er eigenlijk. (Dutch) [Who actually decides?]. Statistica Neerlandica, 20(2):155–169, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1966.tb00469.x.

Begeer:1966:DMO

[646] W. Begeer. Duel met onzekerheid. (Dutch) [Duel with uncertainty]. Statistica Neerlandica, 20(2):171–188, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1966.tb00470.x.

Anonymous:1966:ESA

[647] Anonymous. Het examen Statistisch Analist 1965. (Dutch) [The Statistical Analyst exam 1965]. Statistica Neerlandica, 20(2):189–198, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00471.x.

Runnenburg:1966:OBB

[648] J. Th. Runnenburg and W. Vervaat. Opmerkingen en berekeningen betreffende “Poisson-proces met interrupties” van J. C. A. Zaat. (Dutch) [Comments and calculations regarding “Poisson process with interruptions” by J. C. A. Zaat]. Statistica Neerlandica, 20(2):199– 213, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00472.x. See [578].

Cramer:1966:PVN

[649] J. S. Cramer. Een prijsindex van nieuwe personenauto’s, 1950–1965. (Dutch) [A price index of new passenger cars, 1950–1965]. Statistica Neerlandica, 20(2):215–224, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1966.tb00473.x.

Cramer:1966:LSV

[650] J. S. Cramer. Levensduur en sloop van personenauto’s in Nederland, 1950–1964. (Dutch) [Service life and scrapping of passenger cars in Netherlands, 1950–1964]. Statistica Neerlandica, 20(2):225– 239, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00474.x.

Geerlings:1966:GBV

[651] J. W. H. Geerlings. De gemiddelde bezetting van personenauto’s in interlokaal verkeer. (Dutch) [The average occupancy of passenger cars in long-distance traffic]. Statistica Neerlandica, 20(2):241– 249, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00475.x.

vanNoerssen:1966:FBG

[652] R. F. van Noerssen. De fout bij het gebruik van een tweetal benaderingsformules. (Dutch) [The error in using two approximation formulas]. Statistica Neerlandica, 20(2):251–256, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1966.tb00476.x.

Anonymous:1966:BDBa

[653] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 20(2):257–274, June 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1966.tb00477.x.

Anonymous:1966:ERA

[654] Anonymous. Het examen O.R.-Analist 1965 (2e deel). (Dutch) [The O. R. Analyst exam 1965 (2nd part)]. Statistica Neerlandica, 20(3–4):283– 295, September 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00478.x.

Hogerzeil:1966:SMS

[655] H. H. W. Hogerzeil. Sturen met statistiek. (Dutch) [Steer with statistics]. Statistica Neerlandica, 20(3–4):297–312, September 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1966.
tb00479.x.

Tinbergen:1966:SMS

[656] J. Tinbergen. Sturen met Statistiek. (Dutch) [Steer with Statistics]. Statistica Neerlandica, 20(3–4):313–325, September 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1966.
tb00480.x.

Adriaanse:1966:TTV

[657] R. P. Adriaanse. Tolerantiegebieden en toetsen voor Kwantielen. (Dutch) [Tolerance areas and keys for Quantiles]. Statistica Neerlandica, 20(3– 4):327–334, September 1966. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1966.tb00481.x.

Romp:1966:SMS

[658] Th. P. J. Romp. Sturen met Statistiek in het Bedrijf. (Dutch) [Steer with statistics in the company]. Statistica Neerlandica, 20(3–4):335– 343, September 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00482.x.

Enters:1966:LVI

[659] J. H. Enters. Levensduuronderzoek van industriële produkten. (Dutch) [Lifespan research of industrial products]. Statistica Neerlandica, 20(3– 4):345–355, September 1966. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1966.tb00483.x.

Verstegt:1966:SPS

[660] J. Ch. W. Verstegt. Statistiek en planning in de sociale sector. (Dutch) [Statistics and planning in the social sector]. Statistica Neerlandica, 20 (3–4):357–368, September 1966. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1966.tb00484.x.

Hamming:1966:AMD

[661] G. Hamming. Automatische marktregulering. (Dutch) [Automatic market regulation]. Statistica Neerlandica, 20(3–4):369–384, September 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1966.tb00485.x.

vanderPutten:1966:BVR

[662] A. van der Putten. Bedrijfsmodellen voor rentabiliteitsprognoses. (Dutch) [Business models for profitability forecasting]. Statistica Neerlandica, 20(3–4):385–398, September 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1966.tb00486.x.

Anonymous:1966:BDBb

[663] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 20(3–4):399–414, September 1966. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1966.tb00487.x.

Somermeyer:1967:PSC

[664] W. H. Somermeyer and R. Teekens. Principles of statistical classification. Statistica Neerlandica, 21(1):1–10, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1967.
tb00541.x.

Meisner:1967:EDD

[665] J. Meisner. Estimation of the distribution of diameters of spherical particles from a given grouped distribution of diameters of observed circles formed by a plane section. Statistica Neerlandica, 21(1):11– 30, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00542.x.

Ajgaonkar:1967:HTC

[666] S. G. Prabhu Ajgaonkar. On Horvitz and Thompson’s T_{1} class of linear estimators for Midzuno’s (Ikeda’s generalised) sampling procedure. Statistica Neerlandica, 21(1):31–38, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1967.tb00543.x.

Wessels:1967:SP

[667] J. Wessels. Stochastic programming. Statistica Neerlandica, 21(1):39– 53, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00544.x.

vanderLaan:1967:EDP

[668] P. van der Laan and J. Oosterhoff. Experimental determination of the power functions of the two-sample rank tests of Wilcoxon, van der Waerden and Terry by Monte Carlo techniques: I. Normal parent distributions. Statistica Neerlandica, 21(1):55–68, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1967.
tb00545.x.

Sarndal:1967:DPE

[669] Carl-Erik Särndal. Distribution pairs and equi-probable sequences. Statistica Neerlandica, 21(1):69–79, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1967.tb00546.x.

Peritz:1967:SNT

[670] Eric Peritz. On some nonparametric tests of dispersion. Statistica Neerlandica, 21(1):81–89, March 1967. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1967.tb00547.x.

vanSoest:1967:SER

[671] J. van Soest. Some experimental results concerning tests of normality. Statistica Neerlandica, 21(1):91–97, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1967.
tb00548.x.

vanderVaart:1967:NDN

[672] H. R. van der Vaart. A note on the derivation of the noncentral chi-square density function. Statistica Neerlandica, 21(1):99– 100, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00549.x. See correction [699].

Neudecker:1967:MPO

[673] H. Neudecker. On matrix procedures for optimizing differentiable scalar functions of matrices. Statistica Neerlandica, 21(1):101–107, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00550.x.

Anonymous:1967:BR

[674] Anonymous. Book reviews. Statistica Neerlandica, 21(1):109–113, March 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00551.x.

vanZwet:1967:HOD

[675] W. R. van Zwet. Het hachelijk oordeel. (Dutch) [The handsome judgment]. Statistica Neerlandica, 21(2):117–130, June 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1967.
tb00552.x.

Prins:1967:DBS

[676] H. J. Prins. The distribution of the brightness of stars. A Poissonprocess with “time”-dependent parameter. Statistica Neerlandica, 21(2): 131–149, June 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00553.x.

Flapper:1967:TVN

[677] P. Flapper. Transformatie van niet-normale verdelingen. (Dutch) [Transformation of non-normal distributions]. Statistica Neerlandica, 21(2): 151–155, June 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00554.x.

Mijnheer:1967:OVW

[678] J. L. Mijnheer. Oneindige verwachting en de wet van de grote aantallen. (Dutch) [Infinite expectation and the law of large numbers]. Statistica Neerlandica, 21(2):157–159, June 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1967.tb00555.x.

vandenBeld:1967:VVV

[679] C. A. van den Beld. Vergelijking van voorspelling en realisatie als bijdrage tot economisch bestuur. (Dutch) [Comparison of prediction and realization as a contribution to economic governance]. Statistica Neerlandica, 21 (2):161–174, June 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00556.x.

vdKnaap:1967:BVP

[680] J. v. d. Knaap. Bepaling van de personeelssterkte op kostenbasis. (Dutch) [Determination of staff strength on a cost basis]. Statistica Neerlandica, 21(2):175–198, June 1967. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1967.tb00557.x.

Timman:1967:PMP

[681] R. Timman. Principles of mathematical programming. Statistica Neerlandica, 21(2):199–209, June 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1967.tb00558.x.

Anonymous:1967:BDBa

[682] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 21(2):211–222, June 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1967.tb00559.x.

Hemelrijk:1967:HNC

[683] J. Hemelrijk. The hypergeometric, the normal and chi-squared. Statistica Neerlandica, 21(3–4):225–228, September 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1967.
tb00560.x.

Steutel:1967:RDI

[684] F. W. Steutel. Random division of an interval. Statistica Neerlandica, 21 (3–4):231–244, September 1967. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1967.tb00561.x.

Vervaat:1967:UGP

[685] W. Vervaat. Uitgedund en gerekt Poissonproces. (Dutch) [Thinned and stretched Poisson process]. Statistica Neerlandica, 21(3–4):245–268, September 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1967.tb00562.x.

Fabius:1967:VVM

[686] J. Fabius. Voorwaardelijke verwachtingen en martingalen. (Dutch) [Conditional expectations and martingales]. Statistica Neerlandica, 21(3– 4):269–285, September 1967. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1967.tb00563.x.

Fabius:1967:PBV

[687] J. Fabius. De Poisson benadering voor de binomiale verdeling. (Dutch) [The Poisson approximation for the binomial distribution]. Statistica Neerlandica, 21(3–4):287–289, September 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1967.
tb00564.x.

Boas:1967:NEC

[688] J. Boas. A note on the estimation of the covariance between two random variables using extra information on the separate variables. Statistica Neerlandica, 21(3–4):291–292, September 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1967.
tb00565.x.

Anonymous:1967:BDBb

[689] Anonymous. Boekbesprekingen. (Dutch) [Book reviews]. Statistica Neerlandica, 21(3–4):293–307, September 1967. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1967.tb00566.x.

Kendall:1968:CAU

[690] M. G. Kendall. Certainty about Uncertainty. Statistica Neerlandica, 22 (1):1–12, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00612.x.

Hemelrijk:1968:BLD

[691] J. Hemelrijk. Back to the Laplace definition. Statistica Neerlandica, 22 (1):13–21, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00613.x.

VanZwet:1968:CPF

[692] W. R. Van Zwet. On convexity preserving families of probability distributions. Statistica Neerlandica, 22(1):23–32, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb00614.x.

Kubik:1968:ELS

[693] K. Kubik. On the efficiency of least squares estimators in non-linear models. Statistica Neerlandica, 22(1):33–36, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb00615.x.

Mijnheer:1968:CSA

[694] J. L. Mijnheer. The conduct of the sample average when the first moment is infinite. Statistica Neerlandica, 22(1):37–41, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb00616.x.

VanDerGrinten:1968:UMC

[695] P. M. E. M. Van Der Grinten. Uncertainty in measurement and control. Statistica Neerlandica, 22(1):43–63, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.

tb00617.x. See correction [714].

Elandt-Johnson:1968:OAS

[696] Regina C. Elandt-Johnson. Optimal allocation in J-stage sampling with stratification from finite population. Statistica Neerlandica, 22(1):65– 68, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00618.x.

Neudecker:1968:KMP

[697] H. Neudecker. The Kronecker matrix product and some of its applications in econometrics. Statistica Neerlandica, 22(1):69–82, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00619.x.

Anonymous:1968:BR

[698] Anonymous. Book reviews. Statistica Neerlandica, 22(1):83–89, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00620.x.

Anonymous:1968:CSN

[699] Anonymous. Correction Statistica Neerlandica 21 (1967) p. 89. Statistica Neerlandica, 22(1):90, March 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00621.x. See [672].

Corsten:1968:FDF

[700] L. C. A. Corsten. Over Factoranalyse. (Dutch) [On factor analysis]. Statistica Neerlandica, 22(2):91–96, June 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1968.tb01358.x.

Mijnheer:1968:SUH

[701] J. L. Mijnheer. Steekproeven uit de halve Cauchy verdeling. (Dutch) [Samples from the half-Cauchy distribution]. Statistica Neerlandica, 22 (2):97–101, June 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1968.tb01359.x.

Anthonisse:1968:IDL

[702] Jac. M. Anthonisse. Een indelingsprobleem. (Dutch) [A layout problem]. Statistica Neerlandica, 22(2):103–118, June 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1968.

tb01360.x. See correction [713].

Cramer:1968:AAN

[703] J. S. Cramer. Autoverkopen en autopark in Nederland, 1950–1970. (Dutch) [Car sales and car fleet in Netherlands, 1950–1970]. Statistica Neerlandica, 22(2):119–131, June 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1968.tb01361.x.

Tilanus:1968:RGS

[704] C. B. Tilanus. Het rekenkundig gemiddelde stijgingspercentage — een ondoelmatige groei-indicator. (Dutch) [The arithmetic average rate of increase — an ineffective growth indicator]. Statistica Neerlandica, 22(2): 133–139, June 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1968.tb01362.x.

Anonymous:1968:BDBa

[705] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 22(2):141–149, June 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1968.tb01363.x.

Kriens:1968:VVS

[706] J. Kriens. Het verdelen van steekproeven over subpopulaties bij accountantscontroles. (Dutch) [Distributing samples across subpopulations in audits]. Statistica Neerlandica, 22(3):151–157, September 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00628.x.

Adriaanse:1968:SEV

[707] R. P. Adriaanse and P. Van Der Laan. Statistische eigenschappen van een vlambeveiliging met behulp van een gasontladingsbuis welke gevoelig is voor ultraviolet licht. (Dutch) [Statistical properties of a flame arrester using a gas discharge tube that is sensitive to ultraviolet light]. Statistica Neerlandica, 22(3):159–172, September 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1960.tb00629.x.

VanVeen:1968:AMZ

[708] S. C. Van Veen. Abraham de Moivre en zijn “Doctrine of Chances”. (Dutch) [Abraham de Moivre and his “Doctrine of Chances”]. Statistica Neerlandica, 22(3):173–177, September 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1960.tb00630.x.

Kleijnen:1968:TVI

[709] J. P. C. Kleijnen. Een toepassing van “importance sampling”. (Dutch) [An application of “importance sampling”]. Statistica Neerlandica, 22 (3):179–198, September 1968. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1960.tb00631.x.

Doornbos:1968:AVK

[710] R. Doornbos and J. H. C. Lisman. Afleiding van kwartaalcijfers uit jaartotalen. (Dutch) [Derivation of quarterly figures from annual totals]. Statistica Neerlandica, 22(3):199–205, September 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1960.
tb00632.x.

Tilanus:1968:TNV

[711] C. B. Tilanus. Tabel en nomogram voor het aflezen van jaarlijkse groeipercentages van grafieken op semi-logaritmisch papier. (Dutch) [Table and nomogram for reading annual growth rates from graphs on semi-logarithmic paper]. Statistica Neerlandica, 22(3):207–212, September 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00633.x.

Erkelens:1968:BVL

[712] J. Erkelens. Een berekeningsmethode voor de logistische curve. (Dutch) [A calculation method for the logistic curve]. Statistica Neerlandica, 22 (3):213–217, September 1968. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1960.tb00634.x.

Anthonisse:1968:CBI

[713] Jac. M. Anthonisse. Correctie bij “Een indelingsprobleem”. (Dutch) [Correction to “A classification problem”]. Statistica Neerlandica, 22(3): 219, September 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00636.x. See [702].

VanDerGrinten:1968:CUM

[714] P. M. E. M. Van Der Grinten. Correction to “Uncertainty in measurement and control”. Statistica Neerlandica, 22(3):219, September 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00635.x. See [695].

Anonymous:1968:BDBb

[715] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 22(3):221–233, September 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00637.x.

Anonymous:1968:CDC

[716] Anonymous. Correctie. (Dutch) [Correction]. Statistica Neerlandica, 22(3):233, September 1968. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1960.tb00638.x.

VanSoest:1968:AVV

[717] J. Van Soest. Afschatting van verdelingen van kwadratische vormen met toepassingen op aanpassingstoetsen voor normaliteit en exponentialiteit. (Dutch) [Estimation of distributions of quadratic forms with applications to adjustment tests for normality and exponentiality]. Statistica Neerlandica, 22(4):235–240, December 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00639.x.

VanFrankenhuysen:1968:TVN

[718] J. H. Van Frankenhuysen. Toets voor normaliteit van Shapiro en Wilk. (Dutch) [Test for normality of Shapiro and Wilk]. Statistica Neerlandica, 22(4):241–248, December 1968. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1960.tb00640.x.

Lombaers:1968:ESA

[719] H. J. M. Lombaers. Enige statistische aspecten van simulatie. (Dutch) [Some statistical aspects of simulation]. Statistica Neerlandica, 22(4): 249–255, December 1968. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1960.tb00641.x.

Abrahamse:1968:SME

[720] A. P. J. Abrahamse and J. Koerts. Simultaneous minimization of the error probabilities in a statistical test: an economic example. Statistica Neerlandica, 22(4):257–266, December 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1960.tb00642.x.

Heesterman:1968:FP

[721] A. R. G. Heesterman. Forecasting the past. Statistica Neerlandica, 22 (4):267–271, December 1968. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1960.tb00643.x.

Drion:1968:SNS

[722] E. F. Drion. Steriel of niet steriel, dat is de vraag. (Dutch) [Sterile or not sterile, that is the question]. Statistica Neerlandica, 22(4):273– 278, December 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1960.tb00644.x.

Anonymous:1968:BDBc

[723] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 22(4):279–289, December 1968. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1960.tb00645.x.

Runnenburg:1969:LTS

[724] J. Th. Runnenburg. Limit theorems for stochastic processes occurring in studies of the light-sensitivity of the human eye. Statistica Neerlandica, 23(1):1–17, March 1969. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1969.tb00070.x.

Molenaar:1969:HPP

[725] W. Molenaar. How to poison Poisson (when approximating binomial tails). Statistica Neerlandica, 23(1):19–40, March 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1969.
tb00071.x.

vanSoest:1969:SGF

[726] J. van Soest. Some goodness of fit tests for exponential distributions. Statistica Neerlandica, 23(1):41–51, March 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1969.tb00072.x.

Verdooren:1969:RAM

[727] L. R. Verdooren. Representation of ANOVA models with vectors and vector spaces. Statistica Neerlandica, 23(1):53–65, March 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1969.
tb00073.x.

Runnenburg:1969:AIL

[728] J. Th. Runnenburg and W. Verváat. Asymptotical independence of the lengths of subintervals of a randomly partitioned interval; a sample from S. Ikeda’s work. Statistica Neerlandica, 23(1):67–77, March 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00074.x.

Vervaat:1969:UBD

[729] W. Vervaat. Upper bounds for the distance in total variation between the binomial or negative binomial and the Poisson distribution. Statistica Neerlandica, 23(1):79–86, March 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1969.tb00075.x.

vanRootselaar:1969:BBL

[730] B. van Rootselaar and J. Hemelrijk. Back to “Back to the Laplace definition”. Statistica Neerlandica, 23(1):87–89, March 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1969.
tb00076.x.

Anonymous:1969:BR

[731] Anonymous. Book reviews. Statistica Neerlandica, 23(1):91–95, March 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00077.x.

vanMontfort:1969:IVV

[732] M. A. J. van Montfort. Inleiding over de verdeling van extremen. (Dutch) [Introduction to the distribution of extremes]. Statistica Neerlandica, 23 (2):97–111, June 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00078.x.

vanMontfort:1969:TOO

[733] M. A. J. van Montfort. Toets op ongeassocieerdheid in een 2 × 2-tabel bij trekking zonder teruglegging. (Dutch) [Test for non-association in a 2×2 table for a draw without replacement]. Statistica Neerlandica, 23(2): 113–114, June 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00079.x.

Heiden:1969:HOO

[734] J. A. V. D. Heiden. “How to obtain the optimum from a linear programming problem, when the calculations have to be carried out in a decentralized way”. Statistica Neerlandica, 23(2):115– 149, June 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00080.x.

Derksen:1969:ASM

[735] J. B. D. Derksen. Adjustments for seasonal movements with the Census Method II; recent experiences. Statistica Neerlandica, 23(2):151– 159, June 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00081.x.

Massizzo:1969:OVI

[736] A. I. V. Massizzo, W. Kok, and J. H. C. Lisman. De ontwikkeling van de inkomensongelijkheid gemeten volgens informatietheoretische maatstaven. (Dutch) [The development of income inequality measured according to information-theoretic standards]. Statistica Neerlandica, 23(2): 161–171, June 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00082.x.

Anonymous:1969:BDBa

[737] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 23(2):172–178, June 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1969.tb00083.x.

Hamaker:1969:WKD

[738] H. C. Hamaker. De Wet en de kansspelen. (Dutch) [The Law and games of chance]. Statistica Neerlandica, 23(3):179–191, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00084.x.

Hemelrijk:1969:KBD

[739] J. Hemelrijk. Kans en behendigheid. (Dutch) [Chance and agility]. Statistica Neerlandica, 23(3):193–202, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1969.
tb00085.x.

Hamaker:1969:NWK

[740] H. C. Hamaker. Nogmaals de Wet en de Kansspelen; commentaar op Hemelrijks beschouwingen. (Dutch) [Once again the Law and the Games of Chance; commentary on Hemelrijks reflections]. Statistica Neerlandica, 23(3):203–207, September 1969. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1969.tb00086.x.

Cohen:1969:DBD

[741] Ir. J. W. Cohen. Discussie bijdrage. (Dutch) [Discussion contribution]. Statistica Neerlandica, 23(3):209–212, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1969.
tb00087.x.

Sittig:1969:RBO

[742] J. Sittig. Rechtszekerheid bij onzekerheid. (Dutch) [Legal certainty in the face of uncertainty]. Statistica Neerlandica, 23(3):213–225, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00088.x.

deHaan:1969:SRC

[743] L. de Haan and J. Th. Runnenburg. Some remarks concerning the quotient of sample median and sample range for a sample of size 2n + 1 from a normal distribution. Statistica Neerlandica, 23(3):226– 234, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00089.x.

Bouma:1969:PAS

[744] N. Bouma and A. Vehmeyer. Percentiles and approximations of the sample median over the sample range for samples of size 3, 5, 7 and 9 from a standard normal distribution. Statistica Neerlandica, 23(3):235– 239, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00090.x.

Molenaar:1969:ARH

[745] W. Molenaar. Additional remark on “How to poison Poisson (when approximating binomial tails)”. Statistica Neerlandica, 23(3):241, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00091.x.

deBoer:1969:TVK

[746] J. de Boer. Toepassing van Kosten’s lotingsas in simulaties. (Dutch) [Application of Kosten’s draw axis in simulations]. Statistica Neerlandica, 23(3):243–248, September 1969. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1969.tb00092.x.

deKroon:1969:SBC

[747] J. El. M. de Kroon. Suggestie betreffende het combineren van overschrijdingskansen. (Dutch) [Suggestion regarding combining exceedance probabilities]. Statistica Neerlandica, 23(3):249–251, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00093.x.

Anonymous:1969:BDBb

[748] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 23(3):253–262, September 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1969.tb00094.x.

Willemze:1969:BAV

[749] F. G. Willemze and J. H. C. Lisman. Bij het afscheid van de hoofdredacteur Prof. Dr. J. Hemelrijk. (Dutch) [At the farewell of the Editorin-Chief Prof. Dr. J. Hemelrijk]. Statistica Neerlandica, 23(4):263–264, December 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00095.x.

Lisman:1969:MDV

[750] J. H. C. Lisman. Het “Muzikale dobbelspel” van Mozart. (Dutch) [The “Musical dice game” by Mozart]. Statistica Neerlandica, 23(4):265–267, December 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00096.x.

vanZwet:1969:SZW

[751] W. R. van Zwet. Stemmingen zonder winnaar. (Dutch) [Voting without a winner]. Statistica Neerlandica, 23(4):269–276, December 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00097.x.

Hermans:1969:CKT

[752] J. Hermans. De chi-kwadraat-toets voor aanpassing van continue verdelingen. (Dutch) [The chi-square test for fitting continuous distributions]. Statistica Neerlandica, 23(4):277–285, December 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1969.
tb00098.x.

Voorn:1969:SVP

[753] W. J. Voorn. Schatting van een populatiegemiddelde bij ongelijke trekkingskansen. (Dutch) [Estimation of a population average with unequal draw probabilities]. Statistica Neerlandica, 23(4):287–294, December 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00099.x.

Sander:1969:ISD

[754] P. C. Sander. Invariantie in de schattingstheorie. (Dutch) [Invariance in estimation theory]. Statistica Neerlandica, 23(4):295–300, December 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1969.tb00100.x.

Derksen:1969:IMM

[755] J. B. D. Derksen. Indexcijfers met een meerjarige basisperiode. (Dutch) [Index figures with a multi-year base period]. Statistica Neerlandica, 23(4):301–303, December 1969. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1969.tb00101.x.

Verheyen:1969:IOD

[756] P. A. Verheyen. Investeren en onzekerheid. (Dutch) [Investing and uncertainty]. Statistica Neerlandica, 23(4):305–320, December 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1969.
tb00102.x.

Anonymous:1969:BDBc

[757] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 23(4):321–331, December 1969. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1969.tb00103.x.

Boschloo:1970:RCL

[758] R. D. Boschloo. Raised conditional level of significance for the 2 × 2-table when testing the equality of two probabilities. (Dutch) [Raised conditional level of significance for the 2 × 2-table when testing the equality of two probabilities]. Statistica Neerlandica, 24(1): 1–9, March 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00104.x.

Lancaster:1970:NF

[759] T. Lancaster. Note on forecasting. Statistica Neerlandica, 24(1):37– 39, March 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00105.x.

vanderGiessen:1970:SNL

[760] A. A. van der Giessen. Solving non-linear systems by computer; a new method. Statistica Neerlandica, 24(1):41–50, March 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1970.
tb00106.x.

Sengupta:1970:FPA

[761] J. K. Sengupta and B. C. Sanyal. A fractile programming approach with extreme sample estimates of parameters. Statistica Neerlandica, 24(1): 51–59, March 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00107.x.

Burley:1970:SRA

[762] S. P. Burley. Short-run auction price adjustments. Statistica Neerlandica, 24(2):61–70, June 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00108.x.

Anthonisse:1970:ISLa

[763] Jac. M. Anthonisse. An input system for linear programming problems. Statistica Neerlandica, 24(2):71–81, June 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1970.tb00109.x.

Louter:1970:KTN

[764] A. S. Louter and J. Koerts. On the Kuiper test for normality with mean and variance unknown. Statistica Neerlandica, 24(2):83– 87, June 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00110.x.

Maggu:1970:CTQ

[765] Parkash Lal Maggu. On certain type of queues in series. Statistica Neerlandica, 24(2):89–97, June 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1970.tb00111.x.

Anonymous:1970:BR

[766] Anonymous. Book reviews. Statistica Neerlandica, 24(2):99–108, June 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00112.x.

Gupta:1970:NMW

[767] Y. P. Gupta. A note on the moments of the Wald’s estimator. Statistica Neerlandica, 24(3):109–123, September 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1970.tb00113.x.

Terhal:1970:SVB

[768] P. H. J. J. Terhal. Het schatten van het binnenwerk van een matrix bij gegeven randtotalen. (Dutch) [Estimating the interior of a matrix given marginal totals]. Statistica Neerlandica, 24(3):125–126, September 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00114.x.

Heesterman:1970:AER

[769] A. R. G. Heesterman. On applied econometric research. Statistica Neerlandica, 24(3):127–131, September 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1970.tb00115.x.

Kalmijn:1970:CIB

[770] W. M. Kalmijn. Confidence intervals for batch properties based on both sampling and analytical variance. Statistica Neerlandica, 24(3):133–141, September 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00116.x.

Anthonisse:1970:ISLb

[771] Jac. M. Anthonisse. An input system for linear programming problems *. Statistica Neerlandica, 24(3):143–153, September 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1970.
tb00117.x.

Holtzer:1970:BDB

[772] J. P. A. Holtzer. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 24(3):155–160, September 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1970.tb00118.x.

Anonymous:1970:SDS

[773] Anonymous. De statistische dag STAT25. (Dutch) [The statistical day STAT25]. Statistica Neerlandica, 24(4):161–162, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00119.x.

Enters:1970:HUL

[774] J. H. Enters. Herinneringen uit het losse geheugen. (Dutch) [Memories from loose memory]. Statistica Neerlandica, 24(4):163–165, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00120.x.

Anonymous:1970:EVZ

[775] Anonymous. Eretoespraak van Zijne Excellentie Dr. G. H. Veringa. (Dutch) [Honorary speech by His Excellency Dr. G. H. Veringa]. Statistica Neerlandica, 24(4):167–168, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1970.
tb00121.x.

vanEttinger:1970:SWV

[776] J. van Ettinger. Statistiek in de wereld van nu. (Dutch) [Statistics in today’s world]. Statistica Neerlandica, 24(4):169–181, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00122.x.

Freudenthal:1970:EVB

[777] H. Freudenthal. De ezel van Buridanus. (Dutch) [The donkey of Buridanus]. Statistica Neerlandica, 24(4):183–193, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1970.
tb00123.x.

Meinardi:1970:OVD

[778] J. J. Meinardi. Oorlog en vrede. (Dutch) [War and peace]. Statistica Neerlandica, 24(4):195–208, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1970.tb00124.x.

deJonge:1970:GZP

[779] H. de Jonge. Gezond — ziek: Problemen bij het onderzoek van nieuwe geneesmiddelen. (Dutch) [Healthy — sick: Problems in the research of new drugs]. Statistica Neerlandica, 24(4):209–217, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00125.x.

Somermeijer:1970:ARO

[780] W. H. Somermeijer. Arm — en — rijk oftewel: leren — de dommekracht onder de inkomenslift. (Dutch) [Poor — and — rich in other words: learning — the dumbass under the income elevator]. Statistica Neerlandica, 24(4):219–236, December 1970. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1970.tb00126.x.

Anonymous:1970:VDP

[781] Anonymous. Van Dantzig-prijs. (Dutch) [Van Dantzig prize]. Statistica Neerlandica, 24(4):237–240, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1970.tb00127.x.

Willemze:1970:BVP

[782] F. G. Willemze. Benoeming van Prof. Dr. J. Hemelrijk tot erelid van de vereniging. (Dutch) [Appointment of Prof. Dr. J. Hemelrijk as an honorary member of the association]. Statistica Neerlandica, 24(4):241– 244, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1970.tb00128.x.

Hemelrijk:1970:TWD

[783] J. Hemelrijk. Toeval en ‘wetmatigheid’. (Dutch) [Coincidence and ’lawfulness’]. Statistica Neerlandica, 24(4):245–253, December 1970. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1970.
tb00129.x.

Schaafsma:1971:NPT

[784] Willem Schaafsma. The Neyman–Pearson theory for testing statistical hypotheses. Statistica Neerlandica, 25(1):1–27, March 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1971.
tb00130.x.

Tijms:1971:OIP

[785] H. C. Tijms. The optimality of (s, S) inventory policies in the infinite period model. Statistica Neerlandica, 25(1):29–43, March 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1971.
tb00131.x.

Misra:1971:ECM

[786] P. N. Misra and Y. P. Gupta. Estimation of the covariance matrix of structural disturbances in a complete equation system. Statistica Neerlandica, 25(1):45–55, March 1971. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1971.tb00132.x.

Hopkins:1971:GCP

[787] Carl E. Hopkins and Alan J. Gross. A generalization of Cochran’s procedure for the combining of r × c contingency tables. Statistica Neerlandica, 25(1):57–62, March 1971. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1971.tb00133.x.

Leeflang:1971:PER

[788] P. S. H. Leeflang and B. M. S. van Praag. A procedure to estimate relative powers in binary contacts and an application to Dutch Football League results. Statistica Neerlandica, 25(1):63–84, March 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00134.x.

Anonymous:1971:BDBa

[789] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 25(1):85–87, March 1971. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1971.tb00135.x.

Lisman:1971:ICP

[790] J. H. C. Lisman. The information content of patterns: a tentative investigation on the statistical evaluation of mosaics. Statistica Neerlandica, 25 (2):89–100, June 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00136.x.

Vandaele:1971:RMS

[791] Walter H. Vandaele and S. R. Chowdhury. A revised method of scoring. Statistica Neerlandica, 25(2):101–112, June 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1971.
tb00137.x.

Srinivasan:1971:TN

[792] R. Srinivasan. A test for normality. Statistica Neerlandica, 25(2):113– 115, June 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00138.x.

vanDaal:1971:FOV

[793] J. van Daal. De frequentistische opbouw van de waarschijnlijkheidsrekening volgens R. von Mises. (Dutch) [The frequentist structure of the probability theory according to R. von Mises]. Statistica Neerlandica, 25 (2):117–128, June 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00139.x.

Anonymous:1971:BDBb

[794] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 25(2):129–141, June 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1971.tb00140.x.

Gupta:1971:NMC

[795] R. P. Gupta and D. G. Kabe. A note on the moments of a certain GCL estimator. Statistica Neerlandica, 25(3):143–145, September 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00141.x.

Pierce:1971:RTA

[796] David A. Pierce. On residuals and their autocorrelations in fitted time series models. Statistica Neerlandica, 25(3):147–151, September 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00142.x.

Srinivasan:1971:KTN

[797] R. Srinivasan. On the Kuiper test for normality with mean and variance unknown. Statistica Neerlandica, 25(3):153–157, September 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00143.x.

Wise:1971:SPC

[798] M. E. Wise. Skew probability curves with negative powers of time and related to random walks in series. Statistica Neerlandica, 25(3):159– 180, September 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00144.x.

Anonymous:1971:BDBc

[799] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 25(3):181–189, September 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1971.tb00145.x.

Guenther:1971:UBT

[800] William C. Guenther. On the use of best tests to obtain best βcontent tolerance intervals. Statistica Neerlandica, 25(4):191–202, December 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1971.tb00146.x.

Neeleman:1971:ADC

[801] D. Neeleman. An alternative derivation of the k-class estimators. Statistica Neerlandica, 25(4):203–210, December 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1971.
tb00147.x.

Tijms:1971:SRD

[802] H. C. Tijms. Some results for the dynamic (s, S) inventory model *. Statistica Neerlandica, 25(4):211–226, December 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1971.
tb00148.x.

Anonymous:1971:BDBd

[803] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 25(4):227–231, December 1971. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1971.tb00149.x.

vanderGenugten:1972:DRV

[804] B. B. van der Genugten. The distribution of random variables reduced modulo a. Statistica Neerlandica, 26(1):1–13, March 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00150.x.

deJonge:1972:NDS

[805] C. de Jonge and M. A. J. van Montfort. The null distribution of Spearman’s \underline{S} when n = 12. Statistica Neerlandica, 26(1):15–17, March 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00151.x.

Lisman:1972:NGM

[806] J. H. C. Lisman and M. C. A. van Zuylen. Note on the generation of most probable frequency distributions. Statistica Neerlandica, 26(1): 19–23, March 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00152.x.

Raghunandanan:1972:SEP

[807] K. Raghunandanan and R. Srinivasan. Simplified estimation of parameters in a Cauchy distribution. Statistica Neerlandica, 26(1):25– 27, March 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00153.x.

Stol:1972:PDT

[808] Ph. Th. Stol. Probability distributions of two-component soil samples. Statistica Neerlandica, 26(1):29–44, March 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00154.x.

Scheffer:1972:BR

[809] C. L. Scheffer and M. Keuls. Book review. Statistica Neerlandica, 26 (1):45, March 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00155.x.

Avadhani:1972:EST

[810] M. S. Avadhani and B. V. Sukhatme. Estimation in sampling on two successive occasions. Statistica Neerlandica, 26(2):47–54, June 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00156.x.

Gupta:1972:RES

[811] R. P. Gupta. Reliability estimation of a system comprised of k elements from the same truncated exponential model. Statistica Neerlandica, 26 (2):55–60, June 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00157.x.

vanLennep:1972:SWT

[812] A. W. van Lennep. Een symmetrische Walsh transformatie. (Dutch) [A symmetric Walsh transformation]. Statistica Neerlandica, 26(2):61– 68, June 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00158.x.

Wouters:1972:APE

[813] J. Wouters and M. C. A. van Zuylen. An application of Pitman-efficiency to acceptance sampling procedures. Statistica Neerlandica, 26(2):69– 83, June 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00159.x.

Anonymous:1972:ODAa

[814] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 26(2):85, June 1972. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1972.tb00160.x.

Anonymous:1972:BDBa

[815] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 26(2):87–96, June 1972. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1972.tb00161.x.

Corsten:1972:F

[816] L. C. A. Corsten, R. Doornbos, and J. Wessels. Foreword. Statistica Neerlandica, 26(3):1, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00162.x.

Sieben:1972:THC

[817] J. W. Sieben and F. G. Willemze. Tribute to H. C. Hamaker. Statistica Neerlandica, 26(3):3–7, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00163.x.

Anonymous:1972:PHC

[818] Anonymous. Publications of H. C. Hamaker. Statistica Neerlandica, 26(3):9–13, September 1972. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1972.tb00164.x.

Bosch:1972:UMA

[819] A. J. Bosch. From univariate- to multivariate analysis. Statistica Neerlandica, 26(3):15–20, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00165.x.

Keen:1972:LDG

[820] J. Keen. A life distribution for great tits (Parus major). Statistica Neerlandica, 26(3):21–23, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00166.x.

Mothes:1972:PSE

[821] Jean Mothes. Probabilité subjective et efficacité industrielle, (French) [Subjective probability and industrial efficiency]. Statistica Neerlandica, 26(3):25–28, September 1972. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1972.tb00167.x.

Olmstead:1972:GHL

[822] Paul S. Olmstead. Grouping of high (or low) values in observed data. Statistica Neerlandica, 26(3):29–36, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00168.x.

Terry:1972:ASP

[823] Milton E. Terry. The applied statistician as a professional. Statistica Neerlandica, 26(3):37–41, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1972.tb00169.x.

Stange:1972:ELR

[824] K. Stange. Ueber die einfache lineare Regression (bei positiver Einflussgrösse) mit einer Nebenbedingung. (German) [On simple linear regression (with a positive influencing factor) with a secondary condition]. Statistica Neerlandica, 26(3):43–54, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00170.x.

vanYzeren:1972:BDI

[825] J. van Yzeren. A bivariate distribution with instructive properties as to normality, correlation and dependence. Statistica Neerlandica, 26(3):55– 56, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00171.x.

Adam:1972:EIG

[826] Adolf Adam. Von der Erhaltung der Information. (German) [On the preservation of information]. Statistica Neerlandica, 26(3):57–60, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00172.x.

Corsten:1972:MET

[827] L. C. A. Corsten and A. C. van Eijnsbergen. Multiplicative effects in two-way analysis of variance. Statistica Neerlandica, 26(3):61–68, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00173.x.

Gnanadesikan:1972:MES

[828] R. Gnanadesikan. Methods for evaluating similarity of marginal distributions. Statistica Neerlandica, 26(3):69–78, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00174.x.

Levenbach:1972:RAE

[829] G. J. Levenbach. Residuals analysis — an example. Statistica Neerlandica, 26(3):79–83, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00175.x.

Hogben:1972:OIC

[830] David Hogben. OMNITAB II: a computing system for statistical problem solving using Dutch rather than FORTRAN or ALGOL. Statistica Neerlandica, 26(3):85–99, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00176.x.

Deming:1972:STI

[831] W. Edwards Deming. Some theory on the influence of the inspector and environmental conditions. Statistica Neerlandica, 26(3):101–105, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00177.x.

Hansen:1972:E

[832] Morris H. Hansen and W. Edwards Deming. An example. Statistica Neerlandica, 26(3):107–112, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00178.x.

Bicking:1972:AST

[833] Charles A. Bicking. The application of statistical theory to technology. Statistica Neerlandica, 26(3):113–117, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00179.x.

Mandel:1972:PCA

[834] John Mandel. Principal components, analysis of variance and data structure. Statistica Neerlandica, 26(3):119–129, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00180.x.

Hald:1972:SRS

[835] A. Hald. Some remarks on the sequential probability ratio test. Statistica Neerlandica, 26(3):131–135, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1972.tb00181.x.

Derksen:1972:MTR

[836] J. B. D. Derksen. A method for taking road traffic censuses. Statistica Neerlandica, 26(3):137–143, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1972.tb00182.x.

Doornbos:1972:CTN

[837] R. Doornbos and W. L. M. M. Senden. A comparison of two nonparametric k-sample slippage tests. Statistica Neerlandica, 26(3):145–154, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00183.x.

vanderLaan:1972:EDD

[838] P. van der Laan and J. Prakken. Exact distribution of Durbin’s distribution-free test statistic for Balanced Incomplete Block Designs, and comparison with the Chi-square and F approximation. Statistica Neerlandica, 26(3):155–164, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1972.tb00184.x.

Ott:1972:SSP

[839] Ellis R. Ott and Frank W. Wehrfritz. A special screening program for many treatments. Statistica Neerlandica, 26(3):165–170, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00185.x.

Jaros:1972:EGF

[840] Frantšek Jaroš and Agnes H. ŽAludová. The estimation of guaranteed fatigue life under random loading. Statistica Neerlandica, 26(3):171– 181, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00186.x.

Epstein:1972:HTA

[841] Benjamin Epstein. Hamaker type acceptance tests for the occurrence rate in a Poisson process. Statistica Neerlandica, 26(3):183–189, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00187.x.

Klamkin:1972:APR

[842] M. S. Klamkin and J. H. van Lint. An asymptotic problem in renewal theory. Statistica Neerlandica, 26(3):191–196, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00188.x.

Stommels:1972:ATS

[843] F. J. Stommels. Attuning two successive sampling systems (for instance supplier-customer). Statistica Neerlandica, 26(3):197–201, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00189.x.

Linssen:1972:VCA

[844] H. N. Linssen. The vacuum-cleaner and attachment. Statistica Neerlandica, 26(3):203–209, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00190.x.

Rogers:1972:USL

[845] William H. Rogers and John W. Tukey. Understanding some longtailed symmetrical distributions. Statistica Neerlandica, 26(3):211–226, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00191.x.

Mosteller:1972:DAL

[846] Frederick Mosteller. A data-analytic look at Goldbach counts. Statistica Neerlandica, 26(3):227–242, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1972.tb00192.x.

Wessels:1972:ICU

[847] J. Wessels. Inventory control with unknow demand distribution: a slow-mover case. Statistica Neerlandica, 26(3):243–251, September 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00193.x.

Buhrman:1972:LVK

[848] J. M. Buhrman. Een lading van kraaltjes. (Dutch) [A load of beads]. Statistica Neerlandica, 26(4):97–102, December 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1972.
tb00194.x.

Fase:1972:DLE

[849] M. M. G. Fase. The distribution of lifetime earnings: a problem of intertemporal aggregation. Statistica Neerlandica, 26(4):103–111, December 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00195.x.

Guenther:1972:VSP

[850] William C. Guenther. Variables sampling plans for the Poisson and the binomial. Statistica Neerlandica, 26(4):113–120, December 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00196.x.

Raghunandanan:1972:OSR

[851] K. Raghunandanan and S. A. Patil. On order statistics for random sample size. Statistica Neerlandica, 26(4):121–126, December 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1972.tb00197.x.

Dubbelman:1972:NCD

[852] C. Dubbelman, A. P. J. Abrahamse, and J. Koerts. A new class of disturbance estimators in the general linear model. Statistica Neerlandica, 26(4):127–142, December 1972. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1972.tb00198.x.

Tracy:1972:NMP

[853] Derrick S. Tracy and Rana P. Singh. A new matrix product and its applications in partitioned matrix differentiation. Statistica Neerlandica, 26(4):143–157, December 1972. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1972.tb00199.x.

Anonymous:1972:ODAb

[854] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 26(4):159, December 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00200.x.

Anonymous:1972:BDBb

[855] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 26(4):161–171, December 1972. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1972.tb00201.x.

Hawkins:1973:RTO

[856] Douglas M. Hawkins. Repeated testing for outliers. Statistica Neerlandica, 27(1):1–10, March 1973. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1973.tb00202.x.

Ruymgaart:1973:NNB

[857] F. H. Ruymgaart. Non-normal bivariate densities with normal marginals and linear regression functions. Statistica Neerlandica, 27(1):11–17, March 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00203.x.

Otten:1973:NDS

[858] A. Otten. The null distribution of Spearman’s S when n = 13(1)16. Statistica Neerlandica, 27(1):19–20, March 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1973.tb00204.x.

vanPraag:1973:OPS

[859] B. M. S. van Praag and F. van Doeland. Optimal packing systems. Statistica Neerlandica, 27(1):21–28, March 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1973.tb00205.x.

Quade:1973:PC

[860] Dana Quade. The pair chart. Statistica Neerlandica, 27(1):29–45, March 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00206.x.

Levenbach:1973:GLG

[861] Hans Levenbach. A generalization of the logistic growth function and its estimation. Statistica Neerlandica, 27(1):47–54, March 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1973.
tb00207.x.

Lepage:1973:MTM

[862] Yves Lepage. A maximin test for means. Statistica Neerlandica, 27(1): 55–57, March 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00208.x.

vanBeek:1973:MNI

[863] Paul van Beek. The maximum of a number of independent residual waiting times. Statistica Neerlandica, 27(2):59–67, June 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1973.
tb00209.x.

Lempers:1973:STS

[864] F. B. Lempers and T. Kloek. On a simple transformation for second-order autocorrelated disturbances in regression analysis. Statistica Neerlandica, 27(2):69–75, June 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00210.x.

vanEk:1973:EEM

[865] J. L. P. van Ek and C. B. Tilanu. Enige ervaringen met economischtechnische bedrijfsmodellen. (Dutch) [Some experiences with economictechnical business models]. Statistica Neerlandica, 27(2):77–88, June 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (elec-

tronic).  URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00211.x.

Anonymous:1973:ODAa

[866] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 27(2):89–90, June 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1973.tb00212.x.

Anonymous:1973:BDBa

[867] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 27(2):91–93, June 1973. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1973.tb00213.x.

Sander:1973:A

[868] P. C. Sander. Announcement. Statistica Neerlandica, 27(2):95–96, June 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00214.x.

vanderVaart:1973:EBM

[869] H. Robert van der Vaart. On the existence of bivariate moments of lower order given the existence of moments of higher order. Statistica Neerlandica, 27(3):97–102, September 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1973.tb00215.x.

Guenther:1973:DSS

[870] William C. Guenther. Determination of sample size for tests concerning means and variances of normal distributions. Statistica Neerlandica, 27 (3):103–113, September 1973. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1973.tb00216.x.

Madan:1973:PQS

[871] K. C. Madan. A priority queueing system with service interruptions. Statistica Neerlandica, 27(3):115–123, September 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1973.
tb00217.x.

Buhrman:1973:OSW

[872] J. M. Buhrman. On order statistics when the sample size has a binomial distribution. Statistica Neerlandica, 27(3):125–126, September 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00218.x.

Neudecker:1973:BSR

[873] H. Neudecker. De BLUF-schatter: een rechtstreekse afleiding. (Dutch) [The BLUF estimator: a direct derivation]. Statistica Neerlandica, 27 (3):127–130, September 1973. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1973.tb00219.x.

Anonymous:1973:ODAb

[874] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 27(3):131, September 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1973.tb00220.x.

Anonymous:1973:BDBb

[875] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 27(3):133–138, September 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1973.tb00221.x.

Koerts:1973:SIS

[876] J. Koerts and E. de Leede. Statistical inference and subjective probabilities. Statistica Neerlandica, 27(4):139–161, December 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1973.
tb00222.x.

Theune:1973:CPD

[877] J. A. Theune. Comparison of power for the D’Agostino and the Wilk– Shapiro test of normality for small and moderate samples. Statistica Neerlandica, 27(4):163–168, December 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1973.tb00223.x.

Teekens:1973:REM

[878] R. Teekens, J. Koerts, and A. S. Louter. On the robustness of an estimator of the mean of the dependent variable in a multiplicative model. Statistica Neerlandica, 27(4):171–179, December 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1973.
tb00224.x.

Moors:1973:LBR

[879] J. J. A. Moors. A lower bound for the relative precision of proportional to Neyman allocation. Statistica Neerlandica, 27(4):181–183, December 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00225.x.

Hempenius:1973:NRB

[880] A. L. Hempenius. A note on the relationship between the feasible regions of several types of chance constraints. Statistica Neerlandica, 27(4):185– 186, December 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1973.tb00226.x.

Anonymous:1973:ODAc

[881] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 27(4):187, December 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1973.tb00227.x.

Anonymous:1973:SIU

[882] Anonymous. Statistical inference under order restrictions. Statistica Neerlandica, 27(4):189, December 1973. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1973.tb00228.x.

Cohen:1974:SIM

[883] J. W. Cohen. Some ideas and models in reliability theory. Statistica Neerlandica, 28(1):1–10, March 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1974.tb00229.x.

Vos:1974:SMOa

[884] J. W. E. Vos. Steekproeftrekking met ongelijke kansen. (Dutch) [Unequal probability sampling]. Statistica Neerlandica, 28(1):11–49, March 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1974.tb00230.x.

Corsten:1974:AME

[885] L. C. A. Corsten and A. C. van Eijnsbergen. Addendum to “Multiplicative effects in two-way analysis of variance”. Statistica Neerlandica, 28 (1):51, March 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1974.tb00231.x.

Anonymous:1974:ODAa (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1974.tb00232.x.

Cohen:1974:SAQ

[887] J. W. Cohen. Some aspects of queueing theory. Statistica Neerlandica, 28 (2):55–67, June 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1974.tb00233.x.

Vos:1974:SMOb

[888] J. W. E. Vos. Steekproeftrekking met ongelijke kansen. (Dutch) [Sampling with unequal probabilities]. Statistica Neerlandica, 28(2):69– 108, June 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1974.tb00234.x.

vanMontfort:1974:LRA

[889] M. A. J. van Montfort. Lineaire regressie-analyse, ook bij gegevens in strijd met het model. (Dutch) [Linear regression analysis, even with data that conflicts with the model]. Statistica Neerlandica, 28(2):109– 116, June 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1974.tb00235.x. See comment [912].

Anonymous:1974:ODAb

[890] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 28(2):121–122, June 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1974.tb00236.x.

Anonymous:1974:BDBa

[891] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 28(2):123–125, June 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1974.tb00237.x.

Thomasse:1974:PRS

[892] A. H. Thomasse. Practical recipes to solve the Behrens–Fisher problem. Statistica Neerlandica, 28(3):127–138, September 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1974.
tb00744.x.

Ponstein:1974:TNS

[893] J. Ponstein. Theory and numerical solution of a discrete queueing problem. Statistica Neerlandica, 28(3):139–152, September 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1974.
tb00745.x.

Upperman:1974:OVK

[894] P. M. Upperman and A. M. Dévény. De oplossing van een kwaliteitsprobleem door het gebruik van proefschema’s. (Dutch) [The solution of a quality problem through the use of test schemes]. Statistica Neerlandica, 28(3):153–172, September 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1974.tb00746.x.

Anonymous:1974:ODAc

[895] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 28(3):173–174, September 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1974.tb00747.x.

VanOorschot:1974:CVK

[896] H. J. L. Van Oorschot. Een criterium voor het kiezen van een proefopzet. (Dutch) [A criterion for choosing an experimental design]. Statistica Neerlandica, 28(4):175–207, December 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1974.tb00748.x.

vanHouwelingen:1974:EBR

[897] J. C. van Houwelingen. An empirical Bayes rule for testing simple hypothesis against simple alternative. Statistica Neerlandica, 28(4):209– 221, December 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1974.tb00749.x.

Anonymous:1974:ODAd

[898] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 28(4):223, December 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1974.tb00750.x.

Anonymous:1974:BDBb

[899] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 28(4):225–227, December 1974. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1974.tb00751.x.

Wessels:1975:DSM

[900] J. Wessels and J. A. E. E. van Nunen. Discounted semi-Markov decision processes: linear programming and policy iteration. Statistica Neerlandica, 29(1):1–7, March 1975. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1975.tb00238.x.

vandePanne:1975:SMT

[901] C. van de Panne. Some models of truck allocation for household refuse collection. Statistica Neerlandica, 29(1):9–21, March 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1975.
tb00239.x.

Bruyneel:1975:BTV

[902] G. Bruyneel, H. Muller(-Malek), and S. de Samblanckx. Een bedrijfseconomische toepassing van de theorie der Markov-ketens. (Dutch) [A business economics application of the theory of Markov chains]. Statistica Neerlandica, 29(1):23–27, March 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1975.tb00240.x.

Anonymous:1975:BDBa

[903] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 29(1):29, March 1975. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1975.tb00241.x.

Anonymous:1975:ODAa

[904] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 29(1):31–32, March 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1975.tb00242.x.

Wiggers:1975:BAV

[905] B. G. Wiggers and B. Dorhout. Bij het afscheid van Corsten en Scheffer als redacteuren van Statistica Neerlandica. (Dutch) [On the farewell of Corsten and Scheffer as editors of Statistica Neerlandica]. Statistica Neerlandica, 29(2):33, June 1975. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1975.tb00243.x.

Enters:1975:JJT

[906] J. H. Enters. A. J. de Jong — ter nagedachtenis. (Dutch) [A. J. de Jong — in memory]. Statistica Neerlandica, 29(2):35–37, June 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00244.x.

Snijders:1975:TRB

[907] Tom Snijders. A test for randomness in behaviour. Statistica Neerlandica, 29(2):39–48, June 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00245.x.

Kreijger:1975:ORV

[908] R. G. Kreijger and T. Rouwens. Optimale rantsoenering van aardolie en aardolieprodukten in Nederland; een toepassing van input-outputanalyse en lineaire programmering. (Dutch) [Optimal rationing of petroleum and petroleum products in the Netherlands; an application of inputoutput analysis and linear programming]. Statistica Neerlandica, 29(2): 49–56, June 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00246.x.

Linssen:1975:FPS

[909] H. N. Linssen, J. Hermans, P. van Bemmel, and R. de Graaf. Fit performance of several programs used for non-linear least squares problems. Statistica Neerlandica, 29(2):57–61, June 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1975.tb00247.x.

Hamaker:1975:APT

[910] Hugo C. Hamaker. An alternative procedure for taking a random sample. Statistica Neerlandica, 29(2):63–66, June 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1975.tb00248.x.

Droppert:1975:EEM

[911] A. J. Droppert. Enige ervaringen met kursussen in de OR met behulp van management games. (Dutch) [Some experiences with courses in the OR using management games]. Statistica Neerlandica, 29(2): 67–73, June 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00249.x.

Baaij:1975:COL

[912] J. G. Baaij. Commentaar op “Lineaire regressie, ook bij gegevens in strijd met het model”. (Dutch) [Comment on “Linear regression, even with data that conflicts with the model”]. Statistica Neerlandica, 29(2): 75–76, June 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00250.x. See [889].

Anonymous:1975:BDBb

[913] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 29(2):77–79, June 1975. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1975.tb00251.x.

Albers:1975:EDC

[914] W. Albers. Efficiency and deficiency considerations in the symmetry problem. Statistica Neerlandica, 29(3):81–92, September 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1975.
tb00252.x.

Linssen:1975:NMC

[915] H. N. Linssen. Nonlinearity measures: a case study. Statistica Neerlandica, 29(3):93–99, September 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1975.tb00253.x.

Landis:1975:RSMa

[916] J. Richard Landis and Gary G. Koch. A review of statistical methods in the analysis of data arising from observer reliability studies (Part I). Statistica Neerlandica, 29(3):101–123, September 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00254.x.

Bosch:1975:OBA

[917] A. J. Bosch. Opmerkingen bij een aantal optimaliteitscriteria voor proefopzetten. (Dutch) [Comments on a number of optimality criteria for experimental designs]. Statistica Neerlandica, 29(3):125–126, September 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00255.x.

Anonymous:1975:ODAb

[918] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 29(3):127–128, September 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1975.tb00256.x.

Guenther:1975:IHU

[919] William C. Guenther. The inverse hypergeometric — a useful model. Statistica Neerlandica, 29(4):129–144, December 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1975.
tb00257.x.

Mendieta:1975:ODL

[920] E. J. Mendieta, H. N. Linssen, and R. Doornbos. Optimal designs for linear mixture models. Statistica Neerlandica, 29(4):145–150, December 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00258.x.

Landis:1975:RSMb

[921] J. Richard Landis and Gary G. Koch. A review of statistical methods in the analysis of data arising from observer reliability studies (Part II). Statistica Neerlandica, 29(4):151–161, December 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1975.tb00259.x.

Steutel:1975:ODA

[922] F. W. Steutel. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 29(4):163–164, December 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1975.tb00260.x.

Anonymous:1975:BDBc

[923] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 29(4):165–170, December 1975. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1975.tb00261.x.

Susarla:1976:PSE

[924] V. Susarla. A property of Stein’s estimator for the mean of a multivariate normal distribution. Statistica Neerlandica, 30(1):1–5, March 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1976.tb00262.x.

Verdooren:1976:LBD

[925] L. R. Verdooren. Loten bij delicate vragen; een overzicht van “randomized response“-technieken. (Dutch) [Draw lots for delicate questions; an overview of “randomized response” techniques]. Statistica Neerlandica, 30(1):7–24, March 1976. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1976.tb00263.x.

Lageweg:1976:MML

[926] B. J. Lageweg, J. K. Lenstra, and A. H. G. Rinnooy Kan. Minimizing maximum lateness on one machine: computational experience and some applications. Statistica Neerlandica, 30(1):25–41, March 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1976.tb00264.x.

Anonymous:1976:ODAa

[927] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 30(1):43–49, March 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1976.tb00265.x.

vanderWal:1976:MG

[928] J. van der Wal and J. Wessels. On Markov games. Statistica Neerlandica, 30(2):51–71, June 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1976.tb00266.x.

Wierenga:1976:MBC

[929] B. Wierenga. Merkkeuze bij consumenten. (Dutch) [Brand choice among consumers]. Statistica Neerlandica, 30(2):73–92, June 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1976.
tb00267.x.

Schouten:1976:CC

[930] Hubert J. A. Schouten. On the Continuity Correction. Statistica Neerlandica, 30(2):93–95, June 1976. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1976.tb00268.x.

Hordijk:1976:SPE

[931] A. Hordijk and H. C. Tijms. A simple proof of the equivalence of the limiting distributions of the continuous-time and the embedded process of the queue size in the M/G/1 queue. Statistica Neerlandica, 30(2): 97–100, June 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1976.tb00269.x.

Anonymous:1976:ODAb

[932] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 30(2):101, June 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1976.tb00270.x.

Weima:1976:BAB

[933] Jan Weima. Bacteriological assay: the best motley row. Statistica Neerlandica, 30(3):103–117, September 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1976.tb00271.x.

Shorack:1976:RSL

[934] Galen R. Shorack. Robust Studentization of location estimates. Statistica Neerlandica, 30(3):119–141, September 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1976.tb00272.x.

Gill:1976:MLS

[935] R. D. Gill. The model of Latent Structure Analysis. Statistica Neerlandica, 30(3):143–149, September 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1976.tb00273.x. See comments

[1047].

Anonymous:1976:ODAc

[936] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 30(3):151–158, September 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1976.tb00274.x.

deHaan:1976:SEE

[937] Laurens de Haan. Sample extremes: an elementary introduction. Statistica Neerlandica, 30(4):161–172, December 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1976.
tb00275.x.

Wessels:1976:FFR

[938] J. Wessels and J. A. E. E. van Nunen. FORMASY FOrecasting and Recruitment in MAnpower SYstems. Statistica Neerlandica, 30(4):173– 193, December 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1976.tb00276.x.

Wijngaard:1976:RTV

[939] J. Wijngaard. Regen, een toepassing van dynamische programering. (Dutch) [Rain, an application of dynamic programming]. Statistica Neerlandica, 30(4):195–199, December 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1976.tb00277.x.

Anonymous:1976:ODAd

[940] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 30(4):201–206, December 1976. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1976.tb00278.x.

Groeneboom:1977:BEP

[941] P. Groeneboom and J. Oosterhoff. Bahadur efficiency and probabilities of large deviations. Statistica Neerlandica, 31(1):1–24, March 1977. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1002/stan.1977.31.1.1.

Schaafsma:1977:CDP

[942] Willem Schaafsma and Gerrit N. van Vark. Classification and discrimination problems with applications, part I. Statistica Neerlandica, 31(1): 25–45, March 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1002/stan.1977.31.1.25.

Anonymous:1977:ODAa

[943] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 31(1):47–52, March 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1002/stan.1977.31.1.47.

Doksum:1977:SGM

[944] Kjell A. Doksum. Some graphical methods in statistics: a review and some extensions. Statistica Neerlandica, 31(2):53–68, June 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1977.
tb00752.x.

VanDerGenugten:1977:ISI

[945] B. B. Van Der Genugten. Identification in statistical inference. Statistica Neerlandica, 31(2):69–89, June 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1977.tb00753.x.

Anonymous:1977:ODAb

[946] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 31(2):91–92, June 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1977.tb00754.x.

Anonymous:1977:BDB

[947] Anonymous. Boekbespreking. (Dutch) [Book review]. Statistica Neerlandica, 31(2):93–94, June 1977. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1977.tb00755.x.

VanHouwelingen:1977:MEB

[948] J. C. Van Houwelingen. Monotonizing empirical Bayes estimators for a class of of discrete distributions with monotone likelihood ratio. Statistica Neerlandica, 31(3):95–104, September 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1977.tb00756.x.

Hutchinson:1977:LSM

[949] T. P. Hutchinson. Latent structure models applied to the joint distribution of drivers’ injuries in road accidents. Statistica Neerlandica, 31 (3):105–111, September 1977. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1977.tb00757.x.

Peters:1977:SPP

[950] Robert J. Peters, Klaas Boskma, and Hendrika E. Kupper. Stochastic programming in production planning: a case with none-simple recourse. Statistica Neerlandica, 31(3):113–126, September 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1977.
tb00758.x.

Romanowska:1977:NUB

[951] Małgorzata Romanowska. A note on the upper bound for the distance in total variation between the binomal and the Poisson distribution. Statistica Neerlandica, 31(3):127–130, September 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1977.
tb00759.x.

Anonymous:1977:ODAc

[952] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 31(3):131–138, September 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1977.tb00760.x.

Anonymous:1977:A

[953] Anonymous. Announcement. Statistica Neerlandica, 31(3):139, September 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1977.tb00761.x.

Ketellapper:1977:SEM

[954] R. H. Ketellapper, B. Bos, M. A. Kooyman, and W. Voorhoeve. A simultaneous econometric model for the Dutch economy. Statistica Neerlandica, 31(4):141–159, December 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1977.tb00762.x.

Mellenbergh:1977:DDB

[955] G. J. Mellenbergh, H. Koppelaar, and W. J. Van Der Linden. Dichotomous decisions based on dichotomously scored items: a case study. Statistica Neerlandica, 31(4):161–169, December 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1977.
tb00763.x.

Wallin:1977:OTV

[956] J. F. E. Wallin. Overflow traffic from the viewpoint of renewal theory. Statistica Neerlandica, 31(4):171–178, December 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1977.
tb00764.x.

Boomsma:1977:CAC

[957] A. Boomsma. Comparing approximations of confidence intervals for the product-moment correlation coefficient. Statistica Neerlandica, 31 (4):179–185, December 1977. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1977.tb00765.x.

Anonymous:1977:ODAd

[958] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 31(4):187–192, December 1977. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1977.tb00766.x.

Melis:1977:MPM

[959] P. Melis. Mathematical programming methods. Statistica Neerlandica, 31(4):193, December 1977. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1977.tb00767.x.

Willems:1978:RF

[960] Jan C. Willems. Recursive filtering. Statistica Neerlandica, 32(1):1– 39, March 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1978.tb01382.x.

Otter:1978:DKF

[961] Pieter W. Otter. The discrete Kalman filter applied to linear regression models: statistical considerations and an application. Statistica Neerlandica, 32(1):41–56, March 1978. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1978.tb01383.x.

vanBaaren:1978:CEB

[962] A. van Baaren. On a class of extensions to the Branley–Terry model in paired comparisons. Statistica Neerlandica, 32(1):57–66, March 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1978.tb01384.x.

Anonymous:1978:ODAa

[963] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 32(1):67–72, March 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1978.tb01385.x.

Runnenburg:1978:MMM

[964] J. Th. Runnenburg. Mean, median, mode. Statistica Neerlandica, 32 (2):73–79, June 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1978.tb01386.x.

Stam:1978:DBS

[965] A. J. Stam. Distance between sampling with and without replacement. Statistica Neerlandica, 32(2):81–91, June 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1978.tb01387.x.

Perng:1978:TET

[966] S. K. Perng. A test for equality of two exponential distributions. Statistica Neerlandica, 32(2):93–102, June 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1978.tb01388.x.

Anonymous:1978:ODAb

[967] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 32(2):103–108, June 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1978.tb01389.x.

Noether:1978:DFC

[968] Gottfried E. Noether. Distribution-free confidence intervals. Statistica Neerlandica, 32(3):109–122, September 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1978.tb01390.x.

Hemelrijk:1978:RBS

[969] J. Hemelrijk. Rules for building statistical models. Statistica Neerlandica, 32(3):123–134, September 1978. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1978.tb01391.x.

Anonymous:1978:C

[970] Anonymous. Correction. Statistica Neerlandica, 32(3):135, September 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1978.tb01393.x.

Anonymous:1978:PPC

[971] Anonymous. Prize problem: In connection with the foregoing paper. Statistica Neerlandica, 32(3):135, September 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1978.
tb01392.x.

Hawkins:1978:ATT

[972] Douglas M. Hawkins. Analysis of three tests for one or two outliers. Statistica Neerlandica, 32(3):137–148, September 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1978.
tb01394.x.

deKroon:1978:GKO

[973] J. P. M. de Kroon. Gemiddelde en kwadraatsom opnieuw bekeken. (Dutch) [Mean and sum of squares revisited]. Statistica Neerlandica, 32(3):149–153, September 1978. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1978.tb01395.x.

Anonymous:1978:OO

[974] Anonymous. Opgavenrubriek: Oplossingen. Statistica Neerlandica, 32 (3):155–162, September 1978. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1978.tb01396.x.

teRiele:1978:PRP

[975] H. J. J. te Riele. The proportional representation problem in the Second Chamber: an approach via minimal distances. Statistica Neerlandica, 32(4):163–179, December 1978. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1978.tb01397.x.

Elffers:1978:IPI

[976] Henk Elffers, Jelke Bethlehem, and Richard Gill. Indeterminacy problems and the interpretation of factor analysis results. Statistica Neerlandica, 32(4):181–199, December 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1978.tb01398.x.

Magnus:1978:MPQ

[977] Jan R. Magnus. The moments of products of quadratic forms in normal variables. Statistica Neerlandica, 32(4):201–210, December 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1978.tb01399.x.

Anonymous:1978:ODAc

[978] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 32(4):211–219, December 1978. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1978.tb01400.x.

vanZwet:1979:MMM

[979] W. R. van Zwet. Mean, median, mode II. Statistica Neerlandica, 33 (1):1–5, March 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00657.x.

Fase:1979:NPM

[980] M. M. G. Fase, D. van der Hoeven, and M. van Nieuwkerk. A numerical planning model for a central bank’s bank notes operations. Statistica Neerlandica, 33(1):7–25, March 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1979.tb00658.x.

Sahai:1979:EVP

[981] Ashok Sahai. An efficient variant of the product and ratio estimators. Statistica Neerlandica, 33(1):27–35, March 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1979.
tb00659.x.

Srivenkataramana:1979:RPM

[982] T. Srivenkataramana and D. S. Tracy. On ratio and product methods of estimation in sampling. Statistica Neerlandica, 33(1):37–49, March 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00660.x.

Sacks:1979:NET

[983] S. T. Sacks and S. Selvin. A note on the extension of tables for the Friedman statistic. Statistica Neerlandica, 33(1):51–54, March 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1979.
tb00661.x.

Anonymous:1979:ODAa

[984] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 33(1):55–64, March 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1979.tb00662.x.

Smit:1979:BA

[985] J. C. Smit. Beware of averages! Statistica Neerlandica, 33(2):65– 71, June 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00663.x.

Don:1979:EPQ

[986] F. J. H. Don. The expectation of products of quadratic forms in normal variables. Statistica Neerlandica, 33(2):73–79, June 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1979.
tb00664.x.

Ganesalingam:1979:CST

[987] S. Ganesalingam and G. J. McLachlan. A case study of two clustering methods based on maximum likelihood. Statistica Neerlandica, 33(2): 81–90, June 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00665.x.

Schaafsma:1979:CDP

[988] Willem Schaafsma and Gerrit N. van Vark. Classification and discrimination problems with applications, Part IIa. Statistica Neerlandica, 33 (2):91–126, June 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00666.x.

Anonymous:1979:ODAb

[989] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 33(2):127–132, June 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1979.tb00667.x.

Magnus:1979:EPQ

[990] Jan R. Magnus. The expectation of products of quadratic forms in normal variables: the practice. Statistica Neerlandica, 33(3):131–136, September 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00668.x.

Koopman:1979:TSP

[991] P. A. R. Koopman. Testing symmetry with a procedure, combining the sign test and the signed rank test. Statistica Neerlandica, 33(3):137– 142, September 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00669.x.

Talman:1979:SPO

[992] A. J. J. Talman. A simple proof of the optimality of the best N-policy in the M/G/1 queueing control problem with removable server. Statistica Neerlandica, 33(3):143–150, September 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1979.tb00670.x.

Hall:1979:RCL

[993] W. J. Hall and Jon A. Wellner. The rate of convergence in law of the maximum of an exponential sample. Statistica Neerlandica, 33(3):151– 154, September 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00671.x.

Anonymous:1979:ODAc

[994] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 33(3):155–165, September 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1979.tb00672.x.

Shorack:1979:WEP

[995] Galen R. Shorack. The weighted empirical process of row independent random variables with arbitrary distribution functions. Statistica Neerlandica, 33(4):169–189, December 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1979.tb00673.x.

Kubat:1979:MMN

[996] Peter Kubat. Mean or median? (A note on an old problem). Statistica Neerlandica, 33(4):191–196, December 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1979.tb00674.x.

Winter:1979:ADC

[997] B. B. Winter. An alternate development of conditioning. Statistica Neerlandica, 33(4):197–212, December 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1979.tb00675.x.

Sahai:1979:BVI

[998] Ashok Sahai and J. M. Buhrman. Bounds for the variance of an inverse binomial estimator. Statistica Neerlandica, 33(4):213–215, December 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1979.tb00676.x.

Anonymous:1979:RPD

[999] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 33(4):217–221, December 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1979.tb00677.x.

Anonymous:1979:ODAd

[1000] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 33(4):223–229, December 1979. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1979.tb00678.x.

Runnenburg:1980:BAV

[1001] J. Th. Runnenburg and R. J. Mokken. Bij het afscheid van Schaafsma als hoofdredacteur van Statistica Neerlandica. (Dutch) [On the farewell of Schaafsma as Editor-in-Chief of Statistica Neerlandica]. Statistica Neerlandica, 34(1):1, March 1980. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1980.tb00679.x.

Elffers:1980:IPM

[1002] Henk Elffers. On interpreting the product moment correlation coefficient. Statistica Neerlandica, 34(1):3–11, March 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00680.x.

Kaas:1980:MMM

[1003] R. Kaas and J. M. Buhrman. Mean, median and mode in binomial distributions. Statistica Neerlandica, 34(1):13–18, March 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00681.x.

Cressie:1980:EPU

[1004] Noel Cressie. M-estimation in the presence of unequal scale. Statistica Neerlandica, 34(1):19–32, March 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1980.tb00682.x.

Verbon:1980:MLE

[1005] H. A. A. Verbon. Maximum likelihood estimation of a labour demand system. An application of a model of seemingly unrelated regression equations with the regression errors composed of two components. Statistica Neerlandica, 34(1):33–48, March 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1980.tb00683.x.

Anonymous:1980:RPD

[1006] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 34(1):49–53, March 1980. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1980.tb00684.x.

Anonymous:1980:ODAa

[1007] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 34(1):55–64, March 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1980.tb00685.x.

Anonymous:1980:SMH

[1008] Anonymous. Shewhart medal 1979 for Hugo C. Hamaker. Statistica Neerlandica, 34(2):65, June 1980. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1980.tb00686.x.

Hoekstra:1980:IC

[1009] A. H. Hoekstra and J. Th. Runnenburg. Inequalities compared. Statistica Neerlandica, 34(2):67–82, June 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1980.tb00687.x.

Verdooren:1980:EVC

[1010] L. R. Verdooren. On estimation of variance components. Statistica Neerlandica, 34(2):83–106, June 1980. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1980.tb00688.x.

Rao:1980:LDR

[1011] P. V. Rao and D. D. Wackerly. A large deviation result for the difference in sample medians. Statistica Neerlandica, 34(2):107–116, June 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1980.tb00689.x.

vanderLaan:1980:APT

[1012] P. van der Laan and J. Weima. Asymptotic power of the twosample test of Wilcoxon for logistic shift alternatives, and comparison with simulation results. Statistica Neerlandica, 34(2):117–121, June 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1980.tb00690.x.

Hooghiemstra:1980:BEL

[1013] G. Hooghiemstra. Brownian excursion and limit theorems for the M/G/1 queue. Statistica Neerlandica, 34(2):123–124, June 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00691.x.

Gill:1980:CSI

[1014] R. D. Gill. Censoring and stochastic integrals. Statistica Neerlandica, 34 (2):124, June 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1980.tb00692.x.

Vijn:1980:PIL

[1015] P. Vijn. Prior information in linear models. Statistica Neerlandica, 34 (2):125, June 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1980.tb00693.x.

deKroon:1980:ODA

[1016] R. Doornbosen J. P. M. de Kroon. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 34(2):127–133, June 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00694.x.

Anonymous:1980:VDP

[1017] Anonymous. Van Dantzig-prijs. (Dutch) [Van Dantzig prize]. Statistica Neerlandica, 34(3):135–139, September 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1980.tb00695.x.

Bierens:1980:CSE

[1018] Herman J. Bierens. Consistent selection of explanatory variables. Statistica Neerlandica, 34(3):141–150, September 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00696.x.

Boyer:1980:ACF

[1019] John E. Boyer, Jr. and Dennis C. Cilliland. Admissibility considerations in the finite state compound and empirical Bayes decision problems. Statistica Neerlandica, 34(3):151–159, September 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00697.x.

Anonymous:1980:RDN

[1020] Anonymous. Regression Diagnostics, a new book by D. W. Belsley, E. Kuh, and R. E. Welsch. Statistica Neerlandica, 34(3):161–164, September 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1980.tb00698.x.

Runnenburg:1980:ODA

[1021] J. Th. Runnenburg. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 34(3):165–172, September 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00699.x.

Akkerboom:1980:TLC

[1022] J. C. Akkerboom, P. Schepers, and J. v.d. Werff. Thin layer chromatography, a case study. Statistica Neerlandica, 34(4):173–187, December 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1980.tb00700.x.

Fung:1980:SSB

[1023] Karen Yuen Fung. Small sample behaviour of some nonparametric multisample location tests in the presence of dispersion differences. Statistica Neerlandica, 34(4):189–196, December 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1980.tb00701.x.

Tsao:1980:RPB

[1024] How Jan Tsao. On the risk performance of Bayes empirical Bayes procedures for classification between N(−1, l) and N(1, 1). Statistica Neerlandica, 34(4):197–208, December 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1980.tb00702.x.

Vos:1980:MDR

[1025] J. W. E. Vos. Mixing of direct, ratio, and product method estimators. Statistica Neerlandica, 34(4):209–218, December 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00703.x.

Vos:1980:MOW

[1026] H. H. Vos. Measurement of orientations to work. Statistica Neerlandica, 34(4):219, December 1980. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1980.tb00704.x.

Linssen:1980:FRM

[1027] H. N. Linssen. Functional relationship and minimum sum estimation. Statistica Neerlandica, 34(4):220, December 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1980.
tb00705.x.

Anonymous:1980:ODAb

[1028] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 34(4):221–227, December 1980. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1980.tb00706.x.

Runnenburg:1981:BAV

[1029] J. Th. Runnenburg and R. J. Mokken. Bij het afscheid van Doornbos als hoofdredacteur van Statistica Neerlandica. (Dutch) [On the farewell of Doornbos as Editor-in-Chief of Statistica Neerlandica]. Statistica Neerlandica, 35(1):1, March 1981. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1981.tb00707.x.

Hempenius:1981:FAA

[1030] A. L. Hempenius. Forecast accuracy analysis: Applied to forecasts of the Dutch Central Planning Bureau, 1964–1978. Statistica Neerlandica, 35 (1):3–37, March 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1981.tb00708.x.

Hawkins:1981:TEV

[1031] Douglas M. Hawkins. Testing for equality of variance of correlated normal variables. Statistica Neerlandica, 35(1):39–47, March 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1981.
tb00709.x.

Keilson:1981:UPT

[1032] J. Keilson. On the unimodality of passage time densities in birth-death processes. Statistica Neerlandica, 35(1):49–55, March 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1981.
tb00710.x.

vanDoorn:1981:SMB

[1033] E. van Doorn. Stochastic monotonicity of birth-death processes. Statistica Neerlandica, 35(1):57, March 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1981.tb00711.x.

vanderLaan:1981:SFP

[1034] G. van der Laan. Simplicial fixed point algorithms. Statistica Neerlandica, 35(1):58, March 1981. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1981.tb00712.x.

Talman:1981:VDF

[1035] A. J. J. Talman. Variable dimension fixed point algorithms and triangulations. Statistica Neerlandica, 35(1):59, March 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1981.
tb00713.x.

Stijnen:1981:ABM

[1036] Th. Stijnen. On the asymptotic behaviour of monotonized empirical Bayes rules. Statistica Neerlandica, 35(1):60, March 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1981.
tb00714.x.

Anonymous:1981:ODAa (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1981.tb00715.x.

deRoos:1981:RAN

[1038] Gerlof Th. de Roos and Willem Schaafsma. Is recreation affecting the number of birds’ nests? A case study. Statistica Neerlandica, 35(2): 69–90, June 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1981.tb00716.x.

Kan:1981:CLP

[1039] A. H. G. Rinnooy Kan and J. Telgen. The complexity of linear programming. Statistica Neerlandica, 35(2):91–107, June 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1981.
tb00717.x.

Kremer:1981:AOA

[1040] Erhard Kremer. Asymptotic optimality of all signed rank tests at infinity. Statistica Neerlandica, 35(2):109–113, June 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1981.
tb00718.x.

deBoer:1981:PEI

[1041] P. M. C. de Boer. Price effects in input–output relations. Statistica Neerlandica, 35(2):115, June 1981. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1981.tb00719.x.

vanWinkel:1981:AP

[1042] E. F. G. van Winkel. ARIMA-processen. Statistica Neerlandica, 35 (2):116, June 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1981.tb00720.x.

vanderHoek:1981:RMN

[1043] G. van der Hoek. Reduction methods in nonlinear programming. Statistica Neerlandica, 35(2):117, June 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1981.tb00721.x.

Anonymous:1981:ODAb

[1044] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 35(2):119–128, June 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1981.tb00722.x.

Aase:1981:MRA

[1045] Knut Kristian Aase. Model reference adaptive systems applied to regression analyses. Statistica Neerlandica, 35(3):129–155, September 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1981.tb00723.x.

Best:1981:TCE

[1046] D. J. Best and J. C. W. Rayner. Are two classes enough for the χ^{2} goodness of fit test? Statistica Neerlandica, 35(3):157–163, September 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1981.tb00724.x.

Altham:1981:SCL

[1047] P. M. E. Altham. Some comments on latent structure analysis, arising out of a paper by R. D. Gill. Statistica Neerlandica, 35(3):165– 172, September 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1981.tb00725.x. See [935].

Bierens:1981:RMA

[1048] H. J. Bierens. Robust methods and asymptotic theory in nonlinear econometrics. Statistica Neerlandica, 35(3):173, September 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1981.
tb00726.x.

Wansbeek:1981:QEP

[1049] T. J. Wansbeek. Quantitative effects in panel data modelling. Statistica Neerlandica, 35(3):174, September 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1981.tb00727.x.

Anonymous:1981:ODAc

[1050] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 35(3):175–183, September 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1981.tb00728.x.

Sittig:1981:MJV

[1051] J. Sittig. In memoriam Jan van Ettinger. Statistica Neerlandica, 35 (4):185–188, December 1981. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1981.tb00729.x.

DeKroon:1981:DFT

[1052] J. De Kroon and P. Van Der Laan. Distribution-free test procedures in two-way layouts; a concept of rank-interaction. Statistica Neerlandica, 35(4):189–213, December 1981. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1981.tb00730.x.

Ramanarayanan:1981:ESS

[1053] R. Ramanarayanan. Explicit steady state solution for a Markovian inventory system. Statistica Neerlandica, 35(4):215–219, December 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1981.tb00731.x.

Jewett:1981:TME

[1054] R. I. Jewett and A. E. Ronner. Trimmed means and M-estimates. Statistica Neerlandica, 35(4):221–227, December 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1981.
tb00732.x.

vanHalem:1981:RPD

[1055] C. van Halem. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 35(4):229–230, December 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1981.tb00733.x.

Anonymous:1981:ODAd

[1056] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 35(4):231–238, December 1981. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1981.tb00734.x.

Corsten:1982:LIH

[1057] L. C. A. Corsten. Levensbericht over Ir. S. H. Justesen 1906–1981. Statistica Neerlandica, 36(1):1–2, March 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1982.tb00768.x.

Goria:1982:STS

[1058] M. N. Goria. A survey of two-sample location-scale problem, asymptotic relative efficiencies of some rank tests. Statistica Neerlandica, 36(1): 3–13, March 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00769.x.

Orban:1982:PDF

[1059] John Orban and Douglas A. Wolfe. Properties of a distribution-free two-stage two-sample median test. Statistica Neerlandica, 36(1):15– 22, March 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00770.x.

Bass:1982:TST

[1060] Thomas A. Bass and Karen Yuen Fung. A two-stage test for one-way analysis of variance. Statistica Neerlandica, 36(1):23–30, March 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00771.x.

Davis:1982:RCD

[1061] Richard A. Davis. The rate of convergence in distribution of the maxima. Statistica Neerlandica, 36(1):31–35, March 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1982.
tb00772.x.

Anonymous:1982:ODAa

[1062] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 36(1):37–43, March 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1982.tb00773.x.

Schouten:1982:MPI

[1063] H. J. A. Schouten. Measuring pairwise interobserver agreement when all subjects are judged by the same observers. Statistica Neerlandica, 36 (2):45–61, June 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00774.x.

Roebruck:1982:CFTa

[1064] P. Roebruck. Canonical forms and tests of hypotheses: Part I: the general univariate mixed model. Statistica Neerlandica, 36(2):63– 74, June 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00775.x.

Roebruck:1982:CFTb

[1065] P. Roebruck. Canonical forms and tests of hypotheses Part II: Multivariate mixed linear models. Statistica Neerlandica, 36(2):75–80, June 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00776.x.

Bierens:1982:UWL

[1066] H. J. Bierens. A uniform weak law of large numbers under π-mixing with application to nonlinear least squares estimation. Statistica Neerlandica, 36(2):81–86, June 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00777.x.

Anonymous:1982:BR

[1067] Anonymous. Book review. Statistica Neerlandica, 36(2):87–89, June 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00778.x.

Anonymous:1982:ODAb

[1068] Anonymous. Opgavenrubriek. (Dutch) [Assignment section]. Statistica Neerlandica, 36(2):91–100, June 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1982.tb00779.x.

Anonymous:1982:A

[1069] Anonymous. Announcement. Statistica Neerlandica, 36(3):101, September 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00780.x.

deJong:1982:ACI

[1070] Paul M. de Jong. Asymptotic confidence intervals for the effect of closing an area upon the number of birds’ nests. Statistica Neerlandica, 36(3):103–108, September 1982. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1982.tb00781.x.

Maher:1982:MAF

[1071] M. J. Maher. Modelling association football scores. Statistica Neerlandica, 36(3):109–118, September 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1982.tb00782.x.

Fligner:1982:DFT

[1072] M. A. Fligner and D. A. Wolfe. Distribution-free tests for comparing several treatments with a control. Statistica Neerlandica, 36(3):119– 127, September 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00783.x.

Luschgy:1982:MCT

[1073] H. Luschgy. Minimax character of the two-sample χ^{2}-test. Statistica Neerlandica, 36(3):129–134, September 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1982.tb00784.x.

Runnenburg:1982:ANV

[1074] J. Th. Runnenburg. Asymptotic normality in vacancies on a line. Statistica Neerlandica, 36(3):135–148, September 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1982.
tb00785.x.

Anonymous:1982:RPDa

[1075] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 36(3):149–153, September 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1982.tb00786.x.

Anonymous:1982:PS

[1076] Anonymous. Problem section. Statistica Neerlandica, 36(3):155–163, September 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00787.x.

Kleijnen:1982:SAS

[1077] Jack P. C. Kleijnen. Statistical aspects of simulation: an updated survey. Statistica Neerlandica, 36(4):165–186, December 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1982.
tb00788.x.

Koeze:1982:LLB

[1078] P. Koeze. The life-length of banknotes. Statistica Neerlandica, 36(4):187– 207, December 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1982.tb00789.x.

Deshpande:1982:IIP

[1079] M. N. Deshpande and S. G. Prabhu-Ajgaonkar. An IPPS (inclusion probability proportional to size) sampling scheme. Statistica Neerlandica, 36(4):209–212, December 1982. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1982.tb00790.x.

Anonymous:1982:RPDb

[1080] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 36(4):213, December 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1982.tb00791.x.

Lossers:1982:PS

[1081] O. P. Lossers, A. J. Van Es, R. Doornbos, P. Kuiper, A. A. Jagers, F. W. Steutel, A. J. Bosch, J. C. Smit, J. Th. Runnenburg, B. B. Van der Genugten, and E. E. C. Van Damme. Problem section. Statistica Neerlandica, 36(4):215–220, December 1982. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1982.tb00792.x.

deKroon:1983:GFR

[1082] J. de Kroon and P. van der Laan. A generalization of Friedman’s rank statistic. Statistica Neerlandica, 37(1):1–14, March 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1983.
tb00793.x.

Das:1983:RM

[1083] P. Das and P. G. H. Mulder. Regression to the mode. Statistica Neerlandica, 37(1):15–20, March 1983. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1983.tb00794.x.

Clarkson:1983:ADP

[1084] D. B. Clarkson and D. B. Wolfson. An application of a displaced Poisson process. Statistica Neerlandica, 37(1):21–28, March 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1983.
tb00795.x.

vanHouwelingen:1983:MEB

[1085] J. C. van Houwelingen and Th. Stijnen. Monotone empirical Bayes estimators for the continuous one-parameter exponential family. Statistica Neerlandica, 37(1):29–43, March 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1983.tb00796.x.

Anonymous:1983:PSa

[1086] Anonymous. Problem section. Statistica Neerlandica, 37(1):45–50, March 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00797.x.

Anonymous:1983:DWE

[1087] Anonymous. Dr. W. Edward Deming Medal for Hugo C. Hamaker. Statistica Neerlandica, 37(2):51, June 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1983.tb00798.x.

Games:1983:RSP

[1088] P. A. Games, H. J. Keselman, and J. C. Rogan. A review of simultaneous pairwise multiple comparisons. Statistica Neerlandica, 37(2): 53–58, June 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00799.x.

Engel:1983:AVI

[1089] J. Engel. Analysis of variance on indirectly measured variables. Statistica Neerlandica, 37(2):59–68, June 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1983.tb00800.x.

Schemper:1983:NST

[1090] M. Schemper. A nonparametric k-sample test for data defined by intervals. Statistica Neerlandica, 37(2):69–71, June 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1983.
tb00801.x.

Falk:1983:RED

[1091] M. Falk. Relative efficiency and deficiency of kernel type estimators of smooth distribution functions. Statistica Neerlandica, 37(2): 73–83, June 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00802.x.

vanBueren:1983:RPD

[1092] W. M. Lammerts van Bueren. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 37(2):85, June 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1983.tb00803.x.

Anonymous:1983:PSb

[1093] Anonymous. Problem section. Statistica Neerlandica, 37(2):87–93, June 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00804.x.

VanEs:1983:RNG

[1094] A. J. Van Es, R. D. Gill, and C. Van Putten. Random number generators for a pocket calculator. Statistica Neerlandica, 37(3):95–102, September 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00805.x.

Kardaun:1983:SSA

[1095] O. Kardaun. Statistical survival analysis of male larynx-cancer patients — a case study. Statistica Neerlandica, 37(3):103–125, September 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00806.x.

VanDerGenugten:1983:ABE

[1096] B. B. Van Der Genugten. The asymptotic behaviour of the estimated generalized least squares method in the linear regression model. Statistica Neerlandica, 37(3):127–141, September 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1983.tb00807.x.

Smit:1983:DCS

[1097] J. C. Smit and W. Vervaat. On divergence and convergence of sums of nonnegative random variables. Statistica Neerlandica, 37(3):143–147, September 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00808.x.

Anonymous:1983:PSc

[1098] Anonymous. Problem section. Statistica Neerlandica, 37(3):149–160, September 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00809.x.

deLeeuw:1983:PCA

[1099] J. de Leeuw. On the prehistory of Correspondence Analysis. Statistica Neerlandica, 37(4):161–164, December 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1983.tb00810.x.

Dharmadhikari:1983:MMM

[1100] S. W. Dharmadhikari and K. Joag-dev. Mean, median, mode III. Statistica Neerlandica, 37(4):165–168, December 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1983.
tb00811.x.

Tigelaar:1983:ISS

[1101] H. H. Tigelaar. The informative sample size for dynamic multiple equation systems with moving average errors. Statistica Neerlandica, 37 (4):169–179, December 1983. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1983.tb00812.x.

Dzhaparidze:1983:IPA

[1102] K. O. Dzhaparidze. On iterative procedures of asymptotic inference. Statistica Neerlandica, 37(4):181–189, December 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1983.
tb00813.x.

Anonymous:1983:PSd

[1103] Anonymous. Problem section. Statistica Neerlandica, 37(4):190–196, December 1983. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1983.tb00814.x.

Mukherjee:1984:MER

[1104] D. Mukherjee and D. C. Hurst. Maximum entropy revisited. Statistica Neerlandica, 38(1):1–12, March 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1984.tb01091.x.

Gupta:1984:DOS

[1105] D. Gupta and R. C. Gupta. On the distribution of order statistics for a random sample size. Statistica Neerlandica, 38(1):13–19, March 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01092.x.

Albers:1984:ACI

[1106] Willem Albers and Peter Löhnberg. An approximate confidence interval for the difference between quantiles in a bio-medical problem. Statistica Neerlandica, 38(1):20–22, March 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1984.tb01093.x.

deKok:1984:AMS

[1107] A. G. de Kok. Algorithmic methods for single server systems with repeated attempts. Statistica Neerlandica, 38(1):23–32, March 1984. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1984.
tb01094.x.

Dijkstra:1984:STU

[1108] Jan B. Dijkstra, Tjeu J. M. Rietjens, and Fred W. Steutel. A simple test for uniformity. Statistica Neerlandica, 38(1):33–44, March 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1984.
tb01095.x.

Anonymous:1984:RPDa

[1109] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 38(1):45–47, March 1984. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1984.tb01096.x.

Anonymous:1984:PSa

[1110] Anonymous. Problem section. Statistica Neerlandica, 38(1):48–52, March 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01097.x.

Anonymous:1984:LRR

[1111] Anonymous. List of recent referees. Statistica Neerlandica, 38(1): 53, March 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01098.x.

Snijders:1984:AVM

[1112] Tom A. B. Snijders. Antithetic variates for Monte Carlo estimation of probabilities. Statistica Neerlandica, 38(2):55–73, June 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1984.
tb01099.x.

Brown:1984:HLE

[1113] L. D. Brown, F. H. Ruymgaart, and D. R. Truax. Hodges–Lehmann efficacies for likelihood ratio type tests in curved bivariate normal families. Statistica Neerlandica, 38(2):75–90, June 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1984.tb01100.x.

Oosterloo:1984:CIT

[1114] Sebie Oosterloo. Confidence intervals for test information and relative efficiency. Statistica Neerlandica, 38(2):91–107, June 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1984.
tb01101.x.

Does:1984:ABS

[1115] Ronald J. M. M. Does. The asymptotic behavior of simple linear rank statistics. Statistica Neerlandica, 38(2):109–130, June 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1984.
tb01102.x.

Anonymous:1984:RPDb

[1116] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 38(2):131, June 1984. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1984.tb01103.x.

Anonymous:1984:PSb

[1117] Anonymous. Problem section. Statistica Neerlandica, 38(2):133– 138, June 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01104.x.

Anonymous:1984:PJH

[1118] Anonymous. Professor J. Hemelrijk Emeritus. Statistica Neerlandica, 38 (2):139, June 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01105.x.

Jansen:1984:RAR

[1119] Marlies E. Jansen. Ridit analysis, a review. Statistica Neerlandica, 38 (3):141–158, September 1984. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1984.tb01106.x.

Engel:1984:MRD

[1120] Jan Engel. Models for response data showing extra-Poisson variation. Statistica Neerlandica, 38(3):159–167, September 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1984.
tb01107.x.

Rothe:1984:AAE

[1121] Günter Rothe. Asymptotic absolute efficiency of projection tests with an application to N-ranking tests. Statistica Neerlandica, 38(3):169– 187, September 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01108.x.

deKroon:1984:CPG

[1122] J. de Kroon and P. van der Laan. A comparison of the powers of a generalized Friedman’s rank test and an aligned rank procedure based on simulation. Statistica Neerlandica, 38(3):189–198, September 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01109.x.

Boxma:1984:JAD

[1123] O. J. Boxma. The joint arrival and departure process for the M/M/L queue. Statistica Neerlandica, 38(3):199–208, September 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1984.
tb01110.x.

Anonymous:1984:RPDc

[1124] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 38(3):209–211, September 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1984.tb01111.x.

Anonymous:1984:PSc

[1125] Anonymous. Problem section. Statistica Neerlandica, 38(3):213–217, September 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01112.x.

Odink:1984:TVM

[1126] J. G. Odink and E. van Imhoff. True versus measured Theil inequality. Statistica Neerlandica, 38(4):219–232, December 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1984.
tb01113.x.

Albers:1984:AEB

[1127] W. Albers and H. J. Veldman. Adaptive estimation of binomial probabilities under misclassification. Statistica Neerlandica, 38(4):233–247, December 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01114.x.

Consul:1984:DOS

[1128] P. C. Consul. On the distributions of order statistics for a random sample size. Statistica Neerlandica, 38(4):249–256, December 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01115.x.

Saxena:1984:BEG

[1129] Ashok K. Saxena. Bayesian estimation of T–G ridge model. Statistica Neerlandica, 38(4):257–260, December 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1984.tb01116.x.

Overdijk:1984:EMF

[1130] D. A. Overdijk. Embedded matrices for finite Markov chains. Statistica Neerlandica, 38(4):261–272, December 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1984.tb01117.x.

vanSlijpe:1984:RWR

[1131] A. R. D. van Slijpe. Random walks on regular polyhedra and other distance-regular graphs. Statistica Neerlandica, 38(4):273–292, December 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01118.x.

Anonymous:1984:PSd

[1132] Anonymous. Problem section. Statistica Neerlandica, 38(4):293–298, December 1984. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1984.tb01119.x.

Geurts:1985:SSS

[1133] J. H. J. Geurts. Some small-sample non-proportional hazards results for the Kaplan–Meier estimator. Statistica Neerlandica, 39(1):1– 13, March 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01120.x.

Dik:1985:DGQ

[1134] J. J. Dik and M. C. M. de Gunst. The distribution of general quadratic forms in normal variables. Statistica Neerlandica, 39(1):14– 26, March 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01121.x.

Engel:1985:AVB

[1135] J. Engel. On the analysis of variance for beta-binomial responses. Statistica Neerlandica, 39(1):27–34, March 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1985.tb01122.x.

Rayner:1985:CSN

[1136] J. C. W. Rayner, D. J. Best, and K. G. Dodds. The construction of the simple χ^{2} and Neyman smooth goodness of fit tests. Statistica Neerlandica, 39(1):35–50, March 1985. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1985.tb01123.x.

Prabhu–Ajgaonkar:1985:IUB

[1137] S. G. Prabhu-Ajgaonkar and S. M. Shah. Invariance of the unordered best estimator in the T -class for Midzuno’s and for Ikeda–Sen’s sampling procedures. Statistica Neerlandica, 39(1):51–56, March 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1985.
tb01124.x.

Anonymous:1985:PSa

[1138] Anonymous. Problem section. Statistica Neerlandica, 39(1):57–61, March 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01125.x.

deKroon:1985:GPR

[1139] J. P. M. de Kroon. A grandson pays respects to his grandfather. Statistica Neerlandica, 39(1):63, March 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01126.x.

Doornbos:1985:FYS

[1140] R. Doornbos. Forty years of statistics and operations research in The Netherlands. Statistica Neerlandica, 39(2):65–72, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1985.
tb01127.x.

Stamhuis:1985:VVS

[1141] I. H. Stamhuis. The “Vereeniging Voor De Statistiek”, 1857–1892. Statistica Neerlandica, 39(2):73–79, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01128.x.

vanZwet:1985:VHR

[1142] W. R. van Zwet. Van de Hulst on robust statistics: a historical note. Statistica Neerlandica, 39(2):81–95, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1985.tb01129.x.

Lenstra:1985:C

[1143] J. K. Lenstra. Corrigenda. Statistica Neerlandica, 39(2):96, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (elec-

tronic).  URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01130.x.

Lageweg:1985:SIP

[1144] B. J. Lageweg, J. K. Lenstra, A. H. G. Rinnooy Kan, and L. Stougie. Stochastic integer programming by dynamic programming. Statistica Neerlandica, 39(2):97–113, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01131.x.

Oosterhoff:1985:CCC

[1145] J. Oosterhoff. The choice of cells in chi-square tests. Statistica Neerlandica, 39(2):115–128, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01132.x.

Bekker:1985:EVE

[1146] P. A. Bekker, T. J. Wansbeek, and A. Kapteyn. Errors in variables in econometrics: New developments and recurrent themes. Statistica Neerlandica, 39(2):129–141, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01133.x.

Cohen:1985:SEQ

[1147] J. W. Cohen and O. J. Boxma. A survey of the evolution of queueing theory. Statistica Neerlandica, 39(2):143–158, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1985.
tb01134.x.

Corsten:1985:CSI

[1148] L. C. A. Corsten. Current statistical issues in agricultural research. Statistica Neerlandica, 39(2):159–168, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1985.tb01135.x.

Molenaar:1985:SSB

[1149] I. W. Molenaar. Statistics in the social and behavioral sciences. Statistica Neerlandica, 39(2):169–179, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01136.x.

Runnenburg:1985:PTA

[1150] J. Th. Runnenburg. Probability theory and its applications. Statistica Neerlandica, 39(2):181–189, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01137.x.

vanStrik:1985:BM

[1151] R. van Strik, R. van Stnk, and P. D. Bezemer. Biostatistics in medicine. Statistica Neerlandica, 39(2):191–201, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1985.
tb01138.x.

Upperman:1985:IS

[1152] P. M. Upperman. Industrial statistics. Statistica Neerlandica, 39(2): 203–217, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01139.x.

vanderWal:1985:MDP

[1153] J. van der Wal and J. Wessels. Markov decision processes. Statistica Neerlandica, 39(2):219–233, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01140.x.

Anonymous:1985:RPD

[1154] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 39(2):235–237, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01141.x.

Anonymous:1985:PSb

[1155] Anonymous. Problem section. Statistica Neerlandica, 39(2):239– 246, June 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01142.x.

Nixdorf:1985:CLT

[1156] R. Nixdorf. Central limit theorms in C[0, 1] for a class of estimators of a distribution function. Statistica Neerlandica, 39(3):251–260, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01143.x.

Bron:1985:SMA

[1157] K. Bron, B. Kosse, K. Poortema, W. Schaafsma, M. J. A. de Voigt, and S. Y. Van Der Werf. Some methodological aspects of estimating the first few moments of the distribution of the gamma ray multiplicity. Statistica Neerlandica, 39(3):261–277, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1985.

tb01144.x. See comments [1158].

vdGenugten:1985:CSM

[1158] B. B. v.d. Genugten and H. H. Tigelaar. Comments on: Some methodological aspect of estimating the first few moments of the distribution of the gamma ray multiplicity. Statistica Neerlandica, 39(3):278–281, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01145.x. See [1157].

Hoekstra:1985:DEF

[1159] J. A. Hoekstra. Designing experiments to fit and validate regression models. Statistica Neerlandica, 39(3):283–290, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1985.
tb01146.x.

Arsham:1985:PPA

[1160] H. Arsham and D. R. Miller. A Poisson process approximation for generalized K–S confidence reg. Statistica Neerlandica, 39(3):291–302, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01147.x.

Laubscher:1985:IFU

[1161] N. F. Laubscher. The indicator function: a useful device for determining certain minimal sufficient statistics. Statistica Neerlandica, 39(3):303– 307, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01148.x.

Upadhyaya:1985:EPM

[1162] L. N. Upadhyaya, H. P. Singh, and J. W. E. Vos. On the estimation of population means and ratios using supplementary information. Statistica Neerlandica, 39(3):309–318, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1985.tb01149.x.

Dekker:1985:RPD

[1163] R. Dekker. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 39(3):319–320, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1985.tb01150.x.

Anonymous:1985:PSc

[1164] Anonymous. Problem section. Statistica Neerlandica, 39(3):321–325, September 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01151.x.

Eiselt:1985:EPS

[1165] H. A. Eiselt and C.-L. Sandblom. External pivoting in the simplex algorithm. Statistica Neerlandica, 39(4):327–341, December 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1985.
tb01152.x.

Barnett:1985:BED

[1166] V. Barnett. The bivariate exponential distribution; a review and some new results. Statistica Neerlandica, 39(4):343–356, December 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01153.x.

Dwass:1985:LTB

[1167] M. Dwass. A limit theorem for Bernoulli RV’s and Feller’s shoe problem. Statistica Neerlandica, 39(4):357–360, December 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1985.
tb01154.x.

Clarkson:1985:DPP

[1168] D. B. Clarkson and D. B. Wolfson. A displaced Poisson process applied to screening for breast cancer in stratified samples with false negatives. Statistica Neerlandica, 39(4):361–373, December 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01155.x.

Yang:1985:BCD

[1169] S-S. Yang. On bootstrapping a class of differentiable statistical functionals with applications to L- and M-estimates. Statistica Neerlandica, 39(4):375–385, December 1985. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1985.tb01156.x.

Hombas:1985:CND

[1170] V. C. Hombas. Characterizing the normal density as a solution of a differential equation. Statistica Neerlandica, 39(4):387–388, December 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01157.x.

Anonymous:1985:PSd

[1171] Anonymous. Problem section. Statistica Neerlandica, 39(4):389–393, December 1985. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1985.tb01158.x.

Rothe:1986:LTT

[1172] G. Rothe. Linear trend test versus global test: a comparison. Statistica Neerlandica, 40(1):1–11, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01159.x.

Fidler:1986:ATP

[1173] V. Fidler and N. J. D. Nagelkerke. Agreement on a two-point scale. Statistica Neerlandica, 40(1):13–20, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1986.tb01160.x.

Engel:1986:SPD

[1174] J. Engel. Split-plot design: Model and analysis for count data. Statistica Neerlandica, 40(1):21–33, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01161.x.

Stam:1986:NSR

[1175] A. J. Stam. A note on sampling with and without replacement. Statistica Neerlandica, 40(1):35–38, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01162.x.

deHaan:1986:SPA

[1176] L. de Haan. A stochastic process that is autoregressive in two directions of time. Statistica Neerlandica, 40(1):39–45, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1986.
tb01163.x.

Hempenius:1986:MFC

[1177] A. L. Hempenius and P. G. H. Mulder. Multiple failure causes and observations of time dependent covariables. Statistica Neerlandica, 40(1): 47–63, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01164.x.

Thavaneswaran:1986:MRA

[1178] A. Thavaneswaran. Model reference adaptive system estimates for counting processes. Statistica Neerlandica, 40(1):65–72, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1986.
tb01165.x.

Schriever:1986:RPD

[1179] B. F. Schriever and P. M. de Jong. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 40(1):73–74, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1986.
tb01166.x.

Anonymous:1986:PSa

[1180] Anonymous. Problem section. Statistica Neerlandica, 40(1):75–77, March 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01167.x.

vanStrik:1986:M

[1181] Roel van Strik. In Memoriam. Statistica Neerlandica, 40(2):79– 80, June 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01191.x.

Fidler:1986:CTB

[1182] V. Fidler. Change-over trials with binary data: Estimation. Statistica Neerlandica, 40(2):81–86, June 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01192.x.

Abel:1986:NTA

[1183] U. Abel. A nonparametric test against ordered alternatives for data defined by intervals. Statistica Neerlandica, 40(2):87–91, June 1986. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1986.
tb01193.x.

Chakraborti:1986:DFC

[1184] S. Chakraborti and M. M. Desu. A distribution-free confidence interval for the difference between quantiles with censored data. Statistica Neerlandica, 40(2):93–98, June 1986. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1986.tb01194.x.

Lalitha:1986:PMT

[1185] S. Lalitha and P. C. Joshi. Performance of Murphy’s test for two outliers in a linear model. Statistica Neerlandica, 40(2):99–107, June 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01195.x.

Don:1986:CNL

[1186] F. J. H. Don. Constrained non-linear least squares. Statistica Neerlandica, 40(2):109–115, June 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01196.x.

Consul:1986:UGP

[1187] P. C. Consul and F. Famoye. On the unimodality of generalized Poisson distribution. Statistica Neerlandica, 40(2):117–122, June 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1986.
tb01197.x.

Shorack:1986:ARW

[1188] G. R. Shorack and J. Beirlant. The appropriate reduction of the weighted empirical process. Statistica Neerlandica, 40(2):123–128, June 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01198.x.

Lenstra:1986:C

[1189] J. K. Lenstra. Corrigenda. Statistica Neerlandica, 40(2):129–130, June 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01199.x.

Anonymous:1986:PSb

[1190] Anonymous. Problem section. Statistica Neerlandica, 40(2):131– 133, June 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01200.x.

Groeneveld:1986:SWF

[1191] R. A. Groeneveld. Skewness for the Weibull family. Statistica Neerlandica, 40(3):135–140, September 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01509.x.

Consul:1986:UGN

[1192] P. C. Consul and F. Famoye. On the unimodality of the generalized negative binomial distribution. Statistica Neerlandica, 40(3):141–144, September 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01510.x.

Fidler:1986:GSR

[1193] V. Fidler and N. J. D. Nagelkerke. A generalized signed-rank test for comparison of P treatments. Statistica Neerlandica, 40(3):145–155, September 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01511.x.

Lalitha:1986:PSR

[1194] S. Lalitha and P. C. Joshi. Performance of Studentized range statistic for two outliers in a linear model. Statistica Neerlandica, 40(3):157– 168, September 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01512.x.

Heijmans:1986:FOE

[1195] R. D. H. Heijmans and J. R. Magnus. On the first-order efficiency and asymptotic normality of maximum likelihood estimators obtained from dependent observations. Statistica Neerlandica, 40(3):169–188, September 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01513.x.

Anonymous:1986:RPDa

[1196] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 40(3):189–190, September 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01514.x.

Steutel:1986:PS

[1197] F. W. Steutel. Problem section. Statistica Neerlandica, 40(3):191–203, September 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01515.x.

Keen:1986:SME

[1198] A. Keen, J. T. N. M. Thissen, J. A. Hoekstra, and J. Jansen. Successive measurement experiments. Statistica Neerlandica, 40(4):205–223, December 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01201.x.

Loynes:1986:RRR

[1199] R. M. Loynes. Recursive and related residuals. Statistica Neerlandica, 40(4):225–235, December 1986. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1986.tb01202.x.

Cramer:1986:EPM

[1200] J. S. Cramer. Estimation of probability models from income class data. Statistica Neerlandica, 40(4):237–250, December 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1986.
tb01203.x.

Brunner:1986:RTD

[1201] E. Brunner and N. Neumann. Rank tests in 2×2 designs. Statistica Neerlandica, 40(4):251–272, December 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01204.x.

Goovaerts:1986:ORW

[1202] M. J. Goovaerts, M. Vandebroeck, and R. Kaas. Ordering of risks and weighted compound distributions. Statistica Neerlandica, 40(4):273–282, December 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01205.x.

Anonymous:1986:RPDb

[1203] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 40(4):283–285, December 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1986.tb01206.x.

Anonymous:1986:PSc

[1204] Anonymous. Problem section. Statistica Neerlandica, 40(4):287–292, December 1986. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1986.tb01207.x.

Gerards:1987:PCC

[1205] A. M. H. Gerards and A. W. J. Kolen. Polyhedral combinatorics in combinatorial optimization. Statistica Neerlandica, 41(1):1–25, March 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01168.x.

Chaudhuri:1987:RRT

[1206] A. Chaudhuri and R. Mukherjee. Randomized response techniques: a review. Statistica Neerlandica, 41(1):27–44, March 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1987.
tb01169.x.

Fidler:1987:LLM

[1207] V. Fidler. Log-linear mixed models for categorical data. Statistica Neerlandica, 41(1):45–57, March 1987. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1987.tb01170.x.

Schemper:1987:NEV

[1208] M. Schemper. Nonparametric estimation of variance, skewness and kurtosis of the distribution of a statistic by jackknife and bootstrap techniques. Statistica Neerlandica, 41(1):59–64, March 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1987.
tb01171.x.

Einmahl:1987:RPD

[1209] J. H. J. Einmahl. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 41(1):65–67, March 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1987.tb01172.x.

Anonymous:1987:PSa

[1210] Anonymous. Problem section. Statistica Neerlandica, 41(1):69–75, March 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01173.x.

Wallin:1987:WTD

[1211] J. F. E. Wallin. Waiting time distribution for a semaphone system. Statistica Neerlandica, 41(2):77–87, June 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1987.tb01174.x.

Arnold:1987:EPC

[1212] B. F. Arnold and E. von Col Iani. Economic Process Control. Statistica Neerlandica, 41(2):89–97, June 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1987.tb01175.x.

Schriever:1987:MRS

[1213] B. F. Schriever. Monotonicity of rank statistics in some nonparametric testing problems. Statistica Neerlandica, 41(2):99–109, June 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01176.x.

Hering:1987:PME

[1214] F. Hering, G. Trenkler, and P. Stahlecker. Partial minimax estimation in regression analysis. Statistica Neerlandica, 41(2):111–128, June 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01177.x.

Ahsanullah:1987:DPR

[1215] M. Ahsanullah and B. Holland. Distributional properties of record values from the geometric distribution. Statistica Neerlandica, 41(2):129– 137, June 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01178.x.

Anonymous:1987:RPDa

[1216] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 41(2):139–142, June 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1987.tb01179.x.

Anonymous:1987:PSb

[1217] Anonymous. Problem section. Statistica Neerlandica, 41(2):143– 150, June 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01180.x.

Stam:1987:SPA

[1218] A. J. Stam. Statistical problems in ancient numismatics. Statistica Neerlandica, 41(3):151–174, September 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1987.tb01208.x.

Hooghiemstra:1987:PCV

[1219] G. Hooghiemstra. A path construction for the virtual waiting time of an M/G/1 queue. Statistica Neerlandica, 41(3):175–182, September 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01209.x.

Oosterhoff:1987:NCF

[1220] J. Oosterhoff and B. F. Schriever. A note on complete families of distributions. Statistica Neerlandica, 41(3):183–190, September 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01210.x.

vanPutten:1987:CMP

[1221] B. van Putten. On the construction of multivariate permutation tests in the two-sample case. Statistica Neerlandica, 41(3):191–202, September 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01211.x.

Moors:1987:RPS

[1222] J. J. A. Moors. On the relative precision of several allocations in stratified sampling. Statistica Neerlandica, 41(3):203–210, September 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01212.x.

Anonymous:1987:RPDb

[1223] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 41(3):211–213, September 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1987.tb01213.x.

Anonymous:1987:PSc

[1224] Anonymous. Problem section. Statistica Neerlandica, 41(3):215–222, September 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01214.x.

vanMontfort:1987:REV

[1225] K. van Montfort, A. Mooijaart, and J. de Leeuw. Regression with errors in variables: estimators based on third order moments. Statistica Neerlandica, 41(4):223–238, December 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1987.tb01215.x.

Abouammoh:1987:DU

[1226] A. M. Abouammoh. On discrete ∝-unimodality. Statistica Neerlandica, 41(4):239–244, December 1987. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1987.tb01216.x.

deGunst:1987:DGQ

[1227] M. C. M. de Gunst. On the distribution of general quadratic functions in normal vectors. Statistica Neerlandica, 41(4):245–252, December 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01217.x.

Moothathu:1987:ECS

[1228] T. S. K. Moothathu. Exceptional convergence of the sample variance. Statistica Neerlandica, 41(4):253–256, December 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1987.
tb01218.x.

Hillegers:1987:RPD

[1229] L. T. M. E. Hillegers. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 41(4):257–258, December 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1987.tb01219.x.

Steutel:1987:PS

[1230] F. W. Steutel. Problem section. Statistica Neerlandica, 41(4):259–266, December 1987. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1987.tb01220.x.

deVos:1988:LTB

[1231] J. C. de Vos and W. Vervaat. Local times of Bernoulli walk. Statistica Neerlandica, 42(1):1–16, March 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1988.tb01516.x.

Engel:1988:RTR

[1232] J. Engel. Rank tests and random blocking of classified data. Statistica Neerlandica, 42(1):17–27, March 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1988.tb01517.x.

Scarsini:1988:SSD

[1233] M. Scarsini. On a simple sequential decision problem. Statistica Neerlandica, 42(1):29–35, March 1988. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1988.tb01518.x.

Phillips:1988:SAB

[1234] G. D. A. Phillips and B. P. M. McCabe. Some applications for Basil’s independence theorem in testing econometric models. Statistica Neerlandica, 42(1):37–46, March 1988. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1988.tb01519.x.

Evans:1988:CBB

[1235] M. A. Evans. Choosing a base for the BLUS vector. Statistica Neerlandica, 42(1):47–52, March 1988. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1988.tb01520.x.

Alzaid:1988:FOI

[1236] A. Alzaid and M. Al-Osh. First-order integer-valued autoregressive (INAR(1)) process: Distributional and regression properties. Statistica Neerlandica, 42(1):53–61, March 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1988.tb01521.x.

Anonymous:1988:RPDa

[1237] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 42(1):63–66, March 1988. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1988.tb01522.x.

Steutel:1988:PS

[1238] F. W. Steutel. Problem section. Statistica Neerlandica, 42(1):67– 74, March 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01523.x.

Anonymous:1988:CP

[1239] Anonymous. Call for papers. Statistica Neerlandica, 42(1):75–78, March 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01524.x.

Snijders:1988:EID

[1240] T. A. B. Snijders. Editor’s introduction: Discussion on statistics versus data-analysis. Statistica Neerlandica, 42(1):81, March 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1988.
tb01525.x.

Molenaar:1988:FSI

[1241] I. W. Molenaar. Formal statistics and informal data analysis, or why laziness should be discouraged. Statistica Neerlandica, 42(2):83– 90, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01221.x.

deLeeuw:1988:MT

[1242] J. de Leeuw. Models and techniques. Statistica Neerlandica, 42(2): 91–98, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01222.x.

deGruijter:1988:DAS

[1243] D. N. M. de Gruijter. Data analysis and statistics, report of a discussion. Statistica Neerlandica, 42(2):99–102, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1988.
tb01223.x.

Rousseeuw:1988:RSE

[1244] P. J. Rousseeuw and A. M. Leroy. A robust scale estimator based on the shortest half. Statistica Neerlandica, 42(2):103–116, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01224.x.

Driessen:1988:SMS

[1245] S. Driessen. The sieve method, a sampling method for audit practice. Statistica Neerlandica, 42(2):117–136, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1988.tb01225.x.

Steutel:1988:NDU

[1246] F. W. Steutel. Note on discrete a-unimodality. Statistica Neerlandica, 42 (2):137–140, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01226.x. See correction [1247].

Abouammoh:1988:CDU

[1247] A. M. Abouammoh. Correction to “On discrete a-unimodality”. Statistica Neerlandica, 42(2):141, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1988.tb01227.x. See [1246].

vanderLaan:1988:RPD

[1248] B. S. van der Laan. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 42(2):143–144, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1988.tb01228.x.

Anonymous:1988:PSa

[1249] Anonymous. Problem section. Statistica Neerlandica, 42(2):145– 152, June 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01229.x.

Does:1988:ADG

[1250] R. J. M. M. Does, R. Helmers, and C. A. J. Klaassen. Approximating the distribution of Greenwood’s statistic. Statistica Neerlandica, 42(3):153– 162, September 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01230.x.

Timmer:1988:MDV

[1251] E. Timmer. The De Mortibus distribution and variability of biological magnitudes. Statistica Neerlandica, 42(3):163–179, September 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01231.x.

Kriens:1988:NMP

[1252] J. Kriens and J. Th. van Lieshout. Notes on the Markowitz portfolio selection method. Statistica Neerlandica, 42(3):181–191, September 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01232.x.

Ahsanullah:1988:CPO

[1253] M. Ahsanullah. Characteristic properties of order statistics based on random sample size from an exponential distribution. Statistica Neerlandica, 42(3):193–198, September 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1988.tb01233.x.

Anonymous:1988:C

[1254] Anonymous. Corrigenda. Statistica Neerlandica, 42(3):199, September 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01234.x.

Anonymous:1988:RPDb

[1255] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 42(3):201–204, September 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1988.tb01235.x.

Anonymous:1988:PSb

[1256] Anonymous. Problem section. Statistica Neerlandica, 42(3):205–214, September 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01236.x.

vanHouwelingen:1988:LRR

[1257] J. C. van Houwelingen and S. le Cessie. Logistic regression, a review. Statistica Neerlandica, 42(4):215–232, December 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1988.
tb01237.x.

Engel:1988:PLR

[1258] J. Engel. Polytomous logistic regression. Statistica Neerlandica, 42 (4):233–252, December 1988. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1988.tb01238.x.

Morgan:1988:EMQ

[1259] B. J. T. Morgan. Extended models for quantal response data. Statistica Neerlandica, 42(4):253–272, December 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1988.tb01239.x.

Verhelst:1988:LBP

[1260] N. Verhelst and I. W. Molenaar. Logit based parameter estimation in the Rasch model. Statistica Neerlandica, 42(4):273–295, December 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01240.x.

Cramer:1988:LME

[1261] J. S. Cramer and G. Ridder. The logit model in economics. Statistica Neerlandica, 42(4):297–314, December 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1988.tb01241.x.

Anonymous:1988:RPDc

[1262] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 42(4):315–316, December 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1988.tb01242.x.

Anonymous:1988:PSc

[1263] Anonymous. Problem section. Statistica Neerlandica, 42(4):317–322, December 1988. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1988.tb01243.x.

Buishand:1989:SEC

[1264] T. A. Buishand. Statistics of extremes in climatology. Statistica Neerlandica, 43(1):1–30, March 1989. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1989.tb01244.x.

Aarts:1989:SAI

[1265] E. H. L. Aarts and P. J. M. van Laarhoven. Simulated annealing: an introduction. Statistica Neerlandica, 43(1):31–52, March 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1989.
tb01245.x.

Steutel:1989:IIF

[1266] F. W. Steutel and J. G. F. Thiemann. On the independence of integer and fractional parts. Statistica Neerlandica, 43(1):53–59, March 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01246.x.

Anonymous:1989:RPDa

[1267] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 43(1):61–62, March 1989. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1989.tb01247.x.

Anonymous:1989:PSa

[1268] Anonymous. Problem section. Statistica Neerlandica, 43(1):63–67, March 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01248.x.

Corsten:1989:IOL

[1269] L. C. A. Corsten. Interpolation and optimal linear prediction. Statistica Neerlandica, 43(2):69–84, June 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1989.tb01249.x.

vanOmmeren:1989:WTD

[1270] J. C. W. van Ommeren and R. D. Nobel. On the waiting time distribution in a GI/G/1 queue with a Coxian-2 service time distribution. Statistica Neerlandica, 43(2):85–90, June 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1989.tb01250.x.

Rothe:1989:SRT

[1271] G. Rothe. Some rank tests for one-sided many–one comparisons. Statistica Neerlandica, 43(2):91–108, June 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1989.tb01251.x.

Janssen:1989:LAN

[1272] A. Janssen. Local asymptotic normality for randomly censored models with applications to rank tests. Statistica Neerlandica, 43(2):109– 125, June 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01252.x.

Anonymous:1989:RPDb

[1273] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 43(2):127–132, June 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1989.tb01253.x.

Anonymous:1989:PSb

[1274] Anonymous. Problem section. Statistica Neerlandica, 43(2):133– 139, June 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01254.x.

Linssen:1989:AIM

[1275] H. N. Linssen and L. T. M. E. Hillegers. Approximate inference in multivariate nonlinear functional relationships. Statistica Neerlandica, 43 (3):141–156, September 1989. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1989.tb01255.x.

Bekker:1989:ATA

[1276] P. A. Bekker and H. Neudecker. Albert’s theorem applied to problems of efficiency and MSE superiority. Statistica Neerlandica, 43(3):157–167, September 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01256.x.

vanderLaan:1989:SLP

[1277] P. van der Laan. Selection from logistic populations. Statistica Neerlandica, 43(3):169–174, September 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1989.tb01257.x.

Einmahl:1989:SEP

[1278] J. H. J. Einmahl. On the standarized empirical process. Statistica Neerlandica, 43(3):175–179, September 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1989.tb01258.x.

Anonymous:1989:RPDc

[1279] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 43(3):181–184, September 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1989.tb01259.x.

Anonymous:1989:PSc

[1280] Anonymous. Problem section. Statistica Neerlandica, 43(3):185–190, September 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01260.x.

Wartna:1989:MJH

[1281] J. A. Wartna. In memoriam: Johannes Hubertus Cornelius Lisman (1908–13 April 1989). Statistica Neerlandica, 43(3):191–192, September 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01261.x.

Kleijnen:1989:RSS

[1282] J. P. C. Kleijnen, J. Kriens, H. Timmermans, and H. van den Wildenberg. Regression sampling in statistical auditing: a practical survey and evaluation. Statistica Neerlandica, 43(4):193–209, December 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01262.x.

deVos:1989:RSS

[1283] A. F. de Vos. On regression sampling in statistical auditing: Bad answers to the wrong questions? Statistica Neerlandica, 43(4):211–223, December 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01263.x.

Kleunen:1989:R

[1284] J. P. C. Kleunen, J. Kreins, H. Timmermans, and H. Van Den Wildenberg. Rejoinder. Statistica Neerlandica, 43(4):225, December 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01264.x.

Pollock:1989:MAM

[1285] D. S. G. Pollock. The misspecification of Arma models. Statistica Neerlandica, 43(4):227–244, December 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1989.tb01265.x.

Paardekooper:1989:NPI

[1286] H. C. H. Paardekooper, H. B. A. Steens, and G. van der Hoek. A note on properties of iterative procedures of asymptotic inference. Statistica Neerlandica, 43(4):245–253, December 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1989.tb01266.x.

Anonymous:1989:PSd

[1287] Anonymous. Problem section. Statistica Neerlandica, 43(4):255–258, December 1989. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1989.tb01267.x.

deLeeuw:1990:LTB

[1288] J. de Leeuw, P. G. M. van der Heijden, and P. Verboon. A latent time– budget model. Statistica Neerlandica, 44(1):1–22, March 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1990.
tb01268.x.

Dean:1990:NER

[1289] A. M. Dean and D. A. Wolfe. A note on exchangeability of random variables. Statistica Neerlandica, 44(1):23–27, March 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1990.
tb01269.x.

Alzaid:1990:SRD

[1290] A. A. Alzaid and D. A. Al-Osh. Some results on discrete amonotonicity. Statistica Neerlandica, 44(1):29–33, March 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1990.
tb01270.x.

Anonymous:1990:RPDa

[1291] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 44(1):35–38, March 1990. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1990.tb01271.x.

Anonymous:1990:PSa

[1292] Anonymous. Problem section. Statistica Neerlandica, 44(1):39–43, March 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01272.x.

Anonymous:1990:C

[1293] Anonymous. Corrigenda. Statistica Neerlandica, 44(1):44, March 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (elec-

tronic).  URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01273.x.

deHaan:1990:FAE

[1294] L. de Haan. Fighting the arch-enemy with mathematics. Statistica Neerlandica, 44(2):45–68, June 1990. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1990.tb01526.x.

Windmeijer:1990:ADS

[1295] F. A. G. Windmeijer. The asymptotic distribution of the sum of weighted squared residuals in binary choice models. Statistica Neerlandica, 44(2): 69–78, June 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01527.x.

Einmahl:1990:EDF

[1296] J. H. J. Einmahl. The empirical distribution function as a tail estimator. Statistica Neerlandica, 44(2):79–82, June 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1990.tb01528.x.

Deng:1990:CED

[1297] Lih-Yuan Deng and Raj S. Chhikara. On the characterization of the exponential distribution by the independence of its integer and fractional parts. Statistica Neerlandica, 44(2):83–85, June 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1990.
tb01529.x.

Bristol:1990:DFC

[1298] D. R. Bristol. Distribution–free confidence intervals for the difference between quantiles. Statistica Neerlandica, 44(2):87–90, June 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1990.
tb01530.x.

Anonymous:1990:PSb

[1299] Anonymous. Problem section. Statistica Neerlandica, 44(2):91–101, June 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01531.x.

Anonymous:1990:ORL

[1300] Anonymous. Operations research in logistics: Editorial introduction. Statistica Neerlandica, 44(3):103–104, September 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1990.
tb01274.x.

Boender:1990:BLP

[1301] C. G. E. Boender and A. H. G. Rinnooy. A Bayesian learning procedure for the (s, Q) inventory policy. Statistica Neerlandica, 44(3):105– 114, September 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01275.x.

Dessouky:1990:SIJ

[1302] M. I. Dessouky, B. J. Lageweg, J. K. Lenstra, and S. L. van de Velde. Scheduling identical jobs on uniform parallel machines. Statistica Neerlandica, 44(3):115–123, September 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1990.tb01276.x.

Spieksma:1990:SSS

[1303] F. C. R. Spieksma, K. Vrieze, and A. G. Oerlemans. On the system setup and the scheduling problem in a Flexible Manufacturing System (FMS). Statistica Neerlandica, 44(3):125–138, September 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01277.x.

Savelsbergh:1990:PIH

[1304] M. W. P. Savelsbergh. A parallel insertion heuristic for vehicle routing with side constraints. Statistica Neerlandica, 44(3):139–148, September 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01278.x.

Langenhoff:1990:ATM

[1305] L. J. G. Langenhoff and W. H. M. Zijm. An analytical theory of multiechelon production/distribution systems. Statistica Neerlandica, 44(3): 149–174, September 1990. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1990.tb01279.x.

Steerneman:1990:PS

[1306] A. G. M. Steerneman. Problem section. Statistica Neerlandica, 44 (3):175–183, September 1990. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1990.tb01280.x.

vanderLaan:1990:VDA

[1307] P. van der Laan. The Van Dantzig Award. Statistica Neerlandica, 44 (4):185–193, December 1990. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1990.tb01281.x.

Engel:1990:AUL

[1308] B. Engel. The analysis of unbalanced linear models with variance components. Statistica Neerlandica, 44(4):195–219, December 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01282.x.

Engel:1990:QLI

[1309] J. Engel. Quasi-likelihood inference in a generalized linear mixed model for balanced data. Statistica Neerlandica, 44(4):221–239, December 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01283.x.

Strijbosch:1990:MDD

[1310] L. W. G. Strijbosch, R. J. M. M. Does, and W. Albers. Multiple-dose design and bias-reducing methods for limiting dilution assays. Statistica Neerlandica, 44(4):241–261, December 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1990.tb01284.x.

Shorack:1990:SPR

[1311] G. R. Shorack. A simple proof of a rate of embedding for the partial sum process. Statistica Neerlandica, 44(4):263–266, December 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01285.x.

Anonymous:1990:RPDb

[1312] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 44(4):267–274, December 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1990.tb01286.x.

Anonymous:1990:PSc

[1313] Anonymous. Problem section. Statistica Neerlandica, 44(4):275–283, December 1990. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1990.tb01287.x.

Kyriakoussis:1991:CLS

[1314] A. Kyriakoussis and H. Papageorgiou. Characterizations of logarithmic series distributions. Statistica Neerlandica, 45(1):1–8, March 1991. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1991.
tb01288.x.

Boscher:1991:CLR

[1315] H. Boscher. Contamination in linear regression models and its influence on estimators. Statistica Neerlandica, 45(1):9–19, March 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1991.
tb01289.x.

Ahsanullah:1991:RVL

[1316] M. Ahsanullah. Record values of the Lomax distribution. Statistica Neerlandica, 45(1):21–29, March 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1991.tb01290.x.

Romeijn:1991:SBA

[1317] H. E. Romeijn. Shake-and-bake algorithms for the identification of nonredundant linear inequalities. Statistica Neerlandica, 45(1):31–50, March 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01291.x.

Jansen:1991:CBE

[1318] M. G. H. Jansen and T. A. B. Snijders. Comparisons of Bayesian estimation procedures for two-way contingency tables without interaction. Statistica Neerlandica, 45(1):51–65, March 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1991.tb01292.x.

Anonymous:1991:PS

[1319] Anonymous. Problem section. Statistica Neerlandica, 45(1):67–68, March 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01293.x.

Anonymous:1991:GEP

[1320] Anonymous. Guest Editors’ preface. Statistica Neerlandica, 45(1):71– 72, March 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01294.x.

Kmenta:1991:LVE

[1321] J. Kmenta. Latent variables in econometrics. Statistica Neerlandica, 45 (2):73–84, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01295.x.

Wansbeek:1991:MEP

[1322] T. J. Wansbeek and R. H. Koning. Measurement error and panel data^{1}. Statistica Neerlandica, 45(2):85–92, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1991.tb01296.x.

Friedmann:1991:AAF

[1323] R. Friedmann, H. J. Mittag, and A. Brandtstater. An application of approximate finite sample results to parameter estimation in a linear errors-in-variables model. Statistica Neerlandica, 45(2):93– 106, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01297.x.

Bekker:1991:RAD

[1324] P. A. Bekker, K. van Montfort, and A. Mooijaart. Regression analysis with dichotomous regressors and misclassification. Statistica Neerlandica, 45(2):107–119, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1991.tb01298.x.

Wegge:1991:ILV

[1325] L. L. Wegge. Identification with latent variables. Statistica Neerlandica, 45(2):121–143, June 1991. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1991.tb01299.x.

Schneeweiss:1991:MLV

[1326] H. Schneeweiss. Models with latent variables: LISREL versus PLS. Statistica Neerlandica, 45(2):145–157, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1991.tb01300.x.

Mooijaart:1991:RNT

[1327] A. Mooijaart and P. M. Bentler. Robustness of normal theory statistics in structural equation models. Statistica Neerlandica, 45(2):159– 171, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01301.x.

Satorra:1991:CSA

[1328] A. Satorra, W. E. Saris, and W. M. de Pijper. A comparison of several approximations to the power function of the likelihood ratio test in covariance structure analysis. Statistica Neerlandica, 45(2):173– 185, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01302.x.

Hsiao:1991:SRM

[1329] C. Hsiao and G. Taylor. Some remarks on measurement errors and the identification of panel data models. Statistica Neerlandica, 45(2): 187–194, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01303.x.

Faber:1991:CCA

[1330] J. Faber. Cooperation and conflict among nations: an application of multi-sample confirmatory factor analysis. Statistica Neerlandica, 45(2): 195–206, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01304.x.

Anonymous:1991:RPDa

[1331] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 45(2):207–216, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1991.tb01305.x.

Steerneman:1991:PSa

[1332] Ton Steerneman. Problem section. Statistica Neerlandica, 45(2):217– 225, June 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01306.x.

Chakraborti:1991:LRT

[1333] S. Chakraborti and M. M. Desu. Linear rank tests for comparing treatments with a control when data are subject to unequal patterns of censorship. Statistica Neerlandica, 45(3):227–254, September 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01307.x.

Eilers:1991:NDE

[1334] P. H. C. Eilers. Nonparametric density estimation with grouped observations. Statistica Neerlandica, 45(3):255–269, September 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1991.
tb01308.x.

Lee:1991:RML

[1335] K. R. Lee and C. H. Kapadia. Restricted maximum likelihood estimation under Eisenhart model ill. Statistica Neerlandica, 45(3):271–282, September 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01309.x.

Frank:1991:SAC

[1336] O. Frank. Statistical analysis of change in networks. Statistica Neerlandica, 45(3):283–293, September 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1991.tb01310.x.

Nowicki:1991:ADR

[1337] K. Nowicki. Asymptotic distributions in random graphs with applications to social networks. Statistica Neerlandica, 45(3):295–325, September 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01311.x.

Revfeim:1991:ACI

[1338] K. J. A. Revfeim. Approximation for the cumulative and inverse gamma distribution. Statistica Neerlandica, 45(3):327–331, September 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01312.x.

Anonymous:1991:RPDb

[1339] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 45(3):333–335, September 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1991.tb01313.x.

Steerneman:1991:PSb

[1340] Ton Steerneman. Problem section. Statistica Neerlandica, 45(3):337– 343, September 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01314.x.

Barendregt:1991:SAS

[1341] L. G. Barendregt and M. J. Rottschäfer. A statistical analysis of spatial point patterns a case study. Statistica Neerlandica, 45(4):345–363, December 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01315.x.

deVos:1991:RME

[1342] A. F. de Vos. On randomization, modeling and experimental design; a new example in an old discussion. Statistica Neerlandica, 45(4):365– 377, December 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01316.x. See remarks [1398].

Christmann:1991:SLL

[1343] A. Christmann and U. Gather. Significance levels of linear rank tests using Edgeworth expansions. Statistica Neerlandica, 45(4):379–389, December 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01317.x.

Scherrer:1991:SSL

[1344] W. Scherrer, M. Deistler, M. Kopel, and W. Reitgruber. Solution sets for linear dynamic errors-in-variables models. Statistica Neerlandica, 45 (4):391–404, December 1991. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1991.tb01318.x.

Neudecker:1991:SUR

[1345] H. Neudecker and F. A. G. Windmeijer. R^{2} in seemingly unrelated regression equations. Statistica Neerlandica, 45(4):405–411, December 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01319.x.

Steerneman:1991:PSc

[1346] Ton Steerneman. Problem section. Statistica Neerlandica, 45(4):413– 419, December 1991. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1991.tb01320.x.

Keller:1992:GEP

[1347] W. J. Keller and P. Kooiman. Guest Editors’ preface. Statistica Neerlandica, 46(1):1–2, March 1992. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1992.tb01321.x.

Anonymous:1992:OST

[1348] Anonymous. Opening speech by Tore Dalenius. Statistica Neerlandica, 46(1):3–4, March 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01322.x.

Keller:1992:DPM

[1349] W. J. Keller and J. G. Bethlehem. Disclosure protection of microdata: problems and solutions. Statistica Neerlandica, 46(1):5–19, March 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01323.x.

Skinner:1992:IDP

[1350] C. J. Skinner. On identification disclosure and prediction disclosure for microdata. Statistica Neerlandica, 46(1):21–32, March 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1992.
tb01324.x.

Greenberg:1992:SMR

[1351] B. V. Greenberg and L. V. Zayatz. Strategies for measuring risk in public use microdata files. Statistica Neerlandica, 46(1):33–48, March 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01325.x.

Mokken:1992:DRM

[1352] R. J. Mokken, P. Kooiman, J. Pannekoek, and L. C. R. J. Willenborg. Disclosure risks for microdata. Statistica Neerlandica, 46(1): 49–67, March 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01326.x.

Blien:1992:DRM

[1353] U. Blien, H. Wirth, and M. Muller. Disclosure risk for microdata stemming from official statistics. Statistica Neerlandica, 46(1):69–82, March 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01327.x.

Anonymous:1992:RPDa

[1354] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 46(1):83–84, March 1992. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1992.tb01328.x.

Steerneman:1992:PS

[1355] Ton Steerneman. Problem section. Statistica Neerlandica, 46(1):85– 94, March 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01329.x.

Anonymous:1992:PC

[1356] Anonymous. Production changes. Statistica Neerlandica, 46(1):95, March 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01330.x.

Kroonenberg:1992:MAV

[1357] P. M. Kroonenberg. In memoriam Albert Verbeek (1946–1990). Statistica Neerlandica, 46(2–3):97–105, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1992.tb01331.x.

Verbeek:1992:CGL

[1358] A. Verbeek. The compactification of generalized linear models. Statistica Neerlandica, 46(2–3):107–142, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1992.tb01332.x.

Aboukalam:1992:RDC

[1359] M. A. F. Aboukalam and H. Ai-Nachawati. Robust dynamic clustering. Statistica Neerlandica, 46(2–3):143–152, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1992.
tb01333.x.

vanderLaan:1992:SSL

[1360] P. van der Laan. On subset selection from logistic populations. Statistica Neerlandica, 46(2–3):153–163, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1992.tb01334.x.

Heesterman:1992:CLT

[1361] C. C. Heesterman and R. D. Gill. A central limit theorem for Mestimators by the von Mises method. Statistica Neerlandica, 46(2–3): 165–177, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01335.x.

vanderVaart:1992:ALM

[1362] A. W. van der Vaart. Asymptotic linearity of minimax estimators. Statistica Neerlandica, 46(2–3):179–194, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1992.tb01336.x.

Deng:1992:RSN

[1363] L. Y. Deng and R. S. Chhikara. Robustness of some non-uniform random variate generators. Statistica Neerlandica, 46(2–3):195–207, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01337.x.

Pakes:1992:CGM

[1364] A. G. Pakes. A characterization of gamma mixtures of stable laws motivated by limit theorems. Statistica Neerlandica, 46(2–3):209– 218, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01338.x.

Anonymous:1992:RPDb

[1365] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 46(2–3):219–222, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1992.tb01339.x.

Anonymous:1992:PSa

[1366] Anonymous. Problem section. Statistica Neerlandica, 46(2–3):223– 227, July 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01340.x.

Does:1992:VCC

[1367] R. J. M. M. Does and B. F. Schriever. Variables control chart limits and tests for special causes. Statistica Neerlandica, 46(4):229–245, December 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01341.x.

Weba:1992:NEN

[1368] M. Weba. A note on the estimation of non-differentiable functions of multinomial probabilities. Statistica Neerlandica, 46(4):247–249, December 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01342.x.

Plachky:1992:ELM

[1369] D. Plachky. On the existence of locally mvu estimators. Statistica Neerlandica, 46(4):251–253, December 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1992.tb01343.x.

Majsnerowska:1992:NMP

[1370] M. Majsnerowska. A note on mixed Poisson approximation. Statistica Neerlandica, 46(4):255–258, December 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1992.tb01344.x.

Hoogendoorn:1992:EWU

[1371] A. W. Hoogendoorn. Estimating the Weight Undersize Distribution for the Wicksell problem. Statistica Neerlandica, 46(4):259–282, December 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01345.x.

Rey:1992:DSL

[1372] W. J. J. Rey. Descriptive statistics of large data sets by scatter plots, an exploratory approach. Statistica Neerlandica, 46(4):283–297, December 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01346.x.

Anonymous:1992:PSb

[1373] Anonymous. Problem section. Statistica Neerlandica, 46(4):299–300, December 1992. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1992.tb01347.x.

Grubel:1993:GEP

[1374] R. Grubel. Guest Editor’s preface. Statistica Neerlandica, 47(1):1– 2, March 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01401.x.

Nelder:1993:ASC

[1375] J. A. Nelder. Aspects of statistical computing, past, present and future. Statistica Neerlandica, 47(1):3–8, March 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1993.tb01402.x.

Kendall:1993:CAP

[1376] W. S. Kendall. Computer algebra in probability and statistics. Statistica Neerlandica, 47(1):9–25, March 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1993.tb01403.x.

Rocke:1993:CRE

[1377] D. M. Rocke and D. L. Woodruff. Computation of robust estimates of multivariate location and shape. Statistica Neerlandica, 47(1):27– 42, March 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01404.x.

Hardle:1993:BAA

[1378] W. Härdle and P. Hall. On the backfitting algorithm for additive regression models. Statistica Neerlandica, 47(1):43–57, March 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1993.
tb01405.x.

Embrechts:1993:SAF

[1379] P. Embrechts, R. Grübel, and S. M. Pitts. Some applications of the fast Fourier transform algorithm in insurance mathematics. Statistica Neerlandica, 47(1):59–75, March 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1993.tb01406.x. This paper is

dedicated to Professor W. S. Jewell on the occasion of his 60th birthday.

Dol:1993:RPD

[1380] Wietse Dol. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 47(1):77–80, March 1993. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1993.tb01407.x.

Haneveld:1993:PS

[1381] L. A. Klein Haneveld. Problem section. Statistica Neerlandica, 47(1): 81–88, March 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01408.x.

Poortema:1993:EGS

[1382] K. Poortema, R. Y. J. Tarnminga, and W. Schaafsma. Estimation in a growth study with irregular measurement times. Statistica Neerlandica, 47(2):89–95, June 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01409.x.

leCessie:1993:BLM

[1383] S. le Cessie and J. C. van Houwelingen. Building logistic models by means of a non parametric goodness of fit test: a case study. Statistica Neerlandica, 47(2):97–109, June 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1993.tb01410.x.

vanderGenugten:1993:EIW

[1384] B. B. van der Genugten. Efficiency of iterated WLS in the linear model with completely unknown heteroskedasticity. Statistica Neerlandica, 47 (2):111–125, June 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01411.x.

Boender:1993:BEM

[1385] C. G. E. Boender and H. K. van Dijk. Bayes estimates of multi-criteria decision alternatives using Monte Carlo integration. Statistica Neerlandica, 47(2):127–151, June 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1993.tb01412.x.

Anonymous:1993:PSa

[1386] Anonymous. Problem section. Statistica Neerlandica, 47(2):153– 160, June 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01413.x.

Jansen:1993:APA

[1387] J. Jansen and J. A. Hoekstra. The analysis of proportions in agricultural experiments by a generalized linear mixed model. Statistica Neerlandica, 47(3):161–174, September 1993. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1993.tb01414.x.

Groeneboom:1993:IEM

[1388] P. Groeneboom and H. P. Lopuhaa. Isotonic estimators of monotone densities and distribution functions: basic facts. Statistica Neerlandica, 47(3):175–183, September 1993. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1993.tb01415.x.

Moors:1993:ELM

[1389] J. J. A. Moors and J. C. van Houwelingen. Estimation of linear models with nequal restrictions. Statistica Neerlandica, 47(3):185–198, September 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01416.x.

Kallenberg:1993:TEV

[1390] W. C. M. Kallenberg, G. S. Mudholkar, and P. Subbaiah. Testing equality of variances in the analysis of repeated measurements. Statistica Neerlandica, 47(3):199–219, September 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1993.tb01417.x.

Akahira:1993:AMF

[1391] M. Akahira and K. Takeuchi. On the application of the Minkowski– Farkas theorem to sampling designs. Statistica Neerlandica, 47(3):221– 223, September 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01418.x.

Anonymous:1993:PSb

[1392] Anonymous. Problem section. Statistica Neerlandica, 47(3):225–232, September 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01419.x.

Merckens:1993:ISE

[1393] A. Merckens and P. A. Bekker. Identification of simultaneous equation models with measurement error: a computerized evaluation. Statistica Neerlandica, 47(4):233–244, December 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1993.tb01420.x.

Alamatsaz:1993:DUD

[1394] M. H. Alamatsaz. On discrete α-unimodal distributions. Statistica Neerlandica, 47(4):245–252, December 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1993.tb01421.x.

Mulder:1993:SPA

[1395] P. G. H. Mulder. The simultaneous processes of ageing and mortality. Statistica Neerlandica, 47(4):253–267, December 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1993.
tb01422.x.

Asmussen:1993:ALH

[1396] S. Asmussen and V. Schmidt. The ascending ladder height distribution for a certain class of dependent random walks. Statistica Neerlandica, 47(4):269–277, December 1993. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1993.tb01423.x.

Gilat:1993:QLF

[1397] D. Gilat and T. P. Hill. Quantile-locating functions and the distance between the mean and quantiles. Statistica Neerlandica, 47(4):279–283, December 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01424.x.

Jansen:1993:RME

[1398] M. J. W. Jansen. On randomization, modeling and experimental design: some remarks on a paper by de Vos. Statistica Neerlandica, 47(4):285– 289, December 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01425.x. See [1342].

deVos:1993:R

[1399] A. F. de Vos. Reply. Statistica Neerlandica, 47(4):289–290, December 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01426.x.

Anonymous:1993:PSc

[1400] Anonymous. Problem section. Statistica Neerlandica, 47(4):291–296, December 1993. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1993.tb01427.x.

Engel:1994:SAA

[1401] B. Engel and A. Keen. A simple approach for the analysis of generalized linear mixed models. Statistica Neerlandica, 48(1):1–22, March 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01428.x.

Artalejo:1994:NRR

[1402] J. R. Artalejo. New results in retrial queueing systems with breakdown of the servers. Statistica Neerlandica, 48(1):23–36, March 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1994.
tb01429.x.

Nieuwenhuis:1994:BGB

[1403] G. Nieuwenhuis. Bridging the gap between a stationary point process and its Palm distribution. Statistica Neerlandica, 48(1):37–62, March 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01430.x.

Wilms:1994:AUD

[1404] R. J. G. Wilms and J. J. A. M. Brands. On the asymptotically uniform distribution modulo 1 of extreme order statistics. Statistica Neerlandica, 48(1):63–70, March 1994. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1994.tb01431.x.

vandenDoel:1994:ACN

[1405] I. T. van den Doel and J. F. Kiviet. Asymptotic consequences of neglected dynamics in individual effects models. Statistica Neerlandica, 48 (1):71–85, March 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01432.x.

Anonymous:1994:PSa

[1406] Anonymous. Problem section. Statistica Neerlandica, 48(1):87–96, March 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01433.x.

Wessels:1994:SNN

[1407] J. Wessels. Statistica Neerlandica: a new parent for a middle-aged child. Statistica Neerlandica, 48(2):97–98, July 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1994.tb01434.x.

Doornbos:1994:MPD

[1408] R. Doornbos. In memoriam Prof. dr. H. C. [Hugo Christiaan] Hamaker 1905–1993. Statistica Neerlandica, 48(2):99–102, July 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1994.
tb01435.x.

Drees:1994:OSK

[1409] H. Drees and H. Milbrodt. The one-sided Kolmogorov–Smirnov test in signal detection problems with Gaussian white noise. Statistica Neerlandica, 48(2):103–116, July 1994. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1994.tb01436.x.

Koning:1994:KTI

[1410] A. J. Koning. KMT-type inequalities and goodness-of-fit tests. Statistica Neerlandica, 48(2):117–132, July 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1994.tb01437.x.

Berg:1994:PMV

[1411] S. Berg and D. Lepelley. On probability models in voting theory. Statistica Neerlandica, 48(2):133–146, July 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1994.tb01438.x.

Wierda:1994:MSP

[1412] S. J. Wierda. Multivariate statistical process control — recent results and directions for future research. Statistica Neerlandica, 48(2):147– 168, July 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01439.x.

Windmeijer:1994:GFL

[1413] F. A. G. Windmeijer. Goodness of fit in the linear model without a constant term. Statistica Neerlandica, 48(2):169–181, July 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1994.
tb01440.x.

Anonymous:1994:RPD

[1414] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 48(2):183–186, July 1994. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1994.tb01441.x.

Anonymous:1994:PSb

[1415] Anonymous. Problem section. Statistica Neerlandica, 48(2):187– 200, July 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01442.x.

Wellner:1994:CFM

[1416] J. A. Wellner. Covariance formulas via marginal martingales. Statistica Neerlandica, 48(3):201–207, November 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1994.tb01443.x.

Christmann:1994:SPL

[1417] A. Christmann, U. Gather, and G. Scholz. Some properties of the length of the shortest half. Statistica Neerlandica, 48(3):209–213, November 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01444.x.

Chaudhuri:1994:SDS

[1418] A. Chaudhuri. Small domain statistics: a review. Statistica Neerlandica, 48(3):215–236, November 1994. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1994.tb01445.x.

Meester:1994:UPT

[1419] R. W. J. Meester. Uniqueness in percolation theory. Statistica Neerlandica, 48(3):237–252, November 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1994.tb01446.x.

Thavaneswaran:1994:NMR

[1420] A. Thavaneswaran and B. Abraham. A note on Model Reference Adaptive System (MRAS) estimate with infinite variance. Statistica Neerlandica, 48(3):253–257, November 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1994.tb01447.x.

Indefrey:1994:EWF

[1421] P. Indefrey and H. Baayeny. Estimating word frequencies from lexical dispersion data. Statistica Neerlandica, 48(3):259–270, November 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01448.x.

Windmeijer:1994:GFT

[1422] F. A. G. Windmeijer. A goodness-of-fit test in the multinomial legit model based on weighted squared residuals. Statistica Neerlandica, 48 (3):271–283, November 1994. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1994.tb01449.x.

Anonymous:1994:PSc

[1423] Anonymous. Problem section. Statistica Neerlandica, 48(3):285–293, November 1994. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1994.tb01450.x.

Beirlant:1995:ACI

[1424] J. Beirlant and J. H. J. Einmahl. Asymptotic confidence intervals for the length of the short t under random censoring. Statistica Neerlandica, 49 (1):1–8, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01451.x.

vanderVaart:1995:EID

[1425] A. W. van der Vaart. Efficiency of infinite dimensional Mestimators. Statistica Neerlandica, 49(1):9–30, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1995.
tb01452.x.

Kopocinski:1995:LSS

[1426] B. Kopociński. The largest square of successes covering the origin for Bernoulli trials on the lattice. Statistica Neerlandica, 49(1):31– 39, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01453.x.

Tribouley:1995:PEM

[1427] K. Tribouley. Practical estimation of multivariate densities using wavelet methods. Statistica Neerlandica, 49(1):41–62, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1995.
tb01454.x.

Kroese:1995:DI

[1428] A. H. Kroese, E. A. van der Meulen, K. Poortema, and W. Schaafsma. Distributional inference. Statistica Neerlandica, 49(1):63–82, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01455.x.

Rosler:1995:DSR

[1429] U. Rösler. Distributions slanted to the right. Statistica Neerlandica, 49 (1):83–93, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01456.x.

Willenborg:1995:TLS

[1430] L. C. R. J. Willenborg. Testing the logical structure of questionnaires. Statistica Neerlandica, 49(1):95–109, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1995.
tb01457.x.

Anonymous:1995:SME

[1431] Anonymous. Semiparametric models: an evaluation. Statistica Neerlandica, 49(1):111–125, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1995.tb01458.x.

Anonymous:1995:PSa

[1432] Anonymous. Problem section. Statistica Neerlandica, 49(1):127–132, March 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01459.x.

Marsaglia:1995:RVI

[1433] G. Marsaglia. Random variables with independent integer and fractional parts. Statistica Neerlandica, 49(2):133–137, July 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1995.
tb01460.x.

Willenborg:1995:DEI

[1434] L. C. R. J. Willenborg. Data editing and imputation from a computational point of view. Statistica Neerlandica, 49(2):139–151, July 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01461.x.

Huang:1995:ANN

[1435] J. Huang and J. A. Wellner. Asymptotic normality of the NPMLE of linear functionals for interval censored data, case 1. Statistica Neerlandica, 49(2):153–163, July 1995. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1995.tb01462.x.

Barendregt:1995:EPL

[1436] L. G. Barendregt and M. C. van Pul. On the estimation of the parameters for the Littlewood model in software reliability. Statistica Neerlandica, 49 (2):165–184, July 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01463.x.

Davies:1995:DF

[1437] P. L. Davies. Data features. Statistica Neerlandica, 49(2):185– 245, July 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01464.x.

Anonymous:1995:BR

[1438] Anonymous. Book reviews. Statistica Neerlandica, 49(2):247–249, July 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01465.x.

Anonymous:1995:RPD

[1439] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 49(2):251–255, July 1995. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1995.tb01466.x.

Anonymous:1995:PSb

[1440] Anonymous. Problem section. Statistica Neerlandica, 49(2):257– 266, July 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01467.x.

vanHouwelingen:1995:SIL

[1441] Hans van Houwelingen, Dirk Sikkel, and Tom Wansbeek. Special issue on longitudinal models. Statistica Neerlandica, 49(3):267–268, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01468.x.

Kortram:1995:CIM

[1442] R. A. Kortram, A. C. M. van Rooij, A. J. Lenstra, and G. Ridder. Constructive identification of the mixed proportional hazards model. Statistica Neerlandica, 49(3):269–281, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1995.
tb01469.x.

vandenBerg:1995:DPD

[1443] G. J. van den Berg and M. Lindeboom. Durations in panel data subject to attrition: a note on estimation in the case of a stock sample. Statistica Neerlandica, 49(3):282–293, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1995.tb01470.x.

vanOpstal:1995:TSW

[1444] R. van Opstal and F. J. R. van de Pol. The transition from school to work in 1979–1987 in The Netherlands. Statistica Neerlandica, 49(3):294– 309, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01471.x.

Zwinderman:1995:PNH

[1445] A. H. Zwinderman, J. C. van Houwelingen, and D. Schweitzer. Prediction of the next hypercalcemia free period: application of random effect models with selection on first event. Statistica Neerlandica, 49(3):310– 323, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01472.x.

Charlier:1995:SMS

[1446] E. Charlier, B. Melenberg, and A. H. O. van Soest. A smoothed maximum score estimator for the binary choice panel data model with an application to labour force participation. Statistica Neerlandica, 49(3):324– 342, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01473.x.

vandenDoel:1995:NDP

[1447] I. T. van den Doel and J. F. Kiviet. Neglected dynamics in panel data models; consequences and detection in finite samples. Statistica Neerlandica, 49(3):343–361, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1995.tb01474.x.

Jansen:1995:LLM

[1448] R. A. R. G. Jansen and J. H. L. Oud. Longitudinal LISREL model estimation from incomplete panel data using the EM algorithm and the Kalman smoother. Statistica Neerlandica, 49(3):362–377, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01475.x.

Sikkel:1995:MMP

[1449] D. Sikkel and A. W. Hoogendoorn. Models for monthly penetrations with incomplete panel data. Statistica Neerlandica, 49(3):378– 391, November 1995. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1995.tb01476.x.

Gill:1996:SNG

[1450] R. D. Gill. Statistica Neerlandica Golden Jubilee Editor’s preface. Statistica Neerlandica, 50(1):1–2, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1996.tb01477.x.

Aardal:1996:PTC

[1451] K. Aardal and C. P. M. van Hoesel. Polyhedral techniques in combinatorial optimization I: Theory. Statistica Neerlandica, 50(1):3–26, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01478.x.

Does:1996:ISR

[1452] R. J. M. M. Does and K. C. B. Roes. Industrial statistics and its recent contributions to total quality in The Netherlands. Statistica Neerlandica, 50(1):27–51, March 1996. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1996.tb01479.x.

Donders:1996:OEA

[1453] M. W. M. Donders and A. C. F. Vorst. Options and earnings announcements: an empirical study for the European Options Exchange. Statistica Neerlandica, 50(1):52–68, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1996.tb01480.x.

Geskus:1996:AOE

[1454] R. B. Geskus and P. Groeneboom. Asymptotically optimal estimation of smooth functionals for interval censoring, part 1. Statistica Neerlandica, 50(1):69–88, March 1996. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1996.tb01481.x.

Heesterbeek:1996:CET

[1455] J. A. P. Heesterbeek and K. Dietz. The concept of R_{0} in epidemic theory. Statistica Neerlandica, 50(1):89–110, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1996.
tb01482.x.

vanderHeijden:1996:FYM

[1456] P. G. M. van der Heijden and K. Sijtsma. Fifty years of measurement and scaling in the Dutch social sciences. Statistica Neerlandica, 50(1): 111–135, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01483.x.

denHollander:1996:RP

[1457] F. den Hollander. Random polymers. Statistica Neerlandica, 50(1):136– 145, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01484.x.

Jansen:1996:IPM

[1458] B. Jansen, C. Roos, and T. Terlaky. Interior point methods, a decade after Karmarkar — a survey, with application to the smallest eigenvalue problem. Statistica Neerlandica, 50(1):146–170, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01485.x.

Melenberg:1996:MCC

[1459] B. Melenberg and A. H. O. van Soest. Measuring the costs of children: parametric and semiparametric estimators. Statistica Neerlandica, 50(1): 171–192, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01486.x.

Petersen:1996:VCM

[1460] J. H. Petersen, P. K. Andersen, and R. D. Gill. Variance components models for survival data. Statistica Neerlandica, 50(1):193– 211, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01487.x.

Koning:1996:RPD

[1461] R. H. Koning. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 50(1):212–214, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1996.tb01488.x.

Anonymous:1996:PSa

[1462] Anonymous. Problem section. Statistica Neerlandica, 50(1):215–223, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01489.x.

Maassen:1996:RWV

[1463] H. Maassen and F. W. Steutel. Remembering Wim Vervaat. Statistica Neerlandica, 50(1):225–230, March 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.1996.tb01490.x.

OBrien:1996:RWV

[1464] G. L. O’Brien. The research of Wim Vervaat. Statistica Neerlandica, 50 (2):231–241, July 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01491.x.

OBrien:1996:LDP

[1465] G. L. O’Brien and W. Vervaat. Large deviation principles and loglog laws for observation processes. Statistica Neerlandica, 50(2):242–260, July 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01492.x.

Deheuvels:1996:FLI

[1466] P. Deheuvels. Functional laws of the iterated logarithm for small increments of empirical processes. Statistica Neerlandica, 50(2):261– 280, July 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01493.x.

Gerritse:1996:SMC

[1467] B. Gerritse. Sanov meets Cramér again. Statistica Neerlandica, 50(2): 281–294, July 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01494.x.

Hansen:1996:SSD

[1468] B. G. Hansen. Stability and self-decomposability of semi-group valued random variables. Statistica Neerlandica, 50(2):295–305, July 1996. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1996.
tb01495.x.

Holwerda:1996:LCR

[1469] H. Holwerda and W. Vervaat. Lattices of capacities, and related topologies. Statistica Neerlandica, 50(2):306–324, July 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1996.
tb01496.x.

Anonymous:1996:PSb

[1470] Anonymous. Problem section. Statistica Neerlandica, 50(2):325– 332, July 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01497.x.

Gill:1996:VDA

[1471] Richard D. Gill. The van Dantzig Award 1995. Statistica Neerlandica, 50(3):333–336, November 1996. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1996.tb01498.x.

vanHouwelingen:1996:TYR

[1472] Hans C. van Houwelingen and Editor of Statistica Neerlandica. Thank you Richard Gill. Statistica Neerlandica, 50(3):337–338, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01499.x.

Fung:1996:LIS

[1473] W. K. Fung. On the lack of invariance of some Bayesian outlier models. Statistica Neerlandica, 50(3):339–343, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.1996.

tb01500.x. See correction [1491].

vanLieshout:1996:NMS

[1474] M. N. M. van Lieshout and A. J. Baddeley. A nonparametric measure of spatial interaction in point patterns. Statistica Neerlandica, 50(3):344– 361, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01501.x.

Tajuddin:1996:SMS

[1475] I. H. Tajuddin. A simple measure of skewness. Statistica Neerlandica, 50(3):362–366, November 1996. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.1996.tb01502.x.

Pap:1996:ABS

[1476] G. Pap and M. C. A. van Zuijlen. On the asymptotic behaviour of the Stringer bound. Statistica Neerlandica, 50(3):367–389, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01503.x.

Wesolowski:1996:NCS

[1477] J. Wesolowski. A new conditional specification of the bivariate Poisson conditionals distribution. Statistica Neerlandica, 50(3):390–393, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01504.x.

Koziol:1996:WKS

[1478] J. A. Koziol. A weighted Kuiper statistic for goodness of fit. Statistica Neerlandica, 50(3):394–403, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.1996.tb01505.x.

Schultz:1996:TSS

[1479] R. Schultz, L. Stougie, and M. H. van der Vlerk. Two-stage stochastic integer programming: a survey. Statistica Neerlandica, 50(3):404– 416, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01506.x.

Moors:1996:CSD

[1480] J. J. A. Moors, R. Th. A. Wagemakers, V. M. J. Coenen, R. M. J. Heuts, and M. J. B. T. Janssens. Characterizing systems of distributions by quantile measures. Statistica Neerlandica, 50(3):417–430, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01507.x.

Anonymous:1996:PSc

[1481] Anonymous. Problem section. Statistica Neerlandica, 50(3):431–434, November 1996. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.1996.tb01508.x.

Feldman:1997:LDC

[1482] P. Feldman, S. T. Rachev, and L. Rüschendorf. Limiting distribution of the collision resolution interval. Statistica Neerlandica, 51(1): 1–22, March 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00034.

Hooghiemstra:1997:CLT

[1483] G. Hooghiemstra and M. Keane. A central limit theorem for sums of correlated products. Statistica Neerlandica, 51(1):23–34, March 1997. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00035.

Gill:1997:NEU

[1484] R. D. Gill. Nonparametric estimation under censoring and passive registration. Statistica Neerlandica, 51(1):35–54, March 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00036.

Kuijpers:1997:FMS

[1485] C. M. H. Kuijpers, C. A. J. Hurkens, and J. B. M. Melissen. Fast movement strategies for a step-and-scan wafer stepper. Statistica Neerlandica, 51(1):55–71, March 1997. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00037.

DeBoer:1997:NAM

[1486] P. M. C. De Boer and R. Harkema. A new approach to maximum likelihood estimation of sum-constrained linear models in case of undersized samples. Statistica Neerlandica, 51(1):72–89, March 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00038.

VanBuuren:1997:OTC

[1487] S. Van Buuren. Optimal transformations for categorical autoregressive time series. Statistica Neerlandica, 51(1):90–106, March 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00039.

Sassen:1997:HRR

[1488] S. A. E. Sassen, H. C. Tijms, and R. D. Nobel. A heuristic rule for routing customers to parallel servers. Statistica Neerlandica, 51(1):107– 121, March 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00040.

Anonymous:1997:RPDa

[1489] Anonymous. Recent Ph.D theses in The Netherlands. Statistica Neerlandica, 51(1):122–124, March 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00041.

Anonymous:1997:PSa

[1490] Anonymous. Problem section. Statistica Neerlandica, 51(1):125–128, March 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00042.

Anonymous:1997:CWK

[1491] Anonymous. Correction to W. K. Fung, On the lack of invariance of some Bayesian outlier models. Statistica Neerlandica, 51(1):128, March 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00043. See [1473].

Keen:1997:AMM

[1492] A. Keen and B. Engel. Analysis of a mixed model for ordinal data by iterative re-weighted REML. Statistica Neerlandica, 51(2):129– 144, July 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00044.

Engel:1997:MBK

[1493] Bas Engel. In memoriam Bertus Keen 1945–1996. Statistica Neerlandica, 51(2):145, July 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00045.

Ahrens:1997:SCM

[1494] J. H. Ahrens and U. Rösler. Signed central moments of integer random variables with decreasing density. Statistica Neerlandica, 51(2): 146–154, July 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00046.

Henze:1997:SPA

[1495] N. Henze. A statistical and probabilistic analysis of popular lottery tickets. Statistica Neerlandica, 51(2):155–163, July 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00047.

Viharos:1997:TIE

[1496] L. Viharos. Tail index estimation based on linear combinations of intermediate order statistics. Statistica Neerlandica, 51(2):164–177, July 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00048.

VanDerLaan:1997:NEB

[1497] M. J. Van Der Laan. Nonparametric estimators of the bivariate survival function under random censoring. Statistica Neerlandica, 51(2): 178–200, July 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00049.

Geskus:1997:AOE

[1498] R. B. Geskus and P. Groeneboom. Asymptotically optimal estimation of smooth functionals for interval censoring, part 2. Statistica Neerlandica, 51(2):201–219, July 1997. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00050.

Lopuhaa:1997:AEE

[1499] H. P. Lopuhaä. Asymptotic expansion of S-estimators of location and covariance. Statistica Neerlandica, 51(2):220–237, July 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00051.

Neuts:1997:HDC

[1500] M. F. Neuts, D. E. Rauschenberg, and Jian-Min Li. How did the cookie crumble? Identifying fragmentation procedures. Statistica Neerlandica, 51(2):238–251, July 1997. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00052.

Anonymous:1997:PSb

[1501] Anonymous. Problem section. Statistica Neerlandica, 51(2):252– 256, July 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00053.

He:1997:QRE

[1502] X. He and G. Wang. Qualitative robustness of S∗-estimators of multivariate location and dispersion. Statistica Neerlandica, 51(3):257–268, November 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00054.

Breusch:1997:ENC

[1503] T. S. Breusch, J. C. Robertson, and A. H. Welsh. The emperor’s new clothes: a critique of the multivariate t regression model. Statistica Neerlandica, 51(3):269–286, November 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00055.

Chang:1997:CD

[1504] J. T. Chang and D. Pollard. Conditioning as disintegration. Statistica Neerlandica, 51(3):287–317, November 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00056.

VanDerGenugten:1997:BHC

[1505] B. B. Van Der Genugten. Blackjack in Holland Casino’s: basic, optimal and winning strategies. Statistica Neerlandica, 51(3):318–344, November 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00057.

Liu:1997:KIE

[1506] S. Liu and H. Neudecker. Kantorovich inequalities and efficiency comparisons for several classes of estimators in linear models. Statistica Neerlandica, 51(3):345–355, November 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00058.

Coolen:1997:SBS

[1507] F. P. A. Coolen. On the stopping boundary for sequential selection of the normal population with the largest mean. Statistica Neerlandica, 51(3):356–365, November 1997. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00059.

Gather:1997:NIM

[1508] U. Gather, J. Pawlitschko, and I. Pigeot. A note on invariance of multiple tests. Statistica Neerlandica, 51(3):366–372, November 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.t01-1-00060.

Anonymous:1997:RPDb

[1509] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 51(3):373–377, November 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00061.

Anonymous:1997:PSc

[1510] Anonymous. Problem section. Statistica Neerlandica, 51(3):378–382, November 1997. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00062.

VanDenBrink:1998:IUS

[1511] W. P. Van Den Brink. The improper use of social security: an estimation problem. Statistica Neerlandica, 52(1):1–5, March 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00064.

Jongbloed:1998:EDT

[1512] G. Jongbloed. Exponential deconvolution: two asymptotically equivalent estimators. Statistica Neerlandica, 52(1):6–17, March 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00065.

Stein:1998:AST

[1513] A. Stein. Analysis of space–time variability in agriculture and the environment with geostatistics. Statistica Neerlandica, 52(1):18–41, March 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00066.

Romeijn:1998:GFA

[1514] H. E. Romeijn. A general framework for approximate sampling with an application to generating points on the boundary of bounded convex regions. Statistica Neerlandica, 52(1):42–59, March 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00067.

DeHaan:1998:CTI

[1515] L. De Haan and L. Peng. Comparison of tail index estimators. Statistica Neerlandica, 52(1):60–70, March 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00068.

Menendez:1998:ADD

[1516] M. L. Menéndez, D. Morales, L. Pardo, and I. Vajda. Asymptotic distributions of ϕ-divergences of hypothetical and observed frequencies on refined partitions. Statistica Neerlandica, 52(1):71–89, March 1998. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00069.

Khmaladze:1998:GFT

[1517] E. V. Khmaladze. Goodness of fit tests for “Chimeric” alternatives. Statistica Neerlandica, 52(1):90–111, March 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00070.

Granville:1998:EIP

[1518] V. Granville. Estimation of the intensity of a Poisson point process by means of nearest neighbor distances. Statistica Neerlandica, 52(1):112– 124, March 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00071.

Anonymous:1998:PSa

[1519] Anonymous. Problem section. Statistica Neerlandica, 52(1):125–128, March 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00072.

Moors:1998:SPP

[1520] J. J. A. Moors, R. Smeets, and F. W. M. Boekema. Sampling with probabilities proportional to the variable of interest. Statistica Neerlandica, 52 (2):129–140, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00073.

Ayala:1998:SLB

[1521] G. Ayala and A. Simó. Stochastic labelling of biological images. Statistica Neerlandica, 52(2):141–152, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00074.

Molenberghs:1998:MMD

[1522] G. Molenberghs, B. Michiels, M. G. Kenward, and P. J. Diggle. Monotone missing data and pattern-mixture models. Statistica Neerlandica, 52(2): 153–161, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00075.

Warners:1998:NAC

[1523] J. P. Warners. A nonlinear approach to a class of combinatorial optimization problems. Statistica Neerlandica, 52(2):162–184, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00076.

Krug:1998:RLM

[1524] B. Krug. Reweighted loglikelihood method in robust estimation in a mixed unbalanced model. Statistica Neerlandica, 52(2):185–199, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00077.

Lee:1998:SEE

[1525] M.-J. Lee and H. Kim. Semiparametric econometric estimators for a truncated regression model: a review with an extension. Statistica Neerlandica, 52(2):200–225, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00078.

Meester:1998:CEP

[1526] R. Meester and J. E. Steif. Consistent estimation of percolation quantities. Statistica Neerlandica, 52(2):226–238, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00079.

Chiu:1998:EDD

[1527] S. N. Chiu and D. Stoyan. Estimators of distance distributions for spatial patterns. Statistica Neerlandica, 52(2):239–246, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00080.

Anonymous:1998:RPD

[1528] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 52(2):247–248, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00081.

Anonymous:1998:PSb

[1529] Anonymous. Problem section. Statistica Neerlandica, 52(2):249– 254, June 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00082.

Franses:1998:LDS

[1530] Philip Hans Franses. Large data sets in finance and marketing: introduction by the special issue editor. Statistica Neerlandica, 52(3):255– 257, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00083.

Granger:1998:EIM

[1531] C. W. J. Granger. Extracting information from mega-panels and high-frequency data. Statistica Neerlandica, 52(3):258–272, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00084.

Andersen:1998:TUF

[1532] T. G. Andersen and T. Bollerslev. Towards a unified framework for high and low frequency return volatility modeling. Statistica Neerlandica, 52(3):273–302, November 1998. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00085.

Balasubramanian:1998:MLD

[1533] S. Balasubramanian, S. Gupta, W. Kamakura, and M. Wedel. Modeling large data sets in marketing. Statistica Neerlandica, 52(3):303– 323, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00086.

Caserta:1998:ARR

[1534] S. Caserta, J. Danielsson, and C. G. De Vries. Abnormal returns, risk, and options in large data sets. Statistica Neerlandica, 52(3):324–335, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00087.

Hafner:1998:SAP

[1535] C. M. Hafner and H. Herwartz. Structural analysis of portfolio risk using beta impulse response functions. Statistica Neerlandica, 52(3):336– 355, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00088.

Abdous:1998:MMM

[1536] B. Abdous and R. Theodorescu. Mean, median, mode IV. Statistica Neerlandica, 52(3):356–359, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00089.

Haggstrom:1998:ESA

[1537] O. Häggström and K. Nelander. Exact sampling from anti-monotone systems. Statistica Neerlandica, 52(3):360–380, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00090.

Anonymous:1998:PSc

[1538] Anonymous. Problem section. Statistica Neerlandica, 52(3):381–383, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00091.

Anonymous:1998:CV

[1539] Anonymous. Contents of volume 52. Statistica Neerlandica, 52(3):384– 385, November 1998. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00092.

Senden:1999:IE

[1540] W. J. M. Senden and W. R. Van Zwet. Introducing EURANDOM. Statistica Neerlandica, 53(1):1–4, March 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00093.

Poortema:1999:MOC

[1541] K. Poortema. On modelling overdispersion of counts. Statistica Neerlandica, 53(1):5–20, March 1999. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00094.

Hoogendoorn:1999:DPI

[1542] A. W. Hoogendoorn and D. Sikkel. Description of purchase incidence by multivariate heterogeneous Poisson processes. Statistica Neerlandica, 53 (1):21–35, March 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00095.

Albers:1999:VCS

[1543] W. Albers, W. C. M. Kallenberg, and G. R. J. Arts. A variety of criteria for setting test limits. Statistica Neerlandica, 53(1):36– 54, March 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00096.

Pannekoek:1999:SMS

[1544] J. Pannekoek. Statistical methods for some simple disclosure limitation rules. Statistica Neerlandica, 53(1):55–67, March 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00097.

Marron:1999:EWM

[1545] J. S. Marron. Effective writing in mathematical statistics. Statistica Neerlandica, 53(1):68–75, March 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00098.

Swanson:1999:FSP

[1546] N. R. Swanson. Finite sample properties of a simple LM test for neglected nonlinearity in error-correcting regression equations. Statistica Neerlandica, 53(1):76–95, March 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00099.

Pakes:1999:CMG

[1547] A. G. Pakes. On the convergence of moments of geometric and harmonic means. Statistica Neerlandica, 53(1):96–110, March 1999. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00100.

Bednarski:1999:LUB

[1548] T. Bednarski and W. Florczak. On a local uniform bootstrap validity. Statistica Neerlandica, 53(1):111–121, March 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00101.

Anonymous:1999:PSa

[1549] Anonymous. Problem section. Statistica Neerlandica, 53(1):122–128, March 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00102.

Stougie:1999:IPA

[1550] Leen Stougie. Introducing the paper by Aardal and van Hoesel. Statistica Neerlandica, 53(2):129–130, July 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00103.

Aardal:1999:PTC

[1551] K. Aardal and C. P. M. Van Hoesel. Polyhedral techniques in combinatorial optimization II: applications and computations. Statistica Neerlandica, 53(2):131–177, July 1999. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00104.

Gather:1999:JER

[1552] U. Gather, S. Köhne, and I. Pigeot. Jackknife estimators of a relative risk in 2 × 2 and 2 × 2 × K contingency tables. Statistica Neerlandica, 53(2): 178–196, July 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00105.

Dellaert:1999:SDG

[1553] N. P. Dellaert and W. A. Luijten. Statistical disclosure in general threedimensional tables. Statistica Neerlandica, 53(2):197–221, July 1999. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00106.

Goldie:1999:GDO

[1554] Charles M. Goldie and Ross A. Maller. Generalized densities of order statistics. Statistica Neerlandica, 53(2):222–246, July 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00107.

Wu:1999:CUD

[1555] F. Wu and S. Dharmadhikari. Convolutions of α-unimodal discrete distributions. Statistica Neerlandica, 53(2):247–250, July 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00108.

Anonymous:1999:RPDa

[1556] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 53(2):251, July 1999. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00109.

Anonymous:1999:PSb

[1557] Anonymous. Problem section. Statistica Neerlandica, 53(2):252– 256, July 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00110.

Wong:1999:RAM

[1558] Weng Kee Wong. Recent advances in multiple-objective design strategies. Statistica Neerlandica, 53(3):257–276, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00111.

Munzel:1999:NMP

[1559] U. Munzel. Nonparametric methods for paired samples. Statistica Neerlandica, 53(3):277–286, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00112.

Beirlant:1999:PLD

[1560] J. Beirlant, A. Berlinet, and L. Györfi. On piecewise linear density estimators. Statistica Neerlandica, 53(3):287–308, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00113.

Dey:1999:DDE

[1561] A. K. Dey and F. H. Ruymgaart. Direct density estimation as an illposed inverse estimation problem. Statistica Neerlandica, 53(3):309–326, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00114.

Gather:1999:RES

[1562] U. Gather and V. Schultze. Robust estimation of scale of an exponential distribution. Statistica Neerlandica, 53(3):327–341, November 1999. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00115.

Thabane:1999:BSB

[1563] L. Thabane and M. Safiul Haq. On Bayesian selection of the best normal population using the Kullback–Leibler divergence measure. Statistica Neerlandica, 53(3):342–360, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00116.

Roussanov:1999:SBD

[1564] B. V. Roussanov. Stepwise Bayes distributional inference for a finite population. Statistica Neerlandica, 53(3):361–374, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00117.

Anonymous:1999:RPDb

[1565] Anonymous. Recent Ph.D. theses. Statistica Neerlandica, 53(3):375, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00118.

Anonymous:1999:PSc

[1566] Anonymous. Problem section. Statistica Neerlandica, 53(3):376–382, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00119.

Anonymous:1999:SNI

[1567] Anonymous. Statistica Neerlandica 53, 1999 index. Statistica Neerlandica, 53(3):383–384, November 1999. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00120.

VanHouwelingen:2000:ESN

[1568] Hans C. Van Houwelingen. Editorial: Statistics in the new millennium. Statistica Neerlandica, 54(1):1–2, January 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00121.

Moors:2000:RAC

[1569] J. J. A. Moors, B. B. Van Der Genugten, and L. W. G. Strijbosch. Repeated audit controls. Statistica Neerlandica, 54(1):3–13, January 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00122.

Longford:2000:ADI

[1570] Nicholas T. Longford. An alternative definition of individual bioequivalence. Statistica Neerlandica, 54(1):14–36, January 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00123.

Jones:2000:RRA

[1571] M. C. Jones. Rough-and-ready assessment of the degree and importance of smoothing in functional estimation. Statistica Neerlandica, 54(1):37– 46, January 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00124.

Ahmed:2000:LSP

[1572] S. E. Ahmed and A. K. Basu. Least squares, preliminary test and Steintype estimation in general vector AR(p) models. Statistica Neerlandica, 54(1):47–66, January 2000. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00125.

Falin:2000:BMP

[1573] G. I. Falin and A. Gómez-Corral. On a bivariate Markov process arising in the theory of single-server retrial queues. Statistica Neerlandica, 54(1):

67–78, January 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00126.

Nielsen:2000:RCR

[1574] S. F. Nielsen. Relative coarsening at random. Statistica Neerlandica, 54(1):79–99, January 2000. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00127.

VanEs:2000:CRP

[1575] A. J. Van Es, R. Helmers, and M. Hušková. On a crossroad of resampling plans: bootstrapping elementary symmetric polynomials. Statistica Neerlandica, 54(1):100–110, January 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00128.

Anonymous:2000:PSa

[1576] Anonymous. Problem section. Statistica Neerlandica, 54(1):111–112, January 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00129.

Anonymous:2000:GE

[1577] Anonymous. Guest editorial. Statistica Neerlandica, 54(2):113–115, July 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00130.

VanNunen:2000:JWH

[1578] J. A. E. E. Van Nunen and J. Van Der Wal. Jaap Wessels — his life with stochastic processes. Statistica Neerlandica, 54(2):116– 126, July 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00131.

VanHoutum:2000:RBC

[1579] G. J. Van Houtum and W. H. M. Zijm. On the relationship between cost and service models for general inventory systems. Statistica Neerlandica, 54(2):127–147, July 2000. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00132.

Hordijk:2000:OSC

[1580] A. Hordijk and J. A. Loeve. Optimal static customer routing in a closed queuing network. Statistica Neerlandica, 54(2):148–159, July 2000. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00133.

Wijngaard:2000:FRM

[1581] J. Wijngaard and S. Stidham. Forward recursion for Markov decision processes with skip-free-to-the-right transitions. Part II: non-standard applications. Statistica Neerlandica, 54(2):160–174, July 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00134.

Dellaert:2000:PPS

[1582] N. P. Dellaert, A. G. De Kok, and Wang Wei. Push and pull strategies in multi-stage assembly systems. Statistica Neerlandica, 54(2):175– 189, July 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00135.

Sassen:2000:MGQ

[1583] S. A. E. Sassen and J. Van Der Wal. The M/G/∞ queueing model for optimistic concurrency control. Statistica Neerlandica, 54(2):190– 201, July 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00136.

Adan:2000:MSB

[1584] I. J. B. F. Adan and J. A. C. Resing. Multi-server batch-service systems. Statistica Neerlandica, 54(2):202–220, July 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00137.

Boxma:2000:MMQ

[1585] O. J. Boxma and I. A. Kurkova. The M/M/1 queue in a heavy-tailed random environment. Statistica Neerlandica, 54(2):221–236, July 2000. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00138.

VanDerAalst:2000:ADT

[1586] W. M. P. Van Der Aalst, K. M. Van Hee, and H. A. Reijers. Analysis of discrete-time stochastic Petri nets. Statistica Neerlandica, 54(2):

237–255, July 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00139.

Aarts:2000:NNU

[1587] E. H. L. Aarts and H. M. M. Ten Eikelder. A neural network unification. Statistica Neerlandica, 54(2):256–264, July 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00140.

Dzhaparidze:2000:SAM

[1588] K. Dzhaparidze, P. J. C. Spreij, and J. H. Van Zanten. Some aspects of modeling and statistical inference for financial models. Statistica Neerlandica, 54(3):265–292, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00141.

Stegeman:2000:HTV

[1589] A. Stegeman. Heavy tails versus long-range dependence in selfsimilar network traffic. Statistica Neerlandica, 54(3):293–314, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00142.

Baddeley:2000:EFE

[1590] A. J. Baddeley, M. Kerscher, K. Schladitz, and B. T. Scott. Estimating the J function without edge correction. Statistica Neerlandica, 54(3):315– 328, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00143.

Baddeley:2000:NSP

[1591] A. J. Baddeley, J. Møller, and R. Waagepetersen. Non- and semiparametric estimation of interaction in inhomogeneous point patterns. Statistica Neerlandica, 54(3):329–350, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00144.

Belitser:2000:LMP

[1592] E. Belitser. Local minimax pointwise estimation of a multivariate density. Statistica Neerlandica, 54(3):351–365, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00145.

Allaart:2000:IRM

[1593] P. C. Allaart. Inequalities relating maximal moments to other measures of dispersion. Statistica Neerlandica, 54(3):366–373, November 2000. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00146.

Tuenter:2000:GPD

[1594] H. J. H. Tuenter. On the generalized Poisson distribution. Statistica Neerlandica, 54(3):374–376, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00147.

Anonymous:2000:RPD

[1595] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 54(3):377, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00148.

Anonymous:2000:PSb

[1596] Anonymous. Problem section. Statistica Neerlandica, 54(3):378–382, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.t01-1-00149.

Anonymous:2000:CV

[1597] Anonymous. Contents of volume 54. Statistica Neerlandica, 54(3):383– 384, November 2000. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00150.

Franses:2001:E

[1598] Philip Hans Franses. Editorial. Statistica Neerlandica, 55(1):1, March 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00151.

VanHouwelingen:2001:SIP

[1599] Hans Van Houwelingen. Special issue: Prediction in medical statistics. Statistica Neerlandica, 55(1):2, March 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00152.

Hand:2001:MDA

[1600] D. J. Hand. Measuring diagnostic accuracy of statistical prediction rules. Statistica Neerlandica, 55(1):3–16, March 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00153.

VanHouwelingen:2001:SPL

[1601] J. C. Van Houwelingen. Shrinkage and penalized likelihood as methods to improve predictive accuracy. Statistica Neerlandica, 55(1):17– 34, March 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00154.

Vach:2001:CVC

[1602] W. Vach. Cross validation calibration based shrinkage — revisited for the construction of prognostic indices. Statistica Neerlandica, 55(1):35– 52, March 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00155.

Vach:2001:VSS

[1603] K. Vach, W. Sauerbrei, and M. Schumacher. Variable selection and shrinkage: Comparison of some approaches. Statistica Neerlandica, 55 (1):53–75, March 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00156.

Steyerberg:2001:AST

[1604] E. W. Steyerberg, M. J. C. Eijkemans, and J. D. F. Habbema. Application of shrinkage techniques in logistic regression analysis: a case study. Statistica Neerlandica, 55(1):76–88, March 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00157.

Royston:2001:LDM

[1605] P. Royston. The lognormal distribution as a model for survival time in cancer, with an emphasis on prognostic factors. Statistica Neerlandica, 55(1):89–104, March 2001. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00158.

Anonymous:2001:RPD

[1606] Anonymous. Recent Ph.D. theses in The Netherlands. Statistica Neerlandica, 55(1):105–107, March 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00159.

Eisinga:2001:I

[1607] Rob Eisinga and Philip Hans Franses. Introduction. Statistica Neerlandica, 55(2):109–110, July 2001. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00160.

Beck:2001:TSC

[1608] Nathaniel Beck. Time-series–cross-section data. Statistica Neerlandica, 55(2):111–133, July 2001. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00161.

Rosen:2001:BFI

[1609] Ori Rosen, Wenxin Jiang, Gary King, and Martin A. Tanner. Bayesian and frequentist inference for ecological inference: The R × C case. Statistica Neerlandica, 55(2):134–156, July 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00162.

DeRooij:2001:DAM

[1610] Mark De Rooij. Distance association models for the analysis of repeated transition frequency tables. Statistica Neerlandica, 55(2):157– 181, July 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00163.

Feder:2001:TSA

[1611] Moshe Feder. Time series analysis of repeated surveys: The state–space approach. Statistica Neerlandica, 55(2):182–199, July 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00164.

Oud:2001:QLD

[1612] Johan H. L. Oud. Quasi-longitudinal designs in SEM state space modeling. Statistica Neerlandica, 55(2):200–220, July 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00165.

McLaren:2001:RPT

[1613] C. H. McLaren and D. G. Steel. Rotation patterns and trend estimation for repeated surveys using rotation group estimates. Statistica Neerlandica, 55(2):221–238, July 2001. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00166.

Girma:2001:NDM

[1614] Sourafel Girma. A note on dynamic modelling from short and heterogeneous pseudo panels. Statistica Neerlandica, 55(2):239–248, July 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00167.

Pelzer:2001:ETP

[1615] Ben Pelzer, Rob Eisinga, and Philip Hans Franses. Estimating transition probabilities from a time series of independent cross sections. Statistica Neerlandica, 55(2):249–262, July 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00168.

Kloek:2001:OHT

[1616] Teun Kloek. Obituary: Henri Theil, 1924–2000. Statistica Neerlandica, 55(3):263–269, November 2001. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00169.

DeJong:2001:JDM

[1617] F. De Jong, F. C. Drost, and B. J. M. Werker. A jump-diffusion model for exchange rates in a target zone. Statistica Neerlandica, 55(3):270– 300, November 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00170.

Imhof:2001:ERO

[1618] L. Imhof, J. Lopez-Fidalgo, and W. K. Wong. Efficiencies of rounded optimal approximate designs for small samples. Statistica Neerlandica, 55(3):301–318, November 2001. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00171.

Molenberghs:2001:MCD

[1619] G. Molenberghs and H. Geys. Multivariate clustered data analysis in developmental toxicity studies. Statistica Neerlandica, 55(3):319–345, November 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00172.

Albers:2001:HAP

[1620] C. J. Albers and W. Schaafsma. How to assign probabilities if you must. Statistica Neerlandica, 55(3):346–357, November 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00173.

Mertens:2001:DIR

[1621] B. J. A. Mertens. Downdating: Interdisciplinary research between statistics and computing. Statistica Neerlandica, 55(3):358–366, November 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00174.

Ghosh:2001:ECA

[1622] D. Ghosh. Efficiency considerations in the additive hazards model with current status data. Statistica Neerlandica, 55(3):367–376, November 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00175.

Anonymous:2001:PS

[1623] Anonymous. Problem section. Statistica Neerlandica, 55(3):377–383, November 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00176.

Anonymous:2001:CV

[1624] Anonymous. Content of volume 55. Statistica Neerlandica, 55(3):384– 385, November 2001. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00177.

Franses:2002:E

[1625] Philip Hans Franses. Editorial. Statistica Neerlandica, 56(1):1, February 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.06001.

Paap:2002:WAM

[1626] Richard Paap. What are the advantages of MCMC based inference in latent variable models? Statistica Neerlandica, 56(1):2–22, February 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.

1111/1467-9574.00060. See discussion [1628].

Pelzer:2002:BET

[1627] Ben Pelzer and Rob Eisinga. Bayesian estimation of transition probabilities from repeated cross sections. Statistica Neerlandica, 56(1):23– 33, February 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00063.

Koopman:2002:DMB

[1628] S. J. Koopman. Discussion of ‘MCMC-based inference’ by R. Paap. Statistica Neerlandica, 56(1):34–40, February 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.06002. See

[1626].

Schultze:2002:IOE

[1629] V. Schultze and J. Pawlitschko. The identification of outliers in exponential samples. Statistica Neerlandica, 56(1):41–57, February 2002. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.01600.

Geelhoed:2002:MDA

[1630] B. Geelhoed and H. J. Glass. Multinomial distributions applied to random sampling of particulate materials. Statistica Neerlandica, 56(1): 58–76, February 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.03500.

Knoth:2002:MMV

[1631] S. Knoth and W. Schmid. Monitoring the mean and the variance of a stationary process. Statistica Neerlandica, 56(1):77–100, February 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.03000.

Wellner:2002:CCT

[1632] J. A. Wellner and R. T. Smythe. Computing the covariance of two Brownian area integrals. Statistica Neerlandica, 56(1):101–109, February 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.03300.

Hansen:2002:NBI

[1633] Martin B. Hansen and Steffen L. Lauritzen. Nonparametric Bayes inference for concave distribution functions. Statistica Neerlandica, 56(1):110– 127, February 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.04600.

Einmahl:2002:GE

[1634] John Einmahl and Axel Munk. Guest editorial. Statistica Neerlandica, 56 (2):129–131, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00187.

Csorgo:2002:SDN

[1635] Sándor Csörgő. The smoothing dichotomy in nonparametric regression under long-memory errors. Statistica Neerlandica, 56(2):132– 142, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00188.

Denker:2002:ASL

[1636] Manfred Denker and Susanne Koch. Almost sure local limit theorems. Statistica Neerlandica, 56(2):143–151, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00189.

Leadbetter:2002:SLC

[1637] M. R. Leadbetter and G. V. Spaniolo. On statistics at level crossings by a stationary process. Statistica Neerlandica, 56(2):152–164, May 2002. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00190.

Mair:2002:MCS

[1638] B. A. Mair, Murali Rao, and J. M. M. Anderson. A minimum chi-squared method for indirect parameter estimation from Poisson data. Statistica Neerlandica, 56(2):165–178, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00191.

Brunner:2002:WRS

[1639] Edgar Brunner, Sebastian Domhof, and Madan L. Puri. Weighted rank statistics in factorial designs with fixed effects. Statistica Neerlandica, 56 (2):179–194, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00192.

Revesz:2002:LBL

[1640] P. Révész. Large balls left empty by a critical branching Wiener field. Statistica Neerlandica, 56(2):195–205, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00193.

VanRooij:2002:FIH

[1641] Arnoud Van Rooij. Fourier inversion and the Hausdorff distance. Statistica Neerlandica, 56(2):206–213, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00194.

Serfling:2002:QFM

[1642] Robert Serfling. Quantile functions for multivariate analysis: approaches and applications. Statistica Neerlandica, 56(2):214–232, May 2002. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00195.

Dempfle:2002:NED

[1643] Astrid Dempfle and Winfried Stute. Nonparametric estimation of a discontinuity in regression. Statistica Neerlandica, 56(2):233–242, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00196.

Fiocco:2002:SCP

[1644] Marta Fiocco and Willem R. Van Zwet. Statistics for the contact process. Statistica Neerlandica, 56(2):243–251, May 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00197.

Goovaerts:2002:SPA

[1645] M. J. Goovaerts and R. Kaas. Some problems in actuarial finance involving sums of dependent risks. Statistica Neerlandica, 56(3):253– 269, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.03600.

Moors:2002:TSS

[1646] J. J. A. Moors and L. W. G. Strijbosch. Two-step sequential sampling. Statistica Neerlandica, 56(3):270–284, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.05000.

Tsionas:2002:BIT

[1647] Efthymios G. Tsionas. Bayesian inference in time series models using kernel quasi likelihoods. Statistica Neerlandica, 56(3):285–294, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.04800.

Bouzar:2002:MRD

[1648] Nadjib Bouzar. Mixture representations for discrete Linnik laws. Statistica Neerlandica, 56(3):295–300, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.04900.

Berlinet:2002:GPL

[1649] A. Berlinet, T. Hobza, and I. Vajda. Generalized piecewise linear histograms. Statistica Neerlandica, 56(3):301–313, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.05600.

Fredette:2002:NAP

[1650] Marc Fredette and Jean-François Angers. A new approximation of the posterior distribution of the log–odds ratio. Statistica Neerlandica, 56(3): 314–329, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.t01-1-00057.

Didelez:2002:MSE

[1651] Vanessa Didelez. ML- and semiparametric estimation in logistic models with incomplete covariate data. Statistica Neerlandica, 56(3):330– 345, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.t01-1-00059.

Liang:2002:IEB

[1652] Tachen Liang. An improved empirical Bayes test for positive exponential families. Statistica Neerlandica, 56(3):346–361, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.t01-1-
00070.

Wedel:2002:CVF

[1653] Michel Wedel. Concomitant variables in finite mixture models. Statistica Neerlandica, 56(3):362–375, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.t01-1-00072.

DenHertog:2002:CQA

[1654] Dick Den Hertog, Etienne De Klerk, and Kees Roos. On convex quadratic approximation. Statistica Neerlandica, 56(3):376–385, August 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.t01-1-00075.

Bandyopadhyay:2002:TBS

[1655] Uttam Bandyopadhyay and Atanu Biswas. Test of Bernoulli success probability in inverse sampling for nearer alternatives using adaptive allocation. Statistica Neerlandica, 56(4):387–399, November 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00208.

Biswas:2002:BAD

[1656] Atanu Biswas and Jean-François Angers. A Bayesian adaptive design in clinical trials for continuous responses. Statistica Neerlandica, 56(4):400– 414, November 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00209.

Cai:2002:TSA

[1657] Zongwu Cai. A two-stage approach to additive time series models. Statistica Neerlandica, 56(4):415–433, November 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00210.

Antoniadis:2002:WTS

[1658] A. Antoniadis, D. Leporini, and J.-C. Pesquet. Wavelet thresholding for some classes of non–Gaussian noise. Statistica Neerlandica, 56(4):434– 453, November 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00211.

Loubes:2002:AES

[1659] Jean-Michel Loubes and Sara Van De Geer. Adaptive estimation with soft thresholding penalties. Statistica Neerlandica, 56(4):453–478, November 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.t01-1-00212.

Pak:2002:OTA

[1660] Kevin Pak and Nanda Piersma. Overview of OR techniques for airline revenue management. Statistica Neerlandica, 56(4):479–495, November 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00213.

Franses:2002:FSC

[1661] Philip Hans Franses. From first submission to citation: an empirical analysis. Statistica Neerlandica, 56(4):496–509, November 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00214.

Anonymous:2002:ISN

[1662] Anonymous. Index to Statistica Neerlandica 56, 2002. Statistica Neerlandica, 56(4):510–512, November 2002. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.t01-1-00215.

VanBuuren:2003:E

[1663] Stef Van Buuren and Rob Eisinga. Editorial. Statistica Neerlandica, 57 (1):1–2, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00216.

Rubin:2003:NMI

[1664] Donald B. Rubin. Nested multiple imputation of NMES via partially incompatible MCMC. Statistica Neerlandica, 57(1):3–18, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00217.

Schafer:2003:MIM

[1665] Joseph L. Schafer. Multiple imputation in multivariate problems when the imputation and analysis models differ. Statistica Neerlandica, 57(1): 19–35, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00218.

Brand:2003:TSE

[1666] Jaap P. L. Brand, Stef van Buuren, Karin Groothuis-Oudshoorn, and Edzard S. Gelsema. A toolkit in SAS for the evaluation of multiple imputation methods. Statistica Neerlandica, 57(1):36–45, February 2003. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00219.

Kamakura:2003:LAM

[1667] Wagner A. Kamakura and Michel Wedel. List augmentation with model based multiple imputation: a case study using a mixed-outcome factor model. Statistica Neerlandica, 57(1):46–57, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00220.

Rassler:2003:NIB

[1668] Susanne Rässler. A non-iterative Bayesian approach to statistical matching. Statistica Neerlandica, 57(1):58–74, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00221.

Fitzmaurice:2003:MHD

[1669] Garrett M. Fitzmaurice. Methods for handling dropouts in longitudinal clinical trials. Statistica Neerlandica, 57(1):75–99, February 2003. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00222.

Albert:2003:RET

[1670] Paul S. Albert and Dean A. Follmann. A random effects transition model for longitudinal binary data with informative missingness. Statistica Neerlandica, 57(1):100–111, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00223.

Molenberghs:2003:SAC

[1671] Geert Molenberghs, Herbert Thijs, Michael G. Kenward, and Geert Verbeke. Sensitivity analysis of continuous incomplete longitudinal outcomes. Statistica Neerlandica, 57(1):112–135, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00224.

Groeneboom:2003:DEU

[1672] P. Groeneboom and G. Jongbloed. Density estimation in the uniform deconvolution model. Statistica Neerlandica, 57(1):136–157, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00225.

DeGooijer:2003:CDE

[1673] Jan G. De Gooijer and Dawit Zerom. On conditional density estimation. Statistica Neerlandica, 57(2):159–176, May 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00226.

VanLieshout:2003:CMP

[1674] M. N. M. Van Lieshout and R. S. Stoica. The candy model: properties and inference. Statistica Neerlandica, 57(2):177–206, May 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00227.

Dette:2003:SMA

[1675] Holger Dette and Axel Munk. Some methodological aspects of validation of models in nonparametric regression. Statistica Neerlandica, 57(2): 207–244, May 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00228.

Anevski:2003:EDC

[1676] Dragi Anevski. Estimating the derivative of a convex density. Statistica Neerlandica, 57(2):245–257, May 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00229.

Kundu:2003:DDS

[1677] Debasis Kundu and Swagata Nandi. Determination of discrete spectrum in a random field. Statistica Neerlandica, 57(2):258–284, May 2003. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00230.

Bijleveld:2003:EI

[1678] Catrien Bijleveld and Hans Nijboer. Editorial introduction. Statistica Neerlandica, 57(3):285–288, August 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00231.

VanDerHeijden:2003:ESC

[1679] Peter G. M. Van Der Heijden, Maarten Cruyff, and Hans C. Van Houwelingen. Estimating the size of a criminal population from police records using the truncated Poisson regression model. Statistica Neerlandica, 57(3):289–304, August 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00232.

Bijleveld:2003:LMM

[1680] Catrien C. J. H. Bijleveld and Ab Mooijaart. Latent Markov modelling of recidivism data. Statistica Neerlandica, 57(3):305–320, August 2003. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00233.

vanTulder:2003:EEI

[1681] F. P. van Tulder and B. C. J. van Velthoven. Econom(etr)ics of crime and litigation. Statistica Neerlandica, 57(3):321–346, August 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00234.

Elffers:2003:ANI

[1682] Henk Elffers. Analysing neighbourhood influence in criminology. Statistica Neerlandica, 57(3):347–367, August 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/1467-9574.00235.

Sjerps:2003:SAI

[1683] M. Sjerps and A. D. Kloosterman. Statistical aspects of interpreting DNA profiling in legal cases. Statistica Neerlandica, 57(3):368– 389, August 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00236.

Cordeiro:2003:TCS

[1684] Gauss M. Cordeiro, Denise A. Botter, Lúcia P. Barroso, and Silvia L. P. Ferrari. Three corrected score tests for generalized linear models with dispersion covariates. Statistica Neerlandica, 57(4):391–409, November 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/1467-9574.00237.

Cherchye:2003:MAD

[1685] Laurens Cherchye and Thierry Post. Methodological advances in DEA: a survey and an application for the Dutch electricity sector. Statistica Neerlandica, 57(4):410–438, November 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00238.

Koopman:2003:TSM

[1686] Siem Jan Koopman and Marius Ooms. Time series modelling of daily tax revenues. Statistica Neerlandica, 57(4):439–469, November 2003. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00239.

Klaassen:2003:DS

[1687] Chris A. J. Klaassen and J. Theo Runnenburg. Discrete spacings. Statistica Neerlandica, 57(4):470–483, November 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/1467-9574.00240.

Kippers:2003:ESC

[1688] Jeanine Kippers, Erjen van Nierop, Richard Paap, and Philip Hans Franses. An empirical study of cash payments. Statistica Neerlandica, 57(4):484–508, November 2003. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/1467-9574.00241.

Anonymous:2003:SNI

[1689] Anonymous. Statistica Neerlandica 57, 2003: Index. Statistica Neerlandica, 57(4):509–510, November 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/1467-9574.00242.

Molenaar:2004:AHH

[1690] Ivo W. Molenaar. About handy, handmade and handsome models. Statistica Neerlandica, 58(1):1–20, February 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1046/j.0039-0402.2003.00110.x.

Goeman:2004:PSU

[1691] J. J. Goeman, S. Le Cessie, R. J. Baatenburg de Jong, and S. A. Van De Geer. Predicting survival using disease history: a model combining relative survival and frailty. Statistica Neerlandica, 58(1):21–34, February 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00244.x.

Schafgans:2004:FSP

[1692] Marcia M. A. Schafgans. Finite sample properties for the semiparametric estimation of the intercept of a censored regression model. Statistica Neerlandica, 58(1):35–56, February 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1046/j.0039-0402.2003.00107.x.

Geelhoed:2004:EPS

[1693] B. Geelhoed and H. J. Glass. Estimators for particulate sampling derived from a multinomial distribution. Statistica Neerlandica, 58(1):57– 74, February 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1046/j.0039-0402.2003.00106.x.

Huang:2004:STE

[1694] Kuo-Chung Huang. A survey technique for estimating the proportion and sensitivity in a dichotomous finite population. Statistica Neerlandica, 58(1):75–82, February 2004. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1046/j.0039-0402.2003.00113.x.

Ferger:2004:CMT

[1695] Dietmar Ferger. A continuous mapping theorem for the argmaxfunctional in the non-unique case. Statistica Neerlandica, 58(1):83– 96, February 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1046/j.0039-0402.2003.00111.x.

Resing:2004:MMQ

[1696] Jacques Resing and Ronald Rietman. The M/M/1 queue with gated random order of service. Statistica Neerlandica, 58(1):97–110, February 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1046/j.0039-0402.2003.00112.x.

Shieh:2004:ERC

[1697] Gwowen Shieh and Show-Li Jan. The effectiveness of randomized complete block design. Statistica Neerlandica, 58(1):111–124, February 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1046/j.0039-0402.2003.00109.x.

Maas:2004:EI

[1698] Cora Maas and Jeroen Vermunt. Editorial introduction. Statistica Neerlandica, 58(2):125–126, May 2004. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1046/j.0039-0402.2003.00251.x.

Maas:2004:RIM

[1699] Cora J. M. Maas and Joop J. Hox. Robustness issues in multilevel regression analysis. Statistica Neerlandica, 58(2):127–137, May 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1046/j.0039-0402.2003.
00252.x.

Fox:2004:MRE

[1700] Jean-Paul Fox. Modelling response error in school effectiveness research. Statistica Neerlandica, 58(2):138–160, May 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1046/j.0039-0402.2003.
00253.x.

Ebbes:2004:RRE

[1701] Peter Ebbes, Ulf Böckenholt, and Michel Wedel. Regressor and randomeffects dependencies in multilevel models. Statistica Neerlandica, 58(2): 161–178, May 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1046/j.0039-0402.2003.00254.x.

Kato:2004:THR

[1702] Bernet Sekasanvu Kato and Herbert Hoijtink. Testing homogeneity in a random intercept model using asymptotic, posterior predictive and plug-in p-values. Statistica Neerlandica, 58(2):179–196, May 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1046/j.0039-0402.2003.00255.x.

Candel:2004:PEB

[1703] Math J. J. M. Candel. Performance of empirical Bayes estimators of random coefficients in multilevel analysis: Some results for the random intercept-only model. Statistica Neerlandica, 58(2):197– 219, May 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1046/j.0039-0402.2003.00256.x.

Vermunt:2004:EAE

[1704] Jeroen K. Vermunt. An EM algorithm for the estimation of parametric and nonparametric hierarchical nonlinear models. Statistica Neerlandica, 58(2):220–233, May 2004. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1046/j.0039-0402.2003.00257.x.

vanDuijn:2004:REM

[1705] Marijtje A. J. van Duijn, Tom A. B. Snijders, and Bonne J. H. Zijlstra. p_{2}: a random effects model with covariates for directed graphs. Statistica Neerlandica, 58(2):234–254, May 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1046/j.0039-0402.2003.00258.x.

Corander:2004:BAD

[1706] Jukka Corander and Mattias Villani. Bayesian assessment of dimensionality in reduced rank regression. Statistica Neerlandica, 58(3):255– 270, August 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00108.x.

Lok:2004:ECE

[1707] Judith Lok, Richard Gill, Aad Van Der Vaart, and James Robins. Estimating the causal effect of a time-varying treatment on time-to-event using structural nested failure time models. Statistica Neerlandica, 58(3): 271–295, August 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00123.x.

Wilde:2004:EME

[1708] Joachim Wilde. Estimating multiple equation hybrid models with endogenous dummy regressors. Statistica Neerlandica, 58(3):296–312, August 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00122.x.

Manatunga:2004:EIW

[1709] Amita K. Manatunga and David Oakes. Efficiency of independence working analysis of correlated survival data with general cluster size. Statistica Neerlandica, 58(3):313–321, August 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2004.00124.x.

Brijs:2004:MPM

[1710] Tom Brijs, Dimitris Karlis, Gilbert Swinnen, Koen Vanhoof, Geert Wets, and Puneet Manchanda. A multivariate Poisson mixture model for marketing applications. Statistica Neerlandica, 58(3):322–348, August 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00125.x.

Berkhout:2004:BPC

[1711] Peter Berkhout and Erik Plug. A bivariate Poisson count data model using conditional probabilities. Statistica Neerlandica, 58(3):349–364, August 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00126.x.

Cramer:2004:SBL

[1712] J. S. Cramer. Scoring bank loans that may go wrong: a case study. Statistica Neerlandica, 58(3):365–380, August 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2004.00127.x.

Franses:2004:FYS

[1713] Philip Hans Franses. Fifty years since Koyck (1954). Statistica Neerlandica, 58(4):381–387, November 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2004.00266.x.

Bohm:2004:APE

[1714] S. Böhm, V. Schmidt, and L. Heinrich. Asymptotic properties of estimators for the volume fractions of jointly stationary random sets. Statistica Neerlandica, 58(4):388–406, November 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2004.00267.x.

Papanastassiou:2004:ESS

[1715] Demetrios Papanastassiou and Demetrios Ioannides. The estimation of a state space model by estimating functions with an application. Statistica Neerlandica, 58(4):407–427, November 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2004.00268.x.

Groenen:2004:CRV

[1716] Patrick J. F. Groenen and Jacqueline J. Meulman. A comparison of the ratio of variances in distance-based and classical multivariate analysis. Statistica Neerlandica, 58(4):428–439, November 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00269.x.

Boswijk:2004:IET

[1717] H. Peter Boswijk and Jurgen A. Doornik. Identifying, estimating and testing restricted cointegrated systems: an overview. Statistica Neerlandica, 58(4):440–465, November 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2004.00270.x.

Demirtas:2004:SDI

[1718] Hakan Demirtas. Simulation driven inferences for multiply imputed longitudinal datasets. Statistica Neerlandica, 58(4):466–482, November 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00271.x.

Hobijn:2004:GKT

[1719] Bart Hobijn, Philip Hans Franses, and Marius Ooms. Generalizations of the KPSS-test for stationarity. Statistica Neerlandica, 58(4):483–502, November 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00272.x.

Anonymous:2004:I

[1720] Anonymous. Index. Statistica Neerlandica, 58(4):503–505, November 2004. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2004.00273.x.

Klugkist:2005:EI

[1721] Irene Klugkist and Herbert Hoijtink. Editorial introduction. Statistica Neerlandica, 59(1):1–2, February 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2005.00274.x.

Berger:2005:PMP

[1722] James O. Berger and German Molina. Posterior model probabilities via path-based pairwise priors. Statistica Neerlandica, 59(1):3–15, February 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00275.x.

Chen:2005:CML

[1723] Ming-Hui Chen. Computing marginal likelihoods from a single MCMC output. Statistica Neerlandica, 59(1):16–29, February 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2005.
00276.x.

Chib:2005:ARM

[1724] Siddhartha Chib and Ivan Jeliazkov. Accept–reject Metropolis–Hastings sampling and marginal likelihood estimation. Statistica Neerlandica, 59(1):30–44, February 2005. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2005.00277.x.

vanderLinde:2005:DVS

[1725] Angelika van der Linde. DIC in variable selection. Statistica Neerlandica, 59(1):45–56, February 2005. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2005.00278.x.

Klugkist:2005:BMS

[1726] Irene Klugkist, Bernet Kato, and Herbert Hoijtink. Bayesian model selection using encompassing priors. Statistica Neerlandica, 59(1):57– 69, February 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00279.x.

Maris:2005:PPV

[1727] Gunter Maris. Posterior predictive p-values for classical null hypotheses. Statistica Neerlandica, 59(1):70–81, February 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2005.
00280.x.

Galindo-Garre:2005:TLL

[1728] Francisca Galindo-Garre and Jeroen K. Vermunt. Testing log-linear models with inequality constraints: a comparison of asymptotic, bootstrap, and posterior predictive p-values. Statistica Neerlandica, 59(1): 82–94, February 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00281.x.

Fox:2005:BMI

[1729] Jean-Paul Fox and Cees A. W. Glas. Bayesian modification indices for IRT models. Statistica Neerlandica, 59(1):95–106, February 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00282.x.

Zijlstra:2005:MSR

[1730] Bonne J. H. Zijlstra, Marijtje A. J. van Duijn, and Tom A. B. Snijders. Model selection in random effects models for directed graphs using approximated Bayes factors. Statistica Neerlandica, 59(1):107– 118, February 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00283.x.

Balk:2005:DPQ

[1731] Bert M. Balk. Divisia price and quantity indices: 80 years after. Statistica Neerlandica, 59(2):119–158, May 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2005.00284.x.

Haldrup:2005:LPF

[1732] Niels Haldrup and Peter Lildholdt. Local power functions of tests for double unit roots. Statistica Neerlandica, 59(2):159–179, May 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2005.
00285.x.

VanHoesel:2005:OTN

[1733] Stan Van Hoesel. Optimization in telecommunication networks. Statistica Neerlandica, 59(2):180–205, May 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2005.00286.x.

Masse:2005:NTD

[1734] Jean-Claude Massé and Radu Theodorescu. Neyman type A distribution revisited. Statistica Neerlandica, 59(2):206–213, May 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2005.
00287.x.

Nadarajah:2005:SDA

[1735] Saralees Nadarajah and Samuel Kotz. Sampling distributions associated with the multivariate t distribution. Statistica Neerlandica, 59(2): 214–234, May 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00288.x.

Dehling:2005:EI

[1736] Herold Dehling and Ton Steerneman. Editorial introduction. Statistica Neerlandica, 59(3):235–238, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2005.00300.x.

Bekker:2005:IVE

[1737] Paul A. Bekker and Jan van der Ploeg. Instrumental variable estimation based on grouped data. Statistica Neerlandica, 59(3):239– 267, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00296.x.

Eaton:2005:SPM

[1738] Morris L. Eaton and D. A. S. Fraser. Studentization and prediction in a multivariate normal setting. Statistica Neerlandica, 59(3):268– 276, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00297.x.

Haneveld:2005:MAR

[1739] Willem K. Klein Haneveld. Multilinear approximation on rectangles and the related moment problem. Statistica Neerlandica, 59(3):277– 297, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00295.x.

Lauter:2005:SRS

[1740] Jürgen Läuter, Ekkehard Glimm, and Markus Eszlinger. Search for relevant sets of variables in a high-dimensional setup keeping the familywise error rate. Statistica Neerlandica, 59(3):298–312, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00290.x.

Lee:2005:GTS

[1741] Mei-Ling Ting Lee and Herold G. Dehling. Generalized two-sample U-statistics for clustered data. Statistica Neerlandica, 59(3):313–323, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00298.x.

Mikosch:2005:HMM

[1742] Thomas Mikosch. How to model multivariate extremes if one must? Statistica Neerlandica, 59(3):324–338, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2005.
00289.x.

VanDerHoeven:2005:GSF

[1743] T. Van Der Hoeven and A. G. M. Steerneman. Gas system failure rate. Statistica Neerlandica, 59(3):339–364, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2005.
00294.x.

vanVark:2005:SAM

[1744] G. N. van Vark. Some applications of multivariate statistics to physical anthropology. Statistica Neerlandica, 59(3):365–374, August 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00291.x.

DeBoer:2005:E

[1745] Paul De Boer. Erratum. Statistica Neerlandica, 59(4):375, November 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00306.x.

Nicolau:2005:PVI

[1746] João Nicolau. Processes with volatility-induced stationarity: an application for interest rates. Statistica Neerlandica, 59(4):376–396, November 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00292.x.

Albers:2005:EFU

[1747] C. J. Albers, G. Th. De Roos, and W. Schaafsma. Estimating a frequency unseen: an application to ornithology. Statistica Neerlandica, 59(4):397– 413, November 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00293.x.

Breitung:2005:PUR

[1748] Jörg Breitung and Samarjit Das. Panel unit root tests under crosssectional dependence. Statistica Neerlandica, 59(4):414–433, November 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00299.x.

vandenHout:2005:FDW

[1749] Ardo van den Hout and Gerty J. L. M. Lensvelt-Mulders. On a 2 × 2 factorial design where the use of randomized response is one of the factors. Statistica Neerlandica, 59(4):434–447, November 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00301.x.

Henriques:2005:ERK

[1750] Carla Henriques and Paulo Eduardo Oliveira. Exponential rates for kernel density estimation under association. Statistica Neerlandica, 59 (4):448–466, November 2005. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2005.00302.x.

Huisman:2005:ORP

[1751] Dennis Huisman, Leo G. Kroon, Ramon M. Lentink, and Michiel J. C. M. Vromans. Operations research in passenger railway transportation. Statistica Neerlandica, 59(4):467–497, November 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2005.
00303.x.

deBoer:2005:RRR

[1752] Paul M. C. de Boer and Christian M. Hafner. Ridge regression revisited. Statistica Neerlandica, 59(4):498–505, November 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2005.
00304.x.

Anonymous:2005:I

[1753] Anonymous. Index. Statistica Neerlandica, 59(4):506–508, November 2005. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2005.00307.x.

Koning:2006:MTS

[1754] Ruud H. Koning. In memoriam Ton Steerneman (1954–28 September 2005). Statistica Neerlandica, 60(1):1–2, February 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2006.
00310.x.

Best:2006:TBI

[1755] D. J. Best, P. B. Brockhoff, and J. C. W. Rayner. Tests for balanced incomplete block ranked data with ties. Statistica Neerlandica, 60(1): 3–11, February 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00305.x.

Denteneer:2006:MDT

[1756] Dee Denteneer. Models for data transmission delay in cable networks. Statistica Neerlandica, 60(1):12–45, February 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2006.
00308.x.

Wu:2006:PSS

[1757] Shuo-Jye Wu, Ying-Po Lin, and Yi-Ju Chen. Planning step-stress life test with progressively type I group-censored exponential data. Statistica Neerlandica, 60(1):46–56, February 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2006.00309.x.

vanDijk:2006:TEJ

[1758] Bram van Dijk. Treatment effect of job-training programmes on unemployment duration in Slovakia. Statistica Neerlandica, 60(1):57–72, February 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00311.x.

Nobel:2006:WTP

[1759] R. D. Nobel and H. C. Tijms. Waiting-time probabilities in the M/G/1 retrial queue. Statistica Neerlandica, 60(1):73–78, February 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00312.x.

Franses:2006:EI

[1760] Philip Hans Franses, Patrick J. F. Groenen, and Albert P. M. Wagelmans. Editorial introduction. Statistica Neerlandica, 60(2): 79, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00316.x.

Dreze:2006:FYE

[1761] J. Drèze. Fifty years of econometric institute. Statistica Neerlandica, 60 (2):80–84, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00321.x.

VanDijk:2006:REA

[1762] H. K. Van Dijk, J. F. Kaashoek, and A. P. M. Wagelmans. ‘Rotterdam econometrics’: an analysis of publications of the econometric institute 1956–2004. Statistica Neerlandica, 60(2):85–111, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00320.x.

Wagelmans:2006:RPE

[1763] Albert P. M. Wagelmans. The Rotterdam program in Econometrics & Management Science. Statistica Neerlandica, 60(2):112–123, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00324.x.

Gower:2006:SDA

[1764] J. C. Gower. Statistica data analytica est et aliter. Statistica Neerlandica, 60(2):124–134, May 2006. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2006.00313.x.

Groenen:2006:VYP

[1765] Patrick J. F. Groenen and L. Andries van der Ark. Visions of 70 years of psychometrics: the past, present, and future. Statistica Neerlandica, 60 (2):135–144, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00318.x.

Don:2006:MME

[1766] F. J. H. Don and J. P. Verbruggen. Models and methods for economic policy: 60 years of evolution at CPB. Statistica Neerlandica, 60(2): 145–170, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00323.x.

Malinvaud:2006:DEY

[1767] E. Malinvaud. Disequilibrium econometrics: 25 years later. Statistica Neerlandica, 60(2):171–180, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2006.00319.x.

McCullough:2006:REY

[1768] B. D. McCullough, Charles G. Renfro, and Houston H. Stokes. Re-examining 50-year-old OLS estimates of the Klein–Goldberger model. Statistica Neerlandica, 60(2):181–193, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2006.
00322.x.

Cramer:2006:WCE

[1769] Mars Cramer. The working conditions of econometric scholars, then and now. Statistica Neerlandica, 60(2):194–205, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2006.
00315.x.

Ooms:2006:ESD

[1770] Marius Ooms and Jurgen A. Doornik. Econometric software development: past, present and future. Statistica Neerlandica, 60(2):206– 224, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00317.x.

Bos:2006:DDN

[1771] Frits Bos. The development of the Dutch national accounts as a tool for analysis and policy. Statistica Neerlandica, 60(2):225–258, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00314.x.

Boender:2006:TFY

[1772] Guus Boender and Gerrit Timmer. Twenty-five years of applied Operations Research at ORTEC. Statistica Neerlandica, 60(2):259– 268, May 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00325.x.

Does:2006:EI

[1773] Ronald J. M. M. Does and Jeroen De Mast. Editorial introduction. Statistica Neerlandica, 60(3):269, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2006.00334.x.

DeMast:2006:ISD

[1774] Jeroen De Mast and Ronald J. M. M. Does. Industrial statistics: a discipline with opportunities and challenges. Statistica Neerlandica, 60(3): 270–282, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00329.x.

Baillie:2006:CAS

[1775] David H. Baillie and Chris A. J. Klaassen. Credit to and in acceptance sampling. Statistica Neerlandica, 60(3):283–291, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2006.
00331.x.

Albers:2006:SAC

[1776] Willem Albers and Wilbert C. M. Kallenberg. Self-adapting control charts. Statistica Neerlandica, 60(3):292–308, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2006.
00330.x.

DelCastillo:2006:SPA

[1777] E. Del Castillo. Statistical process adjustment: a brief retrospective, current status, and some opportunities for further work. Statistica Neerlandica, 60(3):309–326, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2006.00328.x.

Koning:2006:MBC

[1778] Alex J. Koning. Model-based control charts in phase 1 statistical process control. Statistica Neerlandica, 60(3):327–338, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00335.x.

Ramaker:2006:PAI

[1779] Henk-Jan Ramaker, Eric N. M. Van Sprang, Johan A. Westerhuis, Stephen P. Gurden, Age K. Smilde, and Frank H. Van Der Meulen. Performance assessment and improvement of control charts for statistical batch process monitoring. Statistica Neerlandica, 60(3):339– 360, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00337.x.

Goos:2006:OVO

[1780] Peter Goos. Optimal versus orthogonal and equivalent-estimation design of blocked and split-plot experiments. Statistica Neerlandica, 60(3):361– 378, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00333.x.

Huele:2006:RSA

[1781] A. Freek Huele and Jan Engel. A response surface approach to tolerance design. Statistica Neerlandica, 60(3):379–395, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2006.
00332.x.

Kenett:2006:ASC

[1782] R. Kenett and P. Thyregod. Aspects of statistical consulting not taught by academia. Statistica Neerlandica, 60(3):396–411, August 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00327.x.

Koning:2006:EI

[1783] Ruud Koning and Siem Heisterkamp. Editorial introduction. Statistica Neerlandica, 60(4):413, November 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2006.00344.x.

Yekutieli:2006:AMI

[1784] Daniel Yekutieli, Anat Reiner-Benaim, Yoav Benjamini, Gregory I. Elmer, Neri Kafkafi, Noah E. Letwin, and Norman H. Lee. Approaches to multiplicity issues in complex research in microarray analysis. Statistica Neerlandica, 60(4):414–437, November 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2006.00343.x.

Franses:2006:MPT

[1785] Philip Hans Franses. On modeling panels of time series. Statistica Neerlandica, 60(4):438–456, November 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2006.00339.x.

Brijs:2006:MIR

[1786] Tom Brijs, Filip Van den Bossche, Geert Wets, and Dimitris Karlis. A model for identifying and ranking dangerous accident locations: a case study in Flanders. Statistica Neerlandica, 60(4):457–476, November 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00341.x.

Zhang:2006:STU

[1787] Biao Zhang. A score test under logistic regression models based on case–control data. Statistica Neerlandica, 60(4):477–496, November 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2006.00336.x.

Sun:2006:AHR

[1788] Liuquan Sun, Zhigang Zhang, and Jianguo Sun. Additive hazards regression of failure time data with covariate measurement errors. Statistica Neerlandica, 60(4):497–509, November 2006. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2006.00340.x.

VanHoesel:2007:EI

[1789] Stan Van Hoesel and Albert Wagelmans. Editorial introduction. Statistica Neerlandica, 61(1):1–3, February 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2007.00363.x.

Kolen:2007:GAP

[1790] Antoon Kolen. A genetic algorithm for the partial binary constraint satisfaction problem: an application to a frequency assignment problem. Statistica Neerlandica, 61(1):4–15, February 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2007.00357.x.

VanHoesel:2007:CPR

[1791] Stan P. M. Van Hoesel and Albert P. M. Wagelmans. On the pcoverage problem on the real line. Statistica Neerlandica, 61(1):16– 34, February 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00347.x.

Oosten:2007:DGR

[1792] Maarten Oosten, Jeroen H. G. C. Rutten, and Frits C. R. Spieksma. Disconnecting graphs by removing vertices: a polyhedral approach. Statistica Neerlandica, 61(1):35–60, February 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2007.00350.x.

Pijls:2007:HES

[1793] Wim Pijls. Heuristic estimates in shortest path algorithms. Statistica Neerlandica, 61(1):61–74, February 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00348.x.

Brauner:2007:MCI

[1794] N. Brauner, Y. Crama, A. Grigoriev, and J. Van De Klundert. Multiplicity and complexity issues in contemporary production scheduling. Statistica Neerlandica, 61(1):75–91, February 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2007.00362.x.

Dorndorf:2007:DMF

[1795] Ulrich Dorndorf, Florian Jaehn, Chen Lin, Hui Ma, and Erwin Pesch. Disruption management in flight gate scheduling. Statistica Neerlandica, 61(1):92–114, February 2007. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2007.00361.x.

Koster:2007:TLB

[1796] Arie M. C. A. Koster and Adrian Zymolka. Tight LP-based lower bounds for wavelength conversion in optical networks. Statistica Neerlandica, 61(1):115–136, February 2007. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2007.00351.x.

Kuijpers:2007:OPD

[1797] Bart H. M. Kuijpers and Peter C. Schotman. Optimal prepayment of Dutch mortgages. Statistica Neerlandica, 61(1):137–155, February 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00349.x.

Lok:2007:NAV

[1798] Reinder B. Lok, Dolores Romero Morales, and Dries Vermeulen. Note on the applicability of the VCG mechanism to capacitated assignment problems and extensions. Statistica Neerlandica, 61(1):156–171, February 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00356.x.

Yu:2007:KBM

[1799] Keming Yu, Jorge Mateu, and Emilio Porcu. A kernel-based method for nonparametric estimation of variograms. Statistica Neerlandica, 61(2): 173–197, May 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00326.x.

Anevski:2007:ITC

[1800] Dragi Anevski. Interarrival times in a counting process and bird watching. Statistica Neerlandica, 61(2):198–208, May 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2007.
00338.x.

Bentkus:2007:CBM

[1801] V. Bentkus, N. Kalosha, and M. C. A. Van Zuijlen. Confidence bounds for the mean in nonparametric multisample problems. Statistica Neerlandica, 61(2):209–231, May 2007. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2007.00342.x.

Tian:2007:ENI

[1802] Guo-Liang Tian, Ming Tan, and Kai Wang Ng. An exact non-iterative sampling procedure for discrete missing data problems. Statistica Neerlandica, 61(2):232–242, May 2007. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2007.00345.x.

Gomes:2007:RBT

[1803] M. Ivette Gomes, Cristina Miranda, and Clara Viseu. Reduced-bias tail index estimation and the jackknife methodology. Statistica Neerlandica, 61(2):243–270, May 2007. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2007.00346.x.

Anonymous:2007:C

[1804] Anonymous. Corrigendum. Statistica Neerlandica, 61(2):271, May 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (elec-

tronic).  URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00367.x.

Tomizawa:2007:ELD

[1805] Sadao Tomizawa, Nobuko Miyamoto, Kouji Yamamoto, and Akinobu Sugiyama. Extensions of linear diagonal-parameter symmetry and quasisymmetry models for cumulative probabilities in square contingency tables. Statistica Neerlandica, 61(3):273–283, August 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2007.
00353.x.

Huang:2007:COR

[1806] Kuo-Chung Huang. Constructing optimal randomized response designs with consideration for the level of privacy protection. Statistica Neerlandica, 61(3):284–291, August 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00352.x.

Ghosh:2007:AMS

[1807] Santu Ghosh, Kalyan Das, and Peter Congdon. Analysis of marginally specified semi-nonparametric models for clustered binary data. Statistica Neerlandica, 61(3):292–304, August 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00355.x.

Jasso:2007:NCD

[1808] Guillermina Jasso and Samuel Kotz. A new continuous distribution and two new families of distributions based on the exponential. Statistica Neerlandica, 61(3):305–328, August 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00358.x.

Bandyopadhyay:2007:ATT

[1809] Uttam Bandyopadhyay, Atanu Biswas, and Shirsendu Mukherjee. Adaptive two-treatment two-period crossover design for binary treatment responses. Statistica Neerlandica, 61(3):329–344, August 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2007.
00360.x.

Paulo:2007:BAE

[1810] M. J. Paulo and A. Otten. A Bayesian approach for exact optimal measurement and cutting times for a Eucalyptus production forest. Statistica Neerlandica, 61(3):345–357, August 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00359.x.

Porcu:2007:CFS

[1811] E. Porcu, J. Mateu, and M. Bevilacqua. Covariance functions that are stationary or nonstationary in space and stationary in time. Statistica Neerlandica, 61(3):358–382, August 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00364.x.

Bodnar:2007:SMB

[1812] Olha Bodnar and Wolfgang Schmid. Surveillance of the mean behavior of multivariate time series. Statistica Neerlandica, 61(4):383–406, November 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00365.x.

Schmid:2007:ELR

[1813] Matthias Schmid, Hans Schneeweiss, and Helmut Küchenhoff. Estimation of a linear regression under microaggregation with the response variable as a sorting variable. Statistica Neerlandica, 61(4):407–431, November 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00366.x.

McHale:2007:MSM

[1814] Ian McHale and Phil Scarf. Modelling soccer matches using bivariate discrete distributions with general dependence structure. Statistica Neerlandica, 61(4):432–445, November 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00368.x.

Alink:2007:DGC

[1815] Stan Alink, Matthias Löwe, and Mario V. Wüthrich. Diversification for general copula dependence. Statistica Neerlandica, 61(4):446–465, November 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00370.x.

Herwartz:2007:TRE

[1816] Helmut Herwartz. Testing for random effects in panel models with spatially correlated disturbances. Statistica Neerlandica, 61(4):466–487, November 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00369.x.

Zeileis:2007:GFT

[1817] Achim Zeileis and Kurt Hornik. Generalized M-fluctuation tests for parameter instability. Statistica Neerlandica, 61(4):488–508, November 2007. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00371.x.

Anonymous:2007:SN

[1818] Anonymous. Statistica Neerlandica 61, 2007. Statistica Neerlandica, 61(4):509–511, November 2007. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2007.00380.x.

Oud:2008:EI

[1819] Johan H. L. Oud and Hermann Singer. Editorial introduction. Statistica Neerlandica, 62(1):1–3, February 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00385.x.

Oud:2008:CTM

[1820] Johan H. L. Oud and Hermann Singer. Continuous time modeling of panel data: SEM versus filter techniques. Statistica Neerlandica, 62(1): 4–28, February 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00376.x.

Singer:2008:NCT

[1821] Hermann Singer. Nonlinear continuous time modeling approaches in panel research. Statistica Neerlandica, 62(1):29–57, February 2008. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2007.
00373.x.

Delsing:2008:ARR

[1822] Marc J. M. H. Delsing and Johan H. L. Oud. Analyzing reciprocal relationships by means of the continuous-time autoregressive latent trajectory model. Statistica Neerlandica, 62(1):58–82, February 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00386.x.

Toharudin:2008:ARB

[1823] Toni Toharudin, Johan H. L. Oud, and Jaak B. Billiet. Assessing the relationships between Nationalism, Ethnocentrism, and Individualism in Flanders using Bergstrom’s approximate discrete model. Statistica Neerlandica, 62(1):83–103, February 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00378.x.

Koopman:2008:ESC

[1824] Siem Jan Koopman, Marius Ooms, André Lucas, Kees van Montfort, and Victor Van Der Geest. Estimating systematic continuous-time trends in recidivism using a non-Gaussian panel data model. Statistica Neerlandica, 62(1):104–130, February 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00375.x.

Chow:2008:CTM

[1825] Sy-Miin Chow and Guangjian Zhang. Continuous-time modelling of irregularly spaced panel data using a cubic spline model. Statistica Neerlandica, 62(1):131–154, February 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00379.x.

Naskar:2008:SAC

[1826] Malay Naskar. Semiparametric analysis of clustered survival data under nonparametric frailty. Statistica Neerlandica, 62(2):155–172, May 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00372.x.

Grillenzoni:2008:SIS

[1827] Carlo Grillenzoni. Statistics for image sharpening. Statistica Neerlandica, 62(2):173–192, May 2008. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2007.00374.x.

Demirtas:2008:ICD

[1828] Hakan Demirtas and Donald Hedeker. Imputing continuous data under some non-Gaussian distributions. Statistica Neerlandica, 62(2):193– 205, May 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00377.x.

Nadarajah:2008:DJK

[1829] Saralees Nadarajah. On a distribution of Jasso and Kotz. Statistica Neerlandica, 62(2):206–207, May 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2007.00381.x.

Horvath:2008:PDA

[1830] Csilla Horváth and Jaap E. Wieringa. Pooling data for the analysis of dynamic marketing systems. Statistica Neerlandica, 62(2):208– 229, May 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2007.00382.x.

Lee:2008:MMV

[1831] Myoung jae Lee. Method-of-moment view of linear simultaneous equation systems. Statistica Neerlandica, 62(2):230–238, May 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2007.
00383.x.

Eisinga:2008:ILT

[1832] Rob Eisinga. Information loss for 2×2 tables with missing cell counts: binomial case. Statistica Neerlandica, 62(2):239–254, May 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2007.
00384.x.

Adan:2008:EI

[1833] Ivo Adan, Alessandro Di Bucchianico, and Remco Van Der Hofstad. Editorial introduction. Statistica Neerlandica, 62(3):255, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00403.x.

DenHollander:2008:EDE

[1834] W. Th. Frank Den Hollander, Wim J. M. Senden, and Willem R. Van Zwet. EURANDOM: a decade of European stochastics. Statistica Neerlandica, 62(3):256–265, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2008.00398.x.

Aalto:2008:RST

[1835] Samuli Aalto and Urtzi Ayesta. Recent sojourn time results for multilevel processor-sharing scheduling disciplines. Statistica Neerlandica, 62(3): 266–282, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00393.x.

Denteneer:2008:IPP

[1836] Dee Denteneer, Sem Borst, Peter van de Ven, and Guido Hiertz. IEEE 802.11s and the Philosophers’ problem. Statistica Neerlandica, 62(3): 283–298, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00396.x.

Janssen:2008:BRM

[1837] A. J. E. M. Janssen and J. S. H. Van Leeuwaarden. Back to the roots of the M/D/s queue and the works of Erlang, Crommelin and Pollaczek. Statistica Neerlandica, 62(3):299–313, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00395.x.

Camia:2008:SLT

[1838] Federico Camia. Scaling limits of two-dimensional percolation: an overview. Statistica Neerlandica, 62(3):314–330, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2008.
00400.x.

VanEnter:2008:GNG

[1839] Aernout C. D. Van Enter, Frank Redig, and Evgeny Verbitskiy. Gibbsian and non-Gibbsian states at Eurandom. Statistica Neerlandica, 62(3): 331–344, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00394.x.

Merkl:2008:MFL

[1840] Franz Merkl, Aniko Öry, and Silke W. W. Rolles. The ‘magic formula’ for linearly edge-reinforced random walks. Statistica Neerlandica, 62(3): 345–363, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00402.x.

Armstrong:2008:CFM

[1841] Nicola J. Armstrong. The changing focus of microarray analysis. Statistica Neerlandica, 62(3):364–373, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2008.00399.x.

Grunwald:2008:ECE

[1842] Peter Grünwald. Entropy concentration and the empirical coding game. Statistica Neerlandica, 62(3):374–392, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2008.
00401.x.

Zhang:2008:CSA

[1843] Jian Zhang and Faming Liang. Convergence of stochastic approximation algorithms under irregular conditions. Statistica Neerlandica, 62(3):393– 403, August 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00397.x.

Demirhan:2008:BEL

[1844] Haydar Demirhan and Canan Hamurkaroğlu. Bayesian estimation of log odds ratios from R×C and 2×2×K contingency tables. Statistica Neerlandica, 62(4):405–424, November 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2008.00387.x.

Jacome:2008:SRP

[1845] M. A. Jácome and R. Cao. Strong representation of the presmoothed quantile function estimator for censored data. Statistica Neerlandica, 62(4):425–440, November 2008. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2008.00388.x.

Bekker:2008:QSS

[1846] R. Bekker, O. J. Boxma, and J. A. C. Resing. Queues with service speed adaptations. Statistica Neerlandica, 62(4):441–457, November 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00389.x.

Riaz:2008:MPM

[1847] Muhammad Riaz. Monitoring process mean level using auxiliary information. Statistica Neerlandica, 62(4):458–481, November 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2008.
00390.x.

Borovkova:2008:CAN

[1848] Svetlana Borovkova, Hendrik P. Lopuhaä, and Budi Nurani Ruchjana. Consistency and asymptotic normality of least squares estimators in generalized STAR models. Statistica Neerlandica, 62(4):482–508, November 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00391.x.

Franses:2008:ES

[1849] Philip Hans Franses, Paul De Boer, and Elli Hoek Van Dijke. Editorial statistics. Statistica Neerlandica, 62(4):509, November 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2008.
00405.x.

Anonymous:2008:SNI

[1850] Anonymous. Statistica Neerlandica 62, 2008 index. Statistica Neerlandica, 62(4):510–512, November 2008. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2008.00406.x.

Gill:2009:SEP

[1851] Richard D. Gill. Statistics, ethics and probiotica. Statistica Neerlandica, 63(1):1–12, February 2009. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2008.00411.x.

Zhang:2009:RNR

[1852] Jin Zhang and Xueren Wang. Robust normal reference bandwidth for kernel density estimation. Statistica Neerlandica, 63(1):13–23, February 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00392.x.

Eisinga:2009:BBC

[1853] Rob Eisinga. The beta-binomial convolution model for 2 × 2 tables with missing cell counts. Statistica Neerlandica, 63(1):24–42, February 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00404.x.

Gabriel:2009:SOA

[1854] Edith Gabriel and Peter J. Diggle. Second-order analysis of inhomogeneous spatio-temporal point process data. Statistica Neerlandica, 63(1): 43–51, February 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00407.x.

VanNierop:2009:WDS

[1855] Erjen Van Nierop. Why do statistics journals have low impact factors? Statistica Neerlandica, 63(1):52–62, February 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2008.
00408.x.

Albers:2009:NCC

[1856] Willem Albers and Wilbert C. M. Kallenberg. Normal control charts with nonparametric safeguard. Statistica Neerlandica, 63(1):63–81, February 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00409.x.

Vanbelle:2009:ABI

[1857] S. Vanbelle and A. Albert. Agreement between an isolated rater and a group of raters. Statistica Neerlandica, 63(1):82–100, February 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00412.x.

Steland:2009:SIV

[1858] Ansgar Steland and Henryk Zähle. Sampling inspection by variables: nonparametric setting. Statistica Neerlandica, 63(1):101–123, February 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2008.00413.x.

Franses:2009:WGT

[1859] Philip Hans Franses. Why is GDP typically revised upwards? Statistica Neerlandica, 63(2):125–130, May 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2008.00410.x.

VanDeVen:2009:RPH

[1860] Peter Van De Ven. Reaction to Philip Hans Franses’ note ‘Why is GDP typically revised upwards?’. Statistica Neerlandica, 63(2):131– 132, May 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00418.x.

Kranendonk:2009:RPH

[1861] Henk Kranendonk and Johan Verbruggen. Reaction to Philip Hans Franses’ note ‘Why is GDP typically revised upwards?’. Statistica Neerlandica, 63(2):133–134, May 2009. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2009.00419.x.

Lopez-Fidalgo:2009:ODC

[1862] J. López-Fidalgo, M. J. Rivas-López, and R. Del Campo. Optimal designs for Cox regression. Statistica Neerlandica, 63(2):135– 148, May 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00415.x.

Ceyhan:2009:CST

[1863] Elvan Ceyhan. Class-specific tests of spatial segregation based on nearest neighbor contingency tables. Statistica Neerlandica, 63(2):149– 182, May 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00414.x.

Ursu:2009:EMA

[1864] Eugen Ursu and Pierre Duchesne. Estimation and model adequacy checking for multivariate seasonal autoregressive time series models with periodically varying parameters. Statistica Neerlandica, 63(2):183– 212, May 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00417.x.

vanMontfort:2009:ESE

[1865] Kees van Montfort, Ab Mooijaart, and Frits Meijerink. Estimating structural equation models with non-normal variables by using transformations. Statistica Neerlandica, 63(2):213–226, May 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2009.
00420.x.

Rueda:2009:NMA

[1866] M. Rueda and J. F. Muñoz. New model-assisted estimators for the distribution function using the Pseudo Empirical Likelihood Method. Statistica Neerlandica, 63(2):227–244, May 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2009.00421.x.

VanGeloven:2009:RAS

[1867] Nan Van Geloven, Eric A. Cator, Hendrik P. Lopuhaä, and Mart P. Janssen. Regression analysis on serial dilution data from virus validation robustness studies. Statistica Neerlandica, 63(3):245–257, August 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00416.x.

Marozzi:2009:CID

[1868] Marco Marozzi. A composite indicator dimension reduction procedure with application to university student satisfaction. Statistica Neerlandica, 63(3):258–268, August 2009. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2009.00422.x.

Nedyalkova:2009:GFR

[1869] Desislava Nedyalkova, Lionel Qualité, and Yves Tillé. General framework for the rotation of units in repeated survey sampling. Statistica Neerlandica, 63(3):269–293, August 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2009.00423.x.

Hafner:2009:TLV

[1870] Christian M. Hafner and Helmut Herwartz. Testing for linear vector autoregressive dynamics under multivariate generalized autoregressive heteroskedasticity. Statistica Neerlandica, 63(3):294–323, August 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00424.x.

Rayner:2009:TBU

[1871] J. C. W. Rayner and Eric J. Beh. Towards a better understanding of correlation. Statistica Neerlandica, 63(3):324–333, August 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00425.x.

Franses:2009:EOV

[1872] Philip Hans Franses, Michael McAleer, and Rianne Legerstee. Expert opinion versus expertise in forecasting. Statistica Neerlandica, 63(3): 334–346, August 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00426.x.

Habibi:2009:CPD

[1873] Reza Habibi. Change point detection in scale family distributions. Statistica Neerlandica, 63(3):347–352, August 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2009.00427.x.

Bandyopadhyay:2009:RAP

[1874] Uttam Bandyopadhyay and Rahul Bhattacharya. Response adaptive procedures with dual optimality. Statistica Neerlandica, 63(3):353– 367, August 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00428.x.

DeBoer:2009:TNN

[1875] Paul De Boer and Richard Paap. Testing non-nested demand relations: linear expenditure system versus indirect addilog. Statistica Neerlandica, 63(3):368–384, August 2009. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2009.00429.x.

Jolink:2009:JTS

[1876] Albert Jolink. Jan Tinbergen’s statistical contribution to economic policy. Statistica Neerlandica, 63(4):385–399, November 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2009.
00441.x.

Lee:2009:TZI

[1877] JungBok Lee, Byoung Cheol Jung, and Seo Hoon Jin. Tests for zero inflation in a bivariate zero-inflated Poisson model. Statistica Neerlandica, 63(4):400–417, November 2009. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2009.00430.x.

Liu:2009:MST

[1878] Wei Liu, Jian Tao, Ning-Zhong Shi, and Man-Lai Tang. Model selection in toxicity studies. Statistica Neerlandica, 63(4):418–431, November 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00431.x.

Douhou:2009:RUA

[1879] Salima Douhou and Jan R. Magnus. The reliability of user authentication through keystroke dynamics. Statistica Neerlandica, 63(4):432– 449, November 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00434.x.

Myllymaki:2009:CHI

[1880] Mari Myllymäki and Antti Penttinen. Conditionally heteroscedastic intensity-dependent marking of log Gaussian Cox processes. Statistica Neerlandica, 63(4):450–473, November 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2009.00433.x.

DeGelder:2009:TDC

[1881] E. R. De Gelder and A. P. M. Wagelmans. The two-dimensional cutting stock problem within the roller blind production process. Statistica Neerlandica, 63(4):474–489, November 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2009.00436.x.

Lam:2009:CIM

[1882] Kar Yin Lam, Alex Koning, and Philip Hans Franses. Confidence intervals for maximal reliability in tau-equivalent models. Statistica Neerlandica, 63(4):490–507, November 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2009.00439.x.

Franses:2009:AES

[1883] Philip Hans Franses, Paul De Boer, and Elli Hoek Van Dijke. Announcement and editorial statistics. Statistica Neerlandica, 63(4):508, November 2009. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00442.x.

Anonymous:2009:SN

[1884] Anonymous. Statistica Neerlandica 63, 2009. Statistica Neerlandica, 63(4):509–511, November 2009. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2009.00443.x.

Stepaniak:2010:MPS

[1885] Pieter S. Stepaniak, Christiaan Heij, and Guus De Vries. Modeling and prediction of surgical procedure times. Statistica Neerlandica, 64(1):1– 18, February 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00440.x.

Comas:2010:TOP

[1886] Carlos Comas, Jorge Mateu, and Aila Särkkä. A third-order point process characteristic for multi-type point processes. Statistica Neerlandica, 64(1):19–44, February 2010. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2009.00437.x.

Balabdaoui:2010:EMD

[1887] Fadoua Balabdaoui and Jon A. Wellner. Estimation of a k-monotone density: characterizations, consistency and minimax lower bounds. Statistica Neerlandica, 64(1):45–70, February 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2009.00438.x.

vanNierop:2010:IYI

[1888] Erjen van Nierop. The introduction of the 5-year impact factor: does it benefit statistics journals? Statistica Neerlandica, 64(1):71–76, February 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00448.x.

Kateri:2010:MSS

[1889] M. Kateri, U. Kamps, and N. Balakrishnan. Multi-sample simple stepstress experiment under time constraints. Statistica Neerlandica, 64(1): 77–96, February 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00444.x.

Ling:2010:GAT

[1890] Shiqing Ling and Michael McAleer. A general asymptotic theory for time-series models. Statistica Neerlandica, 64(1):97–111, February 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00447.x.

Famoye:2010:NBG

[1891] Felix Famoye. A new bivariate generalized Poisson distribution. Statistica Neerlandica, 64(1):112–124, February 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2009.00446.x.

Franses:2010:NCO

[1892] Philip Hans Franses and J. S. Cramer. On the number of categories in an ordered regression model. Statistica Neerlandica, 64(1):125–128, February 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00432.x.

Lenstra:2010:MGL

[1893] Jan Karel Lenstra and Henk Tijms. In memoriam Gijs de Leve (1926– 2009). Statistica Neerlandica, 64(2):129–130, May 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2010.
00455.x.

Adan:2010:MJW

[1894] Ivo Adan, Onno Boxma, and Jan Van Der Wal. In memoriam: Jaap Wessels (1939–2009). Statistica Neerlandica, 64(2):131–132, May 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00456.x.

Todem:2010:SRM

[1895] David Todem, KyungMann Kim, Jason Fine, and Limin Peng. Semiparametric regression models and sensitivity analysis of longitudinal data with non-random dropouts. Statistica Neerlandica, 64(2):133– 156, May 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2009.00435.x.

Hox:2010:EEM

[1896] Joop J. Hox, Cora J. M. Maas, and Matthieu J. S. Brinkhuis. The effect of estimation method and sample size in multilevel structural equation modeling. Statistica Neerlandica, 64(2):157–170, May 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2009.
00445.x.

Ezzahrioui:2010:SAR

[1897] M’hamed Ezzahrioui and Elias Ould Saïd. Some asymptotic results of a non-parametric conditional mode estimator for functional timeseries data. Statistica Neerlandica, 64(2):171–201, May 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2010.
00449.x.

Bandyopadhyay:2010:CAA

[1898] Uttam Bandyopadhyay, Atanu Biswas, and Rahul Bhattacharya. A covariate-adjusted adaptive design for two-stage clinical trials with survival data. Statistica Neerlandica, 64(2):202–226, May 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2010.
00451.x.

You:2010:SIS

[1899] Jinhong You and Xian Zhou. Statistical inference on seemingly unrelated varying coefficient partially linear models. Statistica Neerlandica, 64(2): 227–253, May 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00450.x.

Gugushvili:2010:EI

[1900] Shota Gugushvili, Chris Klaassen, and Peter Spreij. Editorial introduction. Statistica Neerlandica, 64(3):255–256, August 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2010.
00459.x.

Papapantoleon:2010:ONA

[1901] Antonis Papapantoleon. Old and new approaches to LIBOR modeling. Statistica Neerlandica, 64(3):257–275, August 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2010.
00458.x.

Corcuera:2010:PVI

[1902] José Manuel Corcuera and Gergely Farkas. Power variation for Itô integrals with respect to α-stable processes. Statistica Neerlandica, 64(3): 276–289, August 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00463.x.

Comte:2010:NPE

[1903] Fabienne Comte and Valentine Genon-Catalot. Non-parametric estimation for pure jump irregularly sampled or noisy Lévy processes. Statistica Neerlandica, 64(3):290–313, August 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2010.00462.x.

Kappus:2010:ECL

[1904] Johanna Kappus and Markus Reiß. Estimation of the characteristics of a Lévy process observed at arbitrary frequency. Statistica Neerlandica, 64(3):314–328, August 2010. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2010.00461.x.

Podolskij:2010:ULT

[1905] Mark Podolskij and Mathias Vetter. Understanding limit theorems for semimartingales: a short survey. Statistica Neerlandica, 64(3):329– 351, August 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00460.x.

Bingham:2010:MEP

[1906] N. H. Bingham, John M. Fry, and Rüdiger Kiesel. Multivariate elliptic processes. Statistica Neerlandica, 64(3):352–366, August 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00465.x.

Pijls:2010:MFI

[1907] Wim Pijls and Walter A. Kosters. Mining frequent itemsets: a perspective from operations research. Statistica Neerlandica, 64(4):367–387, November 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00452.x.

Jackson:2010:MEC

[1908] Monica C. Jackson, Laura Johansen, Cathy Furlong, Abigail Colson, and Kimberly F. Sellers. Modelling the effect of climate change on prevalence of malaria in western Africa. Statistica Neerlandica, 64(4):388– 400, November 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00453.x.

Litvak:2010:SPA

[1909] Nelly Litvak and Maria Vlasiou. A survey on performance analysis of warehouse carousel systems. Statistica Neerlandica, 64(4):401–447, November 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00454.x.

Kamst:2010:OMS

[1910] Richard Kamst, Gerard H. Kuper, and Gerard Sierksma. The Olympic 500-m speed skating; the inner–outer lane difference. Statistica Neerlandica, 64(4):448–459, November 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2010.00457.x.

Zardasht:2010:EWB

[1911] V. Zardasht and M. Asadi. Evaluation of P (X_{t} > Y_{t}) when both X_{t} and Y_{t} are residual lifetimes of two systems. Statistica Neerlandica, 64(4):460– 481, November 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00464.x.

Janssen:2010:ELR

[1912] Arnold Janssen and Stefan Wellek. Exact linear rank tests for two-sample equivalence problems with continuous data. Statistica Neerlandica, 64 (4):482–504, November 2010. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2010.00466.x.

Nishiyama:2010:MCE

[1913] Yoichi Nishiyama. Moment convergence of M-estimators. Statistica Neerlandica, 64(4):505–507, November 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2010.00469.x.

Franses:2010:ES

[1914] Philip Hans Franses, Paul De Boer, and Elli Hoek Van Dijke. Editorial statistics. Statistica Neerlandica, 64(4):508, November 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2010.
00473.x.

Anonymous:2010:IV

[1915] Anonymous. Index of volume 64. Statistica Neerlandica, 64(4):509– 511, November 2010. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00472.x.

Qu:2011:DAL

[1916] Hao Qu and Feng-Chang Xie. Diagnostics analysis for log-Birnbaum– Saunders regression models with censored data. Statistica Neerlandica, 65(1):1–21, February 2011. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2010.00467.x.

Eisinga:2011:SAM

[1917] Rob Eisinga and Ben Pelzer. Saddlepoint approximations to the mean and variance of the extended hypergeometric distribution. Statistica Neerlandica, 65(1):22–31, February 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2010.00468.x.

Einmahl:2011:UMW

[1918] John H. J. Einmahl and Sander G. W. R. Smeets. Ultimate 100-m world records through extreme-value theory. Statistica Neerlandica, 65(1):32– 42, February 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00470.x.

Scarf:2011:NST

[1919] P. A. Scarf and M. M. Yusof. A numerical study of tournament structure and seeding policy for the soccer world cup finals. Statistica Neerlandica, 65(1):43–57, February 2011. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2010.00471.x.

Gill:2011:MHP

[1920] Richard D. Gill. The Monty Hall problem is not a probability puzzle (it’s a challenge in mathematical modelling). Statistica Neerlandica, 65(1): 58–71, February 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00474.x.

Lee:2011:WTP

[1921] Chihoon Lee and Jianqiang C. Wang. Waiting time probabilities in the M/G/1 + M queue. Statistica Neerlandica, 65(1):72–83, February 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00476.x.

Osei:2011:HBM

[1922] Frank B. Osei, Alfred A. Duker, and Alfred Stein. Hierarchical Bayesian modeling of the space-time diffusion patterns of cholera epidemic in kumasi, Ghana. Statistica Neerlandica, 65(1):84–100, February 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00475.x.

Franses:2011:RCP

[1923] Philip Hans Franses and Richard Paap. Random-coefficient periodic autoregressions. Statistica Neerlandica, 65(1):101–115, February 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2010.00477.x.

tenCate:2011:PAA

[1924] Arie ten Cate. Patterns on an ancient artifact: a coincidence? Statistica Neerlandica, 65(1):116–124, February 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2010.00478.x.

Caporin:2011:TNI

[1925] Massimiliano Caporin and Michael McAleer. Thresholds, news impact surfaces and dynamic asymmetric multivariate GARCH. Statistica Neerlandica, 65(2):125–163, May 2011. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2010.00479.x.

Li:2011:AFT

[1926] Jialiang Li and Mei-Ling Ting Lee. Analysis of failure time using threshold regression with semi-parametric varying coefficients. Statistica Neerlandica, 65(2):164–182, May 2011. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2011.00481.x.

vanLieshout:2011:FIP

[1927] M. N. M. van Lieshout. A J-function for inhomogeneous point processes. Statistica Neerlandica, 65(2):183–201, May 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00482.x.

Slooten:2011:FII

[1928] Klaas Slooten and Ronald Meester. Forensic identification: the island problem and its generalisations. Statistica Neerlandica, 65(2):202– 237, May 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00484.x.

Jonker:2011:ETL

[1929] Marianne Jonker, Aad van der Vaart, Zoltan Bochdanovits, and Dorret Boomsma. Estimating and testing linkage disequilibrium from sib data. Statistica Neerlandica, 65(3):239–258, August 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00480.x.

Greer:2011:BIE

[1930] Brandi A. Greer, James D. Stamey, and Dean M. Young. Bayesian interval estimation for the difference of two independent Poisson rates using data subject to under-reporting. Statistica Neerlandica, 65(3):259– 274, August 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00483.x.

vanEs:2011:CKE

[1931] Bert van Es. Combining kernel estimators in the uniform deconvolution problem. Statistica Neerlandica, 65(3):275–296, August 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00485.x.

Xu:2011:SIS

[1932] Qinfeng Xu, Jinhong You, and Xiaoli Li. Statistical inference on seemingly unrelated non-parametric regression models with serially correlated errors. Statistica Neerlandica, 65(3):297–318, August 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00486.x.

Yin:2011:GVB

[1933] Yuliang Yin. Generalized P -values and Bayesian evidence in the onesided testing problems under exponential distributions. Statistica Neerlandica, 65(3):319–336, August 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2011.00487.x.

Ferrari:2011:DTB

[1934] Silvia L. P. Ferrari, Patricia L. Espinheira, and Francisco Cribari-Neto. Diagnostic tools in beta regression with varying dispersion. Statistica Neerlandica, 65(3):337–351, August 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2011.00488.x.

Heij:2011:CSE

[1935] Christiaan Heij and Philip Hans Franses. Correcting for survey effects in pre-election polls. Statistica Neerlandica, 65(3):352–370, August 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00489.x.

Swartz:2011:BCP

[1936] Tim B. Swartz. Bayesian clustering with priors on partitions. Statistica Neerlandica, 65(4):371–386, November 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2011.00490.x.

Nkurunziza:2011:ESR

[1937] Sévérien Nkurunziza and S. Ejaz Ahmed. Estimation strategies for the regression coefficient parameter matrix in multivariate multiple regression. Statistica Neerlandica, 65(4):387–406, November 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00491.x.

Boonstra:2011:MII

[1938] Harm Jan H. Boonstra, Chris J. de Blois, and Gert-Jan M. Linders. Macro-integration with inequality constraints: an application to the integration of transport and trade statistics. Statistica Neerlandica, 65 (4):407–431, November 2011. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2011.00492.x.

Gromicho:2011:GPE

[1939] Joaquim A. S. Gromicho, Erwin Oudshoorn, and Gerhard Post. Generating price-effective intermodal routes. Statistica Neerlandica, 65(4):432– 445, November 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00493.x.

Batsidis:2011:EMD

[1940] A. Batsidis and K. Zografos. Errors of misclassification in discrimination of dimensional coherent elliptic random field observations. Statistica Neerlandica, 65(4):446–461, November 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2011.00494.x.

Caeiro:2011:ACO

[1941] Frederico Caeiro and M. Ivette Gomes. Asymptotic comparison at optimal levels of reduced-bias extreme value index estimators. Statistica Neerlandica, 65(4):462–488, November 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2011.00495.x.

Kuiper:2011:HHM

[1942] Rebecca M. Kuiper and Herbert Hoijtink. How to handle missing data in regression models using information criteria. Statistica Neerlandica, 65(4):489–506, November 2011. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2011.00496.x.

Anonymous:2011:ES

[1943] Anonymous. Editorial statistics. Statistica Neerlandica, 65(4):507, November 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00502.x.

Anonymous:2011:IV

[1944] Anonymous. Index of volume 65. Statistica Neerlandica, 65(4):508– 510, November 2011. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00503.x.

Bakker:2012:SIM

[1945] Bart Bakker and Johan van Rooijen. This special issue on Methodological challenges of register-based research is compiled by guest editors. Statistica Neerlandica, 66(1):1, February 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2011.00520.x.

Bakker:2012:MCR

[1946] Bart F. M. Bakker and Piet J. H. Daas. Methodological challenges of register-based research. Statistica Neerlandica, 66(1):2–7, February 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00505.x.

Bakker:2012:EVA

[1947] Bart F. M. Bakker. Estimating the validity of administrative variables. Statistica Neerlandica, 66(1):8–17, February 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00504.x.

Berka:2012:CEM

[1948] Christopher Berka, Stefan Humer, Mathias Moser, Manuela Lenk, Henrik Rechta, and Eliane Schwerer. Combination of evidence from multiple administrative data sources: quality assessment of the Austrian register-based Census 2011. Statistica Neerlandica, 66(1):18–33, February 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00506.x.

Holmberg:2012:DAQ

[1949] Anders Holmberg. Discussion on assessing quality of administrative data. Statistica Neerlandica, 66(1):34–40, February 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00507.x.

Zhang:2012:TST

[1950] Li-Chun Zhang. Topics of statistical theory for register-based statistics and data integration. Statistica Neerlandica, 66(1):41–63, February 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00508.x.

Kim:2012:RAU

[1951] Gunky Kim and Raymond Chambers. Regression analysis under probabilistic multi-linkage. Statistica Neerlandica, 66(1):64–79, February 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00509.x.

Karr:2012:DSU

[1952] Alan F. Karr. Discussion on statistical use of administrative data: old and new challenges. Statistica Neerlandica, 66(1):80–84, February 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00510.x.

Cai:2012:PVC

[1953] Zongwu Cai and Huaiyu Xiong. Partially varying coefficient instrumental variables models. Statistica Neerlandica, 66(2):85–110, May 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00497.x.

Pang:2012:ORR

[1954] Johnny Pang and Frits Ruymgaart. Optimal risk rate in a functional inverse problem. Statistica Neerlandica, 66(2):111–120, May 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00498.x.

Mahapatra:2012:SEH

[1955] A. K. Mahapatra, Somesh Kumar, and P. Vellaisamy. Simultaneous estimation of hazard rates of several exponential populations. Statistica Neerlandica, 66(2):121–132, May 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2011.00499.x.

Marra:2012:MSD

[1956] Giampiero Marra, David L. Miller, and Luca Zanin. Modelling the spatiotemporal distribution of the incidence of resident foreign population. Statistica Neerlandica, 66(2):133–160, May 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00500.x.

Choi:2012:BER

[1957] Jin young Choi and Myoung jae Lee. Bounding endogenous regressor coefficients using moment inequalities and generalized instruments. Statistica Neerlandica, 66(2):161–182, May 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2011.00501.x.

Wang:2012:SSA

[1958] Shun-Fang Wang and Xue-Ren Wang. Stratified studies about risk ratios in 2 × 2 tables with structural zero. Statistica Neerlandica, 66(2): 183–202, May 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2011.00511.x.

Fu:2012:ECF

[1959] Lianyan Fu, Wei Gao, and Ning-Zhong Shi. Estimating cell frequencies under inequality constraints based on the Kullback–Leibler information. Statistica Neerlandica, 66(2):203–216, May 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2011.
00513.x.

Wit:2012:AMW

[1960] Ernst Wit, Edwin van den Heuvel, and Jan-Willem Romeijn. ‘All models are wrong … ’: an introduction to model uncertainty. Statistica Neerlandica, 66(3):217–236, August 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2012.00530.x.

Longford:2012:WMW

[1961] Nicholas T. Longford. ‘Which model?’ is the wrong question. Statistica Neerlandica, 66(3):237–252, August 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2011.00517.x.

vanderLinde:2012:BVM

[1962] Angelika van der Linde. A Bayesian view of model complexity. Statistica Neerlandica, 66(3):253–271, August 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2011.00518.x.

Claeskens:2012:FEM

[1963] Gerda Claeskens. Focused estimation and model averaging with penalization methods: an overview. Statistica Neerlandica, 66(3):272– 287, August 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00514.x.

Friel:2012:EER

[1964] Nial Friel and Jason Wyse. Estimating the evidence — a review. Statistica Neerlandica, 66(3):288–308, August 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2011.00515.x.

Hastie:2012:MCU

[1965] David I. Hastie and Peter J. Green. Model choice using reversible jump Markov chain Monte Carlo. Statistica Neerlandica, 66(3):309– 338, August 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00516.x.

Wenmackers:2012:MSM

[1966] Sylvia Wenmackers and Danny E. P. Vanpoucke. Models and simulations in material science: two cases without error bars. Statistica Neerlandica, 66(3):339–355, August 2012. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2011.00519.x.

Knoth:2012:MOC

[1967] Sven Knoth and Marianne Frisén. Minimax optimality of CUSUM for an autoregressive model. Statistica Neerlandica, 66(4):357–379, November 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00512.x.

Ferger:2012:ETC

[1968] Dietmar Ferger, Jens Klotsche, and Ulrike Lüken. Estimation and testing of crossing-points in fixed design regression. Statistica Neerlandica, 66 (4):380–402, November 2012. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/j.1467-9574.2012.00521.x.

Giraldo:2012:HCS

[1969] R. Giraldo, P. Delicado, and J. Mateu. Hierarchical clustering of spatially correlated functional data. Statistica Neerlandica, 66(4):403–421, November 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00522.x.

Headrick:2012:SNN

[1970] Todd C. Headrick and Mohan D. Pant. Simulating non-normal distributions with specified L-moments and L-correlations. Statistica Neerlandica, 66(4):422–441, November 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2012.00523.x.

Singh:2012:TSS

[1971] Sarjinder Singh and Stephen A. Sedory. A true simulation study of three estimators at equal protection of respondents in randomized response sampling. Statistica Neerlandica, 66(4):442–451, November 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00524.x.

Thas:2012:UOT

[1972] O. Thas, D. J. Best, and J. C. W. Rayner. Using orthogonal trend contrasts for testing ranked data with ordered alternatives. Statistica Neerlandica, 66(4):452–471, November 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2012.00525.x.

Moller:2012:ASO

[1973] Jesper Møller and Mohammad Ghorbani. Aspects of second-order analysis of structured inhomogeneous spatio-temporal point processes. Statistica Neerlandica, 66(4):472–491, November 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2012.
00526.x.

Diaz-Garcia:2012:OAM

[1974] José A. Díaz-García and Rogelio Ramos-Quiroga. Optimum allocation in multivariate stratified random sampling: stochastic matrix mathematical programming. Statistica Neerlandica, 66(4):492–511, November 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00527.x.

Anonymous:2012:ES

[1975] Anonymous. Editorial statistics. Statistica Neerlandica, 66(4):512, November 2012. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00539.x.

Mendes:2013:TET

[1976] Mehmet Mendeş and Soner Yiğit. Type I error and test power of different tests for testing interaction effects in factorial experiments. Statistica Neerlandica, 67(1):1–26, February 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2012.00528.x. See com-

ments [1997, 1999].

Chang:2013:RJQ

[1977] Chia-Lin Chang and Michael McAleer. Ranking journal quality by harmonic mean of ranks: an application to ISI statistics & probability. Statistica Neerlandica, 67(1):27–53, February 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2012.
00529.x.

Einbeck:2013:NMR

[1978] Jochen Einbeck and James Taylor. A number-of-modes reference rule for density estimation under multimodality. Statistica Neerlandica, 67(1): 54–66, February 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00531.x.

Saffari:2013:IIE

[1979] Seyed Ehsan Saffari, Robiah Adnan, and William Greene. Investigating the impact of excess zeros on hurdle-generalized Poisson regression model with right censored count data. Statistica Neerlandica, 67(1):67– 80, February 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00532.x.

Ahmad:2013:NMT

[1980] M. Rauf Ahmad, Dietrich von Rosen, and Martin Singull. A note on mean testing for high dimensional multivariate data under nonnormality. Statistica Neerlandica, 67(1):81–99, February 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2012.
00533.x.

Al-Rawwash:2013:RML

[1981] Mohammad Al-Rawwash and Amjad D. Al-Nasser. Repeated measures and longitudinal data analysis using higher-order entropies. Statistica Neerlandica, 67(1):100–111, February 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/j.1467-9574.2012.00534.x.

Tressler:2013:MPC

[1982] Ameina Tressler and Alan Chow. Multiple pairwise comparison procedures based on the Lin and Wang test for right censored survival data. Statistica Neerlandica, 67(1):112–120, February 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/epdf/10.1111/j.1467-9574.2012.
00535.x.

Yu:2013:RBI

[1983] Zhuangxiong Yu, Meijin Wang, and Peng Wang. Residual-based IV estimation of dynamic panel data models with fixed effects. Statistica Neerlandica, 67(2):121–144, May 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/j.1467-9574.2012.00536.x.

Franke:2013:PED

[1984] Brice Franke and Thomas Kott. Parameter estimation for the drift of a time inhomogeneous jump diffusion process. Statistica Neerlandica, 67 (2):145–168, May 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/j.1467-9574.2012.00537.x.

Cuong:2013:WCS

[1985] Nguyen Viet Cuong. Which covariates should be controlled in propensity score matching? Evidence from a simulation study. Statistica Neerlandica, 67(2):169–180, May 2013. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12000.

Odijk:2013:NLM

[1986] Michiel A. Odijk and Pieter H. A. J. M. van Gelder. A note on the length, measured along fixed orientations, of a path connecting any two points relative to their straight line distance. Statistica Neerlandica, 67 (2):181–189, May 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12001.

Eisinga:2013:SAS

[1987] Rob Eisinga, Manfred Te Grotenhuis, and Ben Pelzer. Saddlepoint approximations for the sum of independent non-identically distributed binomial random variables. Statistica Neerlandica, 67(2):190–201, May 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12002.

Brinks:2013:SDE

[1988] Ralph Brinks, Rebecca Fischer-Betz, Sandra Landwehr, Matthias Schneider, and Guido Giani. Stochastic differential equation modeling the prevalence of rare chronic diseases with an application to systemic Lupus erythematosus. Statistica Neerlandica, 67(2):202–210, May 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12003.

Cho:2013:OAM

[1989] Wendy K. Tam Cho, Jason J. Sauppe, Alexander G. Nikolaev, Sheldon H. Jacobson, and Edward C. Sewell. An optimization approach for making causal inferences. Statistica Neerlandica, 67(2):211–226, May 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12004.

Gandolfi:2013:MED

[1990] A. Gandolfi, M. Benelli, A. Magi, and S. Chiti. Moment estimation in discrete shifting level model applied to fast array-CGH segmentation. Statistica Neerlandica, 67(3):227–262, August 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12005.

Baran:2013:SEL

[1991] Jerzy Baran and Agnieszka Stȩpień-Baran. Sequential estimation of a location parameter and powers of a scale parameter from delayed observations. Statistica Neerlandica, 67(3):263–280, August 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12006.

Fok:2013:TEM

[1992] Dennis Fok and Philip Hans Franses. Testing earnings management. Statistica Neerlandica, 67(3):281–292, August 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12007.

Roy:2013:ABD

[1993] Surupa Roy, Kalyan Das, and Angshuman Sarkar. Analysis of binary data with the possibility of wrong ascertainment. Statistica Neerlandica, 67(3):293–310, August 2013. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12008.

Kleinke:2013:MII

[1994] Kristian Kleinke and Jost Reinecke. Multiple imputation of incomplete zero-inflated count data. Statistica Neerlandica, 67(3):311–336, August 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12009.

Langovoy:2013:RAS

[1995] Mikhail Langovoy and Olaf Wittich. Randomized algorithms for statistical image analysis and site percolation on square lattices. Statistica Neerlandica, 67(3):337–353, August 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12010.

Davies:2013:AMC

[1996] Tilman M. Davies and Martin L. Hazelton. Assessing minimum contrast parameter estimation for spatial and spatiotemporal log-Gaussian Cox processes. Statistica Neerlandica, 67(4):355–389, November 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12011.

Brunner:2013:CPT

[1997] Edgar Brunner and Madan L. Puri. Comments on the paper ‘Type I error and test power of different tests for testing interaction effects in factorial experiments’ by M. Mendes and S. Yigit (Statistica Neerlandica, 2013, pp. 1–26). Statistica Neerlandica, 67(4):390–396, November 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12012. See [1976].

Shah:2013:CMY

[1998] Denis A. Shah and Laurence V. Madden. A comment on Mendeş and Yiğit (2013), ‘Type I error and test power of different tests for testing interaction effects in factorial experiments’, Statistica Neerlandica, 67:1–26. Statistica Neerlandica, 67(4):397–399, November 2013. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12013.

Konietschke:2013:CTE

[1999] Frank Konietschke, Xin Gao, and Arne C. Bathke. Comment on ‘Type I error and test power of different tests for testing interaction effects in factorial experiments’. Statistica Neerlandica, 67(4):400–402, November 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12014. See [1976].

Allen:2013:FDA

[2000] David E. Allen, Mohammad A. Ashraf, Michael McAleer, Robert J. Powell, and Abhay K. Singh. Financial dependence analysis: applications of vine copulas. Statistica Neerlandica, 67(4):403–435, November 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12015.

Radice:2013:EOE

[2001] Rosalba Radice, Luca Zanin, and Giampiero Marra. On the effect of obesity on employment in the presence of observed and unobserved confounding. Statistica Neerlandica, 67(4):436–455, November 2013. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12016.

Fiszeder:2013:NLV

[2002] Piotr Fiszeder and Grzegorz Perczak. A new look at variance estimation based on low, high and closing prices taking into account the drift. Statistica Neerlandica, 67(4):456–481, November 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12017.

vanGiersbergen:2013:BCS

[2003] Noud P. A. van Giersbergen. Bartlett correction in the stable secondorder autoregressive model with intercept and trend. Statistica Neerlandica, 67(4):482–498, November 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12018.

Anonymous:2013:ES

[2004] Anonymous. Editorial statistics. Statistica Neerlandica, 67(4):499, November 2013. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12020.

Segers:2014:PDE

[2005] Rene Segers and Philip Hans Franses. Panel design effects on response rates and response quality. Statistica Neerlandica, 68(1):1–24, February 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12019.

Huang:2014:LTR

[2006] Gang Huang, Michel Mandjes, and Peter Spreij. Limit theorems for reflected Ornstein–Uhlenbeck processes. Statistica Neerlandica, 68(1): 25–42, February 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12021.

Safarkhani:2014:ONA

[2007] Maryam Safarkhani, Shahab Jolani, and Mirjam Moerbeek. Optimal number of accrual groups and accrual group sizes in longitudinal trials with discrete-time survival endpoints. Statistica Neerlandica, 68(1):43– 60, February 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12022.

Vink:2014:PMM

[2008] Gerko Vink, Laurence E. Frank, Jeroen Pannekoek, and Stef van Buuren. Predictive mean matching imputation of semicontinuous variables. Statistica Neerlandica, 68(1):61–90, February 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12023.

Ramalho:2014:CLE

[2009] Esmeralda A. Ramalho and Joaquim J. S. Ramalho. Convenient links for the estimation of hedonic price indexes: the case of unique, infrequently traded assets. Statistica Neerlandica, 68(2):91–117, May 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12024.

Li:2014:IMR

[2010] Qing Li. Identifiability of mean-reverting measurement error with instrumental variable. Statistica Neerlandica, 68(2):118–129, May 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12025.

Tokola:2014:DCO

[2011] Kari Tokola, Andreas Lundell, Jaakko Nevalainen, and Hannu Oja. Design and cost optimization for hierarchical data. Statistica Neerlandica, 68(2):130–148, May 2014. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12026.

vanderMaas:2014:FIU

[2012] Jordi van der Maas. Forecasting inflation using time-varying Bayesian model averaging. Statistica Neerlandica, 68(3):149–182, August 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12027.

Tran:2014:APT

[2013] Lanh Tran, Ba Chu, Chunfeng Huang, and Kim P. Huynh. Adaptive permutation tests for serial independence. Statistica Neerlandica, 68(3): 183–208, August 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12028.

Wang:2014:QPG

[2014] Feixing Wang and Yingshuai Wang. Quantum prediction GJR model and its applications. Statistica Neerlandica, 68(3):209–224, August 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12029.

Sedivy:2014:IEI

[2015] Ondřej Šedivý and Antti Penttinen. Intensity estimation for inhomogeneous Gibbs point process with covariates-dependent chemical activity. Statistica Neerlandica, 68(3):225–249, August 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12030.

Kopczewska:2014:MSK

[2016] Katarzyna Kopczewska. L-moments skewness and kurtosis as measures of regional convergence and cohesion. Statistica Neerlandica, 68(4):251– 266, November 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12031.

Murakami:2014:SAD

[2017] Hidetoshi Murakami. A saddlepoint approximation to the distribution of the sum of independent non-identically uniform random variables. Statistica Neerlandica, 68(4):267–275, November 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12032.

Androulakis:2014:TPS

[2018] E. Androulakis, C. Koukouvinos, and F. Vonta. Tuning parameter selection in penalized generalized linear models for discrete data. Statistica Neerlandica, 68(4):276–292, November 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12033.

Tian:2014:NNR

[2019] Guo-Liang Tian. A new non-randomized response model: The parallel model. Statistica Neerlandica, 68(4):293–323, November 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12034.

Chien:2014:RID

[2020] Li-Chu Chien and Tsung-Shan Tsou. Robust influence diagnostics for generalized linear models with continuous responses. Statistica Neerlandica, 68(4):324–343, November 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12044.

Franses:2014:ES

[2021] Philip Hans Franses and Elli Hoek van Dijke. Editorial statistics. Statistica Neerlandica, 68(4):344, November 2014. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12050.

Anonymous:2015:IIa

[2022] Anonymous. Issue information. Statistica Neerlandica, 69(1):i–iii, February 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12040.

Tang:2015:VSC

[2023] Yanlin Tang, Xinyuan Song, and Zhongyi Zhu. Variable selection via composite quantile regression with dependent errors. Statistica Neerlandica, 69(1):1–20, February 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12035.

Balk:2015:MRA

[2024] Bert M. Balk. Measuring and relating aggregate and subaggregate total factor productivity change without neoclassical assumptions. Statistica Neerlandica, 69(1):21–48, February 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12045.

Comas:2015:SOA

[2025] C. Comas, F. J. Rodriguez-Cortes, and J. Mateu. Second-order analysis of anisotropic spatiotemporal point process data. Statistica Neerlandica, 69(1):49–66, February 2015. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12046.

Shan:2015:UTA

[2026] Guogen Shan and Gregory Wilding. Unconditional tests for association in 2 × 2 contingency tables in the total sum fixed design. Statistica Neerlandica, 69(1):67–83, February 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12047.

Anonymous:2015:IIb

[2027] Anonymous. Issue information. Statistica Neerlandica, 69(2):i–iii, May 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12041.

Battauz:2015:FAV

[2028] Michela Battauz. Factors affecting the variability of IRT equating coefficients. Statistica Neerlandica, 69(2):85–101, May 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12048.

Nadarajah:2015:SAD

[2029] Saralees Nadarajah, Xiao Jiang, and Jeffrey Chu. A saddlepoint approximation to the distribution of the sum of independent non-identically beta random variables. Statistica Neerlandica, 69(2):102–114, May 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12051.

Kalwij:2015:TTS

[2030] Adriaan Kalwij. Two tests for strict exogeneity in a correlated random effects panel data tobit model. Statistica Neerlandica, 69(2):115–125, May 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12052.

Lovison:2015:GBD

[2031] Gianfranco Lovison. A generalization of the binomial distribution based on the dependence ratio. Statistica Neerlandica, 69(2):126–149, May 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12053.

Abraham:2015:BRB

[2032] Christophe Abraham and Khader Khadraoui. Bayesian regression with B-splines under combinations of shape constraints and smoothness properties. Statistica Neerlandica, 69(2):150–170, May 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12054.

Longford:2015:IRM

[2033] Nicholas T. Longford. On the inefficiency of the restricted maximum likelihood. Statistica Neerlandica, 69(2):171–196, May 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12055.

Anonymous:2015:IIc

[2034] Anonymous. Issue information. Statistica Neerlandica, 69(3):i–iii, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12042.

Gasparini:2015:CRI

[2035] Mauro Gasparini, Mark Von Tress, and Guenter Heimann. Comparative randomized inverse sampling. Statistica Neerlandica, 69(3):197– 211, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12049.

Amin:2015:SPQ

[2036] Muhammad Amin, Lixin Song, Milton Abdul Thorlie, and Xiaoguang Wang. SCAD-penalized quantile regression for high-dimensional data analysis and variable selection. Statistica Neerlandica, 69(3):212–235, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12056.

Withers:2015:BRW

[2037] Christopher S. Withers and Saralees Nadarajah. Bias reduction when data are rounded. Statistica Neerlandica, 69(3):236–271, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12057.

Rahardja:2015:MLE

[2038] Dewi Rahardja and Ying Yang. Maximum likelihood estimation of a binomial proportion using one-sample misclassified binary data. Statistica Neerlandica, 69(3):272–280, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12058.

Weiss:2015:DOI

[2039] Christian H. Weiß and Sebastian Schweer. Detecting overdispersion in INARCH(1) processes. Statistica Neerlandica, 69(3):281–297, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12059.

Vallejos:2015:CMG

[2040] Ronny Vallejos, Felipe Osorio, and Diego Mancilla. The codispersion map: a graphical tool to visualize the association between two spatial variables. Statistica Neerlandica, 69(3):298–314, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12060.

Zhang:2015:GLS

[2041] Jin Zhang. Generalized least squares cross-validation in kernel density estimation. Statistica Neerlandica, 69(3):315–328, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12061.

Thorlie:2015:MFS

[2042] Milton Abdul Thorlie, Lixin Song, Muhammad Amin, and Xiaoguang Wang. Modeling and forecasting of stock index volatility with APARCH models under ordered restriction. Statistica Neerlandica, 69(3):329– 356, August 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12062.

Anonymous:2015:IId

[2043] Anonymous. Issue information. Statistica Neerlandica, 69(4):i–iii, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12043.

Cator:2015:E

[2044] Eric Cator. Editorial. Statistica Neerlandica, 69(4):357, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12081.

Wellek:2015:NES

[2045] Stefan Wellek. Nearly exact sample size calculation for powerful nonrandomized tests for differences between binomial proportions. Statistica Neerlandica, 69(4):358–373, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12063.

Lorenz:2015:DRM

[2046] E. Lorenz, C. Jenkner, W. Sauerbrei, and H. Becher. Dose–response modelling for bivariate covariates with and without a spike at zero: theory and application to binary outcomes. Statistica Neerlandica, 69(4):374– 398, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12064.

Sultana:2015:CTD

[2047] Papia Sultana, Jialiang Li, and Jianhua Hu. Comparison of threedimensional ROC surfaces for clustered and correlated markers, with a proteomics application. Statistica Neerlandica, 69(4):399–418, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12065.

Abegaz:2015:CGG

[2048] Fentaw Abegaz and Ernst Wit. Copula Gaussian graphical models with penalized ascent Monte Carlo EM algorithm. Statistica Neerlandica, 69 (4):419–441, November 2015. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12066.

Garcia:2015:CIS

[2049] V. J. García, M. Martel, and F. J. Vázquez-Polo. Complementary information for skewness measures. Statistica Neerlandica, 69(4):442–459, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12067.

Hyodo:2015:TBD

[2050] Masashi Hyodo, Nobumichi Shutoh, Takahiro Nishiyama, and Tatjana Pavlenko. Testing block-diagonal covariance structure for highdimensional data. Statistica Neerlandica, 69(4):460–482, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12068.

Slaoui:2015:PBS

[2051] Yousri Slaoui. Plug-in bandwidth selector for recursive kernel regression estimators defined by stochastic approximation method. Statistica Neerlandica, 69(4):483–509, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12069.

Vidoni:2015:EKL

[2052] Paolo Vidoni. Estimating the Kullback–Liebler risk based on multifold cross-validation. Statistica Neerlandica, 69(4):510–540, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12070.

Cator:2015:ES

[2053] Eric Cator and Elli Hoek van Dijke. Editorial statistics. Statistica Neerlandica, 69(4):541, November 2015. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12080.

Anonymous:2016:IIa

[2054] Anonymous. Issue information. Statistica Neerlandica, 70(1):1–3, February 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12074.

Zhu:2016:LIA

[2055] Fukang Zhu, Shuangzhe Liu, and Lei Shi. Local influence analysis for Poisson autoregression with an application to stock transaction data. Statistica Neerlandica, 70(1):4–25, February 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12071.

Knottnerus:2016:NVA

[2056] Paul Knottnerus. On new variance approximations for linear models with inequality constraints. Statistica Neerlandica, 70(1):26–46, February 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12072.

Melo:2016:BSL

[2057] Oscar O. Melo, Carlos E. Melo, and Jorge Mateu. Beta spatial linear mixed model with variable dispersion using Monte Carlo maximum likelihood. Statistica Neerlandica, 70(1):47–76, February 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12078.

Anonymous:2016:IIb

[2058] Anonymous. Issue information. Statistica Neerlandica, 70(2):77–79, May 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12075.

Lee:2016:CCR

[2059] Cheon-Sig Lee, Shu-Ching Su, Katrina Mondragon, Veronica I. Salinas, Monique L. Zamora, Stephen Andrew Sedory, and Sarjinder Singh. Comparison of Cramér–Rao lower bounds of variances for at least equal protection of respondents. Statistica Neerlandica, 70(2):80–99, May 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12073.

Karlis:2016:FMC

[2060] Dimitris Karlis, Purushottam Papatla, and Sudipt Roy. Finite mixtures of censored Poisson regression models. Statistica Neerlandica, 70(2): 100–122, May 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12079.

Maiti:2016:CFC

[2061] Raju Maiti, Atanu Biswas, and Samarjit Das. Coherent forecasting for count time series using Box–Jenkins’s AR(p) model. Statistica Neerlandica, 70(2):123–145, May 2016. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12083.

Safarkhani:2016:ODC

[2062] Maryam Safarkhani and Mirjam Moerbeek. D-optimal designs for a continuous predictor in longitudinal trials with discrete-time survival endpoints. Statistica Neerlandica, 70(2):146–171, May 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12085.

Anonymous:2016:IIc

[2063] Anonymous. Issue information. Statistica Neerlandica, 70(3):173–175, August 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12076.

Bourguignon:2016:PGI

[2064] Marcelo Bourguignon. Poisson–geometric INAR(1) process for modeling count time series with overdispersion. Statistica Neerlandica, 70(3):176– 192, August 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12082.

Withers:2016:CTA

[2065] Christopher S. Withers and Saralees Nadarajah. Cusums for tracking arbitrary functionals. Statistica Neerlandica, 70(3):193–228, August 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12084.

Cerasa:2016:CHG

[2066] Andrea Cerasa. Combining homogeneous groups of preclassified observations with application to international trade. Statistica Neerlandica, 70(3):229–259, August 2016. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12086.

Shahtahmassebi:2016:AGP

[2067] Golnaz Shahtahmassebi and Rana Moyeed. An application of the generalized Poisson difference distribution to the Bayesian modelling of football scores. Statistica Neerlandica, 70(3):260–273, August 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12087.

Anonymous:2016:IId

[2068] Anonymous. Issue information. Statistica Neerlandica, 70(4):275–277, November 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12077.

Yoshida:2016:ASP

[2069] Takuma Yoshida. Asymptotics and smoothing parameter selection for penalized spline regression with various loss functions. Statistica Neerlandica, 70(4):278–303, November 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12088.

Zhang:2016:SIR

[2070] Jun Zhang, Nanguang Zhou, Zipeng Sun, Gaorong Li, and Zhenghong Wei. Statistical inference on restricted partial linear regression models with partial distortion measurement errors. Statistica Neerlandica, 70 (4):304–331, November 2016. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12089.

Mahmood:2016:SBC

[2071] Munir Mahmood and Maxwell L. King. On solving bias-corrected nonlinear estimation equations with an application to the dynamic linear model. Statistica Neerlandica, 70(4):332–355, November 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12090.

Liu:2016:MSQ

[2072] Xiaochun Liu. Markov switching quantile autoregression. Statistica Neerlandica, 70(4):356–395, November 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12091.

Das:2016:SNC

[2073] Kalyan Das, Mohamad Elmasri, and Arusharka Sen. A skew-normal copula-driven GLMM. Statistica Neerlandica, 70(4):396–413, November 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12092.

Nguyen:2016:SCT

[2074] Hien D. Nguyen, Geoffrey J. McLachlan, Jeremy F. P. Ullmann, and Andrew L. Janke. Spatial clustering of time series via mixture of autoregressions models and Markov random fields. Statistica Neerlandica, 70(4):414–439, November 2016. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12093.

Cator:2016:ES

[2075] Eric Cator and Elli Hoek van Dijke. Editorial statistics. Statistica Neerlandica, 70(4):440, November 2016. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12100.

Anonymous:2017:IIa

[2076] Anonymous. Issue information. Statistica Neerlandica, 71(1):1–3, January 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12094.

Al-Momani:2017:PRE

[2077] Marwan Al-Momani, Abdulkadir A. Hussein, and S. E. Ahmed. Penalty and related estimation strategies in the spatial error model. Statistica Neerlandica, 71(1):4–30, January 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12098.

Lan:2017:NPR

[2078] Ling Lan, Dipankar Bandyopadhyay, and Somnath Datta. Nonparametric regression in clustered multistate current status data with informative cluster size. Statistica Neerlandica, 71(1):31–57, January 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12099.

Lopuhaa:2017:SEM

[2079] Hendrik P. Lopuhaä and Eni Musta. Smooth estimation of a monotone hazard and a monotone density under random censoring. Statistica Neerlandica, 71(1):58–82, January 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12101.

Anonymous:2017:IIb

[2080] Anonymous. Issue information. Statistica Neerlandica, 71(2):83–85, May 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12095.

Liu:2017:DAV

[2081] Yonghui Liu, Ruochen Sang, and Shuangzhe Liu. Diagnostic analysis for a vector autoregressive model under Student’s t-distributions. Statistica Neerlandica, 71(2):86–114, May 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12102.

Comte:2017:SRH

[2082] Fabienne Comte, Gwennaelle Mabon, and Adeline Samson. Spline regression for hazard rate estimation when data are censored and measured with error. Statistica Neerlandica, 71(2):115–140, May 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12103.

Cereda:2017:BAL

[2083] Giulia Cereda. Bayesian approach to LR assessment in case of rare type match. Statistica Neerlandica, 71(2):141–164, May 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12104.

Anonymous:2017:IIc

[2084] Anonymous. Issue information. Statistica Neerlandica, 71(3):165–167, August 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12096.

Andrade:2017:BRM

[2085] J. A. A. Andrade and Edward Omey. Bayesian robustness modelling using the O-regularly varying distributions. Statistica Neerlandica, 71(3): 168–183, August 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12105.

Shutoh:2017:BCL

[2086] Nobumichi Shutoh, Takahiro Nishiyama, and Masashi Hyodo. Bartlett correction to the likelihood ratio test for MCAR with two-step monotone sample. Statistica Neerlandica, 71(3):184–199, August 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12106.

Medeiros:2017:SST

[2087] Francisco M. C. Medeiros and Silvia L. P. Ferrari. Small-sample testing inference in symmetric and log-symmetric linear regression models. Statistica Neerlandica, 71(3):200–224, August 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12107.

Mandjes:2017:NCL

[2088] Michel Mandjes and Peter Spreij. A note on the central limit theorem for the idleness process in a one-sided reflected Ornstein–Uhlenbeck model. Statistica Neerlandica, 71(3):225–235, August 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12108.

Anonymous:2017:IId

[2089] Anonymous. Issue information. Statistica Neerlandica, 71(4):237–239, November 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12097.

Kabaila:2017:CAI

[2090] Paul Kabaila, Rheanna Mainzer, and Davide Farchione. Conditional assessment of the impact of a Hausman pretest on confidence intervals. Statistica Neerlandica, 71(4):240–262, November 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12109.

Chatterjee:2017:TPT

[2091] Suryasish Chatterjee and Uttam Bandyopadhyay. Three-period, twotreatment crossover design under long-term carryover effect. Statistica Neerlandica, 71(4):263–285, November 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12110.

Chatterjee:2017:USM

[2092] Moutushi Chatterjee and Ashis Kumar Chakraborty. Unification of some multivariate process capability indices for asymmetric specification region. Statistica Neerlandica, 71(4):286–306, November 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12112.

Cator:2017:ES

[2093] Eric Cator and Elli Hoek van Dijke. Editorial statistics. Statistica Neerlandica, 71(4):307, November 2017. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12122.

Anonymous:2018:IIa

[2094] Anonymous. Issue information. Statistica Neerlandica, 72(1):1–3, February 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12117.

Ly:2018:APP

[2095] Alexander Ly, Maarten Marsman, and Eric-Jan Wagenmakers. Analytic posteriors for Pearson’s correlation coefficient. Statistica Neerlandica, 72(1):4–13, February 2018. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12111.

Hanea:2018:ADD

[2096] A. M. Hanea and G. F. Nane. The asymptotic distribution of the determinant of a random correlation matrix. Statistica Neerlandica, 72(1): 14–33, February 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12113.

Musoro:2018:DPM

[2097] Jammbe Z. Musoro, Aeilko H. Zwinderman, Ameen Abu-Hanna, Rob Bosman, and Ronald B. Geskus. Dynamic prediction of mortality among patients in intensive care using the sequential organ failure assessment (SOFA) score: a joint competing risk survival and longitudinal modeling approach. Statistica Neerlandica, 72(1):34–47, February 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12114.

Wu:2018:DCM

[2098] Sheng-Jhih Wu, Sujit K. Ghosh, Yu-Cheng Ku, and Peter Bloomfield. Dynamic correlation multivariate stochastic volatility with latent factors. Statistica Neerlandica, 72(1):48–69, February 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12115.

Anonymous:2018:IIb

[2099] Anonymous. Issue information. Statistica Neerlandica, 72(2):71–73, May 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12118.

MarquesdaSilva:2018:FIU

[2100] Antonio Hermes Marques da Silva Júnior, Jochen Einbeck, and Peter S. Craig. Fisher information under Gaussian quadrature models. Statistica Neerlandica, 72(2):74–89, May 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12116.

Zheng:2018:STS

[2101] Shengchao Zheng and Degao Li. Semiparametric time series regression modeling with a diverging number of parameters. Statistica Neerlandica, 72(2):90–108, May 2018. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12121.

Muro:2018:CBP

[2102] Alan Muro and Sara van de Geer. Concentration behavior of the penalized least squares estimator. Statistica Neerlandica, 72(2):109–125, May 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12123.

Dattner:2018:AOS

[2103] Itai Dattner and Shota Gugushvili. Application of one-step method to parameter estimation in ODE models. Statistica Neerlandica, 72(2): 126–156, May 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12124.

Zhang:2018:CED

[2104] Qiang Zhang and Ping Chen. Credibility estimators with dependence structure over risks and time under balanced loss function. Statistica Neerlandica, 72(2):157–173, May 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12125.

Anonymous:2018:IIc

[2105] Anonymous. Issue information. Statistica Neerlandica, 72(3):175–177, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12119.

Grzegorczyk:2018:EI

[2106] Marco Grzegorczyk. Editorial introduction. Statistica Neerlandica, 72 (3):178, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12151.

Langrock:2018:SBN

[2107] Roland Langrock, Timo Adam, Vianey Leos-Barajas, Sina Mews, David L. Miller, and Yannis P. Papastamatiou. Spline-based nonparametric inference in general state-switching models. Statistica Neerlandica, 72(3):179–200, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12133.

Eilers:2018:TAE

[2108] Paul H. C. Eilers. The truth about the effective dimension. Statistica Neerlandica, 72(3):201–209, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12131.

Agresti:2018:SWI

[2109] Alan Agresti and Claudia Tarantola. Simple ways to interpret effects in modeling ordinal categorical data. Statistica Neerlandica, 72(3):210– 223, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12130.

Simone:2018:MUR

[2110] Rosaria Simone and Gerhard Tutz. Modelling uncertainty and response styles in ordinal data. Statistica Neerlandica, 72(3):224–245, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12129.

Stefanucci:2018:PBD

[2111] Marco Stefanucci, Laura M. Sangalli, and Pierpaolo Brutti. PCA-based discrimination of partially observed functional data, with an application to AneuRisk65 data set. Statistica Neerlandica, 72(3):246–264, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12137.

Giurghita:2018:SMC

[2112] Diana Giurghita and Dirk Husmeier. Statistical modelling of cell movement. Statistica Neerlandica, 72(3):265–280, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12140.

Kamalabad:2018:IND

[2113] Mahdi Shafiee Kamalabad and Marco Grzegorczyk. Improving nonhomogeneous dynamic Bayesian networks with sequentially coupled parameters. Statistica Neerlandica, 72(3):281–305, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12136.

Paun:2018:MMI

[2114] L. Mihaela Păun, M. Umar Qureshi, Mitchel Colebank, Nicholas A. Hill, Mette S. Olufsen, Mansoor A. Haider, and Dirk Husmeier. MCMC methods for inference in a mathematical model of pulmonary circulation. Statistica Neerlandica, 72(3):306–338, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12132.

Sering:2018:LCM

[2115] Konstantin Sering, Petar Milin, and R. Harald Baayen. Language comprehension as a multi-label classification problem. Statistica Neerlandica, 72(3):339–353, August 2018. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12134.

deMelo:2018:FAC

[2116] Mariana Arozo B. de Melo, Cristiano A. C. Fernandes, and Eduardo F. L. de Melo. Forecasting aggregate claims using score-driven time series models. Statistica Neerlandica, 72(3):354–374, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12139.

Stolfi:2018:SMS

[2117] Paola Stolfi, Mauro Bernardi, and Lea Petrella. The sparse method of simulated quantiles: an application to portfolio optimization. Statistica Neerlandica, 72(3):375–398, August 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12141.

Anonymous:2018:IId

[2118] Anonymous. Issue information. Statistica Neerlandica, 72(4):399–401, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12120.

Chen:2018:BTD

[2119] Baoline Chen, Tommaso Di Fonzo, and Nino Mushkudiani. Benchmarking, temporal disaggregation, and reconciliation of systems of time series. Statistica Neerlandica, 72(4):402–405, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12157.

Daalmans:2018:SBS

[2120] Jacco Daalmans. On the sequential benchmarking of subannual series to annual totals. Statistica Neerlandica, 72(4):406–420, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12163.

Temursho:2018:EBB

[2121] Umed Temursho. Entropy-based benchmarking methods. Statistica Neerlandica, 72(4):421–446, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12160.

Quilis:2018:TDE

[2122] Enrique M. Quilis. Temporal disaggregation of economic time series: The view from the trenches. Statistica Neerlandica, 72(4):447–470, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12150.

Bisio:2018:TDD

[2123] Laura Bisio and Filippo Moauro. Temporal disaggregation by dynamic regressions: Recent developments in Italian quarterly national accounts. Statistica Neerlandica, 72(4):471–494, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12156.

Guerrero:2018:RSR

[2124] Víctor M. Guerrero and Francisco Corona. Retropolating some relevant series of Mexico’s System of National Accounts at constant prices: The case of Mexico City’s GDP. Statistica Neerlandica, 72(4):495–519, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12162.

Brent:2018:MLE

[2125] Geoffrey Brent. Maximum likelihood estimation framework for tablebalancing adjustments. Statistica Neerlandica, 72(4):520–532, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12159.

Chen:2018:SRT

[2126] Baoline Chen, Tommaso Di Fonzo, Thomas Howells, and Marco Marini. The statistical reconciliation of time series of accounts between two benchmark revisions. Statistica Neerlandica, 72(4):533–552, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12154.

Mushkudiani:2018:SLD

[2127] Nino Mushkudiani, Jacco Daalmans, and Reinier Bikker. Solving large-data consistency problems at Statistics Netherlands using macrointegration techniques. Statistica Neerlandica, 72(4):553–573, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12158.

McElroy:2018:SAS

[2128] Tucker McElroy. Seasonal adjustment subject to accounting constraints. Statistica Neerlandica, 72(4):574–589, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12161.

Elezovic:2018:RSA

[2129] Suad Elezović and Yingfu Xie. Reconciliation of seasonally adjusted data with applications to the Swedish quarterly national accounts. Statistica Neerlandica, 72(4):590–602, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12155.

Cator:2018:ES

[2130] Eric Cator. Editorial statistics. Statistica Neerlandica, 72(4):603, November 2018. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12165.

Anonymous:2019:IIa

[2131] Anonymous. Issue information. Statistica Neerlandica, 73(1):1–3, February 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12146.

Young:2019:ACT

[2132] D. S. Young, M. Naghizadeh Qomi, and A. Kiapour. Approximate confidence and tolerance limits for the discrete Pareto distribution for characterizing extremes in count data. Statistica Neerlandica, 73(1):4– 21, February 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12126.

Perez-Sanchez:2019:BAL

[2133] José María Pérez-Sánchez, Román Salmerón-Gómez, and Francisco M. Ocaña-Peinado. A Bayesian asymmetric logistic model of factors underlying team success in top-level basketball in Spain. Statistica Neerlandica, 73(1):22–43, February 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12127.

Koopman:2019:MEI

[2134] Siem Jan Koopman, Rutger Lit, and Thuy Minh Nguyen. Modified efficient importance sampling for partially non-Gaussian state space models. Statistica Neerlandica, 73(1):44–62, February 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12128.

Thiam:2019:REP

[2135] Baba Thiam. Relative error prediction in nonparametric deconvolution regression model. Statistica Neerlandica, 73(1):63–77, February 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12135.

Groeneboom:2019:EMS

[2136] Piet Groeneboom and Kim Hendrickx. Estimation in monotone singleindex models. Statistica Neerlandica, 73(1):78–99, February 2019. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12138.

Belomestny:2019:SIM

[2137] Denis Belomestny, Tatiana Orlova, and Vladimir Panov. Statistical inference for moving-average Lévy-driven processes: Fourier-based approach. Statistica Neerlandica, 73(1):100–117, February 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12142.

Li:2019:TSV

[2138] Han Li, Kai Yang, and Dehui Wang. A threshold stochastic volatility model with explanatory variables. Statistica Neerlandica, 73(1):118– 138, February 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12143.

Vendrig:2019:RUV

[2139] Nadia J. Vendrig, Lia Hemerik, Ilona J. Pinter, and Cajo J. F. ter Braak. Relating ultrasonic vocalizations from a pair of rats to individual behavior: a composite link model approach. Statistica Neerlandica, 73(1):139– 156, February 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12144.

Anonymous:2019:IIb

[2140] Anonymous. Issue information. Statistica Neerlandica, 73(2):157– 159, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12147.

Andreis:2019:RWR

[2141] F. Andreis, P. L. Conti, and F. Mecatti. On the role of weights rounding in applications of resampling based on pseudopopulations. Statistica Neerlandica, 73(2):160–175, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12145.

Gaunt:2019:NDP

[2142] Robert E. Gaunt. A note on the distribution of the product of zeromean correlated normal random variables. Statistica Neerlandica, 73(2): 176–179, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12152.

Lopuhaa:2019:CLT

[2143] Hendrik P. Lopuhaä and Eni Musta. A central limit theorem for the Hellinger loss of Grenander-type estimators. Statistica Neerlandica, 73 (2):180–196, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12153.

Fang:2019:ORS

[2144] Longxiang Fang and Tingtao Xu. Ordering results of the smallest and largest order statistics from independent heterogeneous exponentiated gamma random variables. Statistica Neerlandica, 73(2):197–210, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12164.

Barreto-Souza:2019:BID

[2145] Wagner Barreto-Souza and Rodrigo B. Silva. A bivariate infinitely divisible law for modeling the magnitude and duration of monotone periods of log-returns. Statistica Neerlandica, 73(2):211–233, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12166.

Qu:2019:CAM

[2146] Lianqiang Qu and Liuquan Sun. The Cox–Aalen model for recurrentevent data with a dependent terminal event. Statistica Neerlandica, 73 (2):234–255, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12167.

Ayyad:2019:TNG

[2147] C. Ayyad, J. Mateu, M. Omidi, I. Tamayo-Uria, and M. Mohammadzadeh. Trivariate non-Gaussian copulas to analyze the spatial behavior of rat sightings. Statistica Neerlandica, 73(2):256–273, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12168.

Meester:2019:NLC

[2148] Ronald Meester and Timber Kerkvliet. A new look at conditional probability with belief functions. Statistica Neerlandica, 73(2):274–291, May 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12169.

Zhang:2019:NRM

[2149] Jun Zhang, Yujie Gai, Bingqing Lin, and Xuehu Zhu. Nonlinear regression models with single-index heteroscedasticity. Statistica Neerlandica, 73(2):292–316, May 2019. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12170.

Anonymous:2019:IIc

[2150] Anonymous. Issue information. Statistica Neerlandica, 73(3):317–319, August 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12148.

Nogales:2019:NRB

[2151] A. G. Nogales and P. Pérez. A note on the relationship between conditional and unconditional independence and its extensions for Markov kernels. Statistica Neerlandica, 73(3):320–332, August 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12171.

Chipperfield:2019:WAM

[2152] James Chipperfield. A weighting approach to making inference with probabilistically linked data. Statistica Neerlandica, 73(3):333–350, August 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12172.

Marsman:2019:BEE

[2153] Maarten Marsman, Lourens Waldorp, Fabian Dablander, and Eric-Jan Wagenmakers. Bayesian estimation of explained variance in ANOVA designs. Statistica Neerlandica, 73(3):351–372, August 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12173.

Sun:2019:PSB

[2154] Ruili Sun, Tiefeng Ma, and Shuangzhe Liu. Portfolio selection based on semivariance and distance correlation under minimum variance framework. Statistica Neerlandica, 73(3):373–394, August 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12174.

Mao:2019:NIE

[2155] Lu Mao. Nonparametric identification and estimation of current status data in the presence of death. Statistica Neerlandica, 73(3):395– 413, August 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12175.

Schmitt:2019:CFP

[2156] Fernando O. Schmitt, Gustavo L. Gilardoni, and J. A. A. Andrade. A class of flat prior distributions for the Poisson-gamma hierarchical model. Statistica Neerlandica, 73(3):414–433, August 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12176.

Famoye:2019:BEE

[2157] Felix Famoye. Bivariate exponentiated-exponential geometric regression model. Statistica Neerlandica, 73(3):434–450, August 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12177.

Anonymous:2019:IId

[2158] Anonymous. Issue information. Statistica Neerlandica, 73(4):451–453, November 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12149.

Mainassara:2019:DRA

[2159] Yacouba Boubacar Maïnassara and Abdoulkarim Ilmi Amir. Distribution of residual autocorrelations for multiplicative seasonal ARMA models with uncorrelated but nonindependent error terms. Statistica Neerlandica, 73(4):454–474, November 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12178.

Malov:2019:NEC

[2160] Sergey V. Malov. Nonparametric estimation for a current status rightcensored data model. Statistica Neerlandica, 73(4):475–495, November 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12180.

Bansal:2019:SCM

[2161] Shilpa Bansal and Nitin Gupta. Stochastic comparisons of multivariate reversed frailty model. Statistica Neerlandica, 73(4):496–513, November 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12185.

Lu:2019:INA

[2162] Jiannan Lu. Improved Neymanian analysis for 2^{K} factorial designs with binary outcomes. Statistica Neerlandica, 73(4):514–523, November 2019.

CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12186.

Cator:2019:ES

[2163] Eric Cator. Editorial statistics. Statistica Neerlandica, 73(4):524, November 2019. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12194.

Anonymous:2020:IIa

[2164] Anonymous. Issue information. Statistica Neerlandica, 74(1):1–3, February 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12181.

vanGeloven:2020:E

[2165] Nan van Geloven and Maja Pohar Perme. Editorial. Statistica Neerlandica, 74(1):4, February 2020. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12195.

Gasparini:2020:MEM

[2166] Alessandro Gasparini, Keith R. Abrams, Jessica K. Barrett, Rupert W. Major, Michael J. Sweeting, Nigel J. Brunskill, and Michael J. Crowther. Mixed-effects models for health care longitudinal data with an informative visiting process: a Monte Carlo simulation study. Statistica Neerlandica, 74(1):5–23, February 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12188.

Weber:2020:ALH

[2167] Susanne Weber and Martin Wolkewitz. Accounting for length of hospital stay in regression models in clinical epidemiology. Statistica Neerlandica, 74(1):24–37, February 2020. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12187.

Pajouheshnia:2020:ATD

[2168] Romin Pajouheshnia, Noah A. Schuster, Rolf H. H. Groenwold, Frans H. Rutten, Karel G. M. Moons, and Linda M. Peelen. Accounting for timedependent treatment use when developing a prognostic model from observational data: a review of methods. Statistica Neerlandica, 74(1): 38–51, February 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12193.

Park:2020:PLA

[2169] Eunyoung Park, Sookhee Kwon, Jihoon Kwon, Richard Sylvester, and Il Do Ha. Penalized h-likelihood approach for variable selection in AFT random-effect models. Statistica Neerlandica, 74(1):52–71, February 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12179.

Ochoa:2020:PMC

[2170] Lizbeth Burgos Ochoa, Judith J. M. Rijnhart, Brenda W. Penninx, Klaas J. Wardenaar, Jos W. R. Twisk, and Martijn W. Heymans. Performance of methods to conduct mediation analysis with time-to-event outcomes. Statistica Neerlandica, 74(1):72–91, February 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12191.

Anonymous:2020:IIb

[2171] Anonymous. Issue information. Statistica Neerlandica, 74(2):93–95, May 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12182.

Chipperfield:2020:BIU

[2172] James Chipperfield. Bootstrap inference using estimating equations and data that are linked with complex probabilistic algorithms. Statistica Neerlandica, 74(2):96–111, May 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12189.

Han:2020:ETT

[2173] David Han. Expected termination times of progressively Type-I censored step-stress accelerated life tests under continuous and interval inspections. Statistica Neerlandica, 74(2):112–124, May 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12190.

Ranciati:2020:BSM

[2174] Saverio Ranciati, Ernst C. Wit, and Cinzia Viroli. Bayesian smooth-and-match inference for ordinary differential equations models linear in the parameters. Statistica Neerlandica, 74(2):125–144, May 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12192.

Ciginas:2020:LSD

[2175] Andrius Čiginas. Linking of sample data to small areas. Statistica Neerlandica, 74(2):145–158, May 2020. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12199.

Blagus:2020:MSE

[2176] Rok Blagus and Jelle J. Goeman. Mean squared error of ridge estimators in logistic regression. Statistica Neerlandica, 74(2):159–191, May 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12201.

Sedory:2020:UQM

[2177] Stephen A. Sedory, Sarjinder Singh, Oluwaseun L. Olanipekun, and Colin Wark. Unrelated question model with two decks of cards. Statistica Neerlandica, 74(2):192–215, May 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12202.

Anonymous:2020:IIc

[2178] Anonymous. Issue information. Statistica Neerlandica, 74(3):217–219, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12183.

Vinciotti:2020:SNS

[2179] Veronica Vinciotti and Ernst C. Wit. Statistica Neerlandica special issue on Statistical Network Science. Statistica Neerlandica, 74(3):220– 221, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12212.

Britton:2020:EMS

[2180] Tom Britton. Epidemic models on social networks — with inference. Statistica Neerlandica, 74(3):222–241, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12203.

Hansson:2020:EIA

[2181] Disa Hansson and Susanne Strömdahl. Estimating individual action dispositions using binary and frequency egocentric sexual network data. Statistica Neerlandica, 74(3):242–260, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12207.

Richter:2020:IGC

[2182] Francisco Richter, Bart Haegeman, Rampal S. Etienne, and Ernst C. Wit. Introducing a general class of species diversification models for phylogenetic trees. Statistica Neerlandica, 74(3):261–274, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12205.

Fritz:2020:TVH

[2183] Cornelius Fritz, Michael Lebacher, and Göran Kauermann. Tempus volat, hora fugit: a survey of tie-oriented dynamic network models in discrete and continuous time. Statistica Neerlandica, 74(3):275–299, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12198.

Snijders:2020:NDN

[2184] Tom A. B. Snijders, Malick Faye, and Julien Brailly. Network dynamics with a nested node set: Sociability in seven villages in Senegal. Statistica Neerlandica, 74(3):300–323, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12208.

DAngelo:2020:MNH

[2185] Silvia D’Angelo, Marco Alfò, and Thomas Brendan Murphy. Modeling node heterogeneity in latent space models for multidimensional networks. Statistica Neerlandica, 74(3):324–341, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12209.

Schweinberger:2020:SIC

[2186] Michael Schweinberger. Statistical inference for continuous-time Markov processes with block structure based on discrete-time network data. Statistica Neerlandica, 74(3):342–362, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12196.

Pavlovic:2020:MSS

[2187] Dragana M. Pavlović, Bryan R. L. Guillaume, Soroosh Afyouni, and Thomas E. Nichols. Multi-subject stochastic blockmodels with mixed effects for adaptive analysis of individual differences in human brain network cluster structure. Statistica Neerlandica, 74(3):363–396, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12219.

Scutari:2020:BNM

[2188] Marco Scutari. Bayesian network models for incomplete and dynamic data. Statistica Neerlandica, 74(3):397–419, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12197.

Petrakis:2020:OMG

[2189] Nikolaos Petrakis, Stefano Peluso, Dimitris Fouskakis, and Guido Consonni. Objective methods for graphical structural learning. Statistica Neerlandica, 74(3):420–438, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12211.

Sadeghi:2020:HMI

[2190] Kayvan Sadeghi and Alessandro Rinaldo. Hierarchical models for independence structures of networks. Statistica Neerlandica, 74(3):439– 457, August 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12200.

Anonymous:2020:IId

[2191] Anonymous. Issue information. Statistica Neerlandica, 74(4):459–461, November 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12184.

Zhang:2020:LCM

[2192] Jun Zhang, Yiping Yang, and Gaorong Li. Logarithmic calibration for multiplicative distortion measurement errors regression models. Statistica Neerlandica, 74(4):462–488, November 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12204.

Nik:2020:CPF

[2193] Simon Nik and Christian H. Weiß. CLAR(1) point forecasting under estimation uncertainty. Statistica Neerlandica, 74(4):489–516, November 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12206.

Zhang:2020:NBT

[2194] Qingchun Zhang, Dehui Wang, and Xiaodong Fan. A negative binomial thinning-based bivariate INAR(1) process. Statistica Neerlandica, 74 (4):517–537, November 2020. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12210.

Pinheiro:2020:NAA

[2195] Hildete P. Pinheiro, Pranab K. Sen, Aluísio Pinheiro, and Samara F. Kiihl. A nonparametric approach to assess undergraduate performance. Statistica Neerlandica, 74(4):538–558, November 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12217.

Durot:2020:EGT

[2196] Cécile Durot and Eni Musta. On the L_{p} error of the Grenander-type estimator in the Cox model. Statistica Neerlandica, 74(4):559–591, November 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12218.

Hossain:2020:SEE

[2197] Shakhawat Hossain and Shahedul A. Khan. Shrinkage estimation of the exponentiated Weibull regression model for time-to-event data. Statistica Neerlandica, 74(4):592–610, November 2020. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12220.

Bourguignon:2020:RIG

[2198] Marcelo Bourguignon and Diego I. Gallardo. Reparameterized inverse gamma regression models with varying precision. Statistica Neerlandica, 74(4):611–627, November 2020. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12221.

Anonymous:2021:IIa

[2199] Anonymous. Issue information. Statistica Neerlandica, 75(1):1–3, February 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12213.

Kabaila:2021:EDW

[2200] Paul Kabaila, Davide Farchione, Samer Alhelli, and Nathan Bragg. The effect of a Durbin–Watson pretest on confidence intervals in regression. Statistica Neerlandica, 75(1):4–23, February 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12222.

Bellio:2021:MLE

[2201] Ruggero Bellio and Nicola Soriani. Maximum likelihood estimation based on the Laplace approximation for p_{2} network regression models. Statistica Neerlandica, 75(1):24–41, February 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12223.

Rachdi:2021:NNL

[2202] Mustapha Rachdi, Ali Laksaci, Zoulikha Kaid, Abbassia Benchiha, and Fahimah A. Al-Awadhi. k-nearest neighbors local linear regression for functional and missing data at random. Statistica Neerlandica, 75(1): 42–65, February 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12224.

Das:2021:SIV

[2203] Ujjwal Das and Kalyan Das. Selection of influential variables in ordinal data with preponderance of zeros. Statistica Neerlandica, 75(1): 66–87, February 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12225.

Ristic:2021:ESa

[2204] Miroslav Ristic, Marijtje van Duijn, and Nan van Geloven. Editorial statistics. Statistica Neerlandica, 75(1):88, February 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12230.

Anonymous:2021:IIb

[2205] Anonymous. Issue information. Statistica Neerlandica, 75(2):89–91, May 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12214.

Biffignandi:2021:LBD

[2206] Silvia Biffignandi and Alessandro Zeli. Longitudinal business data construction and quality: Two different approaches. Statistica Neerlandica, 75(2):92–114, May 2021. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12228.

Koley:2021:APT

[2207] Arnab Koley and Debasis Kundu. Analysis of progressive Type-II censoring in presence of competing risk data under step stress modeling. Statistica Neerlandica, 75(2):115–136, May 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12226.

Rocha:2021:RLI

[2208] Suelena S. Rocha, Patrícia L. Espinheira, and Francisco Cribari-Neto. Residual and local influence analyses for unit gamma regressions. Statistica Neerlandica, 75(2):137–160, May 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12229.

Groeneboom:2021:EIT

[2209] Piet Groeneboom. Estimation of the incubation time distribution for COVID-19. Statistica Neerlandica, 75(2):161–179, May 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12231.

Guha:2021:NTS

[2210] Apratim Guha, Atanu Biswas, and Abhik Ghosh. A nonparametric twosample test using a general ϕ-divergence-based mutual information. Statistica Neerlandica, 75(2):180–202, May 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12232.

Laumen:2021:SSL

[2211] Benjamin Laumen and Erhard Cramer. k-step stage life testing. Statistica Neerlandica, 75(2):203–233, May 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12227.

Anonymous:2021:IIc

[2212] Anonymous. Issue information. Statistica Neerlandica, 75(3):235–237, August 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12215.

Arab:2021:CTO

[2213] Idir Arab, Paulo Eduardo Oliveira, and Tilo Wiklund. Convex transform order of beta distributions with some consequences. Statistica Neerlandica, 75(3):238–256, August 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12233.

Sadoun:2021:LAE

[2214] Mohamed Sadoun and Mohamed Bentarzi. Locally asymptotically efficient estimation for parametric PINAR(p) models. Statistica Neerlandica, 75(3):257–289, August 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12234.

Wiley:2021:ILR

[2215] Briceön Wiley, Chris Elrod, Phil D. Young, and Dean M. Young. An integrated-likelihood-ratio confidence interval for a proportion based on underreported and infallible data. Statistica Neerlandica, 75(3):290– 298, August 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12235.

Cancho:2021:BSM

[2216] Vicente G. Cancho, Gladys D. C. Barriga, Gauss M. Cordeiro, Edwin M. M. Ortega, and Adriano K. Suzuki. Bayesian survival model induced by frailty for lifetime with long-term survivors. Statistica Neerlandica, 75(3):299–323, August 2021. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12236.

Pal:2021:EDC

[2217] Suvra Pal and Souvik Roy. On the estimation of destructive cure rate model: a new study with exponentially weighted Poisson competing risks. Statistica Neerlandica, 75(3):324–342, August 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12237.

Neumann:2021:BIV

[2218] Michael H. Neumann. Bootstrap for integer-valued GARCH(p, q) processes. Statistica Neerlandica, 75(3):343–363, August 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12238.

Zhang:2021:MCM

[2219] Jun Zhang. Model checking for multiplicative linear regression models with mixed estimators. Statistica Neerlandica, 75(3):364–403, August 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12239.

Anonymous:2021:IId

[2220] Anonymous. Issue information. Statistica Neerlandica, 75(4):405–407, November 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12216.

Balabdaoui:2021:PLS

[2221] Fadoua Balabdaoui, Cécile Durot, and Christopher Fragneau. On the population least-squares criterion in the monotone single index model. Statistica Neerlandica, 75(4):408–436, November 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12240.

vanZwet:2021:SFW

[2222] Erik W. van Zwet and Eric A. Cator. The significance filter, the winner’s curse and the need to shrink. Statistica Neerlandica, 75(4):437– 452, November 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12241.

Bayer:2021:IKR

[2223] Fábio M. Bayer, Francisco Cribari-Neto, and Jéssica Santos. Inflated Kumaraswamy regressions with application to water supply and sanitation in Brazil. Statistica Neerlandica, 75(4):453–481, November 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12242.

Lee:2021:RAR

[2224] Woojoo Lee, Hans-Peter Piepho, and Youngjo Lee. Resolving the ambiguity of random-effects models with singular precision matrix. Statistica Neerlandica, 75(4):482–499, November 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12244.

Atkinson:2021:IAR

[2225] Andrew Atkinson, Suzie Cro, James R. Carpenter, and Michael G. Kenward. Information anchored reference-based sensitivity analysis for truncated normal data with application to survival analysis. Statistica Neerlandica, 75(4):500–523, November 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12250.

Ristic:2021:ESb

[2226] Miroslav Ristic, Marijtje van Duijn, and Nan van Geloven. Editorial statistics. Statistica Neerlandica, 75(4):524, November 2021. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12257.

Anonymous:2022:IIa

[2227] Anonymous. Issue information. Statistica Neerlandica, 76(1):1–3, February 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12245.

Chattopadhyay:2022:CPA

[2228] Subhankar Chattopadhyay, Raju Maiti, Samarjit Das, and Atanu Biswas. Change-point analysis through integer-valued autoregressive process with application to some COVID-19 data. Statistica Neerlandica, 76(1):4– 34, February 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12251.

Aleksandrov:2022:GFT

[2229] Boris Aleksandrov, Christian H. Weiß, and Carsten Jentsch. Goodnessof-fit tests for Poisson count time series based on the Stein–Chen identity. Statistica Neerlandica, 76(1):35–64, February 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12252.

Krause:2022:RPD

[2230] Joscha Krause, Jan Pablo Burgard, and Domingo Morales. Robust prediction of domain compositions from uncertain data using isometric logratio transformations in a penalized multivariate Fay–Herriot model. Statistica Neerlandica, 76(1):65–96, February 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12253.

Briz-Redon:2022:ICG

[2231] Álvaro Briz-Redón, Jorge Mateu, and Francisco Montes. Identifying crime generators and spatially overlapping high-risk areas through a nonlinear model: a comparison between three cities of the Valencian region (Spain). Statistica Neerlandica, 76(1):97–120, February 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12254.

Anonymous:2022:IIb

[2232] Anonymous. Issue information. Statistica Neerlandica, 76(2):121– 123, May 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12246.

Saulo:2022:LSQ

[2233] Helton Saulo, Alan Dasilva, Víctor Leiva, Luis Sánchez, and Hanns de la Fuente-Mella. Log-symmetric quantile regression models. Statistica Neerlandica, 76(2):124–163, May 2022. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12243.

Orsi:2022:CNB

[2234] Carlo Orsi. On the conditional noncentral beta distribution. Statistica Neerlandica, 76(2):164–189, May 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12249.

Sathish:2022:AMA

[2235] Vurukonda Sathish, Siuli Mukhopadhyay, and Rashmi Tiwari. Autoregressive and moving average models for zero-inflated count time series. Statistica Neerlandica, 76(2):190–218, May 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12255.

Ariyo:2022:BMS

[2236] Oludare Ariyo, Emmanuel Lesaffre, Geert Verbeke, Martijn Huisman, Martijn Heymans, and Jos Twisk. Bayesian model selection for multilevel mediation models. Statistica Neerlandica, 76(2):219–235, May 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12256.

Nogales:2022:BED

[2237] Agustín G. Nogales. On Bayesian estimation of densities and sampling distributions: The posterior predictive distribution as the Bayes estimator. Statistica Neerlandica, 76(2):236–250, May 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12258.

Anonymous:2022:IIc

[2238] Anonymous. Issue information. Statistica Neerlandica, 76(3):251–253, August 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12247.

Zhang:2022:TCT

[2239] Shibin Zhang and Xin M. Tu. Tests for comparing time-invariant and time-varying spectra based on the Anderson–Darling statistic. Statistica Neerlandica, 76(3):254–282, August 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12259.

Menssen:2022:PIA

[2240] Max Menssen and Frank Schaarschmidt. Prediction intervals for all of M future observations based on linear random effects models. Statistica Neerlandica, 76(3):283–308, August 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12260.

Hunsberger:2022:RCI

[2241] Sally Hunsberger, Lori Long, Sarah E. Reese, Gloria H. Hong, Ian A. Myles, Christa S. Zerbe, Pleonchan Chetchotisakd, and Joanna H. Shih. Rank correlation inferences for clustered data with small sample size. Statistica Neerlandica, 76(3):309–330, August 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12261.

Etievant:2022:SLP

[2242] Lola Etiévant and Vivian Viallon. On some limitations of probabilistic models for dimension-reduction: Illustration in the case of probabilistic formulations of partial least squares. Statistica Neerlandica, 76(3):331– 346, August 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12262.

Wang:2022:AOL

[2243] Wei Wang, Linjian Li, Sheng Li, Fei Yin, Fang Liao, Tao Zhang, Xiaosong Li, Xiong Xiao, and Yue Ma. Average ordinary least squares-centered penalized regression: a more efficient way to address multicollinearity than ridge regression. Statistica Neerlandica, 76(3):347–368, August 2022. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12263.

Anonymous:2022:IId

[2244] Anonymous. Issue information. Statistica Neerlandica, 76(4):369–371, November 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12248.

Reinikainen:2022:BSS

[2245] Jaakko Reinikainen and Juha Karvanen. Bayesian subcohort selection for longitudinal covariate measurements in follow-up studies. Statistica Neerlandica, 76(4):372–390, November 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12264.

Larsson:2022:BCI

[2246] Rolf Larsson. Bartlett correction of an independence test in a multivariate Poisson model. Statistica Neerlandica, 76(4):391–417, November 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12265.

Wang:2022:OSM

[2247] Tianzhen Wang and Haixiang Zhang. Optimal subsampling for multiplicative regression with massive data. Statistica Neerlandica, 76(4):418– 449, November 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12266.

Gaunt:2022:BDT

[2248] Robert E. Gaunt. The basic distributional theory for the product of zero mean correlated normal random variables. Statistica Neerlandica, 76(4):450–470, November 2022. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12267.

Ristic:2022:ES

[2249] Miroslav Ristic, Marijtje van Duijn, and Nan van Geloven. Editorial statistics. Statistica Neerlandica, 76(4):471, November 2022. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12279.

Anonymous:2023:IIa

[2250] Anonymous. Issue information. Statistica Neerlandica, 77(1):1–3, February 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12269.

Chang:2023:TEC

[2251] Chih-Hao Chang, Kam-Fai Wong, and Wei-Yee Lim. Threshold estimation for continuous three-phase polynomial regression models with constant mean in the middle regime. Statistica Neerlandica, 77(1):4– 47, February 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12268.

Campisi:2023:ASF

[2252] Giovanni Campisi, Luca La Rocca, and Silvia Muzzioli. Assessing skewness in financial markets. Statistica Neerlandica, 77(1):48–70, February 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12273.

Zhou:2023:ABC

[2253] Jing Zhou and Gerda Claeskens. Automatic bias correction for testing in high-dimensional linear models. Statistica Neerlandica, 77(1):71– 98, February 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12274.

Fang:2023:USO

[2254] Longxiang Fang, Narayanaswamy Balakrishnan, Wenyu Huang, and Shuai Zhang. Usual stochastic ordering of the sample maxima from dependent distribution-free random variables. Statistica Neerlandica, 77(1): 99–112, February 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12275.

BenElouefi:2023:IPW

[2255] Rim Ben Elouefi and Foued Saâdaoui. Inverse-probability-weighted logrank test for stratified survival data with missing measurements. Statistica Neerlandica, 77(1):113–129, February 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12276.

Anonymous:2023:IIb

[2256] Anonymous. Issue information. Statistica Neerlandica, 77(2):131– 133, May 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12270.

Battauz:2023:TDC

[2257] Michela Battauz. Testing for differences in chain equating. Statistica Neerlandica, 77(2):134–145, May 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12277.

Calatayud:2023:PMC

[2258] Julia Calatayud, Marc Jornet, and Jorge Mateu. A phenomenological model for COVID-19 data taking into account neighboring-provinces effect and random noise. Statistica Neerlandica, 77(2):146–155, May 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12278.

Boamah-Addo:2023:DTZ

[2259] Kwame Boamah-Addo, Tomasz J. Kozubowski, and Anna K. Panorska. A discrete truncated Zipf distribution. Statistica Neerlandica, 77(2): 156–187, May 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12280.

Yang:2023:BIM

[2260] Kai Yang, Qingqing Zhang, Xinyang Yu, and Xiaogang Dong. Bayesian inference for a mixture double autoregressive model. Statistica Neerlandica, 77(2):188–207, May 2023. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12281.

Roy:2023:ILL

[2261] Soumya Roy and Biswabrata Pradhan. Inference for log-location-scale family of distributions under competing risks with progressive Type-I interval censored data. Statistica Neerlandica, 77(2):208–232, May 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12282.

Bose:2023:PES

[2262] Meghna Bose, Jean-François Angers, and Atanu Biswas. Prior effective sample size in phase II clinical trials with mixed binary and continuous responses. Statistica Neerlandica, 77(2):233–248, May 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12283.

Anonymous:2023:IIc

[2263] Anonymous. Issue information. Statistica Neerlandica, 77(3):249–251, August 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12271.

Wit:2023:E

[2264] Ernst C. Wit. Editorial. Statistica Neerlandica, 77(3):252–253, August 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12307.

Kanga:2023:APN

[2265] S.-H. Arnaud Kanga, Ouagnina Hili, Sophie Dabo-Niang, and Assi N’Guessan. Asymptotic properties of nonparametric quantile estimation with spatial dependency. Statistica Neerlandica, 77(3):254–283, August 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12284.

Zhu:2023:DSM

[2266] Xuwen Zhu and Xiang Zhang. Discretized skew-t mixture model for deconvoluting liquid chromatograph mass spectrometry data. Statistica Neerlandica, 77(3):284–303, August 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12285.

Wang:2023:ERE

[2267] Pei Wang, Erin L. Abner, Changrui Liu, David W. Fardo, Frederick A. Schmitt, Gregory A. Jicha, Linda J. Van Eldik, and Richard J. Kryscio. Estimating random effects in a finite Markov chain with absorbing states: Application to cognitive data. Statistica Neerlandica, 77(3): 304–321, August 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12286.

Kolassa:2023:IPL

[2268] John E. Kolassa and Juan Zhang. Inference in the presence of likelihood monotonicity for proportional hazards regression. Statistica Neerlandica, 77(3):322–339, August 2023. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12287.

Li:2023:TCR

[2269] Zhiming Li, Changxing Ma, and Keyi Mou. Testing the common risk difference of proportions for stratified uni- and bilateral correlated data. Statistica Neerlandica, 77(3):340–364, August 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12288.

Almeida:2023:BSM

[2270] Frederico M. Almeida, Vinícius D. Mayrink, and Enrico A. Colosimo. Bayesian solution to the monotone likelihood in the standard mixture cure model. Statistica Neerlandica, 77(3):365–390, August 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12289.

Borquez:2023:PTI

[2271] Ricardo Bórquez. A portmanteau test for the iid hypothesis. Statistica Neerlandica, 77(3):391–402, August 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12290.

Anonymous:2023:IId

[2272] Anonymous. Issue information. Statistica Neerlandica, 77(4):403–405, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12272.

Wit:2023:NAT

[2273] Ernst-Jan Camiel Wit. New article types in Statistica Neerlandica. Statistica Neerlandica, 77(4):406–407, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12324.

Cox:2023:PPV

[2274] Kyle Cox and Benjamin Kelcey. A partial posterior p value test for multilevel mediation. Statistica Neerlandica, 77(4):408–428, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12291.

Allison:2023:LL

[2275] James S. Allison, Bruno Ebner, and Marius Smuts. Logistic or not logistic? Statistica Neerlandica, 77(4):429–443, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12292.

Burger:2023:RME

[2276] Divan A. Burger, Sean van der Merwe, Emmanuel Lesaffre, Peter C. le Roux, and Morgan J. Raath-Krüger. A robust mixed-effects parametric quantile regression model for continuous proportions: Quantifying the constraints to vitality in cushion plants. Statistica Neerlandica, 77(4):444–470, November 2023. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12293.

Prasad:2023:EFM

[2277] E. Hari Prasad and N. Balakrishna. Estimating function method for nonnegative autoregressive models. Statistica Neerlandica, 77(4):471– 496, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12294.

Kundu:2023:SCL

[2278] Pradip Kundu, Amarjit Kundu, and Biplab Hawlader. Stochastic comparisons of largest claim amounts from heterogeneous portfolios. Statistica Neerlandica, 77(4):497–515, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12296.

Cruz:2023:ACE

[2279] Nelson Alirio Cruz, Luis Alberto López Pérez, and Oscar Orlando Melo. Analysis of cross-over experiments with count data in the presence of carry-over effects. Statistica Neerlandica, 77(4):516–542, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12295.

Pries:2023:OII

[2280] Joris Pries, Etienne van de Bijl, Jan Klein, Sandjai Bhulai, and Rob van der Mei. The optimal input-independent baseline for binary classification: The Dutch Draw. Statistica Neerlandica, 77(4):543–554, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12297.

Jang:2023:NCF

[2281] Yu-Hyeong Jang, Jun Zhao, Hyoung-Moon Kim, Kyusang Yu, Sunghoon Kwon, and SungHwan Kim. New closed-form efficient estimator for multivariate gamma distribution. Statistica Neerlandica, 77(4):555–572, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12299.

Micheloud:2023:ARS

[2282] Charlotte Micheloud, Fadoua Balabdaoui, and Leonhard Held. Assessing replicability with the sceptical p-value: Type-I error control and sample size planning. Statistica Neerlandica, 77(4):573–591, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12312.

vandenHeuvel:2023:ES

[2283] Edwin van den Heuvel, Veronica Vinciotti, and Ernst-Jan Camiel Wit. Editorial statistics. Statistica Neerlandica, 77(4):592, November 2023. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12325.

Anonymous:2024:IIa

[2284] Anonymous. Issue information. Statistica Neerlandica, 78(1):1–3, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12300.

Karlis:2024:MSI

[2285] Dimitris Karlis, Azmi Chutoo, Naushad Mamode Khan, and Vandna Jowaheer. The multilateral spatial integer-valued process of order 1. Statistica Neerlandica, 78(1):4–24, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12298.

Zhang:2024:LRM

[2286] Jun Zhang, Bingqing Lin, and Yan Zhou. Linear regression models with multiplicative distortions under new identifiability conditions. Statistica Neerlandica, 78(1):25–67, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12304.

Rayner:2024:OCB

[2287] J. C. W. Rayner and G. C. Livingston, Jr. Orthogonal contrasts for both balanced and unbalanced designs and both ordered and unordered treatments. Statistica Neerlandica, 78(1):68–78, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12305.

Sun:2024:TJR

[2288] Yucheng Sun. Testing for jumps with robust spot volatility estimators. Statistica Neerlandica, 78(1):79–104, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12306.

Singh:2024:POC

[2289] Kundan Singh, Amulya Kumar Mahto, and Yogesh Mani Tripathi. On partially observed competing risks model for Chen distribution under generalized progressive hybrid censoring. Statistica Neerlandica, 78 (1):105–135, February 2024. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12308.

Sen:2024:CSG

[2290] Abhibasu Sen and Karabi Dutta Choudhury. A case study of Gulf Securities Market in the last 20 years: a long short-term memory approach. Statistica Neerlandica, 78(1):136–166, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12309.

Isilar:2024:RLT

[2291] Melike Işılar and Y. Murat Bulut. Robust Liu-type estimator based on GM estimator. Statistica Neerlandica, 78(1):167–190, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12310.

Wang:2024:ASC

[2292] Antai Wang, Ziyan Guo, Yilong Zhang, and Jihua Wu. The analysis of semi-competing risks data using Archimedean copula models. Statistica Neerlandica, 78(1):191–207, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.

wiley.com/doi/epdf/10.1111/stan.12311.

Li:2024:PAM

[2293] Sheng Li, Wei Wang, Menghan Yao, Junyu Wang, Qianqian Du, Xuelin Li, Xinyue Tian, Jing Zeng, Ying Deng, Tao Zhang, Fei Yin, and Yue Ma. Poisson average maximum likelihood-centered penalized estimator: a new estimator to better address multicollinearity in Poisson regression. Statistica Neerlandica, 78(1):208–227, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12313.

Gzyl:2024:JPU

[2294] Henryk Gzyl and Silvia Mayoral. Joint probabilities under expected value constraints, transportation problems, maximum entropy in the mean. Statistica Neerlandica, 78(1):228–243, February 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12314.

Nguyen-Trang:2024:EAC

[2295] Thao Nguyen-Trang, Tai Vo-Van, and Ha Che-Ngoc. An efficient automatic clustering algorithm for probability density functions and its applications in surface material classification. Statistica Neerlandica, 78 (1):244–260, February 2024. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12315.

Anonymous:2024:IIb

[2296] Anonymous. Issue information. Statistica Neerlandica, 78(2):261– 263, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12301.

Iwashita:2024:GTS

[2297] Toshiya Iwashita and Bernhard Klar. A gamma tail statistic and its asymptotics. Statistica Neerlandica, 78(2):264–280, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12316.

Chen:2024:CRR

[2298] Xinyuan Chen, Denise Esserman, and Fan Li. Competing risks regression for clustered survival data via the marginal additive subdistribution hazards model. Statistica Neerlandica, 78(2):281–301, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12317.

Aguirre-Hamilton:2024:FRR

[2299] Christopher Aguirre-Hamilton, Stephen A. Sedory, and Sarjinder Singh. Franklin’s randomized response model with correlated scrambled variables. Statistica Neerlandica, 78(2):302–309, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12318.

Wang:2024:IEA

[2300] Jun Wang and Yahe Yu. Improved estimation of average treatment effects under covariate-adaptive randomization methods. Statistica Neerlandica, 78(2):310–333, May 2024. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12319.

Andres:2024:YCM

[2301] Antonio Martín Andrés, María Álvarez Hernández, and Francisco Gayá Moreno. The Yates, Conover, and Mantel statistics in 2 × 2 tables revisited (and extended). Statistica Neerlandica, 78(2):334–356, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12320.

Abraham:2024:IPD

[2302] Christophe Abraham. An informative prior distribution on functions with application to functional regression. Statistica Neerlandica, 78(2): 357–373, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12322.

Zheng:2024:SRC

[2303] Siming Zheng, Alan T. K. Wan, and Yong Zhou. Semiparametric recovery of central dimension reduction space with nonignorable nonresponse. Statistica Neerlandica, 78(2):374–396, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12321.

Cai:2024:TCI

[2304] Zongwu Cai, Ying Fang, Ming Lin, and Shengfang Tang. Testing conditional independence in casual inference for time series data. Statistica Neerlandica, 78(2):397–426, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12323.

Genest:2024:ACN

[2305] Christian Genest and Frédéric Ouimet. Asymptotic comparison of negative multinomial and multivariate normal experiments. Statistica Neerlandica, 78(2):427–440, May 2024. CODEN ???? ISSN 0039-0402 (print),

1467-9574 (electronic). URL https://onlinelibrary.wiley.com/doi/
epdf/10.1111/stan.12328.

Ozdemir:2024:FPM

[2306] Ozancan Ozdemir and Ceylan Yozgatligil. Forecasting performance of machine learning, time series, and hybrid methods for low- and highfrequency time series. Statistica Neerlandica, 78(2):441–474, May 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12326.

Anonymous:2024:IIc

[2307] Anonymous. Issue information. Statistica Neerlandica, 78(3):475–477, August 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12302.

Kittani:2024:GVP

[2308] Hilmi Fadel Kittani. Generalized k-variate proportional hazard function for censored survival data. Statistica Neerlandica, 78(3):478–484, August 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12327.

Yang:2024:CBT

[2309] Weichao Yang, Xu Guo, Niwen Zhou, and Changliang Zou. Connections between two classes of estimators for single-index models. Statistica Neerlandica, 78(3):485–490, August 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12329.

Spyropoulou:2024:SPI

[2310] Maria-Zafeiria Spyropoulou and James Bentham. Scaling priors for intrinsic Gaussian Markov random fields applied to blood pressure data. Statistica Neerlandica, 78(3):491–504, August 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://

onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12330.

Buric:2024:AAE

[2311] Tomislav Burić and Lenka Mihoković. Asymptotic approximations of expectations of power means. Statistica Neerlandica, 78(3):505–522, August 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12331.

Ghosh:2024:MLL

[2312] Sayan Ghosh and P. Vellaisamy. Marginal log-linear parameters and their collapsibility for categorical data. Statistica Neerlandica, 78(3): 523–543, August 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574

(electronic). URL https://onlinelibrary.wiley.com/doi/epdf/10.
1111/stan.12332.

Kabaila:2024:CSC

[2313] Paul Kabaila and A. H. Welsh. The concept of sufficiency in conditional frequentist inference. Statistica Neerlandica, 78(3):544–562, August 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL

https://onlinelibrary.wiley.com/doi/epdf/10.1111/stan.12333.

Jones:2024:DCT

[2314] M. C. Jones. Duals of convolution thinned relationships. Statistica Neerlandica, 78(3):563–568, August 2024. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https://onlinelibrary.wiley.

com/doi/epdf/10.1111/stan.12337.